A Complete Bibliography of Publications in the Journal of Computational Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

29 May 2024 Version 1.04

Title word cross-reference

(2 + 1) [1297]. / [976]. 1 [409, 1751]. 1000× [997]. 2 [1703, 1818, 409, 1773, 927, 373, 178, 1505, 1661]. 3 [1254, 664, 663, 529, 1588, 409, 1017, 1061, 385, 231, 1326, 1693, 369, 1139, 1547, 1099, 378, 672, 1106, 1129, 1126, 373, 566, 150, 1690, 1437, 481, 1275, 1661]. 4 [1661]. 5 [1458]. ^{+} [1278]. ^{T} [1931]. _{2} [500, 412, 1123, 889, 918, 498, 1409]. _{3} [78, 918]. _{4} [498]. _{6} [498]. α [548]. β [29, 918]. C [258]. H [1067, 1038]. H_{∈} [1404]. H_{∞} [280]. d [539, 692]. F_{order} [948]. h [1866]. I [487]. K [1391, 682, 1002]. L^{2} [1390]. LU [1038]. N [513]. P [249, 1170, 9]. ϕ [1874]. π [406]. Q [1477, 617]. QR [216]. r [412]. Re_{τ} = 5300 [1408]. BaH^{+} [1291]. SrH^{+} [1291]. σ [290]. V [1729]. Z [409].

-2 [409]. -5 [409]. -benzofuran- [409]. -body [513]. -Centroids [412]. -cycle [1729]. -D [481, 1139, 1099]. -fractional [1874]. -hole [290]. -In [918]. -learning [1477]. -link [1458]. -locus [378]. -LU [1067]. -means [1391]. -MPs [539, 692]. -one [409]. -reachable [682]. -refined [1866]. -skeletons [29]. -stable [548]. -tridiagonal [1002].

/s/ [128].

1 [401]. 101938 [1859]. 19

[1563, 1463, 1670, 1731, 1795, 1827, 1506].

2 [307]. 2010 [32, 40, 22, 12]. 2011 [90, 52]. 2012 [99]. 2013 [211, 286]. 2014 [410, 229, 302, 285, 329, 322]. 2015 [330, 339, 346, 421, 384]. 2016 [457, 468, 526, 502, 742]. 2017 [635, 596, 611, 711, 681]. 2018 [735, 874, 774, 821, 963, 916]. 2019 [1109, 1101, 1035, 1001, 980, 1057, 1019, 1046, 1086, 1234]. 2020 [1148, 1125, 1112, 1167, 1136, 1157, 1228, 1191, 1338]. 2021 [1290, 1258, 1248, 1346, 1268, 1417, 1407, 1388]. 2022 [1479, 1439, 1425, 1555, 1457, 1516, 1657, 1636, 1597]. 2023 [1722, 1837, 1679, 1757, 1777, 1706, 1744, 1816, 1815]. 2024 [1891, 1858, 1914, 1938, 1879]. 24th [553]. 29 [1364].

3 [209]. 3-dihydro- [409]. 32 [1715]. 33258 [402]. 3D [1299].

4-nitrophenyl [409]. 4.0 [1526].

5 [344]. 5.0 [1762]. 51 [1335]. 5G [857].

6-dimethyl- [409]. 67 [1859].

74 [500].

8 [889].

P̄ANDA [418].

AA [834]. AA-stacked [834]. Ab-initio [1581, 1376]. abatement [614]. abdominal [809]. ABFIA [1500]. Abnormal [880]. Above [1427]. Above-ground [1427]. ABS [598, 1078]. ABS-MindBurnout [1078]. ABS-MindHeart [598]. abstraction [1031]. Academia [797]. academic [1256, 1305]. Acari [701]. accelerate [580, 1144]. Accelerated [883, 1091, 448, 1911, 1261, 1715, 352, 1747, 1900, 1592]. Accelerating [1263, 845, 882, 1032, 1547, 129, 1903, 1447, 1007, 411]. Acceleration [602, 1507, 1400]. accelerators [997]. accelerometer [1138]. access [1164, 441, 1460]. accessible [1397]. accommodating [1385]. according [66]. accuracy [260, 882, 380, 1916, 1129, 1358, 1523, 1852, 1050]. accurate [825, 111, 557, 705, 1619]. Accurately [1163]. achieving [763]. acid [1245, 98]. acoustic [642, 1514, 1129, 128]. ACQC [1634]. acquired [1286]. across [10, 1336]. ACT [189]. ACT-R [189]. action [929]. actionable [1843]. active [430, 1037, 1441, 1603]. activity [841, 1149, 1186, 1107]. activity-based [1107]. Adams [513]. Adaptation [506, 1050, 1503, 969, 349, 974, 1364, 95, 178]. adaption [1301]. Adaptive [697, 294, 469, 1296, 1383, 1692, 958, 1648, 1501, 231, 380, 795, 177, 1233, 779, 1745, 422, 241, 883, 1759, 512, 1475, 987, 1664, 373, 1782, 1448, 281, 1360, 335, 9, 433, 438, 1549, 1632, 178, 1525, 1846, 1905, 766, 173, 1062, 982, 1237]. adaptivity [175, 373, 15, 93]. ADAS [1562]. ADCS [272]. adder [687]. adders [504]. additional [1673]. Addressing [1881]. adhesion [445]. ADI [1091]. adiabatic [927]. Adjoint [1192, 232, 754, 1666]. adjoint-based [232]. Adjoints} [1044]. adjuvant [1497]. ADM [1007]. ADME [401]. ADME-T [401]. admissibility [1908]. adolescents [646]. Adomian [1007]. Adsorption [1245, 408, 500, 1131, 417, 1844]. adsorption-desorption [1131]. Advanced [30, 123, 116, 393, 780, 1272]. Advances [100, 391, 1075, 115, 573, 720]. advancing [1803]. advection [1585, 1864, 1311, 141, 1628, 1178, 1569]. advection-diffusion [1628]. advection-dominated [1864]. adversarial [1936]. advertisement [666]. Aerial [1054, 1865, 1435, 974, 1364]. Aerodynamic [389, 1199, 1856]. AFDX [911]. affecting [290]. affine [591]. affinity [891]. AFM [1803]. after [1901]. against [401]. age [616, 415]. age-related [616]. Agent [1507, 640, 433, 886, 968, 1107, 1812, 969, 1240, 1483, 598, 1078, 1617, 177, 775, 1143, 673, 624, 437, 1175, 534, 637, 920, 1043, 268, 1283, 1885, 86, 1015, 1059, 1264, 1716, 1613, 546, 173, 1724, 1521]. Agent- [1107]. Agent-based [1507, 640, 433, 886, 968, 1812, 969, 1483, 1078, 1617, 775, 1143, 624, 437, 1175, 534, 637, 1885, 1015, 1613, 546, 1724, 1521]. agents [401, 678, 407, 672, 432]. aggregate [14]. aggregation [680, 1132, 219, 1855, 686, 863]. aggregation-breakage [1855]. agricultural [900]. agroecosystem [204]. AHP [605]. aided [1813]. air [276, 1919, 235]. air-conditioning [235]. air-water [276]. aircraft [1520, 1165]. airflow [1366, 1900]. airfoil [1856]. airfoils [17]. airway [1849]. AIS [1868]. ALE [1272]. ALE-mixed [1272]. aleatory [716]. AlexNet [984]. algebra [416, 787, 1644]. Algebraic [574, 1518, 152, 148, 280, 1087]. alginate [1861]. alginate-based [1861]. AlgoPy [198]. Algorithm [1733, 722, 27, 806, 1436, 1848, 818, 1428, 1703, 1430, 714, 887, 1252, 312, 182, 1394, 1254, 310, 778, 854, 577, 702, 1513, 1683, 1501, 650, 1122, 602, 1277, 1284, 1500, 1391, 1669, 1103, 1625, 898, 247, 1654, 1825, 959, 1799, 801, 1405, 1553, 351, 1599, 1605, 427, 348, 900, 779, 803, 1113, 600, 729, 1745, 772, 354, 241, 1120, 597, 206, 243, 507, 461, 1931, 223, 987, 1615, 770, 653, 655, 246, 1465, 1542, 1130, 281, 150, 657, 205, 1413, 1360, 86, 1371, 1393, 1587, 731, 232, 613, 1802, 785, 1418, 660, 252, 1647, 1902, 1840, 1335, 1287, 218]. algorithm [1455, 1351, 1631, 1626, 1002, 1282, 254, 82, 366, 1170, 1348, 1675, 1009, 436, 869, 1144, 1062, 649, 853, 424, 1835, 581, 1822, 138, 1004, 1629, 1232, 253]. Algorithmic [1466, 198, 1192, 1676, 1663, 753]. Algorithmic-gradient [1466]. Algorithms [391, 145, 212, 477, 1910, 75, 727, 1327, 986, 184, 1197, 1451, 777, 1139, 1095, 925, 1030, 543, 1347, 1450, 1535, 1274, 1361, 651, 433, 1690, 1539, 578, 1427, 582, 690, 881, 622, 982, 1213]. algorithms-based [925]. alighting [1204]. aligners [786]. Alignment [1193, 1895]. alkanes [1245]. all-hexahedral [1905]. Allen [1877, 1808, 1888]. alliances [1769]. allocating [1005]. Allocation [27, 1066, 1435, 587, 900, 1800, 1334, 1627, 1280, 86, 343, 1188, 1246, 860, 606, 366]. allocations [861]. alloys [840]. along [1044]. alternating [1719, 1205]. alternative [267]. altitude [1054, 1822]. Alya [479]. American [632, 1518]. AMG [87]. ammonia [1863]. among [72, 263]. amounts [1095]. Ampere [1715]. analogue [1211]. analogues [402]. analogy [568, 43]. analyse [1559]. analyses [1754, 1552, 1844]. Analysing [207, 1073]. Analysis [841, 1528, 1662, 1841, 70, 1560, 1155, 1577, 815, 1830, 1667, 941, 961, 654, 102, 1442, 1459, 709, 18, 1778, 905, 1775, 788, 823, 875, 431, 318, 197, 1366, 716, 1054, 1426, 1767, 1176, 958, 396, 1332, 135, 1022, 1131, 28, 1717, 1908, 704, 1620, 1915, 1571, 1746, 698, 1318, 272, 1833, 1372, 1127, 279, 544, 369, 1660, 1137, 1605, 811, 738, 1251, 378, 718, 757, 750, 672, 1464, 1532, 1751, 1672, 270, 508, 1016, 1627, 658, 387, 1280, 1540, 797, 585, 896, 1821, 323, 1851, 550, 269, 1076, 126, 1831, 1918, 1285, 218, 578, 1614, 1854, 1783]. analysis [1052, 1545, 1568, 1081, 1134, 191, 1808, 1810, 303, 1521, 1497, 1720, 533, 816, 952, 1395, 35, 361, 712, 494, 459]. analytic [1712, 333]. Analytical [331, 1899, 1704, 1580]. Analytically [50]. analytics [907, 865]. analyze [236]. Analyzing [1910, 440, 1873, 1894, 637, 1283]. Anarchy [1466]. aneurysm [878]. aneurysms [92, 1106]. ANFIS [1237]. angle [565, 1849]. angular [877]. anisotropic [1818, 1503, 1674, 1020, 1201, 15, 93, 178, 1601]. ANN [429, 1923]. annealing [1688, 582, 1835]. announcements [636]. annual [673]. annular [159]. anode [1604]. anodization [453]. anomalies [726]. anomalous [1337, 792]. Anomaly [788, 627, 1795, 1787]. Anomaly-based [788]. Ansys [1834]. Ansys-Fluent [1834]. Ant [666, 717, 790, 1927, 671, 1435, 777, 779, 1005, 343, 1141]. antenna [1828, 1902]. Antennae [1458]. anti [1673, 407, 1827, 1203, 350]. anti-diabetes [1203]. anti-inflammatory [407]. anti-periodic [350]. anti-predator [1673]. anti-viral [1827]. anticipation [1801]. Anticipative [628]. antipodally [88]. Antlion [975]. aorta [745]. aortic [745, 1301]. Apc [748]. Apollonius [1634]. Application [564, 1713, 1379, 716, 1828, 1284, 746, 1197, 1451, 1712, 177, 74, 1518, 1329, 548, 651, 1539, 126, 1647, 879, 204, 713, 1144, 1066, 490, 668, 494, 778, 641, 28, 201, 1512, 1799, 1605, 970, 1664, 901, 341, 450, 1607, 281, 1533, 1310, 924, 580, 1307, 1840, 876, 582, 353, 1348, 830, 1062, 132]. application-based [1607]. Application-driven [746]. application-specific [1310]. Applications [239, 1795, 672, 927, 1110, 1452, 684, 444, 1581, 1733, 1051, 721, 725, 1342, 1220, 60, 760, 1221, 1184, 1208, 1048, 238, 1298, 87, 579, 56, 1493, 36, 277, 1212, 988, 1492, 236, 557, 1736, 585, 762, 103, 159, 1339, 1180, 856, 227, 859]. Applied [391, 1252, 1104, 1155, 1301, 1590, 623, 305, 533, 1505]. Applying [233, 1617, 605, 1368]. approach [1436, 34, 993, 665, 1003, 552, 1538, 819, 809, 741, 1172, 942, 1008, 1269, 310, 1265, 716, 42, 1462, 503, 1466, 1717, 1240, 1484, 306, 392, 1434, 1298, 478, 187, 1787, 576, 523, 153, 1041, 1321, 79, 1421, 1410, 1372, 1470, 133, 992, 728, 1378, 402, 381, 811, 799, 1518, 1206, 729, 262, 1800, 825, 951, 1084, 1672, 1231, 1527, 112, 1336, 1453, 1473, 495, 618, 341, 764, 1748, 253, 86, 1058, 213, 1210, 1548, 1886, 532, 972, 1083, 492, 1614, 1525, 190, 37, 1168, 1568, 1790, 594, 414, 816, 712]. Approaches [1728, 1528, 1342, 1220, 264, 1806, 1132, 1590, 314, 585, 904, 1697, 1222]. appropriate [1640]. approved [407]. approximability [1361]. approximate [153, 1041, 1609, 1142, 863]. Approximating [760, 842]. approximation [1910, 169, 175, 1098, 1024, 74, 1916, 612, 1450, 1903, 1925]. approximations [1930]. April [1035, 1148, 1290, 1479, 1722, 1891]. AQP [863]. Arabic [872, 905]. arbitrarily [1696]. arbitrary [776, 140, 1077]. arc [522, 1025]. Architecture [1212, 1904, 76, 706, 1344, 1470, 181, 509, 213, 586, 643]. architecture-proof [586]. architectures [18, 703, 234, 56, 693, 45, 592, 1371, 1262, 1459, 846, 1839, 411, 622]. archiving [1068]. area [299, 110, 697, 1489]. areas [1882, 119]. argon [1401]. argon-neon [1401]. arising [1149, 1337]. arithmetic [928, 102, 1626, 1093]. armband [1286]. ARMD [616]. aromatic [406]. Array [249, 261, 372, 1902]. arrays [560]. arrival [700]. arsenic [1574]. art [840]. arterial [1436, 1525]. arteries [368]. artery [1750, 1382]. articulated [193]. Artificial [656, 1243, 315, 1500, 898, 1825, 1249, 348, 245, 730, 511, 982, 701, 429]. ascending [745]. Asian [928]. ASL [1436]. asleep [1443]. aspect [905]. aspect-based [905]. aspects [188, 1754]. assay [1314]. assembling [1392]. assembly [354, 205]. assertion [238]. Assessing [601, 1730, 1918, 1900, 10, 1523]. Assessment [402, 1574, 1877, 375, 363, 1014, 582, 919]. asset [758, 1586]. asset-liability [758]. Assigning [838]. assignment [855, 1180, 766]. assimilated [639]. Assimilation [1431, 1824, 1183, 1341, 1402, 744, 1870, 1667, 1649, 1347, 1782, 1300, 1354, 1831, 879, 1296, 1303, 1327]. assistant [849]. assistants [1281]. Assisted [1772, 1375, 1451, 1723, 1780, 461, 1613, 1803, 458]. associated [1591]. associative [350]. assortativity [137]. assumption [291].

Assuring [298]. asteroid [1615]. asymmetric [787, 1906]. asymmetrical [976]. Asynchronous [1736, 768, 318, 832, 1299]. atmospheric [1294, 357]. atomistic [76, 45, 568]. atoms [278]. atria [929]. atrial [1288]. attack [819, 723, 1602, 1781]. attacks [442]. Attempts [1728, 296]. attention [1646]. attitude [1113, 272]. Attractive [1305]. attribute [1881, 293, 790, 1386]. attributes [1606]. auctions [533]. augment [1563, 1314]. augmentation [1375, 996]. augmented [1619, 892]. August [32]. Australian [1566]. Auto [416, 1093, 1270, 1266, 481]. Auto-tuning [416, 1093, 1270, 1266, 481]. autocorrelation [1889]. Autoencoder [1904]. Automata [42, 769, 444, 445, 504, 47, 1841, 453, 318, 447, 71, 687, 319, 452, 767, 1512, 676, 1032, 1326, 1928, 772, 828, 920, 764, 450, 677, 449, 762, 514, 1384, 766, 365, 1224, 110, 515, 763, 1488, 1028, 1441, 768]. automata-based [676, 772, 762]. Automated [616, 1023, 1704, 1508, 1821, 1852, 1779, 381, 592, 1589, 973, 652, 1392, 713, 878, 1526]. Automatic [1255, 664, 650, 1377, 618, 113, 1526, 125, 1259, 1514, 627, 584, 601, 1410, 656]. automating [1924]. automation [503]. automaton [225, 826, 765]. AutoMF [1827]. autonomic [706]. autonomous [1617, 1331, 1894, 1707]. AutoPas [1270, 1266]. autotuning [1419]. AUVs [1082]. availability [1070]. average [412]. averaged [1573]. averaging [667]. avian [1418]. AVIDA [1719]. avoiding [1889]. awake [1443]. aware [1594, 221, 936, 1599, 763, 183, 1343, 214, 58, 606, 933, 861]. Axial [1395]. Axisymmetric [1201, 554].

B [434, 141, 1759, 1616]. B-spline [434, 1759, 1616]. B-splines [141]. back [835, 1350, 961, 505]. backbone [1399]. backhaul [1188]. backward [576]. backward/forward [576]. bacterial [250, 657, 1637]. Bagging [895]. BaH [1278]. balance [94, 1822]. Balanced [255]. balancer [676]. balancing [736, 1496, 832, 324, 93, 1304]. ballet [191]. ballistic [1340]. BAM [350, 1531]. band [1403]. band-limited [1403]. banking [906]. bar [1703]. Barcelona [1919]. Bare [1232]. Bare-Bones [1232]. barriers [708]. base [796]. Based [1713, 806, 1906, 1565, 1232, 1455, 1243, 1472, 1581, 1634, 504, 1809, 1695, 905, 788, 736, 435, 1454, 680, 809, 110, 1128, 741, 1862, 942, 1342, 668, 666, 662, 1008, 1812, 703, 669, 38, 1727, 1826, 1215, 19, 358, 778, 1040, 958, 969, 466, 1670, 1702, 1738, 1022, 401, 1895, 1240, 702, 1591, 1132, 1327, 374, 148, 1823, 1483, 1683, 717, 975, 1499, 1561, 650, 1434, 116, 1102, 1460, 1004, 1398, 428, 1500, 1669, 1507, 394, 1412, 1861, 523, 1625, 706, 1061, 1638, 184, 1789, 219, 923, 1680, 1072, 640, 598, 1078, 676, 56, 1617, 189, 293]. based [1726, 177, 1629, 1378, 313, 775, 1415, 1564, 1553, 1365, 226, 351, 1331, 1294, 1359, 1010, 921, 748, 1143, 146, 209, 440, 120, 1099, 1049, 248, 36, 912, 738, 803, 600, 1309, 718, 1161, 772, 1419, 925, 906, 1030, 297, 829, 624, 943, 422, 1178, 437, 562, 527, 607, 909, 1544, 1420, 1640, 460, 1773, 1930, 17, 1175, 1536, 543, 833, 1781, 1205, 1541, 534, 1504, 1684, 1859, 507, 637, 1006, 1646, 1686, 1276, 984, 1347, 223, 508, 1615, 1114, 920, 1043, 990, 268, 1467, 603, 1476, 1473, 771, 618, 341, 1274, 1502, 1624, 686, 450, 1283, 1465, 1492, 85, 1130]. based [1607, 1389, 1448, 1885, 176, 848, 1097, 1300, 657, 236, 688, 998, 651, 1573, 244, 1396, 433, 870, 1015, 1358, 1058, 84, 208, 1437, 232, 762, 397, 1138, 269, 924, 726, 967, 1088, 864, 1548, 785, 134, 1716, 1658, 660, 894, 1083, 904, 1307, 514, 796, 1519, 1335, 1287, 999, 1811, 1526, 1014, 1868, 1613, 555, 1631, 782, 491, 1633, 1905, 1304, 1374, 1907, 546, 732, 606, 694, 1792, 1081, 1170, 1253, 1808, 1675, 1813, 997, 1724, 1517, 869, 1521, 830, 911, 1087, 1732, 1766, 1934, 1586, 1497, 1720, 1579, 913, 649, 880, 891, 868, 871, 867, 952, 954, 1395]. based [1929, 1807, 1275, 1701, 332, 908, 948, 1791, 1498, 886, 968, 1107, 712, 325, 399]. basis [978, 1567, 1301, 1006, 1590, 907, 756, 224, 517, 550, 1401, 305, 1616]. Bat [722, 247, 987, 1675]. batch [621, 845]. Batched [844]. battery [1604, 1821, 1658]. Bayes [892]. Bayesian [895, 936, 1141, 537, 760, 1461, 1158, 998, 1241, 1226, 1455]. BBDTC [633]. be [1507]. beads [136]. beam [1805]. bearings [1249]. bed [294]. bee [315, 1500, 898, 246, 245, 982]. Beetle [1458]. behavior [966, 466, 780, 1880, 647, 1333, 1326, 1415, 811, 1200, 939, 1780, 637, 1214, 1028, 1138, 317, 1525, 798, 1021, 365, 976, 465]. behaviors [938, 120, 173]. behaviour [625, 1673, 271, 920, 1488, 218, 886]. behaviours [1107]. Beijing [968, 1107]. BEK [1775]. Bell [1887]. Bénard [1201]. benchmark [557]. Benchmarking [978, 234]. bend [231]. benefit [914, 1280]. benthic [673]. benzofuran [409]. berry [1252]. Best [832, 485, 778, 653]. beta [47, 424]. beta-skeleton [47]. Better [1748, 1550, 534, 1134]. between [1865, 1560, 44, 1238, 486, 1536, 1475, 1016, 1173, 258, 1025, 791, 976, 118]. Beyond [62, 1211, 694]. Bézier [1434, 806, 1308]. BFO [250]. BFS [1080]. BGK [1202]. bi [1483, 845, 1774, 913]. bi-diagonalization [845]. bi-level [1483, 1774, 913]. bias [427]. Biased [261, 834, 949]. biclusters [794]. bidirectional [350]. bidomain [202, 1503]. bifurcation [977, 1540, 1382, 1531]. Big [1463, 633, 839, 1501, 428, 959, 909, 837, 865, 861, 862, 864, 1196]. Bikmeans [948]. bilayer [834]. billets [1253]. bin [612]. binary [1848, 1121, 668, 1254, 1889, 803]. binding [402, 1392]. Bio [239, 648, 60, 904]. Bio-Inspired [239, 904]. bio-interactome [60]. BIOCOMP [257]. bioconvection [1892]. bioinformatics [490, 1342]. Bioinspired [727]. biological [930, 1895, 593, 261, 240, 891]. biologically [1640]. biomarkers [1083, 1642]. biomass [1427]. biomedical [127, 428, 1037, 633, 1783, 633]. biometric [814]. biophysical [1921]. biosensor [1365]. Biped [1004, 1458, 309]. bird [1062]. bit [1096, 599]. bit-flip [1096]. Bitcoin [1766]. black [1727, 224, 517, 1613, 1306, 528]. black-box [224, 517]. black-oil [1727]. blade [1537]. blasting [1901]. Block [1362, 1728, 34, 145, 662, 434, 1625, 1622, 1781, 324, 1039, 896]. block-diagonal [434]. Block-enhanced [1362]. block-tridiagonal [145]. blog [871]. blogs [72]. blood [736, 381, 1750, 1532, 1624, 1771, 1548, 1382]. Bloom [134]. Blue [419]. bluff [1273]. bluff-and-fix [1273]. BMMFOA [803]. board [601, 11, 21, 31, 39, 51, 64, 80, 89, 91, 114, 121, 142, 157, 170, 179, 194, 210, 222, 228, 256, 284, 301, 321, 328, 338, 345, 355, 383, 420, 456, 467, 476, 488, 501, 525, 588, 589, 590, 610, 634, 679, 710, 733, 734, 773, 820, 873, 915, 962, 979, 1000, 1018, 1034, 1045, 1056, 1085, 1100, 1108, 1111, 1124, 1135, 1147, 1156, 1166, 1190, 1209, 1227, 1247, 1257, 1267, 1289, 1325, 1345, 1387, 1406, 1416, 1424, 1438, 1456, 1478, 1515, 1554, 1596, 1635, 1656, 1836, 1678]. Board [1705, 1721, 1743, 1756, 1776, 1814, 1857, 1878, 1890, 1913, 1937]. boarding [1204]. body [299, 1751, 270, 136, 513, 1618]. Boltzmann [563, 572, 359, 505, 463, 564, 1128, 565, 1199, 1259, 1255, 1752, 1297, 1116, 1194, 306, 1311, 1201, 56, 1363, 207, 313, 1711, 1178, 562, 961, 1195, 566, 1179, 1832, 1542, 557, 136, 554, 571, 1401, 1639, 1036, 1623, 567, 922, 411, 1380, 558, 1312, 1236, 1592, 1530, 1643, 1357]. Boltzmann-multiparticle [1179]. bond [396, 759, 273]. bonding [78]. Bone [667, 1933]. bones [1023, 1232]. book [761].

Boolean [259, 836]. boost [1050]. Boosting [472, 1695]. borer [1252]. Born [385]. borne [812]. Boston [524, 386]. bot [700]. both [314]. bottleneck [320]. bottlenecks [1073]. Bottom [1924]. Bottom-up [1924]. Bounce [569, 1350, 961, 505, 283]. bounce-back [1350, 961]. bound [741, 1751, 1618]. boundaries [505]. Boundary [1534, 1775, 463, 1171, 719, 1828, 306, 1620, 385, 1363, 1817, 369, 784, 1356, 934, 1711, 1350, 568, 1930, 505, 1755, 782, 1639, 1643, 1505, 642, 411]. boundary-lattice [463]. boundary-penalized [1620]. bounded [164]. Bounding [185]. Boussinesq [159]. Boussinesq-type [159]. box [224, 517, 1613]. Brain [1212, 746, 155, 795, 1149, 129, 984, 837, 996, 1720]. brain-computer [1720]. branch [741, 1465]. Branch-and-Fix [1465]. Brazilian [1172]. breakage [1629, 1855]. breakdown [144]. breakdowns [296]. breakup [1540]. breast [1461, 1730, 1589, 1497, 880]. bridge [1723, 633, 1117]. Bridging [319, 631, 840, 883, 833]. bright [609]. Brinkman [1421]. broad [1672]. broadcast [297]. Brownian [417, 1788, 1236]. Brusselator [1540]. BSC [485]. BSense [1461]. bubble [1785]. bubble-sort [1785]. Buckley [160]. budget [721, 233, 1011]. Buffered [624]. build [248]. build-or-buy [248]. Building [1317, 971, 199, 788]. built [1207]. Bulk [1736]. Burgers’ [1549, 139]. burnout [1078]. Burr [1411]. burst [740]. burst-switched [740]. bursting [947]. business [792]. butenolide [401]. butterfly [731]. butterfly-inspired [731]. buy [248].

C [1263, 1004, 498, 967]. C-means [967]. CA [454, 1716]. CA-based [1716]. cache [1115, 1151, 1043]. cache-clustering [1151]. cache-partitioning [1151]. cadaveric [1460]. cadaveric-based [1460]. Cahn [1847, 1877, 1808, 1888]. calciner [1863]. calculated [1294, 283]. calculating [412]. Calculation [1715, 1298, 1685, 1691, 580, 1519]. calculations [1278, 288, 1250, 78, 1447, 622, 1291]. Calibration [1358, 1120, 1006, 998, 208, 754, 332, 361]. Call [30, 724]. camera [1006]. Can [403, 1507, 940]. Canadian [370]. Cancer [1220, 1461, 1126, 1360, 1730, 1429, 1838, 1876, 1497, 1572]. cancerisation [1143]. cancerous [1370]. cantilever [1745]. capabilities [1529, 563]. capacitance [1422]. capacity [1469, 1543, 1052]. capillary [1727]. capture [560]. Capturing [1079, 1864]. Caputo [1872]. carbon [13, 496, 109]. carcinogenesis [748]. carcinoma [654, 656, 618, 1642, 658]. card [895, 377]. cardiac [202, 1503, 1761, 1104, 1020, 744, 1318, 1641, 1314, 1358, 15, 93]. CardioRisk [363]. cardiovascular [743, 363, 1831, 1918, 1272]. Care [30, 123, 1798]. Carlo [374, 1712, 155, 1044, 498, 218, 622, 283, 1275]. carotid [1382]. carry [687]. Cartesian [161]. cascade [785]. cascades [945]. cascading [1378]. Case [1908, 127, 968, 978, 625, 72, 1183, 1463, 1288, 815, 1769, 352, 221, 739, 608, 403, 615, 1627, 240, 1607, 932, 357, 487, 1384, 796, 1526, 866, 199, 791, 235, 197]. cases [1763, 451, 557]. Casson [1893, 1854]. caste [1792]. caste-based [1792]. casting [1253]. catalogs [1081]. catalyst [561]. catalysts [1650]. catalytic [454]. categorical [1873, 1924]. categories [1336]. category [936, 82]. category-aware [936]. Cattaneo [1861]. Cauchy [385, 776]. Causal [851]. causes [1793]. causing [831]. cavity [1366, 1537, 1660, 1239, 1260]. CAVS [404]. CBD [866]. CDS [1381]. CE [965, 1320]. Cell [296, 1459, 132, 736, 151, 560, 380, 1326, 1143, 1488, 1358, 1429, 1771, 438, 1384, 311, 876, 1876, 1770, 1282, 1642, 902]. cell-based [736]. cells [504, 1198, 1017, 1441, 1188]. Cellular [453, 515, 1441, 768, 444, 445, 1224, 504, 1841, 318, 447, 71, 687, 319, 452, 767, 1512, 676, 1032, 1326, 83, 678, 748, 772, 451, 828, 920, 450, 677, 225, 449, 826, 514, 1384, 1876, 326, 1134, 365, 110, 1488, 1028]. cement [1863]. center [763, 860]. centered [1601]. Central [160, 1727, 1592, 166]. centrality [1155, 1591, 1746, 544, 718, 842, 1519, 1545, 1668, 1519]. centralized [795]. centre [740]. Centrifugal [425]. Centroid [1561]. centroids [49, 412]. CEO [636]. cerebral [92, 401, 878]. cesium [1844]. CESM [233]. CEV [1518]. CFD [1051, 958, 977, 1911, 87, 56, 168]. CFD-DPM [1911]. CFD/RBD [977]. CFLP [1544]. CH [498]. Chain [1193, 323, 1762, 1483, 914, 1372, 274, 608, 906, 530]. challenge [1801]. Challenges [844, 696, 214, 1213, 654, 143]. change [1469, 1359, 1871, 281, 1548, 716]. change-point [1359]. changes [291, 1336, 1437]. changing [1491]. channel [1621, 1781, 1408, 1820]. channels [1286]. Chaos [1647, 1761, 163, 1328, 1273]. Chaospy [426]. Chaotic [247, 1783, 348, 6, 1141]. Characterisation [739]. characteristic [1643]. Characteristics [1861, 1499, 429, 1712, 1537, 943, 128, 491, 766, 976]. Characterization [972, 295, 409, 1437]. Characterizing [233, 1665, 1163, 1591]. charged [347, 1750]. charging [625, 1251, 458]. chatter [1453]. Chebyshev [1886, 1367]. checking [1181]. checkpointing [955, 1666]. chemical [57, 499, 737, 1398, 78, 291, 1788, 543, 497, 1392]. chemically [1437]. chemically-induced [1437]. Chemistry [286, 400, 1337, 1725]. chemotherapeutic [1370]. China [775, 944, 235, 968, 1107]. Chinese [952]. chip [878]. chirality [496]. choices [1185]. Cholesky [621, 1097, 846]. Christi [1820]. Christov [1861]. chromatic [1481]. chunk [1298]. Churn [1686]. Cipher [1728, 668]. ciphers [1781]. Circle [1634]. Circle-based [1634]. Circuit [1828, 446]. circular [1020, 912, 1893, 1671]. circular-shaped [1020]. circulation [1750]. Citations [72]. cities [739, 1175, 999]. citizen [1054]. city [116, 115, 1189, 1919]. city-scale [1189]. claims [605]. class [1581, 1177, 25, 1484, 360, 847, 1595]. classical [191]. Classification [1567, 1484, 1789, 1121, 1454, 609, 823, 875, 1023, 761, 1501, 650, 1279, 545, 799, 509, 1527, 658, 1612, 1589, 1411, 967, 996, 1083, 847, 1647, 1868, 1286, 1689, 1433]. classifier [1121, 360, 1626]. classifiers [470]. classifying [1443]. CLAVIRE [156]. clay [492]. clay/epoxy [492]. climate [357, 1533, 84, 1548]. climatology [1172]. climbing [315]. Clinical [124, 1220, 969, 1318, 1451, 743, 1730, 130]. cliques [1654]. clogging [1022]. closed [1363, 768]. closest [49]. Cloud [752, 665, 725, 854, 1680, 587, 74, 1599, 1608, 1906, 1922, 1690, 459, 860, 1631, 1070, 1741, 913, 1584, 694]. clouds [956, 859, 1498]. CLSVoF [1842]. Cluster [336, 151, 771, 853]. cluster-based [771]. Cluster-discovery [336]. Clustering [264, 1292, 818, 1809, 1862, 431, 242, 1669, 1746, 184, 1151, 220, 133, 605, 1664, 1624, 1868, 1619, 1634, 769]. clusterization [1502]. clusterization-based [1502]. clusters [28, 786, 7, 149, 1876, 1521]. CNN [1563, 996, 1813]. CNNs [1493]. Co [376, 271, 1834, 1517, 251, 500, 412, 1123, 889, 498, 1409]. Co-evolution [376]. co-matrix [1517]. co-simulation [271, 1834]. co-swarm [251]. Coarse [87, 362, 864]. coarse-grained [362, 864]. Coarse-grid [87]. Coastal [994, 375]. coatings [1133]. code [902, 1259, 1255, 1254, 958, 221, 1710, 46, 899, 59, 1369, 1304]. coded [803, 253]. codes [882, 1508]. coding [1483, 957, 912]. Coefficient [1731, 1746]. coefficient-dependent [1746]. coefficients [25, 412, 1628]. Coevolution [1709]. coevolving [1704]. coffee [1252]. cognition [432]. cognitive [849, 186, 1792]. cognitively [671]. Coherent [1644, 342]. cohesion [248]. cohesive [531]. coins [455]. cold [965, 1365]. collaboration [486, 1790]. Collaborative [633, 466, 914, 146, 209, 331, 794, 888, 985]. collection [295, 1610, 1795, 954]. Collective [1011, 939, 670, 1787, 1053, 316, 344, 1026]. Collision [708, 1179]. Collision-free [708]. collisions [569]. collocation [168, 1886, 1307, 341]. colon [748, 1143]. colonial [312]. colony [315, 671, 1435, 1500, 898, 777, 779, 790, 245, 343, 1927, 982, 1141]. Color [535, 250, 807]. colorectal [748]. coloring [223, 770]. column [1271, 1622]. column-level [1271]. combat [101]. combination [1269, 647, 96, 578]. Combinatorial [131]. Combined [997, 463, 1402, 1688, 417, 1855, 1370, 1689, 614]. Combining [1183, 1342, 1869, 1017, 1835]. COMBO [131]. COMBO-FISH [131]. combustion [737, 1239]. come [136]. commerce [700, 173]. commercial [791]. commitment [803]. commodity [245]. common [552, 1130]. CoMMpass [1285]. Communication [1264, 220, 271, 855, 1480, 1073, 481]. Communication-efficient [1264, 220]. communications [703, 324]. communities [72, 688, 998, 1490]. Community [1699, 1910, 1169, 71, 1512, 801, 772, 752, 1316, 521, 785, 134, 1305, 475]. comovement [1652]. compact [1363, 1930, 1551, 1087]. Compacting [1159]. companies [758]. Comparative [318, 550, 1545, 718, 585, 894]. comparing [544]. Comparison [1030, 603, 174, 1812, 1155, 1402, 739, 1475]. Comparisons [976]. compartment [1798]. compartmental [1662]. compatibility [1548]. compensation [1284]. compensator [1553]. compensatory [1774]. competence [796]. Competing [622]. competition [1734]. Competitive [13, 184, 312, 1140, 946, 205]. compilation [935, 100]. complete [790, 1383]. completion [721]. Complex [42, 542, 313, 316, 344, 1581, 391, 396, 1131, 727, 675, 187, 1333, 898, 1712, 385, 544, 693, 772, 637, 1449, 1114, 112, 1467, 623, 323, 1088, 1899, 67, 398, 1658, 532, 1490, 122, 999, 1639, 1168, 1595, 1545, 366, 582, 798, 783, 1383, 1883, 1026, 1701, 670]. complex-networks [999]. Complexity [1866, 850, 1269, 1881, 1712, 174, 1790]. Component [494, 875, 746, 248, 557, 1471, 1720]. components [196, 538, 1401, 1632]. Composing [120]. composite [185, 1784]. composites [387]. composition [1680, 1924, 1876, 913, 227]. compositional [1162, 1440]. compound [1376, 263]. compounds [404, 405]. compressible [1040, 1514, 1575, 924, 843, 1312, 1896]. compression [1341, 1483, 591, 896]. compressive [109]. compulsive [1083]. Comput [1364, 209, 1859, 1335, 344, 1067]. computable [1924]. Computation [1920, 481, 16, 1895, 983, 1216, 1127, 738, 1055, 1536, 824, 1903, 1846, 104, 1661]. Computational [410, 809, 239, 489, 1234, 1806, 1717, 1221, 1184, 1208, 536, 1404, 1318, 1119, 1754, 1239, 407, 1550, 1532, 1422, 1508, 1126, 593, 225, 258, 286, 400, 549, 1851, 81, 33, 311, 109, 753, 8, 841, 65, 24, 1269, 950, 723, 1767, 69, 499, 1880, 1911, 1512, 55, 1529, 486, 639, 349, 370, 26, 1294, 661, 1095, 1206, 803, 906, 1900, 840, 1219, 1338, 1796, 422, 415, 793, 1645, 1120, 389, 10, 379, 889, 604, 342, 1495, 9, 995, 981, 1210, 1523, 126, 438, 487, 1246, 23, 143, 810, 1014, 190, 326, 720, 204, 173, 1790]. computational [594, 1207, 465, 100, 477, 620]. computationally [719, 461, 474]. computations [1192, 147, 1630, 223, 1039, 899, 1050, 1682, 58, 1592]. compute [786, 1867]. Computer [334, 813, 1324, 723, 1313, 523, 1072, 693, 592, 1323, 320, 920, 1459, 1720]. computers [1522]. Computing [30, 239, 1323, 1110, 665, 1051, 490, 212, 683, 1164, 1269, 307, 529, 1027, 123, 1717, 1075, 702, 1385, 1031, 127, 428, 478, 1321, 1825, 26, 1617, 221, 1065, 68, 156, 1565, 277, 437, 1800, 73, 615, 1334, 1397, 1834, 324, 464, 112, 1173, 849, 1832, 240, 1922, 752, 651, 1821, 213, 200, 186, 459, 904, 573, 1070, 1618, 546, 865, 1118, 694]. concentration [1863, 751]. concept [1496, 1843, 1175, 657]. Concepts [658, 1843]. concerning [1531]. concrete [462, 1745]. Concurrent [154, 721, 437, 1839, 1616]. concurrently [1931]. condition [1022, 1350, 568, 909, 1548]. condition-based [909]. conditioned [630]. conditioning [1932, 235]. conditions [885, 719, 1828, 1245, 784, 1711, 1489, 1185]. Condon [412]. conductance [77]. conducting [379]. conduction [1818, 1685, 1205]. conductivity [1818, 1568]. Conference [410, 1234, 553, 742, 1338, 286, 400, 549, 489]. confidentiality [1070]. configuration [738, 1447]. configurations [267]. confined [1204, 1236]. confinement [570]. conflict [1806, 1779, 1206, 1210]. conflicts [1865]. conformations [4]. conforming [434]. congenital [1318]. congested [119]. conical [54]. conjugate [576, 1363, 305, 353]. connected [1090, 1930, 1729]. Connecting [599]. connection [1906]. connectivity [520, 837]. conquer [478, 414]. consensus [1640, 1510]. conservable [1553]. conservation [1403]. conservative [1877, 1414]. conserved [1063]. Considering [1243, 1132, 1680, 1564, 473, 1854, 1224, 1491]. consistency [1181, 298, 912]. consistent [897, 1408]. consolidation [763]. constant [1518, 1390, 1490, 1440]. constitutive [1761, 970]. constrained [1403, 715, 600, 1005, 1231, 1130, 343, 1690, 517, 1655, 856, 1896, 251, 1677]. constrained-MLCS [1130]. Constraint [1150, 1196, 857, 1543, 304, 1295]. constraints [1174, 721, 201, 855, 1887, 1180, 1367]. Constructing [1741]. Construction [829, 1654, 1186]. Consumption [1163, 665, 579, 951, 791]. Contact [1175, 1805, 565, 1415, 1368]. contagion [1368]. container [1059]. containers [1122]. containing [1250, 334]. contamination [1574]. Contemporary [139]. content [965, 1115, 183, 1929]. contents [441]. context [1051, 1672]. contextual [666]. continuation [1833]. continuity [1933, 1390]. continuous [1667, 1723, 624, 1811, 254, 1170, 1253, 1741]. Continuum [1106, 1486, 109]. contour [1037]. contours [652]. contracting [1449]. contraction [744, 750, 340, 1117]. contrast [1468]. contribution [1256]. Control [986, 974, 1364, 272, 695, 1243, 1003, 1692, 1714, 182, 641, 1753, 201, 1917, 13, 153, 352, 1137, 1458, 697, 63, 1649, 987, 890, 1782, 193, 1784, 1887, 1852, 1755, 1671, 836, 1905, 1374, 1595, 798, 713, 303, 1367, 614, 1677]. controlled [1149]. controllers [798]. Controlling [584, 459]. convection [1008, 1864, 25, 1823, 639, 1201, 1660, 1030, 1677, 1530]. convection-diffusion [25]. convection-diffusion-reaction [639]. convective [1775, 1893]. Convergence [850, 1833, 1847, 1755, 1007, 1675, 1883, 865]. conversational [1281]. conversion [1165]. Conversive [392]. convex [1773, 1630]. ConvLSTM [1738]. convolution [1534, 1614]. convolutional [1054, 1443, 650, 619, 880, 917]. Conway [225]. cooperation [1477]. cooperative [707, 992, 395, 450, 474, 1642]. coordinate [1749]. coordination [1099, 252]. Coprocessor [581]. Core [1715, 918, 18, 145, 28, 1197, 56, 1331, 129, 149, 606, 1665]. Core-level [918]. coronary [1014]. Corpus [1820]. correction [1436, 1402, 422, 84, 217]. correlated [1564]. Correlation [837, 1215, 1169, 1764, 1715, 811, 1091]. corridor [1488, 1934]. Corrigendum [1364, 1859, 524, 1335, 344]. corrosion [1326]. cortex [279, 545]. Cosine [1232, 1799, 1405, 1360]. cosmic [1294]. cosmic-rays [1294]. Cost [895, 859, 861, 1121, 1599, 1544, 1800, 236, 438, 305, 1568, 1694, 1275, 858]. Cost-aware [861]. cost-effective [1275]. cost-efficient [1599, 1800]. Cost-sensitive [895, 1121]. costs [614]. count [1559]. counter [1464]. counterintuitive [406]. countermeasures [844]. counting [295]. coupled [885, 1311, 522, 87, 154, 83, 1486, 828, 1179, 1468, 1732, 1842]. Coupling [307, 1299, 1036, 248, 747, 1834, 273, 782, 1595, 976]. coupling/decoupling [273]. course [315]. Cov19Net [1670]. Covariance [1050]. Cover [679, 710, 733, 716, 1428]. coverage [1200, 1141]. COVID [1563, 1463, 1670, 1731, 1795, 1506, 1827]. COVID-19 [1563, 1463, 1670, 1731, 1795, 1506, 1827]. CPS [1344]. CPU [145, 234, 1931, 1935, 1015, 622]. CPU/GPU [1015]. CPUs [1197]. crack [1394, 369, 98]. create [181]. creative [1483]. credit [895, 235]. Crediting [1256]. creeping [387]. crime [1464]. crimes [1084]. criminal [1778]. crisis [521]. criterion [1653]. critical [1735, 163, 698, 946, 37, 1545]. CRO [543]. Cross [1169, 1715, 1871, 23, 1117, 1807]. cross-bridge [1117]. Cross-correlation [1169, 1715]. cross-disciplinary [23]. cross-project [1871]. crossroad [732]. crow [715, 1360]. crowd [257, 1032, 1074, 1043, 1028, 1052, 867, 332, 361]. crowded [1332]. crowds [193]. cryo [1659]. cryo-electron [1659]. crypt [748, 1143]. Cryptanalyzing [1728]. cryptocurrencies [1169]. crystals [631, 1377]. CT [809]. ctsfeatures [1873]. Cubature [1312]. cubic [4, 1819]. cuboids [300]. Cuckoo [1473, 1473]. CUDA [1747, 7, 357, 305]. CUDA-accelerated [1747]. cues [1492]. culture [1869, 380]. cumulative [981]. current [334, 252]. curriculum [10, 1495]. curvature [1888]. Curve [806, 1251, 1846]. Curved [505, 738]. curves [1434, 1356, 1741]. Curvilinear [994, 1077]. custom [76]. custom-FPGA [76]. customer [1686]. customers [1047]. Customizable [1759]. cut [1023, 1804]. cuts [1041]. cutting [1459, 908]. cutting-edge [1459]. cyber [698, 1344, 906, 398, 866]. cyber-physical [1344, 398, 866]. cyber-security [698]. cyberinfrastructure [1317]. cycle [196, 1465, 1729, 1829]. Cyclic [449, 538]. cyclic-mission [538]. Cylinder [1369, 1893]. Cylinder-lamina [1369]. cylinders [1671].

D [664, 481, 1703, 1865, 1818, 1254, 663, 529, 1588, 1670, 127, 1017, 1061, 385, 231, 1326, 1693, 369, 1139, 1547, 1099, 1906, 672, 1773, 927, 1106, 1129, 1751, 1126, 373, 566, 150, 1690, 1437, 178, 1275, 1505, 1661]. D-bar [1703]. D-Cov19Net [1670]. D-FFP [1865]. D-Grid [127]. D-RES [1906]. D3Q19 [1194]. DAEs [105]. DAG [853]. dam [627]. damage [462, 1520, 749, 492, 1868]. damaged [1412]. damped [1092]. damped-wave [1092]. damping [270]. Darcy [1861]. Data [1862, 164, 1431, 138, 295, 639, 475, 620, 1212, 1782, 1494, 1904, 633, 1595, 1824, 1243, 1153, 712, 1920, 823, 902, 1164, 1220, 431, 197, 664, 1183, 1353, 719, 1068, 1463, 1598, 1806, 1652, 1341, 1155, 1293, 1484, 1501, 1229, 1719, 1764, 574, 1460, 482, 428, 1507, 1402, 744, 627, 1870, 1451, 1789, 1610, 219, 851, 220, 1279, 1795, 133, 959, 1667, 1359, 740, 1486, 763, 1608, 697, 156, 1055, 262, 360, 909, 1158, 759, 1650, 1649, 1347, 837, 497, 828, 1329, 1552, 1664, 1627, 771, 907, 1935, 150, 1300, 1354, 677, 208, 1429, 996, 1831, 300, 1285]. data [860, 1647, 1526, 518, 1868, 1052, 1286, 1427, 1304, 879, 1296, 1374, 1559, 839, 1810, 199, 1303, 1521, 1087, 1665, 130, 865, 954, 1642, 861, 377, 862, 1661, 1327, 864, 491, 1196]. Data-assimilated [639]. data-based [1304]. Data-Brain [1212]. Data-Driven [138, 1153, 1862, 1595, 712, 1220, 1460, 1451, 156, 1650, 1300, 677, 1052, 1374]. data-flow [1486]. databases [1294, 1408]. Datafying [1767]. dataset [1138]. datasets [835, 264, 1362, 1746, 113]. Day [277]. DEA [1132, 659, 608]. deadline [1385, 606]. deadlock [178]. Dealing [1174, 264, 646]. Death [528]. debugging [238]. December [40, 90, 1837]. decentralised [1053]. decentralized [1762, 1774]. Decision [805, 617, 1174, 125, 470, 391, 1881, 523, 706, 743, 146, 209, 333, 124, 1774, 1748, 932, 789, 1386, 1637, 394]. decision-making [391, 1881, 1774, 1748]. decisions [267]. declustering [1081]. Decomposition [543, 288, 817, 431, 447, 1215, 152, 1598, 1104, 1080, 187, 1524, 738, 460, 1205, 1903, 1898, 1480, 150, 1097, 541, 807, 1580, 1002, 1007, 1517, 1611, 1063]. Decomposition-based [543]. decoupling [273]. Deep [872, 1645, 1658, 1563, 1379, 1023, 1412, 1593, 1781, 1834, 1343, 1360, 1589, 1827, 996, 1335, 1287, 1626, 1689, 1436, 905]. Deep-ASL [1436]. Deep-learning [1658]. deepening [1820]. deeper [1708]. default [1381, 757]. defect [408]. deficit [1120]. defined [696, 740]. definite [149]. definition [266, 326]. deformation [1674, 923, 1786, 1780, 387, 1117, 1643]. degeneration [616]. degradation [537, 1017]. degrading [388, 814]. degree [1746, 595, 415, 1114, 469, 140, 1701]. degrees [1165]. dehydrogenase [287]. DEIM [604]. delay [1853, 1470]. delayed [946, 1732]. delays [174, 350, 1531]. deletion [1829]. delivery [1445, 20, 119]. demand [1753, 775, 1907, 540, 1357]. demands [1047]. demixing [704]. Democracy [1772]. democratic [1477]. demographic [1200, 1336]. Demonstrating [899, 1485]. Dempster [1269, 1058]. denoising [581]. dense [787, 857, 1094, 218]. Density [332, 1669, 923, 1359, 440, 335, 1395]. Density-based [332, 1669, 923]. dental [124]. departments [637]. dependence [1916]. dependency [436]. dependent [1818, 642, 1192, 1829, 116, 1746, 1119, 49, 63, 387, 1343, 1925, 1117, 1568, 1888]. depicting [491]. Deployment [419, 1924]. deposition [499, 1133]. depth [1271, 1573, 1134]. depth-averaged [1573]. depths [1723]. derivation [592]. derivative [16, 341, 1851, 1418]. derivative-free [1418]. derivatives [1864, 401, 291, 1867]. derived [645]. Deriving [1087]. descent [1749]. described [1356]. description [789]. desert [885]. Design [966, 1713, 662, 1237, 958, 466, 687, 1772, 1690, 896, 578, 699, 57, 969, 55, 196, 1799, 17, 390, 340, 1650, 593, 246, 1885, 205, 1784, 1548, 446, 1217, 432, 1282, 1185, 1790, 1720]. designed [46]. designing [401, 426]. desirability [1564]. desorption [1131]. desynchronized [1027]. detachment [445]. detail [1379]. detailed [737, 1355]. detect [718, 46, 1464, 1084]. Detecting [1225, 508, 999, 792, 503, 971, 1723]. Detection [1795, 1563, 895, 788, 609, 1910, 494, 875, 955, 664, 1054, 723, 1670, 805, 1561, 1787, 1512, 627, 1410, 1377, 992, 1629, 801, 1493, 1359, 336, 772, 1602, 1672, 984, 261, 281, 1411, 1699, 1241, 1226, 785, 134, 113, 807, 1868, 839, 1305, 997, 952, 945, 948]. Detention [1243]. Determination [272, 163, 825, 531, 1562]. determine [1627]. determined [550]. Determining [480]. deterministic [756]. developed [903]. Developing [75, 923, 1677]. Development [1513, 1197, 1575, 828, 268, 1695, 69, 1321, 79, 1801, 750, 1223, 1230, 586, 810, 877, 1672]. developments [950]. Deviations [1795]. device [429, 273]. devices [1579]. DEVS [471, 1333]. DEVS-M [471]. Devs/Hla [397]. DFT [978, 288, 405]. DGIRM [1276]. diabetes [907, 1203]. diabetic [609]. Diagnosis [1360, 654, 652, 1689, 1572]. Diagnostics [10]. diagonal [434]. diagonalization [845]. diagram [107, 1275, 132]. diagrams [1920, 293]. Dial [1222]. Dial-a-Ride [1222]. diatomic [1202]. dictionary [281]. diffeomorphic [989]. difference [903, 1089, 592, 1095, 1178, 1129, 1551, 1651, 1414, 1395, 1077]. different [1528, 974, 1364, 1074, 1781, 922]. Differential [1556, 43, 255, 1644, 361, 1121, 1874, 1714, 1177, 135, 1469, 1098, 1853, 319, 645, 1936, 1830, 1710, 1331, 1855, 1609, 1233, 608, 1419, 1120, 1930, 1760, 1768, 1329, 629, 1413, 314, 1452, 870, 1925, 1539, 1802, 252, 1307, 1579, 1835]. differentiation [1192, 1514, 198]. difficult [1617]. diffraction [167]. diffuse [155]. diffusion [632, 1585, 1864, 25, 1311, 1080, 141, 1628, 639, 1201, 941, 935, 427, 747, 1788, 946, 67, 785, 1811, 1092, 1569, 626, 1170, 1601, 1551, 1236, 1414, 1677, 1530]. diffusions [949]. diffusive [175, 927]. diffusivity [561, 1568]. digit [101]. digital [125, 192, 19, 1211, 1412, 1437]. Digitally [1772]. digraphs [1735]. dihydro [409]. dilute [567]. Dimension [1844, 1920, 242, 176, 1661]. dimensional [565, 642, 902, 423, 54, 42, 1652, 169, 522, 1235, 1855, 738, 1711, 162, 1532, 776, 1413, 1039, 438, 1042, 1927, 1860, 558, 1551, 1517, 1586, 861]. Dimensionality [835]. dimensions [1297, 44, 1600]. dimethyl [409]. diode [1846]. diodes [537]. diplomacy [1767, 1806, 1880, 1790]. dipole [1278, 74, 1291]. direct [887, 1359, 776, 1401, 1866]. directed [1591, 675, 1726, 1741]. directional [1865, 1867, 1699, 252]. directions [1212]. directly [1924]. disagreeing [1510]. disaster [1021]. disc [619]. disciplinary [23]. disciplines [296]. discontinuities [1601]. discontinuity [927]. Discontinuous [176, 161, 1583, 166, 1276]. discoveries [1318]. Discovering [689, 1642]. discovery [1451, 1704, 336, 1329, 300]. Discrete [1177, 1688, 553, 74, 1374, 1128, 1399, 1176, 471, 1676, 267, 556, 1015, 1522, 1595, 326, 1586, 853, 1357, 1093]. discrete-event [326]. Discrete-time [1374, 1595]. discrete-velocity [1128]. discretely [928]. discretization [434, 1710, 1419]. discretizations [1077]. discrimination [44]. Discriminative [359]. disease [470, 831, 1917, 1318, 1640, 1336, 35, 132]. disease- [35]. diseases [24, 1567, 358, 1102, 812]. dish [380]. disinformation [1464]. Disintegrating [1490]. disjoint [1785]. dislocation [206]. dismantling [1160]. disorder [1083]. disorders [552]. dispatching [1499]. dispersal [303]. dispersion [1768]. dispersions [1711]. dissimilar [1218]. dissimilarity [1787]. dissipative [1541]. dissolution [403]. Distance [1279, 1909, 999, 1822]. distance-based [999]. distances [1536]. distorted [775, 1601]. Distributed [307, 957, 1797, 1428, 966, 1430, 1187, 703, 54, 1122, 1298, 219, 1698, 1139, 855, 50, 1800, 1397, 1043, 1834, 86, 1264, 200, 826, 713, 856, 227, 954, 858, 852]. distributed-memory [1298]. distributed-order [1698]. Distributing [304]. Distribution [1066, 1115, 380, 220, 183, 88, 1901, 1923, 1280, 1442]. distributions [1771]. disturbance [1374]. disturbance-observer [1374]. disturbances [1383]. divergence [434, 166]. divergence-conforming [434]. divergence-free [166]. divergences [1536]. diverse [1049]. diverse-PSO [1049]. Diversified [1473, 1472]. Diversity [1799, 893, 473, 263, 1180]. Diversity-enhanced [1799]. diverters [92]. diverting [1106]. divide [414]. divide-and-conquer [414]. DJaya [1848]. DLDP [1579]. DLDP-FL [1579]. DNA [264, 402, 496]. DNN [1670]. Docker [1122]. docking [402]. document [818]. does [496]. doilies [1612]. domain [391, 152, 1104, 1710, 738, 1650, 776, 1646, 1903, 159, 1433, 1063, 586]. domain-driven [1650]. domains [1308, 1930, 430]. dominant [1130]. Dominated [946, 714, 1864, 243]. domination [682]. doors [390]. doped [408]. Dory [1920]. dot [504, 687]. dot-cellular [687]. double [1749, 415]. double-degree [415]. Dover [97]. down [215]. downhill [506]. download [709]. downselection [1562]. downtown [119]. DPM [1911]. drag [970, 1671]. drag-reducing [970]. drainage [367]. dried [98]. DRIHM [487]. drilling [147, 1794]. Driven [138, 1153, 1862, 1220, 1265, 1385, 1460, 746, 1451, 1718, 156, 1900, 909, 1800, 825, 340, 1650, 268, 1260, 1922, 1300, 677, 1052, 1374, 1595, 839, 530, 712]. driver [920, 1488, 365]. DRLFluent [1834]. drones [966, 1053]. Droplet [570, 1849, 1195, 1842, 569]. dropout [917]. drug [401, 1314]. drugs [1370, 407, 1827]. dry [616]. dryer [603]. dryer/furnace [603]. DSM [868]. DSM-based [868]. DSMC [559, 1260]. Dual [1828, 1867, 1458]. duality [107]. duck [1251]. due [14]. during [1660, 1780, 521, 178, 1506]. Dutch [415]. DWT [1624]. Dynamic [806, 767, 1870, 1610, 556, 768, 300, 303, 1579, 1394, 310, 423, 1176, 396, 1881, 14, 1499, 472, 1333, 1405, 1786, 351, 1544, 1420, 828, 1480, 771, 651, 1690, 1856, 613, 924, 785, 1042, 1246, 807, 514, 93, 1222, 326, 798, 353, 804, 1611, 861, 945, 886, 1733, 138]. dynamical [1474, 1183, 1341, 1676, 279, 105, 1168]. Dynamically [193, 1218, 454, 373]. Dynamics [553, 1508, 1737, 172, 1366, 57, 1169, 306, 522, 374, 1673, 1872, 319, 1917, 1911, 882, 1061, 631, 1769, 639, 1032, 570, 1566, 1704, 1113, 1145, 1074, 729, 1900, 1645, 927, 389, 206, 1373, 1195, 684, 1523, 317, 1876, 530, 1611, 816, 569, 613].

E-commerce [173, 700]. e-health [646]. e-infrastructure [130]. e-infrastructures [227]. E-science [289, 156]. ear [972]. Early [724, 945, 1493, 1044, 1906, 386, 524]. early-exercise [1044]. earth [627, 1638, 233, 1073]. earthquake [1197, 1745, 1487, 1081]. earthquake-resistant [1745]. East [110]. eavesdropping [442]. EB [1333]. EB-DEVS [1333]. echo [1242]. eco [396, 1737]. eco-bond [396]. eco-epidemiological [1737]. ecological [893, 1934]. economic [423, 1731, 1740]. economical [1745]. economics [73]. economies [1042]. ecosystem [475, 208, 485]. ecosystems [1758, 1762]. Ecuador [451]. eddy [737, 970]. Edge [1785, 1334, 1829, 1588, 1702, 1800, 1114, 1467, 1476, 1459, 1601, 1579]. edge-based [1702]. edge-centered [1601]. Edge-disjoint [1785]. Edinburgh [553]. Editorial [356, 344, 11, 21, 31, 39, 51, 64, 80, 89, 91, 114, 121, 142, 157, 170, 179, 194, 210, 222, 228, 256, 284, 301, 321, 328, 338, 345, 355, 383, 420, 456, 467, 476, 488, 501, 525, 588, 589, 590, 610, 634, 679, 710, 733, 734, 773, 820, 873, 915, 962, 979, 1000, 1018, 1034, 1045, 1056, 1085, 1100, 1108, 1111, 1124, 1135, 1147, 1156, 1166, 1190, 1209, 1227, 1247, 1257, 1267, 1289, 1325, 1345, 1387, 1406, 1416, 1424, 1438, 1456, 1478, 1515, 1554, 1596, 1635, 1656, 1836]. Editorial [1678, 1705, 1721, 1743, 1756, 1776, 1814, 1857, 1878, 1890, 1913, 1937, 1234]. EDM [948]. EDM-JBW [948]. Education [484, 1071, 725, 1072, 415, 275]. educational [192, 608]. Edwards [1711]. EEG [910, 44, 617, 1689]. Effect [1488, 1260, 128, 1382, 1025, 751, 1051, 1673, 385, 1239, 1541, 519, 1336, 1310, 303]. Effective [1713, 812, 1471, 665, 1469, 148, 568, 607, 561, 275, 810, 1790, 1275]. Effectiveness [1697, 725]. Effects [1370, 561, 834, 276, 500, 1727, 1765, 1078, 748, 1126, 1028, 1854, 1876, 1286, 1820, 1185]. Efficiency [1503, 1244, 1926, 59, 940]. Efficient [902, 152, 163, 1577, 595, 1403, 1786, 934, 183, 422, 203, 111, 981, 1846, 58, 989, 465, 325, 993, 788, 1865, 1430, 1259, 1255, 182, 719, 1598, 956, 1435, 1513, 781, 975, 1162, 1610, 1319, 1825, 220, 1400, 1685, 1599, 297, 624, 1800, 1640, 1129, 1921, 1371, 1582, 1264, 1441, 1852, 1802, 1755, 126, 705, 1418, 1603, 1296, 606, 353, 1616, 1651, 251, 859, 1063]. efficiently [1867, 940]. effort [832, 633]. EGA [1539]. EigenBlock [281]. eigensolver [884, 1447]. eigenvalue [75, 884, 105]. eight [1582]. eighth [1849]. eighth-generation [1849]. eikonal [1033]. Einstein [412, 516]. elastic [463, 850, 1702, 306, 1395]. elastic-plastic [1702]. elasticity [1342, 1674, 1518]. elastostatic [1394]. electoral [85]. Electric [1828, 1066, 1025, 458]. electrical [885, 1605, 1251, 1422, 495, 43]. electricity [1907, 533]. electro [556]. electro-hydrodynamics [556]. electroanalytical [1337]. electrochemical [1017, 1658]. electrokinetics [1569]. electromagnetic [1588, 231, 1537]. electron [278, 1659]. Electronic [1278, 1250, 1291, 409, 78, 405, 654, 283]. electroosmotic [1528]. electrophoretic [887, 1133]. electrophysiology [1104, 1641, 15, 93]. electrospinning [1442]. electrostatic [902, 290, 550, 1442]. electrostatics [934, 1747]. Element [435, 642, 112, 1375, 1864, 1702, 1469, 1197, 1261, 231, 1797, 1693, 1723, 369, 934, 390, 1780, 1760, 828, 1231, 9, 972, 1884, 1272, 140, 436, 1275, 1828]. element-by-element [1197, 1261]. elementary [452]. elements [233, 1588, 1600, 973, 140]. ellipsometric [1229]. elliptic [1098]. Elzaki [1855]. embarrassingly [597]. embedded [957, 161]. Embedding [8, 1708, 1594, 811, 948, 991]. embeddings [965]. embodied [396]. EMD [704]. Emden [1755, 1802, 1886, 1580, 1840]. emergency [628, 637, 1488, 1224]. emergent [966, 1333]. EMG [604, 1286]. emission [13, 1863]. emissions [1123]. emitting [537]. Emotion [1076, 792]. emotional [188]. emotions [661]. Empirical [1871, 1607, 938]. employees [317]. employing [520, 574]. emulate [1183]. emulating [1374]. emulation [1598, 1915, 688, 998, 695]. emulator [901]. enable [1767]. enabled [1509, 513]. enables [1658]. Enabling [1762]. encapsulation [889]. encoder [1936]. Encoding [817, 982]. ended [1537]. eNelator [35]. Energy [197, 956, 214, 1163, 396, 374, 579, 1610, 94, 380, 857, 951, 1504, 1280, 1310, 1088, 622, 856]. Energy-aware [214]. energy-based [1088]. Energy-efficient [956, 1610]. engine [1520, 1415, 837]. Engineering [489, 536, 950, 57, 1022, 55, 1799, 1492, 557, 479, 720, 813]. English [638]. Enhanced [245, 1653, 854, 591, 1362, 1680, 1799, 1058, 208, 1629]. Enhancement [324, 1436, 669, 657, 1893]. Enhancing [1200, 870, 690, 898, 765]. ENN [908]. ENN-based [908]. ENO [164]. enroute [1107]. Ensemble [1738, 1484, 1649, 246, 1152, 1638, 381, 1461, 1453, 1097, 1300, 1455, 879, 1296, 1279]. ensemble-based [1638]. Enterprise [914]. Entropy [910, 1660, 719, 60, 897, 1132, 1058, 1860]. entropy-consistent [897]. entropy-sparsified [1860]. Enumeration [1612, 539, 1654]. envelopment [1627]. environment [1430, 1385, 647, 94, 207, 1599, 1099, 1049, 1334, 1740, 1223, 1230, 1015, 586, 1631, 1584, 130, 862, 1071, 307]. environmental [1011, 1590, 1189, 1185]. Environments [1824, 854, 613, 762, 398, 174, 1196, 458, 1883]. EPDE [1329]. Epidemic [24, 226, 351, 1704, 1885, 691, 685, 1724]. epidemics [77]. epidemiological [1737]. epidemiology [69]. Epileptic [44, 1149, 1689]. Epiphany [372]. epistasis [378, 997]. epistatic [352]. epistemic [716]. epithelial [1869]. epoxy [492]. EPR [1740]. equality [201, 715]. Equation [1472, 1308, 25, 135, 139, 141, 319, 1628, 1847, 1797, 63, 1129, 776, 1768, 314, 1819, 1802, 1840, 1401, 160, 1551, 1414, 1888, 1530, 1505, 1357, 1033]. Equation-Based [1472]. Equations [43, 1874, 202, 529, 152, 1864, 163, 175, 1620, 1676, 1853, 1936, 1830, 1667, 1710, 1331, 1855, 1609, 1233, 1356, 280, 1419, 1877, 1329, 629, 430, 1413, 1452, 161, 554, 1925, 1539, 1886, 1549, 1307, 1580, 159, 1808, 166, 1677, 1835]. Equatorial [1172]. equilibration [362]. equilibrium [1717, 1688, 1350, 1611]. equivalent [1595]. era [683, 197]. eras [1211]. ERECT [956]. ergodic [1081]. Erlang [546]. erosion [1349]. Erratum [1067, 209]. Error [135, 1436, 955, 216, 1402, 63, 261, 311, 1810, 217]. errors [1812, 1564]. escape [1863]. essay [486]. Establishing [1771, 1770]. Establishment [1934]. estate [640]. estimates [1273]. Estimating [744, 1923, 1632]. Estimation [1764, 1489, 662, 54, 1362, 153, 696, 1359, 1557, 603, 1624, 335, 1831, 266, 1427]. estimators [719]. Estuary [1182]. ethylamine [288]. ethylene [1239]. ETI [854]. ETo [1739]. Euler [897, 1452, 161, 567]. Eulerian [1674]. European [632]. evacuation [678, 320, 1200, 1028]. evaluate [1336]. Evaluating [1361, 1244, 117, 605]. Evaluation [240, 1121, 703, 1696, 434, 1229, 1041, 1151, 26, 1933, 185, 693, 740, 1095, 883, 1646, 1373, 828, 541, 651, 1264, 1902, 311, 699, 366, 830, 867, 235, 1505]. Evanescent [784]. evaporation [1842]. evaporator [798]. Evapotranspiration [1739]. event [1176, 471, 336, 267, 326, 911, 948]. events [1520, 1221, 1184, 1208]. every [599]. Evidential [990, 1909]. evoked [1720]. Evolution [255, 1121, 1714, 645, 1299, 608, 1120, 1556, 1768, 376, 638, 870, 1429, 1658, 252, 1384, 361, 48]. Evolutionary [545, 907, 1050, 1716, 1351, 1282, 714, 310, 510, 1122, 1451, 1127, 1897, 1113, 729, 906, 1030, 437, 461, 548, 314, 1626, 1009, 816, 332, 337]. Evolutionary-Statistical [337]. evolved [1499]. Evolving [382, 1349]. Exact [83, 50, 754, 1144]. examine [1770]. Examining [46]. example [1317]. exascale [484, 477, 683, 1211, 213, 215, 1217, 479]. excavation [1901]. exchange [1105]. excitable [29]. excitation [918]. Exciting [81]. executed [579]. execution [1342, 786, 592, 783, 227]. exemplar [800]. exercise [1044]. exhaust [929]. exhaustive [352, 378]. exhibit [137]. Existence [350]. expansion [1761, 1446, 1683]. expansions [1273]. Expected [534, 1653]. expediency [768]. expenditures [1740]. expensive [422, 461, 474, 224]. Experience [386, 524, 1807]. Experiential [1758]. experiment [418, 1281]. experimental [1758, 1365, 1882, 791]. experiments [969, 1460, 1521]. expert [1672]. expiratory [128]. explainable [1593, 1521]. explanations [1433]. explicit [504, 903, 958, 1095, 1389, 204, 1414]. exploitation [1013]. Exploiting [1038, 482, 516, 864, 944, 56, 1860, 1067]. Exploration [1475, 1901, 458].

Exploration-oriented [1475]. Exploring [431, 1673, 1139, 888, 1185]. Explosive [1218]. exponential [1024, 223, 350, 1822]. export [1627]. exposure [431]. exposures [938]. expression [802]. expressions [926, 804]. extend [76]. Extended [1529, 1202, 1809, 1727, 1718, 757, 543]. Extending [1600, 1273, 516]. extensible [181, 1212, 199]. extension [879]. extensive [996, 1159]. External [1809]. extracellular [747, 751]. Extraction [661, 664, 939, 804, 985, 891]. extracts [263]. extrapolation [103]. Extreme [628, 335, 212, 1221, 1184, 1208, 983, 575, 1586].

Fabrizio [1872]. Face [926, 1335, 1287, 664]. Facilitating [1801, 218]. factor [1381, 1455]. Factorization [1773, 621, 1067, 1038, 216, 1622, 925, 1931, 1827, 846, 690, 838, 1773]. factorizations [844]. factorizing [1931]. Factors [290, 412, 814]. failure [462, 1520, 538]. failures [1378, 1707]. fake [1464]. fall [600]. false [1176]. false-positives [1176]. families [185]. family [1582]. FARE [726]. farm [14]. Farmer [174]. Fast [621, 1692, 1044, 1773, 362, 860, 531, 1619, 436, 1303, 1661, 539, 675, 1017, 801, 629, 134, 1002, 82, 1253, 1470]. faster [997, 1303]. fault [494, 875, 482, 219, 1246, 879, 1595]. fault-tolerant [1246, 879, 1595]. FC [1883]. FC-RRT* [1883]. FD [1149, 1916, 1105, 1092, 1060]. FDA [1303]. FDD [726]. FDD-based [726]. FDM [1754]. Fe [978]. Fe- [978]. fear [1673]. feasibility [1281]. FEAST [884]. Feature [1022, 891, 1848, 818, 788, 666, 1426, 1571, 583, 627, 1789, 1688, 728, 1544, 833, 1650, 1607, 1734, 967, 1808, 804, 985]. Features [1420, 616, 910, 724, 941, 1011, 814, 249, 960]. February [1001, 1125, 1258, 1439]. Federated [1707, 1579]. feedback [1054, 1804]. feeding [673]. felodipine [403]. FEM [524, 141, 385, 386, 1923, 178, 1616].

FEM-ANN [1923]. FENE [970]. FENE-P [970]. FEniCS [1600]. fermentation [890]. Feynman [104]. FFP [1865]. FFT [1470, 481]. FG [647]. FGMs [1818]. fiber [387]. Fibonacci [1500]. fibre [1126]. fibres [1020]. fibrillation [1288]. fiction [638]. fidelity [1529, 1261, 1301, 17, 389, 461, 1926]. Field [806, 1775, 435, 903, 1865, 1674, 631, 279, 822, 784, 480, 580, 1603, 1442, 1682, 555, 558, 1813, 1143]. fields [1536, 495]. Filling [755]. film [1573]. films [1229]. filter [1738, 1327, 1400, 1113, 1347, 1097]. filtering [510, 1648, 1652, 1486, 794, 888]. filters [1022, 583, 1870]. filtrations [1920]. finance [753]. financial [760, 851, 527, 521, 1851]. Finder [418]. Finding [1681, 1073, 1009, 595, 168, 1570, 1393]. findings [1758, 73]. finds [1130]. Fine [1096, 1736]. Fine-grained [1096, 1736]. fingerprinting [1502]. Finite [435, 1732, 903, 1375, 1352, 1558, 958, 1864, 1702, 1104, 1089, 1674, 1197, 1261, 231, 1797, 1693, 1723, 369, 592, 1095, 1928, 934, 1178, 1129, 1780, 1600, 828, 9, 973, 972, 1549, 1884, 1272, 140, 436, 1414, 1395, 1275, 1077]. finite-difference [1129, 1077]. finite-difference-based [1178]. finite-element [1197, 1261]. Finite-time [1732]. finite-volume [1558, 958]. FinOps [1922]. FinOps-driven [1922]. fire [38, 337, 737, 1221, 1184, 1208, 138, 781, 1493, 480, 877]. firefighter [1221, 1184, 1208]. firefighting [1053]. firefly [777, 869]. firewall [726]. First [1250, 78, 291, 1728, 1031, 599]. First-principles [1250, 78]. FISH [131]. Fisher [1536]. fisheries [582]. fitness [1822]. fitness-distance [1822]. fits [534]. fitted [1930]. fitting [1507, 1789]. five [1612]. fix [1273, 1465]. fixed [1753]. FL [1579]. flame [1493, 803, 1694]. flash [1440]. flattening [1251]. flawed [1740]. fleet [1445]. Flexible [227, 1486, 1928, 1850, 398]. Flight [975, 641, 977, 1165, 1062]. flip [1096]. floating [1093]. floating-point [1093]. flocks [1352]. flood [386, 524, 393]. floods [146, 209, 375]. FLOPs [1400]. FLOPs-efficient [1400]. Flow [1349, 1813, 1932, 1528, 1775, 463, 736, 564, 92, 642, 1560, 1040, 1514, 1332, 1311, 522, 1571, 1861, 1041, 1421, 1302, 1363, 515, 1660, 381, 1486, 388, 1575, 1711, 1788, 1892, 320, 354, 568, 1029, 1532, 970, 1106, 692, 920, 1488, 1898, 464, 563, 1260, 572, 128, 1893, 1382, 518, 1854, 1180, 1894, 567, 1511, 1408, 1655, 365, 976, 871, 1224, 1491, 1409, 702]. FlowNet [1813]. flows [1775, 347, 1297, 1116, 396, 1194, 306, 1513, 559, 994, 927, 1573, 165, 1468, 1639, 558, 843, 1592]. flowshop [312]. fluctuations [14]. Fluent [1834]. Fluid [553, 1508, 1529, 1421, 639, 1301, 313, 1660, 1900, 1645, 927, 389, 988, 464, 1523, 924, 1369, 1272, 1854, 555, 782, 567, 1568, 411, 326, 1236, 1842]. fluid-structure [1301]. Fluidic [922]. fluorescence [131]. flux [897, 1861, 1553, 603]. fluxes [673]. fMRI [837]. FOA [1334]. focal [910, 617]. foci [1838]. Focused [327]. focusing [709]. fog [1825, 1599]. foils [334]. follicular [126]. follow [1647]. following [750, 563]. food [1673, 1372]. foraging [250, 657]. foraminifera [673]. Force [506, 1865, 675, 353]. force-directed [675]. forces [1460, 1249]. Forchheimer [1861]. forcing [1752]. forecast [424]. Forecasting [1713, 1164, 1090, 1621, 1738, 1221, 1184, 1208, 527, 460, 393, 245, 904, 1907, 1444, 1912, 1729]. Forecasts [1713]. foreign [1105]. forensics [19]. forest [337, 138, 1225, 1602, 480, 1427, 1791, 628]. form [403]. Formal [1181, 1333, 236, 1185]. formalization [192]. Formation [613, 684, 1838]. formations [1468]. forming [840, 1780, 1471, 131]. formulation [1864, 957, 1363, 1867, 1087]. formulations [1828, 1866]. FORTRAN [357, 382]. Forward [1393, 756, 576, 1630]. fossil [1013]. Fostering [1790]. found [452]. Four [1381, 1612, 1260]. Four-factor [1381]. four-sided [1260]. Fourier [1470, 412, 1682]. fourth [1886]. fourth-order [1886]. fovea [619]. Fowler [1802, 1755, 1886, 1580]. fox [1584]. FPGA [76, 997]. FPGAs [378]. Fractal [1808, 44, 597, 1661]. Fractional [1568, 1835, 1932, 1874, 741, 1864, 25, 135, 808, 1872, 1917, 1628, 1731, 1830, 1698, 1609, 1233, 1774, 691, 1887, 1307, 1092, 1651, 1367, 1531]. Fractional-order [1835, 1917, 1731, 1531]. fracture [1399, 45]. fractures [1791]. fragile [1641]. frame [462, 591]. Framework [1577, 125, 965, 172, 470, 1812, 197, 57, 1027, 396, 471, 499, 1520, 1355, 1412, 1333, 349, 1302, 786, 117, 516, 822, 592, 794, 909, 1921, 1834, 1329, 1535, 379, 889, 364, 240, 236, 244, 314, 995, 1189, 827, 1523, 685, 126, 1868, 546, 783, 1081, 204, 863, 1655, 1517, 1842, 1637, 332, 1159]. frameworks [1730, 695, 199]. Franck [412]. fraud [895]. free [1003, 1008, 1194, 522, 26, 708, 1231, 1453, 1418, 1562, 1036, 1583, 166, 919, 465]. free-surface [1036]. freedom [1165]. freight [993, 119]. frequencies [1328]. frequency [850, 1404, 1787, 1610]. frequency-limited [1404]. Friction [445, 1236]. frog [660]. frog-leaping [660]. Front [679, 710, 733, 1803]. frontal [9, 1839]. fuel [448, 1013, 876]. full [662, 1828]. full-search [662]. full-wave [1828]. Fully [1589, 1729, 903, 1090, 1576, 828, 1898, 430, 1312]. Fully-automated [1589]. fully-connected [1090]. Function [351, 1061, 1318, 1715, 1564, 1306, 907, 224, 517, 1611]. functional [428, 1410, 1372, 1605, 855, 1903, 269]. functioning [1604]. functions [701, 1301, 291, 1590, 335, 1390, 756, 254, 305, 1616, 1835]. fundamental [1220]. fundus [125]. furan [1245]. furnace [603]. fusion [805, 1621, 535, 667, 523, 794, 106, 1420, 1556, 266, 1929]. future [257, 1550, 115]. Fuzzy [494, 1237, 201, 1576, 1578, 964, 967, 819, 741, 1265, 805, 808, 1032, 333, 1608, 1906, 1745, 890, 1774, 1748, 677, 904, 1626, 1374, 1595, 1584, 798]. Fuzzy-PID [798]. fuzzy-rules [1595]. FV [1729]. FV-MgNet [1729].

G [397]. G-Devs [397]. G-Devs/Hla [397]. GA [854, 1229]. GA-ETI [854]. GA-gradient [1229]. gain [1460, 1541]. gain-dissipative [1541]. galaxies [113]. galaxy [113]. Galerkin [1276, 176, 161, 1887, 1884, 166]. game [1717, 957, 914, 857, 992, 1800, 1627, 1584, 533, 674, 225, 82]. games [192, 1718, 1028, 275, 594, 1642, 1709]. GAN [1738, 1638]. GAN-ConvLSTM [1738]. gap [319, 633, 1472]. gaps [755, 840]. Gas [1558, 737, 559, 144, 1249, 1260, 6, 1896, 1409, 1134]. gas-like [6]. gas-phase [737]. gases [1202]. GAT [1646]. gates [455, 922]. Gauss [650, 1403]. Gaussian [760, 1150, 1536, 1354, 688, 756, 1632, 1144]. Gaussianity [417]. Gaze [1492]. GBSA [402]. GCCDC [1746]. GCCOM [994]. GCR [102]. GEMFsim [685]. gene [1127, 1640, 243, 259, 836, 1642, 1473]. gene-therapy [243]. general [551, 1089, 927, 1223, 1230, 1002, 1888, 991, 994]. generalised [1199]. Generalization [1090, 1359]. Generalized [1048, 163, 667, 44, 280, 685, 1549, 863, 1583, 892, 1701]. generated [16, 523]. Generating [613, 190, 1803, 544, 50, 88, 837]. Generation [668, 293, 1494, 257, 1259, 1255, 1908, 522, 1710, 1660, 1849, 1508, 1449, 1280, 699, 1905, 713]. Generator [968, 1841, 1936, 1926, 1033]. generator-encoder [1936]. generators [1850, 1882]. generic [547, 270, 554]. genes [358, 831]. Genetic [138, 806, 547, 657, 366, 854, 1499, 986, 925, 243, 655, 1361, 1570, 254, 424, 253]. genome [258, 491]. genomic [690]. gentle [1681]. genus [263]. geo [858]. geo-distributed [858]. geomechanical [1302]. geometrical [520]. geometries [161, 1811]. geometry [1326, 1239]. geoscience [1319]. gestational [907]. gesture [392, 881]. get [1737]. GGWO [1492]. given [625]. glacier [165]. glass [1365, 120]. Gliomas [795]. Global [1198, 63, 778, 701, 591, 1746, 13, 1801, 120, 750, 415, 1232, 1556, 1475, 1844, 254, 1170, 350, 1472, 24, 1672]. global-best [778]. globally [305]. GLR [894]. GMRES [850]. GO [866]. goaf [1906]. Goal [1578, 921]. GoD [358]. good [136]. goods [117]. Gordon [1698, 1651, 163]. governance [1790, 118]. governing [1566]. GPGPU [148, 216, 45, 1441]. GPGPU-based [148]. GpRr [831]. GPS [971]. GPU [145, 1171, 448, 1040, 234, 1327, 144, 1507, 758, 595, 352, 1900, 951, 1931, 1347, 324, 1935, 108, 150, 899, 1015, 1262, 513, 1253, 997, 436]. GPU-accelerated [448, 352, 1900]. GPU-based [1040]. GPU-CPU [1931]. GPU-enabled [513]. GPU/multi [1031]. GPUs [621, 844, 418, 845, 1715, 74, 1482, 1763, 129, 1487, 7, 1447, 1832]. grade [996]. gradient [1695, 1466, 1229, 576, 305, 353]. gradients [616]. grading [1162]. Graft [511]. grained [1096, 362, 864, 1736]. grammar [293, 1615]. Grammatical [96]. granular [1415, 904]. granularity [584, 300]. Graph [1826, 1908, 1331, 1615, 1614, 841, 1910, 1621, 293, 718, 1606, 1646, 223, 273, 842, 586, 491, 713, 1732, 991, 1708, 1895, 1527]. Graph-based [1331]. Graph-grammar [1615]. graph-theoretic [1732]. graphene [834, 455]. Graphic [781]. graphical [981]. graphics [202]. graphs [396, 682, 1785, 1591, 1292, 62, 436, 132]. gravity [1162]. Greedy [1749]. Greeks [1044]. green [366, 1919, 1934, 1611]. Grey [1081, 535, 898, 1492, 1396, 354]. Grid [1713, 412, 1385, 127, 87, 1321, 1326, 1930, 1246, 578, 1894, 1697, 1592, 1912, 127]. grids [1164, 579, 373, 771, 161, 924, 1866, 1077]. grinding [638]. groove [1249]. Ground [283, 1427]. Ground-state [283]. groundwater [1574, 614]. group [1049, 1518, 333, 443]. group-based [1049]. grouping [1526]. growing [780, 618, 967]. growth [38, 374, 1370, 1126, 134, 1382, 1770]. growth-based [134]. Guadiana [1182]. guanine [261]. guanine-rich [261]. guarantees [842]. Guelph [277]. guidance [1643]. Guided [1707, 1526, 778, 1292]. gyroscopic [1286]. gyrotactic [1892].

H [498]. H1N1 [459]. Haar [135, 1628, 691]. habitat [295]. HAFA [777]. hail [1882]. half [1452]. half-linear [1452]. halt [638]. Hamiltonian [412, 1829, 63, 1465]. Hand [101, 881]. Hand-to-hand [101]. Hankel [1773]. happiness [137]. Haptic [1460]. harbour [563]. hard [1415]. hardware [574, 1373]. harmful [945]. harmonic [152, 1576, 972]. harmonizations [1700]. harmony [241, 913]. Harnessing [1053]. harsh [885]. Harvis [1154]. hashing [1462]. hashtags [947]. having [1864, 1047]. Hawkes [1557]. hazard [448]. HCN [406]. HCS [1807]. HCS-R-HER [1807]. head [438]. Health [30, 123, 893, 1798, 646]. healthcare [605, 73]. Healthy [44]. heart [1318, 598, 751]. heat [1818, 1585, 1469, 522, 1861, 154, 1685, 1205, 603, 899, 1893, 1854, 237]. heating [438]. HEI [608]. Helastic [1342]. Hellinger [1279]. Helmholtz [1432, 1768]. HemeLB [207]. hemodynamic [988, 1154]. hemodynamics [1171, 745, 368]. hepatitis [258]. hepatocellular [1642]. HER [1807]. Heston [1091]. heterogeneity [678, 1088, 1770]. Heterogeneous [960, 151, 1765, 956, 1435, 1445, 959, 1010, 1129, 1373, 1126, 1871, 1016, 657, 532, 411, 856, 1021, 853, 861, 1118]. heteroscedastic [1793]. HetNets [857]. heuristic [1910, 1683, 801, 1599, 359, 343, 731, 1335, 1287]. hexahedral [1905]. hexanoic [1245]. HF [783]. Hidden [392, 530]. Hierarchic [547, 548, 1050]. Hierarchical [1807, 780, 1693, 1245, 1924, 1264, 300, 446]. hierarchies [847]. hierarchy [72, 333, 1361, 300]. High [288, 1027, 1261, 1301, 380, 378, 277, 694, 490, 1585, 1727, 423, 1588, 1652, 1104, 71, 1385, 1823, 1328, 428, 478, 744, 1529, 595, 1139, 1065, 1575, 1129, 1834, 1832, 1922, 1389, 1316, 1468, 1042, 1060, 1860, 1551, 1517, 350, 1395, 197, 837, 1508]. High-accuracy [380]. high-contrast [1468]. high-dimensional [423, 1652, 1042, 1860, 1517]. High-fidelity [1261, 1529]. high-order [1585, 1588, 1104, 1389, 1060, 1551, 350, 1395]. High-performance [1027, 1065, 1834, 1832, 1922, 1316, 837, 1508]. high-resolution [1727, 1823]. high-risk [71]. High-speed [378, 1575]. Higher [120, 1390, 725, 1331, 1600, 1867, 530, 649]. higher-order [1331, 530]. Highly [529, 546, 57, 693, 297, 578]. hill [315]. Hilliard [1847, 1877, 1808]. hindsight [1807]. hip [1471]. hippocampal [1662]. histogram [616]. HIV [1198, 1483]. HIV-1 [1198]. Hla [397]. HMV [470]. Hoechst [402]. holds [257]. hole [290]. Holzapfel [1761]. homogeneous [1308]. Homogenization [1131, 1022]. Homotopy [1560, 1833]. Hooke [582]. Hopper [1715]. horizontal [1893]. Hot [110, 1471, 308]. hotel [905]. HotML [868]. hours [1507]. housing [775, 886]. hp [177, 435, 231, 178]. hp-adaptive [177, 231, 178]. hp}-Finite [435]. HPC [55, 1319, 1486, 1645, 1016, 1316, 485, 879]. HRS [965]. HRS-CE [965]. HTCondor [1320]. HTCondor-CE [1320]. hubbiness [70]. hubs [70]. Human [669, 742, 1366, 1463, 1143, 1206, 1700, 1314, 258, 1875, 368, 1210, 438, 972, 190, 751]. human-level [1700]. human-like [190]. humanoid [190]. humanoids [1748]. Humidity [1444, 1158]. hummingbird [1825]. Hunting [1733]. hurricane [1200, 1820]. Hybrid [628, 27, 1288, 1229, 1302, 1599, 1570, 730, 580, 1335, 1287, 1905, 993, 788, 965, 942, 1013, 416, 780, 28, 702, 216, 1500, 1861, 898, 1617, 777, 857, 271, 333, 925, 183, 527, 1684, 1859, 1931, 1923, 149, 572, 1360, 1015, 967, 1827, 1246, 904, 1455, 846, 863, 869, 1530, 862, 1835]. hybridisation [131]. Hybridization [723, 258]. Hybridized [1791, 315, 1748]. HybridOctree_{H}ex[1905].hydrodynamic[602].hydrodynamics[148, 556, 1179].hydrogel[1392].hydrogen[406, 876].hydrological[1566].hydromagnetic[1892].hype[1211].hyperbolic[169, 1620, 1644].hypergraph[1654].hyperparameter[1461].hyperplane[991].hypersonic[1817].hyperspace[1918].hyperspectral[823, 799].hypervolume[1653].hypothesis[1676].Hypre[1262].hysteresis[565].

i-CLSVoF [1842]. I/O [1163]. I/O-intensive [1163]. IA [1344]. IA-CPS [1344]. IB [364]. IB-LBM [364]. IBJA [1848]. IBM [849]. ICAM [401]. ICAM-1 [401]. ICCS [230]. ice [1915, 1377, 74]. ID [948]. ideal [924, 166, 1513]. ideas [427]. identifiability [1730, 1918]. Identification [552, 795, 1694, 1805, 724, 1394, 1399, 1023, 188, 1731, 348, 1114, 656, 730, 1658, 1525, 1144, 1732, 880, 917]. identifier [246]. Identifying [1742, 1504, 1138, 1285, 1793, 942, 1745, 1467, 1545, 1637, 1701]. identity [724]. if [1630]. IGA [1616]. IGA-FEM [1616]. III [1582]. ill [630]. illumination [926]. image [1483, 667, 1410, 648, 799, 674, 607, 1082, 1012, 1037, 1411, 1811, 1014, 989, 649, 878, 960]. image-based [1811]. Images [964, 1563, 125, 250, 669, 1271, 1410, 545, 618, 1589, 1437, 126, 113, 807, 652, 98]. Imaging [1323, 823, 1422]. Imbalance [1243]. Imbalanced [1121, 1484, 1279]. IMENSE [130]. IMF [704]. immersed [463, 1171, 306, 988, 505, 782, 1639, 1643, 411]. Immersed-Boundary [411]. Immune [1012, 900, 730]. immunity [1102]. immunization [473]. immunohistochemistry [126]. Impact [745, 1849, 20, 1734, 1317, 71, 296, 775, 1074, 1788, 1650, 371, 1740, 1182, 658, 1894]. impacts [117, 1294, 440, 106]. IMPEC [1440]. imperfect [1740]. imperfection [716]. imperialist [205]. Implementation [1711, 108, 150, 1662, 662, 1031, 523, 758, 1564, 324, 1097, 1441, 282, 1091, 305, 581]. implementations [234]. Implementing [95, 1254, 1080]. implications [1023]. Implicit [848, 903, 1261, 923, 1389, 1468, 1159]. implicits [305]. importance [1646]. important [689, 1261, 1207, 1701]. imprecise [1372]. improve [818, 1654, 1898, 1442, 1790]. Improved [722, 44, 987, 1584, 1848, 642, 897, 1683, 650, 1825, 1405, 900, 653, 655, 1477, 1707, 1682, 282, 1631, 1644, 1694, 1842, 1586]. Improvement [1602, 363, 153, 1365, 1653]. improvements [1310]. improves [1358]. Improving [1713, 196, 243, 1794, 533, 862, 337]. impulsive [303]. Imputation [1494, 1904]. in-depth [1134]. in-field [903]. In-silico [1366, 1370]. In-situ [1552]. in-stent [42, 749]. incentive [1448]. inclined [1560, 1892]. including [565, 276, 1569]. inclusions [1308]. incomplete [1466]. incompressible [1513, 558]. Inconsistencies [759, 999]. inconvenient [1068]. incorporated [1563]. incorporating [1727, 737, 937]. incorporation [1421]. Incremental [360, 1780, 1471]. independent [1152, 881]. index [1809, 1237, 1909]. India [1574]. Indian [608]. indicator [1500, 1863]. indicators [1850]. indices [1591]. individual [1714, 688, 998]. individual-based [688, 998]. individuals [295]. indoor [1900, 1502]. Induced [895, 1051, 1861, 1120, 1437, 438]. induction [1629, 244]. Inductive [1778]. industrial [1627]. Industry [444, 914, 825, 1627, 1526]. industry-driven [825]. inequality [201, 1887, 1367]. Inertial [567, 276]. infected [1737]. Infection [1368, 1198, 172]. infections [1507]. infectious [24, 1102]. infer [1928]. Inference [895, 1237, 1429, 1158, 1241, 1226]. inferred [382]. Inferring [831]. inflammatory [407]. inflow [745]. Influence [625, 1326, 678, 814, 1168, 940, 1841, 598, 94, 1537, 946, 937, 1770]. Influential [1114, 1742, 1467, 1694]. influentials [1009]. Influenza [459, 1175, 1637]. INFO [1605]. InfoMol [497]. Information [1352, 935, 371, 122, 1027, 760, 1466, 535, 775, 1140, 941, 388, 1606, 829, 1909, 519, 990, 1389, 323, 67, 785, 266, 518, 1168, 626, 1021, 940, 852]. informative [252]. informed [1496, 1242, 1936, 1817, 1422, 1725]. infrared [839]. Infrastructure [1397, 665, 625, 54, 1512, 156, 752, 458, 130]. infrastructures [698, 289, 37, 227]. infusion [746]. ingrained [660]. inherent [1691]. inhibition [287]. inhibitory [404]. inhomogeneous [1592]. initio [1581, 288, 1376]. inpainting [1412]. input [335, 1918]. insertions [1625]. insight [287, 1838]. Insights [1245, 1220]. Inspired [239, 671, 1483, 648, 731, 1418, 904, 1213]. Instability [1817, 1775, 386, 524]. Instability-wave [1817]. Instance [1484, 189]. instance-based [189]. Institute [1323]. institutional [1790]. institutions [1495]. insurance [758]. integer [1851]. Integral [1573, 1308, 1119, 1356, 1505]. integrals [101, 824, 104]. integrate [965]. integrated [1132, 1786, 493, 1930, 1627, 269]. integrated-RBF [1930]. Integrating [1431, 659, 1606, 1021, 1719]. integration [1874, 1013, 1460, 1331, 1309, 232, 1285, 436, 1616, 130, 960]. integrator [513]. integrity [1070]. integro [1830, 1307]. integro-differential [1830, 1307]. Intel [622, 581]. Intelligence [1243, 809, 942, 723, 1053, 906, 1684, 1859, 792, 933]. intelligence-based [942, 1684, 1859]. Intelligent [1739, 1344, 793, 707, 439, 440, 661, 763, 829, 495, 651, 888, 713, 1238]. intense [111]. intensity [1095]. intensive [1163]. intention [819]. inter [1646, 324, 1304]. inter-block [324]. inter-code [1304]. inter-domain [1646]. interaction [504, 645, 1301, 1127, 369, 378, 162, 1447, 61, 1369, 1272]. interactions [1352, 352, 1606, 406, 1876]. Interactive [823, 7, 19, 675, 1154]. interactome [60]. Interconnected [1234]. interconnects [1509]. interdependence [1686]. interdependent [1504]. Interdisciplinary [1495, 370, 1397, 1790]. interest [936, 833, 1105]. Interestingness [1843]. interests [1680]. interface [1363, 1511]. interface-conjugate [1363]. interfaces [1272, 843, 1720]. interference [857]. intergranular [1326]. interior [1860]. interior-point [1860]. intermetallic [1376]. intermittent [417]. intermodal [993]. internal [1526]. International [1234, 553, 1338, 1767, 410, 489]. internet [827, 1538, 1610, 1602, 1281, 1224, 960]. Internet-of-Things [1281]. internuclear [412]. interoperability [397]. interoperable [268]. Interpolating [518]. interpolation [1277, 177, 168, 1432, 341, 1092, 1296]. interpolation-FD [1092]. interpretable [1729]. intersection [1055, 1123]. interval [875, 1740]. interval-valued [1740]. intervention [431]. interventions [1078]. intra [83]. intra-cellular [83]. intracranial [1106]. Intraday [460]. intrinsic [403]. introducing [324]. Introduction [1146]. intrusion [788, 992]. intrusive [1328, 1064]. intuitionistic [1774, 1578]. Intuitive [406, 934]. invariant [379]. invasion [1838]. invasive [1460]. inventory [1753, 201, 1740]. inverse [34, 1805, 1628, 45, 1422, 1650, 1630, 603, 548]. inversion [218, 690]. investigate [172]. Investigating [500, 271, 836, 775]. Investigation [385, 1573, 1818, 1560, 1260, 749]. investing [886]. investors [944]. ion [1017, 1604, 334, 1821]. ionic [889, 498]. ions [278]. IoT [697, 1185]. IP [181]. Irradiance [885]. irradiative [603]. Irreducible [452]. irregular [767, 585, 1062]. irregularity [1088]. irreversibility [67]. irrigation [1739]. ischemia [973]. ischemic [751]. Ising [7, 1541]. island [1030]. Isogeometric [1098, 1620, 1276, 436, 712, 1751]. isothermal [1162]. isotropic [1195, 1845]. Issue [239, 265, 286, 549, 1324, 230, 497, 1026, 753]. issues [514, 646]. IT2FCM [1608]. items [965]. Iterated [494]. iteration [577, 1609, 1622, 1205]. iterations [579]. iterative [34, 152, 832, 1302, 1447, 149, 1866, 1440, 58]. IV [403]. ivga [675]. IVPs [203]. IWO [853].

J [1364, 209, 1859, 1335, 344, 1067]. Jacobian [1760, 223, 1905]. January [99, 229, 330, 457, 735, 980, 1112, 1248, 1425, 1679, 1858]. Java [110]. JBW [948]. Jeeves [582]. Jeffrey [1568]. jellyfish [1680]. jet [556, 1401]. jets [919]. Jharkhand [1574]. JIT [312]. jitter [704]. job [1078]. jobs [1385]. Joint [1188, 615, 1923, 1471]. journal [1249, 549, 239, 286]. journey [1320]. JPEG [234]. JS [948]. JS-ID [948]. Julia [934]. July [302, 635, 874, 1057, 1167, 1346, 1555, 1777, 1938]. jump [632, 757]. jump-diffusion [632]. jumps [1381, 1060]. jumps-at-default [1381]. June [22, 1757, 1914]. Jungle [694].

K-nearest [653]. k-shell [1114]. kaleidoscopic [33]. Kalman [1738, 1113, 1097]. Kepler [623]. Kernel [951, 494, 1197, 1830, 1557, 357]. key [668, 1909, 1504]. kill [906]. kinase [287]. kind [1367]. kinematic [191]. Kinetic [1128, 1558, 1401]. kinetics [1841, 737, 1117]. KL [1536]. KL-divergences [1536]. Klein [163, 1698, 1651]. kMC [1080]. kNN [443]. knowledge [1451, 599, 422, 827, 789, 871, 991]. knowledge-based [422]. Knudsen [564, 561]. Knuth [1254]. Kolmogorov [1696, 1195]. kriging [325, 1474]. kriging-based [325]. Krylov [1299, 1403]. Kutta [1403].

label [507, 785, 1868]. labeling [1436]. labor [1614]. LACE [1429]. LAD [1795]. laden [1849]. Lagrange [140]. Lagrangian [1403, 342, 1485, 1619, 1644]. Laguerre [1802]. lahars [451]. lake [165]. lakes [159]. Lamé [1356]. lamina [1369]. LAMMPS [782]. Land [475, 1221, 1184, 1208, 716]. landmark [664]. landscape [1934]. lane [1491, 1755, 1840]. language [503, 1710, 934, 638, 1058, 1207, 586]. Laplace [776, 1505]. Laplacian [717, 1504, 1691]. Large [1713, 1008, 1068, 428, 1065, 1920, 70, 719, 1027, 1696, 737, 577, 1513, 1404, 1571, 1362, 1284, 1746, 1115, 801, 1718, 280, 1487, 970, 1205, 509, 1043, 464, 246, 1690, 368, 1539, 1716, 113, 860, 103, 531, 879, 227, 35, 945, 1107, 1795]. large-eddy [970]. Large-Scale [1713, 1008, 428, 1065, 1027, 1404, 1571, 1362, 1284, 1718, 1487, 1043, 1690, 1539, 103, 35, 945, 1107]. large-scaled [1205]. laser [1051, 111]. laser-induced [1051]. latency [19, 1180]. latency-based [19]. latent [1936]. Lattice [56, 505, 463, 564, 1128, 565, 1199, 1259, 1255, 1752, 1297, 1116, 1194, 306, 1201, 1363, 207, 1843, 1711, 1178, 562, 961, 1195, 1179, 1832, 1542, 557, 136, 554, 300, 571, 1481, 1036, 1623, 567, 922, 1380, 558, 1312, 1236, 1592, 1530, 1643, 1311, 313, 566, 563, 572, 1639, 411]. Lattice-Boltzmann [56, 1128, 207, 1311, 411]. lattices [4, 1177, 571]. law [645]. laws [1087]. Layer [1399, 504, 1775, 470, 441, 1031, 454, 505, 626, 880]. Layer-wise [1399]. layered [1040, 1790]. layers [1817]. Layout [221]. Layout-aware [221]. layouts [902]. LBM [559, 1350, 364]. lbmpy [1259, 1255]. LCMARS [930]. LCS [342]. LDDMM [1403]. LDL [1931]. lead [1738]. lead-time [1738]. leader [1647]. Leadership [1733]. leaf [374]. leaping [660]. learn [136, 981, 594]. Learning [1563, 1071, 1431, 1565, 764, 768, 1707, 1824, 1538, 872, 1379, 1812, 1567, 1826, 1183, 1353, 1765, 1869, 1022, 1131, 1908, 1484, 1764, 428, 767, 957, 1402, 1512, 706, 189, 1718, 1593, 1145, 772, 1574, 1645, 1422, 1781, 1901, 1534, 984, 1214, 1871, 10, 1834, 1590, 771, 1274, 1502, 1624, 450, 1343, 1492, 1922, 1607, 281, 376, 1794, 314, 1589, 1734, 275, 762, 973, 1477, 1827, 1725, 1042, 1658, 660, 174, 1455, 1631, 1803, 766, 1689, 1860, 881, 1303, 1766, 982, 1586, 1497, 1572, 1720, 1579, 868, 1637, 1807, 1659, 769, 763, 1533]. learning-assisted [1803]. learning-based [706, 1781, 1502, 1492]. learning-informed [1422]. Learning-to-augment [1563]. least [577, 1094, 1749, 1142, 483, 245, 1522, 424]. least-squares [577, 1749]. Lees [1711]. Legendre [1307]. length [668, 795, 514, 1384, 1117]. length-dependent [1117]. Lens [620]. lesions [609]. Lessoning [780]. levee [627, 386, 524]. Level [1842, 288, 1714, 955, 462, 1271, 737, 1717, 1483, 523, 56, 297, 1574, 1182, 1774, 918, 1700, 335, 1270, 1266, 913]. level-set [737]. Leveraging [1831, 1742, 1547]. Leverett [160]. Lévy [975, 1024]. lexical [638]. LHC [197]. Li [1821]. Li-ion [1821]. liabilities [757]. liability [758]. libraries [16, 787]. Library [307, 1823, 1316, 200, 59, 1907]. lid [1260]. LiDAR [1935, 1427]. Lie [1518]. life [1753, 1423, 1488, 1418, 1840, 674, 225]. life-saving [1418]. lifetimes [412]. lifted [495]. lifting [487]. light [537, 278, 784, 566, 1134, 1392]. light-emitting [537]. light-powered [1392]. lighten [1630]. lighting [1284, 713]. lightweight [1244, 1901, 200]. LiH [283]. like [6, 190]. Liley [279]. limit [761]. limitations [1014]. limited [1404, 1764, 1403, 1011]. line [1354, 205, 217, 1144]. line-search [1354]. linear [1436, 34, 145, 741, 1585, 955, 416, 787, 577, 808, 1263, 1404, 1277, 1299, 1302, 1372, 1331, 1749, 1747, 413, 1541, 1142, 1774, 1452, 1539, 438, 1902, 1846, 251, 630, 880, 1295]. linearization [1702]. liner [1514]. linguistic [797]. linguistics [661]. Link [599, 764, 1778, 310, 1726, 1458, 1544, 1420, 507, 990, 469, 1476, 1804, 532, 1160, 1348, 892]. link-cut [1804]. Linked [612]. linking [724, 1117]. links [508]. Lion [717]. lip [250]. LiPb [1376]. liquid [1013, 889]. liquids [568, 362, 498]. list [430]. Listwise [936]. literature [658]. lithium [1017, 1604]. lithium-ion [1017, 1604]. little [66]. liver [809, 511]. livestock [1138]. living [1186]. LMIs [890]. Load [736, 1713, 1496, 832, 1870, 676, 20, 93, 1304, 1912]. load/propagate/store [1870]. loadability [651]. loading [462, 109]. loads [1445]. Lobatto [1582]. Local [1005, 1623, 1547, 120, 1486, 1306, 624, 390, 1556, 1037, 1371, 1481, 1092, 754, 1579, 581]. locality [1462]. localization [966, 1049, 331, 1502, 1300, 1810]. Localized [1553, 978, 1428]. Locally [166]. Locating [414]. location [466, 119, 954]. locations [689]. locomotion [188]. locus [378]. LOD [1551]. logic [446]. logical [789]. logistic [1789, 997]. logistics [116, 394, 914, 115, 1526]. long [885, 1738, 1933, 697, 829, 1423]. long-term [885]. longest [1130]. longitudinal [977, 1429]. look [1708, 924]. look-up [924]. lookup [622]. loop [1054, 106, 104]. Lorenz [1183]. loss [1763, 483]. losses [1731]. Low [1450, 662, 1889, 787, 897, 1261, 1712, 697, 340, 1542, 896, 213, 305, 1568, 1408, 891, 1822, 1773]. low-altitude [1822]. low-autocorrelation [1889]. low-cost [1568]. low-cost-memory [305]. low-order [1261]. low-power [662, 787]. low-rank [896, 891]. low-speed [340]. low-storage [1408]. low-time [1712]. ls1 [1270, 1266]. LTE [927]. LU [1067]. Ludo [515]. lumpy [1917]. lung [1670, 1849, 1360]. Luo [1020]. LUSI [110]. LVOT [745]. lymphoma [126].

M [471]. M&S [265]. Mach [897].

Machine

[1431, 472, 1718, 1145, 1901, 1274, 1042, 1613, 1824, 1538, 905, 1812, 1567, 1183, 1764, 428, 1402, 527, 1574, 1422, 413, 483, 1871, 1590, 1922, 973, 1725, 881, 1303, 1586, 868, 1637, 1659, 1533]. Machine-assisted [1613]. machines [1499, 627, 763, 245, 359]. Macro [802, 1140, 111, 61, 311, 976]. macro-scale [311]. Macro-to-micro [802]. Macroscopic [388, 1410, 1161, 1117]. macular [616]. magic [231]. magic-T [231]. MAGMA [621]. Magnetic [964, 1775, 76]. magnetized [1750]. magnetotelluric [435]. mainstream [487]. maintenance [909, 796]. make [743]. Makespan [771]. making [1174, 391, 1881, 1774, 1748, 932, 1386, 1637]. malaria [401]. Malware [473, 709]. MaMiCo [1486]. mammograms [967]. mammographic [669]. MAMS [643]. Management [1146, 1865, 1221, 1184, 1208, 1313, 394, 758, 1344, 1897, 857, 1189, 732]. Managing [1332, 1281, 665]. MANET [1187]. MANETs [471]. Manifold [1572, 1856, 1497, 1720]. manifold-based [1497]. manipulator [1003, 986]. manoeuvres [1418]. mantle [1008]. manufacturing [1680, 909, 913]. many [1197, 56, 129]. many-core [1197, 56, 129]. manycore [574, 112, 1262]. map [1379, 799, 480, 236]. Mapping [837, 1373, 1462, 559, 812, 1856]. MapReduce [859]. maps [85, 1690]. March [99, 52, 339, 468, 596, 774, 1019, 1136, 1268, 1457, 1706, 1879]. marching [1866]. Marden [1819]. mardyn [1270, 1266]. Margin [1172]. Marine [208, 395]. maritime [1433]. markers [1285]. market [41, 775, 759, 521, 1614, 944, 235]. markets [527, 534, 6, 533, 886]. Markov [1717, 323, 540, 530]. Markovian [392, 612]. marks [1023]. Mashup [643]. masking [669]. mass [133, 1063, 712]. mass-conserved [1063]. mass-spring [712]. masses [1380]. massive [835]. Massively [1271, 747, 437, 368, 1266, 887, 206, 1270, 1770, 1839]. Massively-parallel [1271]. Master [415]. matching [523, 1936, 912, 1633, 649]. material [462, 647, 1509, 1658, 1845, 1896]. materials [1581, 1145, 497, 464, 1784, 1217, 825]. Mathematical [1322, 1838, 237, 477, 1198, 1102, 26, 1140, 1774, 1875, 571]. mathematics [212]. matrices [844, 145, 845, 1764, 1002]. Matrix [1773, 1050, 1008, 1362, 696, 1482, 1718, 1622, 925, 1142, 1231, 324, 1827, 218, 1583, 217, 1517, 891, 838]. matrix-free [1008, 1231, 1583]. matter [1919]. matters [195, 1724]. maturing [547]. max [819, 666]. maxima [550]. maximal [1654, 692]. maximization [1910, 946, 940]. maximize [1453]. Maximizing [721, 856]. maximum [668, 651]. Maxwell [152]. May [12, 285, 346, 611, 821, 1046, 1157, 1516, 1744]. MBBAT [728]. MCDM [1174, 705]. MCSA [1629]. MDC [1502]. MDP [199]. Mean [1775, 279, 1576, 1037, 1888]. mean-field [279]. Means [1728, 1293, 1391, 1486, 967, 1846, 581]. measure [672, 382, 1088]. measurement [1332, 608, 1168]. measurements [435, 1562]. measures [1809, 1155, 544, 718, 1545, 540].

Measuring [1071, 804, 377]. mechanical [647, 744, 1693, 730, 1618]. mechanically [83]. mechanics [1761, 1760]. mechanism [587, 297, 1448]. mechanisms [189, 987, 749, 1203]. mechanistic [1220]. mechanobiological [749]. media [1932, 689, 1131, 812, 1200, 1129, 1349, 566, 561, 797, 1440, 1811, 792, 1021, 367]. medical [470, 1410, 1012, 1637]. Medicine [30, 123]. MediGRID [127]. medium [1892, 1854, 533]. melanoma [1410]. Melbourne [866]. melody [1700]. melt [1915]. membrane [551, 1693, 324, 240, 1358, 1659]. membranes [463, 306]. memetic [1139, 624, 770, 86, 1015, 1050]. memories [861]. Memory [1598, 1673, 47, 1013, 1298, 1095, 106, 829, 1028, 1343, 1542, 1459, 305, 350, 1423]. Memory-efficient [1598]. memoryless [1845]. memristor [429]. meniscus [380]. meniscus-shaped [380]. Merged [1476]. Merging [715, 1029]. MERO [1923]. Mesh [1449, 1503, 522, 923, 1301, 556, 512, 15, 93, 178, 1623, 1905, 140, 1579]. mesh-free [522]. meshes [1884, 1803, 1601, 1033]. Meshfree [1546, 1573]. meshing [1588]. meshless [1040, 1754, 1641, 1925, 1522]. meso [462]. meso-damage [462]. Mesoscale [1845, 1217]. Mesoscopic [274]. message [579, 601]. messages [336]. meta [544, 359, 731, 1335, 1287]. meta-analysis [544]. meta-heuristic [359, 731, 1335, 1287]. metacomputing [1924]. metagenomic [262]. Metaheuristic [1445, 1240, 1745, 546, 732]. metaheuristic-based [732]. metaheuristics [542]. metal [978, 1453, 889]. metallization [1282]. Metamodeling [327, 1471]. metamorphic [1398]. metaprogramming [1263]. Method [435, 72, 818, 1932, 463, 1171, 609, 1874, 1681, 1805, 1199, 1752, 1297, 1116, 242, 38, 337, 1215, 1040, 25, 135, 737, 1702, 1881, 1469, 201, 522, 1048, 831, 141, 238, 1244, 1229, 1719, 675, 1628, 576, 1638, 1261, 1680, 272, 1201, 1833, 385, 1797, 1933, 857, 1726, 1667, 1863, 369, 1493, 1786, 1855, 1609, 1718, 1524, 811, 912, 779, 1749, 738, 168, 1356, 943, 562, 1546, 1930, 1422, 243, 692, 776, 1646, 1760, 105, 1768, 1114, 1467, 566, 1774, 1578, 387, 1179, 1641, 450, 108, 1542, 176, 1300, 1354, 430, 1573, 1393, 1734, 161, 554, 1468, 924, 1887, 1755, 102, 134, 1522]. method [571, 1549, 1307, 1580, 1884, 514, 1603, 796, 1105, 1519, 1272, 1632, 1455, 782, 1639, 491, 1623, 1633, 1803, 1304, 1060, 305, 567, 922, 1007, 1644, 1408, 140, 353, 1160, 1253, 1810, 1668, 436, 1551, 754, 911, 1367, 1732, 1592, 109, 1579, 913, 985, 867, 1395, 1295, 361, 908, 235, 1643, 1822, 1357, 642, 1560, 412, 112, 563, 1682]. methodologies [865]. Methodology [1713, 1869, 1509, 1590, 215]. methods [1174, 978, 1695, 477, 680, 1096, 1128, 1259, 1255, 1585, 431, 164, 723, 163, 1194, 1328, 1103, 1402, 426, 1789, 1847, 1299, 605, 1685, 1294, 168, 934, 1419, 1178, 1686, 1475, 603, 654, 658, 686, 848, 541, 225, 550, 705, 894, 103, 174, 159, 1092, 1185, 1312, 1616, 166, 804, 1572, 100, 505]. methyladenosine [1225]. methylidene [409]. Metric [964, 1746, 1649, 989]. metrics [1771]. Metropolis [374, 1663]. MGA [1539]. MgMOF [500]. MgMOF-74 [500]. MgNet [1729]. MgO [408]. MHD [1660, 1893, 1854, 166]. MHT [1502]. MHT-MDC [1502]. MIC [622]. Michigan [770]. Micro [871, 802, 111, 311, 804, 976]. Micro-blog [871]. micro-expression [802]. micro-expressions [804]. micro-scale [311]. microarray [264]. microbial [688, 998]. microdroplets [565]. microfluidic [347]. microorganism [1892]. micropolar [1788, 1892]. microRNAs [261, 258]. microrobots [1392]. microscopic [515]. microscopy [618]. microsimulation [97]. microstructure [1029, 1384, 976]. microstructures [1017]. mid [159]. mid-sized [159]. middle [972]. migrating [1876]. migration [1126, 765]. MILC [221]. milk [1265]. milk-run [1265]. milling [1453]. MIMO [1595]. Min [666, 819]. min-max [819]. MindBurnout [1078]. mindfulness [598, 1078]. mindfulness-based [1078]. MindHeart [598]. mine [1906]. minihubs [119]. minima [550]. minimal [1041]. minimally [1460]. minimization [577, 374, 1276, 1694]. minimize [480]. minimum [760]. Mining [682, 1843, 601, 360, 1909, 244]. mirror [304]. misfit [1630]. Missing [1494, 1904, 1824, 508]. mission [538]. mistakes [46]. mitigating [448]. mitigation [1856, 1014]. mixed [564, 574, 298, 1660, 102, 1272, 1491]. mixing [1239]. mixture [1401, 1632, 1144]. MLCS [1130]. MM [402]. MM/GBSA [402]. MMBAIS [722]. MMRF [1285]. MMRF-CoMMpass [1285]. MNERLP [1476]. MNERLP-MUL [1476]. mobile [299, 1434, 1684, 1859, 1189, 626, 830]. mobility [1293, 439, 1888, 377, 24]. Modal [270, 1682, 1700, 722]. mode [704, 460, 1480, 1039, 807, 1611, 659]. mode-oblivious [1039]. Model [994, 750, 1419, 1120, 359, 1243, 1003, 1051, 788, 1198, 903, 110, 1662, 24, 1841, 453, 1503, 1474, 1727, 716, 1183, 1761, 969, 760, 737, 780, 1022, 1762, 983, 1469, 1591, 71, 201, 808, 1132, 14, 1404, 1917, 1020, 1402, 1861, 1512, 1017, 914, 746, 744, 1638, 1712, 1372, 279, 1693, 775, 1140, 1723, 515, 678, 226, 351, 748, 1381, 1518, 802, 1509, 388, 1593, 1461, 757, 1906, 1161, 1306, 1849, 1202, 1030, 1212, 759, 1422, 883, 970, 1106, 1781, 461, 637, 111, 451, 828, 1740, 1488, 324, 405, 1627, 1782, 1641, 1540, 85, 1448, 1798, 688, 998, 1736, 6]. model [1037, 1730, 259, 1875, 1358, 1058, 691, 84, 208, 1441, 372, 269, 1827, 1918, 1246, 1481, 528, 904, 1384, 122, 796, 1844, 646, 1117, 1793, 1014, 1204, 1632, 555, 782, 1052, 1091, 1036, 1569, 1846, 1168, 1595, 1697, 1081, 626, 173, 558, 1813, 892, 1303, 303, 530, 1586, 1497, 649, 365, 414, 1530, 1912, 1929, 1224, 1491, 332, 361, 948, 475, 506, 1573, 1368]. Model-based [1419, 1022, 226, 555]. model-free [1003]. modeled [267]. Modeling [993, 930, 24, 276, 391, 188, 94, 439, 693, 1356, 729, 106, 1750, 750, 1123, 240, 1389, 1310, 973, 317, 876, 700, 1525, 1133, 1919, 644, 48, 1163, 978, 537, 885, 1220, 1237, 462, 423, 1761, 1588, 396, 1806, 1332, 1880, 1828, 1221, 1184, 1208, 1483, 1102, 882, 1333, 1451, 639, 1032, 380, 1378, 1564, 1149, 473, 1786, 45, 287, 943, 1029, 1200, 17, 1759, 1175, 442, 1780, 534, 111, 1475, 1926, 386, 1343, 334, 273, 1533, 9, 1570, 995, 1899, 685, 438, 289, 1440, 1845, 1613, 643, 204, 445, 1665, 367, 391]. Modelling [931, 429, 1047, 398, 841, 318, 42, 760, 522, 1674, 1676, 1288, 1731, 207, 784, 568, 1055, 1126, 524, 393, 1590, 603, 61, 492, 514, 1838, 1340, 783, 594, 237, 130, 886, 118]. Models [392, 1824, 1153, 1695, 564, 1379, 632, 1812, 260, 1908, 1311, 781, 319, 1507, 1017, 631, 1445, 185, 298, 544, 1415, 1294, 743, 1143, 1249, 1928, 1704, 331, 422, 1532, 527, 1158, 556, 389, 390, 7, 1924, 1274, 604, 1885, 1064, 848, 749, 557, 1794, 314, 1073, 511, 981, 1429, 266, 1811, 1770, 1060, 582, 1724, 754, 1144, 1444, 712, 1159, 1409]. MoDeNa [493]. modern [234, 56, 1839]. modes [621]. Modification [148, 1154]. Modified [806, 662, 1254, 385, 1799, 1605, 803, 889, 1097, 1902, 1340, 160, 1883]. modular [817, 95, 541, 199]. modularity [1910]. module [320]. modules [903, 942, 1821]. modus [1084]. molecular [57, 882, 631, 83, 402, 1486, 1245, 1145, 1373, 550, 289, 1380, 569]. molecular-continuum [1486]. molecules [1278, 1250, 412, 1291]. MOLF [741]. moment [1278, 1592, 1291]. moments [649]. momentum [1403, 1573]. monitored [928]. Monitoring [386, 524, 1610, 393, 1189, 459, 894]. mono [1235]. mono-dimensional [1235]. monodomain [1641]. monopole [1828]. Monte [374, 1712, 155, 1044, 498, 218, 622, 283, 1275]. Monte-Carlo [1275]. mood [41, 636]. Moreopt [921]. morphogenesis [1355]. morphological [1877]. morphologies [403]. Morton [1039]. Morton-ordered [1039]. Morven [314]. mosquito [812]. mosquito-borne [812]. moth [803, 1694]. motion [463, 887, 662, 417, 912, 615, 555, 1888]. motivating [780]. motivation [1071]. motor [1629]. motors [83]. mounted [1025]. Moutafis [680]. movements [474, 190]. Movie [833, 921]. moving [1301, 1272]. MPI [28, 149, 1316, 1073, 372, 282]. MPI/OpenMP [28]. MPICH [1315]. MPs [539, 692]. MRI [1436, 795]. msf [1907]. mud [110]. mudflow [110]. MUL [1476]. Multi [1171, 955, 462, 787, 1621, 1710, 1605, 673, 17, 1780, 1043, 1612, 517, 1477, 996, 1440, 1117, 18, 1932, 145, 680, 448, 741, 714, 1013, 470, 850, 391, 1040, 499, 1435, 28, 1762, 746, 631, 1680, 56, 1363, 177, 1863, 1564, 545, 822, 1331, 1855, 1143, 1053, 1576, 1487, 390, 615, 243, 776, 1423, 461, 270, 1449, 920, 1488, 1926, 149, 1774, 1179, 686, 1465, 205, 541, 557, 1798, 1413, 9, 1039, 86, 505, 1059, 1264, 1699, 864, 1716, 860, 514, 999, 1455, 1868, 1839, 732, 606, 798, 626, 140, 173, 1653, 1790, 1732, 1896, 1822, 722, 1578, 1031]. multi- [56]. Multi-agent [673, 1043, 177, 920, 86, 1264, 1716, 173]. multi-algorithm [1455]. multi-body [270]. multi-cell [1488]. Multi-channel [1621]. multi-compartment [1798]. multi-component [746, 557]. multi-core [18, 145, 28, 1331, 149, 606]. multi-crypt [1143]. multi-dimensional [1855, 776, 1413, 1039]. Multi-discretization [1710]. multi-domain [391]. multi-element [390]. Multi-fidelity [17, 461, 1926]. multi-fractional-step [1932]. multi-frequency [850]. multi-frontal [9, 1839]. Multi-GPU [1171, 1031]. multi-GPU/multi-node [1031]. Multi-grade [996]. multi-indicator [1863]. multi-label [1868]. multi-layer [470, 505, 626]. multi-layered [1040, 1790]. Multi-level [955, 56]. multi-material [1896]. multi-mesh [140]. Multi-Modal [722]. multi-node [1031]. Multi-objective [517, 448, 741, 714, 1013, 1435, 1680, 545, 1576, 243, 1774, 1465, 205, 1653, 1578]. multi-parallel [499]. multi-parameter [1762]. Multi-phase [1440]. multi-physics [822, 1487, 9]. multi-projection [680, 686]. Multi-qubit [1612]. multi-query [864]. multi-reflection [1363]. multi-relational [1699]. multi-resolution [1179]. multi-resource [860]. multi-response [1564]. Multi-robot [1477, 1053]. Multi-scale [462, 1117, 391, 1423, 1449, 999]. Multi-stage [1780, 798]. multi-state [541]. Multi-strategy [1605, 1822]. Multi-threaded [787]. multi-variant [732]. multi-vehicle [615]. multi-weighted [1732]. MULTIBAT [1017]. multibody [154]. multiclass [847]. multicomponent [1570, 1380, 1236]. multicore [787, 520, 1522, 846, 58, 582]. multicore-aware [58]. multiday [377]. multidimensional [70, 1353, 1348]. Multidisciplinary [265]. Multifidelity [1856]. multifocal [1838]. Multifractal [98]. multigrid [529, 152, 148, 574]. Multilayer [1895, 1512, 1218, 446, 1668]. multilayered [532]. Multilevel [509, 1578, 1885, 366, 1583]. multilinear [985]. multilingual [661]. Multimedia [181, 959, 294]. multimicrocracks [61]. multimodal [701, 241, 254]. multimode [538]. Multimodel [825]. Multiobjective [547, 1127, 1251, 543]. multiparticle [1179]. Multiphase [1674, 1571, 495, 554, 1603, 558]. Multiphase-field [1674]. Multiphysics [479, 347, 887, 1546, 1924, 586]. Multiple [1252, 38, 1482, 1049, 266, 1689, 917, 1681, 264, 1194, 938, 203, 1487, 1214, 1624, 1285, 1272, 1907, 1531, 414, 1386, 861]. Multiple-precision [1482]. multiple-relaxation [1194]. multiplex [1293, 949, 1378, 1476]. multiplication [1482, 217]. multiply [1930]. multiply-connected [1930]. multipoint [1754]. multipurpose [1570]. Multiscale [683, 1128, 307, 42, 499, 1022, 1131, 349, 1667, 381, 1754, 493, 629, 1195, 1064, 1468, 492, 1569, 130, 307]. multispeed [1350]. multisplitting [1299]. multistage [1666, 1159]. multistate [1041]. multistep [1585, 1103]. Multithreaded [16, 583, 1703]. Multivariate [875, 1621, 1795, 1903, 1433, 1912]. municipality [118]. muscle [750, 1449, 1117, 307]. museums [1332]. Musubi [313]. mutable [246]. mutation [987, 548]. mutual [757, 1686, 937]. MVAPICH [1316]. MVO [1099]. MVO-based [1099]. myeloma [1285]. Myo [1286].

N [412, 970]. N-parallel [970]. N6 [1225]. N6-methyladenosine [1225]. nacelle [1165]. naive [892]. nano [631, 385]. nano-crystals [631]. nanoadditives [1750]. Nanocomposite [825]. nanocomposites [492]. nanofluid [1528, 1560, 1861, 1892, 1893]. nanomaterial [1788]. nanoparticle [1650]. nanophotonic [196]. nanoporous [561]. nanoribbons [455]. nanorods [1395]. nanostructured [429]. Nanotools [493]. nanotubes [496, 408, 109]. nasal [1366]. nasopharyngeal [654, 656, 618, 658]. native [621, 4]. natural [929, 503, 545, 263, 1893, 1203]. Nature [1213]. Nature-inspired [1213]. Navier [572, 1932, 529, 1558, 1421, 1667, 1356]. navigation [1082]. Nb [1581]. Nb-based [1581]. Near [107, 1682, 839]. Near-pole [107]. Nearest [794, 719, 1726, 653]. nearest-neighbors [1726]. nearest-neighbour [719]. nearly [88]. neck [1849]. need [486]. needed [743]. needs [1630]. negative [718, 1643]. negativity [1235]. negligible [564]. negotiation [173]. negotiations [1767, 173]. neighbor [653]. neighborhood [925, 1114, 1467, 1160]. neighborhood-based [925]. neighbors [1726, 870]. neighbour [719]. neon [1401]. nerve [1120]. NESSie.jl [934]. nested [1067, 1038]. Net [1493, 1912].

NetLogo [1483]. nets [267, 766]. Network [429, 67, 459, 968, 1778, 819, 905, 299, 1910, 942, 724, 1164, 1054, 60, 1621, 1446, 1594, 1753, 1895, 1293, 650, 1398, 957, 1412, 1746, 1041, 1410, 645, 1400, 544, 1140, 740, 1249, 333, 943, 1640, 1541, 1423, 1534, 1646, 1627, 907, 1280, 1448, 932, 657, 1784, 651, 1360, 511, 1548, 1335, 1287, 699, 619, 695, 1203, 1305, 1348, 1810, 1790, 1665, 1934, 1579, 367, 880, 944, 917, 1642, 1929, 599, 1473]. network-based [1640]. networking [827]. networks [930, 70, 1399, 137, 310, 1090, 723, 441, 510, 1443, 1652, 682, 760, 805, 520, 1155, 949, 761, 1242, 1915, 675, 428, 1512, 1115, 1936, 94, 1897, 1726, 1378, 1149, 938, 935, 1218, 1704, 718, 772, 697, 708, 183, 800, 1532, 1544, 442, 939, 946, 692, 1504, 1672, 96, 508, 837, 1114, 990, 1467, 1476, 656, 764, 1389, 259, 1368, 937, 323, 1699, 1088, 1899, 67, 785, 134, 532, 1490, 531, 999, 1687, 836, 1633, 1374, 1545, 626, 866, 1160, 1383, 1668, 1009, 1517, 644, 350, 1531, 1732, 1896, 1021, 77, 414, 865, 868, 867, 35, 1701, 945, 858, 182, 1895]. Networks [296, 1527]. Neural [1895, 429, 1557, 1527, 1784, 819, 905, 1910, 1090, 1054, 1621, 1446, 1753, 761, 1915, 650, 1398, 1412, 1400, 1817, 1149, 1249, 333, 800, 96, 656, 246, 1360, 511, 1335, 1287, 619, 1531, 1896, 1934, 1444, 880, 917, 1929, 991]. Neuro [1237, 1626]. Neuro-Fuzzy [1237, 1626]. neurodegenerative [552]. Neuroevolutionary [1765]. neuromorphic [1662]. neuromuscular [1921]. neuron [1662]. neuronal [509]. neuroscience [1120, 1495]. neurovascular [747]. neutron [622]. neutrosophic [809]. neworks [1420]. news [893, 1464, 948]. Newton [1403]. Newtonian [1660, 1750]. Next [257, 936, 1058]. NFV [696]. NIE [1646]. NIE-GAT [1646]. nine [880]. nine-layer [880]. nitrogen [406]. nitrophenyl [409]. NLP [872]. NMR [1773]. no [291, 82]. no-rejection [82]. Node [425, 1538, 977, 1606, 1646, 1527, 1476, 1925, 1270, 1266, 1031]. node-level [1270, 1266]. nodes [1735, 1591, 1742, 718, 1909, 1504, 1467, 1545, 1701]. noise [1563, 323, 1793, 1822, 118]. noise-perturbed [323]. noise-robust [1563]. noisy [1563, 1183]. NOMA [1188]. Non [1328, 220, 1486, 1557, 1773, 1481, 34, 564, 910, 668, 714, 1567, 1308, 760, 559, 392, 1235, 1847, 417, 1660, 912, 1750, 1350, 927, 1930, 538, 243, 1630, 961, 387, 1354, 1851, 924, 1281, 700, 251, 1611, 581]. non-binary [668]. non-boundary-fitted-grid [1930]. non-coding [912]. Non-convex [1773, 1630]. non-diffusive [927]. non-dominated [714, 243]. non-equilibrium [1350, 1611]. non-Gaussian [760, 1354]. non-Gaussianity [417]. non-homogeneous [1308]. non-ideal [924]. non-integer [1851]. Non-intrusive [1328]. non-iterative [34]. non-linear [251]. Non-local [1486, 1481, 581]. non-Markovian [392]. non-negativity [1235]. non-negligible [564]. non-Newtonian [1660, 1750]. non-obstructive [1567]. non-overlapping [1847]. Non-parametric [1557]. non-physical [961]. non-rarefied [559]. non-reinforced [387]. non-repairable [538]. non-stationary [700]. non-trivial [1281]. Non-uniform [220]. Non-uniformly [1773]. Nonbistable [1541]. nonconservative [169]. noncovalent [406]. nondeterministic [1928]. nonhydrostatic [994]. Nonlinear [1917, 616, 1379, 577, 1853, 139, 1261, 884, 1685, 1524, 1754, 1113, 162, 1205, 111, 1337, 1413, 691, 798, 1551, 1063]. Nonlocal [492, 934, 1395]. nonnegative [1622]. Nonstandard [1089]. Nonstationary [1824, 148]. nonuniform [1884]. norm [1718, 176]. normalization [881]. Northshield [1819]. notation [516]. Note [3, 680, 1142]. notion [1235]. Novel [477, 404, 1140, 1609, 582, 1436, 903, 714, 391, 1881, 687, 1823, 1917, 1746, 1319, 728, 748, 1143, 1233, 407, 776, 497, 1898, 1467, 314, 827, 1886, 1626, 1675, 816, 867, 1912, 948]. November [329, 421, 526, 711, 963, 1109, 1228, 1417, 1657, 1816]. NOx [1863]. nuclear [1447, 1079]. number [564, 668, 1841, 897, 1336, 1867, 1481, 1632, 1286, 1521]. numbers [1265, 1311, 922]. Numeric [516]. Numerical [1727, 1560, 737, 977, 1469, 808, 141, 1604, 1933, 1698, 1309, 757, 1892, 162, 562, 988, 1337, 61, 691, 1840, 1340, 877, 1643, 1152, 1198, 1874, 1172, 1818, 887, 169, 135, 702, 478, 144, 1162, 1830, 1419, 1029, 1536, 111, 1882, 1768, 341, 1274, 1540, 474, 1413, 1533, 1802, 1755, 1671, 1580, 514, 1644, 1586]. Numerically [842]. NURBS [1410, 1356]. nutrient [1872]. NVIDIA [1715]. NVRAM [1043]. NVRAM-based [1043].

O-intensive [1163]. obesity [94]. object [1054, 780, 545, 86, 765, 807]. object-oriented [780]. objective [448, 741, 714, 1013, 1435, 1680, 545, 1576, 243, 1774, 1465, 205, 517, 1927, 1653, 1578]. objects [107]. oblivious [1039]. Observation [1341]. observational [1758]. observations [1183, 1119, 1916]. observer [890, 1374]. obsessive [1083]. obstacle [560]. obstacles [1116, 1421, 1074]. obstructive [1567]. occupancy [117, 324]. occurrences [1176]. Ocean [1172, 994, 1533]. oceanography [341]. October [1101, 1407, 1636]. octree [1905]. octree-based [1905]. ODE [1310]. ODM [1524]. odor [1049]. ODROID [523]. off [1312]. off-lattice [1312]. offender [1084]. offline [1402]. Ogden [1761]. oil [1727, 495]. oligonucleotides [131]. Omp [551]. On-line [217]. on-street [920]. one [844, 169, 1484, 409, 360, 1532, 1475, 1432, 1768, 176, 1586]. one-class [1484, 360]. one-dimensional [169, 1532, 1586]. one-sided [844]. one-stage [1475]. one-way [1432, 1768]. Online [1696, 418, 666, 893, 1402, 601, 587, 1140, 935, 718, 932, 521, 1059, 1222, 865, 867, 858]. Ontario [277]. ontological [37]. ontologies [358]. ontology [959]. Open [279, 1051, 1256, 1648, 1520, 1031, 1823, 426, 1151, 1537, 1811, 1213]. open-ended [1537]. Open-source [279, 1051, 1648, 1520, 1823, 1151, 1811]. OpenACC [1447, 357]. OpenCL [74, 581]. OpenDiHu [1921]. OpenFOAM [1513, 923, 1575, 1552, 1369, 919]. Opening [1613]. OpenMP [28, 149, 581]. OpenSBLI [592]. operandi [1084]. operation [1692, 803, 399]. operational [879]. operations [1715, 1794, 585]. Operative [395]. operator [177, 118, 1741]. operators [1177, 1872, 1098, 1817, 1396]. opportunistic [1633]. opportunities [696, 864, 214]. Opposition [717, 660, 975, 1902, 1631]. opposition-based [1631]. optic [619]. optical [409, 144, 78, 1712, 155, 740, 702]. optics [111]. Optimal [248, 390, 1280, 252, 1633, 1521, 614, 1666, 250, 1254, 201, 1404, 1150, 1917, 153, 1833, 1609, 1524, 855, 1458, 1649, 96, 814, 1627, 655, 246, 176, 1570, 613, 1887, 1580, 1296, 1367, 1929, 861, 1560, 1546]. Optimisation [1146, 1196, 993, 443, 562, 1015, 1141, 200, 1141]. Optimization [92, 27, 1004, 547, 1728, 1295, 852, 8, 818, 1733, 490, 1454, 809, 448, 741, 1862, 1375, 250, 714, 1013, 1739, 671, 510, 778, 1332, 1435, 327, 701, 975, 1115, 974, 1364, 1617, 1564, 1403, 715, 226, 351, 1137, 648, 900, 803, 600, 926, 624, 422, 241, 17, 389, 543, 1232, 243, 461, 96, 1663, 1231, 1926, 1173, 790, 603, 450, 1922, 474, 1885, 150, 1354, 657, 623, 343, 730, 84, 208, 1856, 224, 517, 1264, 864, 1548, 1471, 1716, 1418, 904, 1927, 1632, 1455, 1626, 1489, 254, 1894, 1697, 1792, 622, 1170, 1653, 1694, 1813, 869, 251, 237, 1213, 1472]. optimization [913, 1677, 1275, 960, 1791, 712, 325, 701, 1565]. optimization-based [603]. optimizations [1935, 1459]. optimize [1869, 495, 1303]. Optimized [1180, 846, 1594, 1080, 959, 1461]. Optimizer [1081, 809, 315, 975, 898, 1680, 1688, 1492, 474, 1396, 869, 1584, 717, 354]. optimizes [1282]. Optimizing [666, 1514, 1194, 579, 1113, 1262, 1626, 1153, 858, 1342, 799]. option [1119, 1060]. options [632, 1024, 1044, 1518, 928, 756, 1105]. optoelectronic [1581]. orbital [978]. orbits [5]. orchestration [1122, 397]. orchestrations [1181]. order [632, 1585, 1588, 1104, 1620, 761, 1404, 1917, 1017, 1261, 1731, 1830, 352, 1698, 1786, 1331, 291, 1343, 1389, 1390, 1452, 1867, 995, 1582, 1886, 1580, 1525, 1060, 1408, 1551, 1651, 530, 350, 1531, 1395, 1835, 1077]. ordered [1265, 1625, 1039, 1603]. ordered-fuzzy-numbers-driven [1265]. orders [817]. ordinal [715]. ordinary [1539]. organ [1548]. organic [537, 503, 889]. organising [1053]. organism [1186]. organizations [73]. organizing [799, 85]. oriented [616, 666, 780, 79, 181, 1475, 1935, 1813]. original [1369]. orthogonal [431, 1205, 1354]. orthogonalization [219]. Oscillatory [1540, 824]. osseous [535]. osteocyte [1693]. OTM [1546]. our [1790]. outbreaks [1175, 1724]. outburst [165]. outcomes [1730]. outer [551]. outlier [1241, 1226, 839]. output [335]. OutPyR [1241, 1226]. ovarian [1572]. over-current [252]. Over-sampling [1484]. overdetermined [1094]. overlap [481]. Overlapping [38, 1847, 1691]. Overview [1324, 551]. oxidative [1083]. oximetry [1443].

P [970]. P2P [1897]. Pace3D [822]. package [358, 293, 1873, 1223, 1230, 1559]. packet [150]. packets [20]. packing [612]. PAD [1787]. Padasip [1648]. Page [679, 710, 733]. Pages [32, 40, 22, 12, 90, 52, 99, 211, 229, 302, 285, 329, 322, 330, 339, 346, 421, 384, 457, 468, 526, 502, 635, 596, 611, 711, 681, 735, 874, 774, 821, 963, 916, 1035, 1001, 980, 1057, 1019, 1046]. pairing [1538, 1188]. Palabos [782]. PalaCell2D [1355]. paleoanthropological [1023]. PANDA [1393]. pandemic [1506, 459]. panels [885]. pansharpening [1379]. pantograph [1853, 1802, 1835]. pantograph-type [1853]. PAPA [849]. paper [30, 743]. papers [1234, 1338]. parabolic [1585, 1413, 1884, 1063]. paradigm [177, 275, 888, 765, 122]. paradigms [865]. paragaraph2vec [1594]. Parallel [145, 347, 1889, 1080, 1127, 738, 1565, 1546, 147, 1016, 373, 364, 686, 150, 677, 9, 93, 1153, 1118, 736, 75, 1013, 887, 1259, 1255, 312, 1503, 1269, 337, 1254, 529, 57, 54, 1192, 1588, 1271, 499, 1075, 702, 1298, 87, 1103, 758, 153, 584, 923, 1654, 1151, 1299, 1302, 786, 220, 313, 822, 592, 1461, 747, 800, 1120, 597, 206, 970, 1142, 1043, 1898, 1480, 1535, 1664, 1173, 849, 1590, 149, 108, 1130, 1097, 1371, 368, 232, 1310, 1691, 1270, 1266, 481, 573, 282, 782, 1770, 1839, 1653, 856, 581, 1033, 680]. parallel-in-time [1103]. parallelism [1067, 1038, 472, 56, 1736]. Parallelization [509, 958, 781, 583, 512, 149, 623, 58, 582]. Parallelized [18, 430]. Parameter [1714, 1915, 348, 241, 54, 1762, 1852, 1755, 1831, 1918, 1658, 1884, 1603, 481, 1697, 237]. parameterization [1403]. parameterize [781]. parameterized [1507, 1343]. parameters [1805, 163, 1194, 1137, 1358, 84, 1141, 1383]. Parametric [1308, 1178, 1004, 1372, 1557, 1356, 1630, 623, 880, 917, 1505]. parametrisation [1915]. parametrization [289, 712]. parareal [602]. park [914]. parking [615, 920, 1107]. Parkinson [1640]. parSRA [786]. Partial [1329, 43, 1436, 817, 1830, 1331, 1419, 1373, 790, 1413, 1925]. partial-complete [790]. partially [555]. participating [566]. Participatory [1772]. Particle [151, 27, 1061, 224, 1459, 818, 809, 1366, 1870, 1137, 1711, 926, 1849, 96, 1898, 474, 1270, 1266, 1902, 869, 1236, 251, 1822, 902]. Particle-based [1061]. Particle-In-Cell [1459, 151, 902]. particles [347, 1252, 983, 1911, 74, 1845, 1357]. particles-on-demand [1357]. particulate [1898, 567]. partition [1024]. partitioned [1304]. partitioning [1496, 1151, 480]. partnerships [486]. parts [385]. pass [657]. passenger [1204]. passing [1785, 579]. passive [1761, 627]. past [1892, 1854]. patch [1498]. patch-based [1498]. patches [303]. Path [806, 1683, 1434, 1099, 779, 1684, 1859, 1690, 1477, 300, 1180, 1052, 1591]. Pathfinder [1629, 1631]. Pathological [984]. paths [1681, 613, 1737]. pathways [831, 1451]. patient [745, 1106, 1730, 1793, 130]. patient-specific [745, 1106, 1793]. patients [49, 1336, 1624, 1497]. pattern [1779, 1766, 952]. pattern-based [952]. patterned [367]. patterns [795, 739, 1175, 1540, 258, 1138, 860, 540, 1724, 713, 1665, 377]. PBBCache [1151]. PCA [1132, 894]. PCA-based [894]. PDE [702, 1403, 1231, 1498]. PDE-based [702]. PDE-LDDMM [1403]. PDEs [1192, 1343]. PDHK [287]. PDLPP [835]. peak [431, 947]. Péclet [1311]. peculiar [113]. pedagogical [1072]. pedestrian [319, 388, 320]. PEDISWESA’13 [292]. Pegasus [1313]. penalization [1116, 1421, 469]. penalized [1620]. penetration [1794, 1822]. people [594]. PEPC [59]. perceptions [725]. Performance [1121, 703, 28, 434, 591, 45, 740, 1605, 1095, 1902, 978, 490, 1027, 1385, 428, 882, 472, 1451, 898, 575, 207, 185, 1365, 120, 1065, 608, 1763, 106, 277, 1602, 1834, 1898, 1832, 364, 1922, 236, 1573, 1316, 896, 1015, 1310, 1286, 694, 540, 862, 837, 1508]. performances [296]. perimeter [1489]. periodic [719, 560, 5, 1711, 350]. periods [885]. periphery [1665]. perishable [1753]. peristalsis [1528]. permitting [249]. permutation [253]. permutation-coded [253]. pernigraniline [405]. Persian [1412]. persistence [1920]. persistent [1441]. person [1561]. Personalised [1831]. Personalized [936, 744, 953]. perspective [1322, 1897, 1909]. Perspectives [410, 391, 1772, 400, 815]. Perturbation [1473, 1022, 1131]. perturbed [323, 1884]. pest [303, 174]. Petascale [843, 419, 59]. Petersburg [209, 146]. Petri [380, 267, 766]. Petrov [176]. PETSc [1231]. PETSc/TAO [1231]. PGAS [1736]. Ph.D [370]. pharmaceutical [318]. pharmacology [1203]. pharmacophore [287]. Phase [555, 125, 778, 737, 1469, 631, 1553, 822, 568, 1578, 165, 1468, 1440, 558, 1808]. Phase-field [555, 631, 822, 558]. phenols [260]. phenomena [931, 1676, 1216, 981]. phenomenon [780, 1788, 261]. phenotype [748]. phenotypes [1079]. Phenotypic [1838, 729]. pheromone [1927]. Phi [581]. PHOG [616]. phone [438]. phonon [1511]. photovoltaic [1051, 903, 885, 1763, 1846]. phylogenetic [18]. Physarum [1543]. physical [1598, 1022, 1376, 1344, 271, 1294, 1422, 961, 273, 398, 866]. Physics [1242, 1936, 1415, 1725, 1896, 1817, 822, 1487, 17, 9, 1655, 197]. physics-based [17]. Physics-constrained [1896, 1655]. Physics-informed [1242, 1936, 1817]. Physiological [742, 1875]. physiology [783]. pick [953]. pick-up [953]. pickup [1445]. PID [798]. piece [1390]. piece-wise [1390]. piecewise [1846]. PIES [1308]. pipe [1911]. pipeline [1676, 800, 1589, 1811, 1659]. pipelining [855]. pipes [237]. pitch [1787]. Pitchy [1787]. pixelwise [1561]. placement [1150, 1115, 1608]. planar [927, 1803]. plane [978, 1481]. plane-wave [978]. planet [1796]. plankton [1872]. Planning [806, 327, 1683, 1434, 1099, 779, 803, 1458, 615, 1684, 1859, 268, 1690, 1477, 1052, 732]. plant [954]. plantation [1444]. plants [1695, 1763, 1633]. plasma [151, 1365]. plasmon [1365]. plastic [1702, 387]. plasticity [206]. plate [61, 1025, 1568]. plates [647, 1671]. platform [299, 698, 1397, 342, 1392]. platforms [416, 1385, 428, 867]. Playdoh [200]. players [1709]. playing [594]. PLINK [997]. plume [110]. PMP [680]. pneumonia [1637]. POD [1394, 1474, 1486, 604]. POD-DEIM [604]. POD-RBF [1394]. Point [808, 1040, 1277, 1093, 936, 1359, 618, 1130, 1690, 1549, 1687, 1845, 1854, 1092, 1860, 1741, 1498]. point-based [1130]. point-cloud [1741]. point-of-interest [936]. points [912, 50, 88, 304, 953, 824, 1641]. Poiseuille [463, 564]. Poisson [645, 1797]. Pokémon [866]. polar [107]. polarization [500, 278]. pole [107]. policies [1151, 1283, 366]. policy [1885, 1264]. pollution [614, 1677]. polyaniline [405]. polyhedral [1681, 1687]. Polymer [825]. polymeric [362]. polymorph [403, 918]. polymorphic [779]. Polynomial [1761, 163, 1328, 595, 1273, 103]. polynomials [164, 1802, 1307]. polyurethanes [493]. pond [1915]. pooling [880]. popularity [947]. population [645, 600, 729, 253, 491, 303, 1791]. population-based [1791]. populations [729, 672]. Pore [520, 367, 1437]. porous [1932, 1131, 1892, 1349, 464, 1440, 1811, 1854, 367]. porous-media [1131]. Port [1243, 97]. portability [899, 215]. portfolio [120]. ports [395]. Position [1472, 1862, 1561, 440, 1467, 810]. position-based [440]. position-regulated [1862]. positive [149]. positive-definite [149]. positives [1176]. possibility [201, 13]. post [18, 259, 1021]. post-analysis [18]. post-disaster [1021]. post-transcriptional [259]. potassium [751]. potato [1444]. potent [432]. potential [163, 401, 1251, 1464, 1114, 1467, 550, 1820]. potentials [290, 1720]. pottery [1412]. Potts [1481]. Powell [361]. Power [1713, 1446, 1066, 1163, 665, 1695, 662, 787, 14, 1284, 1470, 1053, 1763, 803, 1309, 697, 246, 273, 213, 606]. power-efficient [606]. Power-Grid [1713]. powered [849, 1392]. powerful [238]. Practical [1442]. practice [1313]. Practices [485]. pre [1932, 1294]. pre-calculated [1294]. pre-conditioning [1932]. precision [1739, 1362, 1093, 1482, 102]. Preconditioned [1592, 680, 576, 686]. preconditioner [1583, 630]. preconditioners [434]. predator [1089, 1673, 376]. Predators [1737]. predict [628, 1606, 403, 493, 1455, 1907, 1021, 1611]. Predicting [1353, 941, 637, 1348, 947, 1791]. Prediction [1249, 1778, 110, 1379, 537, 470, 1237, 38, 337, 310, 716, 138, 1561, 1398, 1451, 601, 1726, 1817, 1863, 1461, 527, 909, 1544, 1420, 17, 1901, 1423, 507, 1686, 1871, 990, 469, 1476, 658, 1274, 764, 1730, 258, 1058, 937, 511, 480, 532, 1707, 690, 1348, 892, 365, 944, 1929]. Predictive [1695, 260, 1288, 941, 812, 1782, 1794, 1570]. predictors [609]. predicts [41]. preemptive [587]. Preface [230, 1880, 292, 180, 171, 286, 1, 2, 158, 573]. preferences [85]. preliminary [73]. prescribed [1785]. presence [1116]. preserving [1823, 1235, 17, 830, 985, 1077]. pressure [1932, 1405, 1624, 1643]. pressure-projection [1932]. Prevalence [654]. Preventing [178]. prevention [146, 209, 1882]. prey [1089, 1673, 376]. price [460, 245, 1466]. Pricing [1105, 632, 1119, 1518, 756, 1060, 1586]. primarily [1495]. primary [368]. Principal [494, 1720, 875, 1426]. Principles [1320, 1250, 1558, 1240, 78]. printing [1275]. prioritising [1640]. prioritization [358]. Priority [766]. prism [373]. privacy [830, 1579]. privacy-preserving [830]. private [587]. Probabilistic [847, 1824, 1593, 800, 981, 1368]. Probabilities [1243]. probe [1460]. Probing [323, 1531]. Problem [27, 1193, 1146, 1196, 993, 1428, 75, 1265, 1889, 391, 54, 808, 1132, 1499, 187, 1628, 144, 1421, 1445, 1047, 1576, 354, 1751, 612, 1276, 105, 770, 603, 653, 655, 1465, 1130, 1798, 1396, 343, 826, 1755, 174, 1369, 1272, 1618, 1583, 1063, 1118, 1222]. problem-solving [391]. problems [1805, 741, 1818, 642, 850, 1585, 1308, 510, 778, 1829, 1702, 1469, 148, 1263, 975, 1434, 1328, 1512, 639, 1299, 884, 1617, 777, 1799, 1564, 715, 1685, 1524, 1749, 600, 1356, 162, 241, 1930, 1205, 776, 1142, 461, 509, 373, 1774, 1578, 450, 1492, 474, 176, 9, 1887, 103, 1884, 1840, 1927, 1092, 1601, 1367, 1213, 1295, 1505, 325]. procedural [395]. procedure [1233]. Proceedings [277, 549, 286]. Process [894, 1131, 1322, 1398, 79, 333, 1178, 147, 509, 890, 1337, 688, 1471, 1442, 700, 367, 35]. process-specific [35]. Processes [1150, 499, 522, 1024, 746, 1143, 427, 1557, 1368, 398, 266]. Processing [781, 1051, 202, 661, 262, 497, 1935, 1411, 122, 863, 199, 199]. processor [372, 606]. processors [510, 787, 112]. product [1874, 1600, 1618, 1203]. production [201, 815, 1740, 1526]. production-inventory [201]. products [1753]. Professional [485]. profiler [901]. profiles [745]. profiling [261]. prognostic [1285]. prognostics [1423]. program [1310]. Programmability [215]. programme [415]. programming [808, 1499, 921, 36, 1012, 849, 372, 215, 913, 861, 1159, 1578]. programs [584, 598, 413]. progress [1816, 389, 816]. Project [1850, 197, 127, 1315, 1082, 497, 1871, 1316, 363]. projected [1855, 1336]. Projection [1741, 1932, 680, 177, 686, 630]. projection-based [177]. projections [1390, 985]. projective [1006]. proliferation [1252, 129]. promises [257]. proof [1496, 1175, 586]. proof-of-concept [1496, 1175]. propagate [1870]. propagation [850, 1399, 1027, 781, 1731, 1140, 473, 784, 63, 507, 1534, 1389, 1064, 480, 785, 1844, 644, 1843]. propagator [1432]. Proper [431, 1205]. properties [834, 1376, 78, 184, 493, 825, 405, 283, 1791]. property [1509]. proposed [808]. prospect [534, 464]. prostate [652]. prosthesis [745]. protection [1096, 827, 1340]. protein [4, 551, 1139, 934, 1747, 1606, 35]. protein-protein [1606]. proteome [1186]. protocol [1865, 297, 1663, 657]. protocols [440, 828]. prototype [1281]. prototyping [1151]. providers [116]. provides [1838]. Proximal [527]. pruning [1400, 297]. pseudo [1841, 1299, 1292]. pseudo-guided [1292]. pseudo-linear [1299]. pseudo-random [1841]. pseudoinverse [1142]. pseudonyms [1238]. pseudopotential [562]. pseudopotential-based [562]. Pseudospectral [159, 1048]. PSGformer [1912]. PSO [1859, 1010, 1049, 1608, 309, 1602, 1684, 1477]. PSO-based [1010]. PSO-SA [1859, 1684]. PSO-search [1602]. psycho [797]. psycho-linguistic [797]. psychological [646]. PU [1789]. public [859]. publication [841]. Publisher [680, 3]. Publishing [195]. pull [1256]. pulmonary [1567]. pulsating [237]. pulsations [334]. pulse [1443]. pulses [111]. pumps [495]. purchasing [886]. purpose [1223, 1230]. pursuit [70, 765]. put [1518]. pyramid [616, 649]. pyramidal [1662]. pyrimidine [1250]. pyroligneous [98]. pyROM [995]. pyruvate [287]. Python [197, 1648, 1552, 1223, 1230, 200, 1559, 198]. PythonFOAM [1552].

QCA [1470, 446]. QMBlender [1061]. QML [314]. QML-Morven [314]. QoS [1599, 1800, 20]. QSAR [287]. Quadpack [104]. quadrature [564]. quadrilateral [1803, 1601]. quadrotor [1113]. quadtree [1598]. qualitative [314]. Quality [266, 490, 1919, 235]. Quantification [1339, 1152, 233, 1474, 1328, 426, 1274, 1223, 1230, 1523, 1014, 1409]. quantify [1771]. Quantifying [893, 77, 1088]. quantitative [510, 575]. quantitatively [637]. quantities [163]. Quanto [1381]. Quantum [1634, 455, 1728, 504, 1211, 687, 1080, 1061, 1751, 612, 497, 1450, 1535, 1618, 283, 1081]. quantum-dot [504]. quartic [1819]. Quasi [438, 1354, 1902, 109, 367]. quasi-continuum [109]. Quasi-linear [438]. quasi-orthogonal [1354]. quasi-static [367]. qubit [1612]. queries [863]. query [1430, 864, 300, 863, 443]. queue [1548]. queuing [766]. QuOp MPI [1535]. quotes [759]. QuPAT [102].

R [358, 1807, 189, 1873]. R&D [1740].

R-package [358]. race [1506]. radial [1301, 1590, 907, 756, 224, 517, 1092, 305]. Radiant [1828]. radiation [1861, 1788, 1854]. radiative [233, 412, 1712, 899, 1893]. Rain [600]. Rain-fall [600]. rainbow [1807]. RainbowCrack [282]. raindrop [348]. Ramp [483]. RAMSAS [272]. Random [1904, 1841, 1594, 949, 1292, 46, 1225, 1602, 597, 1536, 828, 1354, 870, 1481, 1427, 1697, 1062, 1791, 1591, 425]. randomized [219, 1094]. range [164, 1311, 592, 697]. range-bounded [164]. Rank [1773, 1450, 896, 880, 891]. rank-based [880]. Ranking [718, 1132, 842, 1668, 936]. ranks [1612]. RANS [919]. Rao [1536, 1351]. Rapid [79, 1462, 1151, 168, 837, 1883]. rare [560, 1633]. Rarefaction [263, 1260]. rarefied [1297, 559, 1260, 1401]. Rate [387, 721, 560, 598, 697, 403, 1740, 1453, 128, 1794]. rates [1105]. ratio [1359]. Rational [1432]. rationalization [1658]. rationalized [1628]. ray [1563]. Rayleigh [1201]. rays [1294]. RBD [977]. RBF [1394, 1024, 1149, 1916, 1930, 1573, 1105, 1060]. RBF-FD [1149, 1916, 1105, 1060]. reachable [682]. reaching [190, 1510]. reaction [288, 639, 1249, 1245, 454, 543, 1337, 1811, 1569, 1551]. reactions [1569]. reactive [931, 956, 1284, 1575, 1788, 927]. reactivity [561]. read [1462, 786]. readmission [628]. Real [1189, 308, 732, 878, 1703, 431, 682, 956, 1676, 1746, 1638, 640, 56, 1016, 1329, 386, 524, 1794, 208, 762, 1840, 465, 1409]. real-estate [640]. real-life [1840]. Real-time [1189, 308, 732, 878, 1703, 431, 956, 1638, 56, 1016, 386, 524]. real-world [682, 1746, 762, 465]. reality [1211, 1154]. reasonable [49]. receptor [401]. reciprocal [137]. reclaimer [399]. recognition [819, 666, 1779, 392, 802, 926, 1335, 1287, 881, 1203, 1766, 891, 960, 908]. recognize [496]. recommendation [936, 829, 833, 888, 871]. recommendations [838]. recommender [965, 466, 921, 925, 830, 933]. Recommending [953]. reconfigurable [878]. Reconstruction [1306, 418, 1703, 663, 591, 375, 1773, 1473, 1371, 353, 1498]. recovery [576, 1119, 1381]. recrystallization [1841, 828, 514, 1384]. rectangular [1592]. rectified [1541, 880]. recurrent [761, 905]. Recursive [1327, 1347]. red [1771, 1584, 879, 1296]. Reduce [236]. Reduced [632, 1480, 1343, 629, 995]. Reduced-communication [1480]. Reducing [1768, 951, 970, 324]. reduction [1436, 242, 1404, 1017, 746, 1470, 1786, 427, 1509, 1882, 270, 790, 771, 604, 585, 1437, 1485, 1671, 1844]. Redundancy [27]. redundant [1742]. REEDS [1058]. reference [912, 1739]. refined [1866]. Refinement [1386, 556, 512, 1623]. reflection [1363, 162, 304]. Reformulated [1682]. reformulation [1582]. Refreshing [1472]. Refreshing-Gap [1472]. refuelling [1107]. regarding [1916]. regime [1671, 1060, 569]. regime-switching [1060]. region [618, 848, 967, 807]. regional [204]. regionally [751]. regions [387, 406]. registration [1082, 989]. regression [1436, 1277, 1789, 1249, 1593, 1330, 1860, 997, 424]. regression-based [997]. regular [7]. regularity [1925]. Regularization [1357, 1759, 1505]. Regularized [824, 1530]. regulated [1528, 1862]. regulation [750, 1283, 259]. Regulatory [1473, 259, 836, 1642]. reinforced [462, 1745, 708, 387]. reinforcement [1869, 900, 1645, 1834, 771, 376, 1803, 1807]. reinforcement-immune [900]. reinforcement-learning [771]. reinsertion [1222]. reisz [667]. rejection [82]. related [616, 1775]. relational [1699]. relationship [1389, 791]. relationships [1348]. relative [1235]. relativistic [1297]. relaxation [1096, 1752, 1194, 1235, 1363, 1178]. relaxing [1506]. relays [252]. release [354]. relevance [1399, 1476, 1058]. Reliability [536, 272, 538, 537, 1041, 1180, 856, 838]. reliable [1853]. ReLU [917]. Remaining [1423]. remote [1164]. removal [1453]. renting [886]. reordering [1040]. repairable [538]. repeated [296]. replay [1807]. Replication [1187]. reply [137]. report [816]. repositioning [407]. Repository [1473]. Repost [937]. representation [482, 1214, 324, 1846, 985, 789]. representations [1765]. reproducible [379]. reputation [1897]. requests [1256]. required [1460, 557]. requirements [458]. RES [1906]. resale [775]. rescheduling [399]. Research [277, 444, 911, 1256, 1068, 1806, 1315, 464, 658, 1316, 633, 720, 132]. researchers [143]. researches [391]. reserve [1013]. reservoir [1302, 373, 1587, 1852]. residual [577, 522, 1276]. residue [668]. Resilience [521, 1899]. resilient [1132, 216]. resistance [1358]. resistances [1687]. resistant [1745]. resolution [1727, 1823, 744, 568, 1179, 726, 1014, 1919, 1655]. resolve [509]. resolved [1194, 1900, 1898, 1408]. Resonance [964, 1365, 1541]. resource [900, 1800, 1922, 1188, 1246, 860, 791]. Resources [1066, 1016]. respect [137, 358, 77]. response [1372, 1801, 1564, 673, 422, 17, 84, 972]. responses [1662]. restaurant [888]. restenosis [42, 749]. restore [1933]. restoring [1412]. Restricted [359]. restrictions [520, 1506]. resulting [1277]. Results [1082, 647, 1365, 49]. retaining [1745]. retinal [619]. retinopathy [609]. retrieval [959]. RetroFitting [1708]. returning [1021]. returns [636]. retweeting [938]. reused [864]. reused-based [864]. revealing [1203]. Reverse [443]. Review [719, 658, 103, 139, 389, 932]. reviews [905]. revisited [291]. Reynolds [922]. RhythmCount [1559]. rhythmicity [1559]. RIBIB [895]. Riccati [280]. rich [261]. Ride [1222]. Riemann [1896]. rigid [136]. ring [800]. Rips [1920]. RISC [372]. rise [1182]. risk [628, 71, 757, 1906, 363, 424, 895]. risks [375, 1070]. river [393, 684, 613]. RNA [1225, 1241, 1226, 1642]. RNA-Seq [1241, 1226]. RNACache [1462]. RNN [1713]. RNN-Based [1713]. road [971, 117, 519, 1393, 23, 1224]. robbery [1084]. Robin [1468]. Robin-coupled [1468]. robot [1003, 1683, 1004, 1053, 1005, 309, 1684, 1859, 1477, 190]. robotic [503, 986, 974, 1364, 1010]. Robotics [1110, 1477]. robots [1434, 1458]. Robust [993, 494, 663, 1928, 1563, 1862, 1753, 1162, 1458, 246, 1548, 1884, 325]. robustness [1523, 1899, 67, 77]. rock [1901, 1791]. Role [647, 417, 408, 1673, 1384]. Rolling [1044, 641, 308, 976]. rolls [308]. Rome [1293]. root [595, 1793]. Rössler [1141]. rotating [1775, 988]. Rotation [628]. rotor [1629, 1165]. Rough [607]. round [919]. roundabout [1123]. roundness [983]. Route [477]. routines [416]. routing [1265, 116, 777, 1047, 440, 20, 1646, 653, 655]. rovibrational [1278, 1291]. RPF [297]. RPF-based [297]. RRKM [288]. RRT* [1683, 1883]. rubrics [10]. Rudy [1020]. rule [192, 667, 1269]. rules [1499, 244, 1374, 1595, 1312]. Run [1397, 1265, 795]. Run-Time [1397]. runaway [1821]. Runge [1403]. Running [623]. runtime [958]. Russia [1769]. Russian [415, 1175, 143]. Russian-Dutch [415]. RUSTICO [1411].

S [918, 128]. SA [1859, 1684]. saddle [977]. saddle-node [977]. safe [1832]. safety [1221, 1184, 1208, 269, 1455]. SAIRF [819]. salesman [1396]. Salesperson [1829]. Salient [807]. sampled [1773]. samples [1696, 1607]. sampling [1692, 1484, 1683, 1501, 1011, 1475, 425]. sandbag [1340]. Sandpile [676]. sandwich [647]. Sanitized [1565]. sarcomere [1117]. sarin [408]. saturated [1660]. savant [1118]. save [687]. saving [1418]. scaffolds [1933]. Scala [1065]. scalability [899, 690]. Scalable [423, 1122, 219, 1487, 212, 477, 887, 778, 1462, 79, 800, 1921, 1231, 232, 481, 578, 546]. Scale [1713, 1724, 212, 835, 1008, 462, 391, 1027, 1068, 76, 535, 1404, 1571, 1362, 428, 478, 1284, 575, 1400, 45, 1547, 1718, 1065, 1900, 1487, 206, 883, 1423, 1449, 1043, 246, 430, 1533, 1690, 1189, 1539, 860, 103, 514, 999, 311, 1117, 1919, 877, 227, 465, 35, 945, 1107]. scale-bridging [883]. scale-free [465]. scale-resolved [1900]. scaled [369, 1205]. scales [417, 350]. Scaling [59, 638, 218, 98]. scans [623]. scattered [518]. scattering [1537, 1783]. SCBM [796]. scenario [1327, 187, 1564, 1347, 1548, 1519]. scenario-based [1564, 1548]. scenarios [974, 1364, 1044, 1204, 366]. scheduler [676, 879]. scheduling [312, 1265, 854, 956, 1499, 1298, 815, 1825, 1599, 1251, 1608, 354, 951, 771, 1631, 1134, 911, 853, 1584, 859, 1118]. scheme [1128, 1752, 1727, 805, 1104, 1853, 559, 1235, 1830, 1797, 1667, 1099, 248, 50, 88, 1129, 512, 1898, 961, 1195, 1440, 1618, 1601, 1144, 1414, 853, 1063, 1701]. schemes [903, 1089, 1823, 1363, 1178, 1866, 1091, 160, 1651, 1666]. Scholes [1306]. Schrödinger [63]. Schwarz [1847]. Sci [1364, 209, 1859, 1335, 344, 1067]. Science [410, 489, 1234, 536, 156, 1055, 549, 81, 8, 1324, 841, 1758, 950, 1054, 1313, 851, 1072, 1323, 1206, 840, 1219, 1338, 1550, 415, 10, 379, 633, 1210, 487, 289, 23, 33, 720, 813, 477, 1234, 1338, 549, 239, 286]. sciences [486, 370, 186, 143]. Scientific [186, 1163, 721, 854, 1221, 1184, 1208, 1075, 238, 79, 26, 221, 36, 95, 1608, 46, 1330, 195, 487, 573, 1323]. sclerosis [917]. Scotland [553]. scour [1723]. scrambling [674]. screen [1282]. screening [616, 125, 1740, 432]. sea [1915, 1182, 879, 1296]. seal [563]. seaport [1059]. Search [722, 1371, 662, 1254, 1680, 974, 1364, 352, 715, 407, 624, 241, 1602, 692, 1556, 497, 1354, 1360, 1050, 366, 582, 1144, 913, 1458, 1473]. Searching [539, 1804, 790]. seasonal [1907]. second [1620, 1830, 1452, 1408, 1651]. second-order [1620, 1830, 1452, 1408, 1651]. secondary [1071, 435, 451]. secondhand [431]. Section [536]. secure [1538, 1822]. security [1237, 1463, 698, 657, 810, 1934, 867, 1709]. seed [618]. seeding [253]. Seeking [1735, 1748]. seepage [1022]. Segmentation [1454, 250, 652, 619, 809, 1271, 1410, 795, 1012, 656, 618, 658, 1037, 1589, 807, 878, 1659]. segments [971]. segregation [1826, 1010]. SEIR [691]. seismic [462, 627, 1261, 150]. seizure [1689]. seizures [1149]. select [1141, 1827]. Selected [1234, 1338, 552, 690]. Selecting [119]. Selection [1499, 1243, 1848, 818, 788, 666, 1484, 404, 1132, 583, 627, 1789, 1688, 728, 1405, 120, 659, 248, 855, 803, 333, 800, 1030, 1650, 618, 359, 1730, 131, 705, 871, 1929]. selective [889]. selectiviey [938]. selectivity [1139]. Selexol [498]. Self [182, 1053, 1006, 1448, 255, 167, 1889, 1298, 13, 799, 162, 987, 85, 1716, 1846, 530]. Self-adaptive [1448, 987, 1846]. self-avoiding [1889]. Self-calibration [1006]. self-control [13]. self-optimization [1716]. Self-organising [1053]. self-organizing [85]. self-scheduling [1298]. Self-similar [167, 162]. Self-stabilizing [182]. self-updating [530]. Semantic [1594, 959, 1181, 943, 939, 1672]. Semantic-aware [1594]. Semi [1403, 1064, 680, 1826, 1712, 1214, 686, 1755, 1766, 1497]. semi-aggregation [680, 686]. semi-analytic [1712]. Semi-intrusive [1064]. Semi-Lagrangian [1403]. semi-numerical [1755]. semi-supervised [1826, 1214, 1766, 1497]. semiconductor [1051]. sensing [155]. sensitive [895, 1121, 1462, 937, 1160]. Sensitivity [260, 716, 1137, 1915, 750, 1523, 1831, 1918, 1844, 361]. Sensor [182, 805, 1150, 36, 1189, 1623]. sensor-based [36]. sentiment [905, 1076]. separation [1401]. separations [412]. September [211, 322, 384, 502, 681, 916, 1086, 1191, 1388, 1597, 1815]. Seq [1241, 1226]. Sequence [761, 267, 668, 1423, 953]. sequences [1889]. sequencing [257, 1429, 1642].

Sequential [1158, 593]. serial [1084]. Series [1713, 1577, 1090, 1621, 1446, 1764, 1795, 133, 1359, 1566, 1873, 904, 709, 1907, 1433, 1444, 1729]. serious [275, 594]. server [791]. serverless [1342]. Service [253, 490, 628, 885, 703, 116, 1680, 1047, 181, 705, 643, 913]. service-oriented [181]. Services [127, 486, 185, 1181, 697, 827, 487]. sessile [1842]. sessions [623]. set [978, 809, 737, 333, 607, 335, 1087, 1804, 1842]. sets [1920, 682, 509, 550]. setting [1730]. Several [540]. severe [276]. Shafer [1058]. shallow [175, 1036, 1557, 505]. shape [663, 1377, 1356, 17, 389, 390, 1856]. shape-preserving [17]. shaped [1020, 1321, 380]. shaping [323]. SHARCNET [277]. Sharing [1238]. Sharpness [964]. shear [463, 306, 1415, 1711, 1900, 976, 1791]. shear-driven [1900]. sheet [1892, 1780, 1471, 1854]. sheets [1869]. shelf [1753]. shell [738, 1114]. Ship [1820, 1243, 1454]. Shock [1882, 1864, 162]. shock-capturing [1864]. shocks [521, 167]. shoot [1643]. shooting [1681]. shop [354]. Short [1713, 1423, 786, 829, 943, 111, 387, 131, 1907]. Short-Term [1713, 1423, 829, 1907]. shortcut [682]. shortest [1681]. shrinking [1854]. shuffled [660]. Siamese [1904]. sibilant [128]. Sickle [132]. side [1781]. sided [844, 1260]. Sidoarjo [110]. Siebeck [1819]. Sifting [503]. SigmaEva [320]. signal [1783]. signaling [867]. Signals [44, 1443, 617, 1689, 191]. signatures [699]. silent [955]. silicene [1604]. silico [1366, 404, 432, 1869, 552, 1370, 1314]. silver [1528, 1365]. silver-based [1365]. silver-water [1528]. SIMD [1197, 203]. similar [162, 167]. similarities [1877]. Similarity [1069, 819, 1809, 1895, 1726, 672, 1771]. Simple [1542, 729, 88, 1397]. simplified [192, 1639]. Simplifying [1527]. simulate [1194, 1149, 1763, 1186, 122, 37]. simulated [1688, 922, 582, 1835]. Simulating [374, 800, 563, 15, 1163, 1116, 1204]. Simulation [463, 53, 391, 560, 1020, 698, 553, 1365, 1294, 146, 209, 1120, 340, 1401, 798, 458, 399, 966, 1198, 435, 564, 931, 110, 276, 172, 885, 151, 1008, 1252, 462, 57, 1869, 396, 471, 641, 737, 1520, 520, 1211, 14, 1483, 144, 1162, 394, 1115, 746, 1333, 1197, 1261, 1604, 640, 1302, 154, 56, 117, 155, 207, 1801, 271, 475, 1415, 274, 715, 439, 381, 1065, 1486, 1509, 1309, 331, 129, 1711, 454, 320, 673, 506, 1178, 1532, 1546, 1487, 1123, 493, 988, 970, 1106, 1397, 294, 451, 1126, 1449, 920, 1043, 1834, 268, 1535, 373]. simulation [1173, 629, 1336, 566, 364, 1641, 1283, 677, 688, 1587, 1154, 528, 876, 218, 308, 555, 1165, 1639, 1511, 326, 479, 1133, 643, 204, 1253, 1521, 365, 976, 35, 1275, 1643, 1107, 118]. Simulation-assisted [458]. Simulation-based [146, 209, 399]. Simulation-driven [340]. simulations [736, 1171, 565, 1199, 347, 887, 1027, 42, 1040, 551, 76, 1080, 1571, 602, 1911, 55, 1017, 1529, 1072, 231, 313, 83, 74, 1547, 1245, 1145, 747, 1419, 1900, 597, 206, 1882, 1450, 1373, 1615, 1921, 1898, 464, 961, 572, 1533, 433, 449, 165, 1358, 498, 368, 924, 1852, 1270, 1266, 513, 1603, 1569, 567, 411, 1380, 1655, 843, 1896, 919, 391]. simulator [1151, 598, 1078, 685]. simulators [1460]. Simultaneous [1176, 615, 623]. Sinc [341]. Sinc-Collocation [341]. Sine [1232, 163, 1799, 1405, 1360]. sine-Gordon [163]. Singapore [625]. single [504, 1752, 1178, 1358, 1429, 1927, 619, 1846, 1511, 1642]. single-cell [1358, 1429, 1642]. single-diode [1846]. single-interface [1511]. single-objective [1927]. single-relaxation-time [1178]. singular [163, 1830, 1002]. singularities [1866]. singularity [986]. singularly [1884]. sink [1854]. SIR [1102, 414]. SIR-based [1102]. SIRQV [351]. SIS [1731, 226]. site [700]. sites [1225, 932, 408, 709]. situ [1571, 1552, 1485, 1244, 131]. Six [1165]. size [271, 1916, 1901, 1384]. sized [1094, 159]. sizes [1876]. skeleton [47, 1693]. skeletons [29]. sketches [523]. skiing [506]. skills [477]. skin [1917]. skyline [1430]. sleep [1443]. slide [126]. slip [961]. slope [1455]. slow [629]. sludge [98]. slugging [276]. small [966, 845, 729, 7, 206, 1188, 1899, 877, 296]. small-scale [877]. small-world [7, 1899]. smallest [105]. Smart [459, 1762, 779, 246, 398, 1912, 377]. smart-card [377]. SmartSim [1533]. Smirnov [1696]. smog [1882]. smoke [431, 929, 1493]. smoothed [1702]. smoothly [755]. SMS [644]. SOARANN [1748]. SOARANN-fuzzy [1748]. Social [65, 968, 724, 310, 689, 1880, 71, 94, 1378, 1140, 938, 935, 812, 718, 772, 183, 943, 1200, 73, 939, 946, 1672, 508, 990, 764, 797, 932, 937, 827, 785, 792, 1633, 866, 1160, 1305, 1348, 1009, 644, 1021, 414, 865, 944, 868, 867, 858, 506, 459]. social-aware [183]. Society [1772, 841]. Socio [671, 1792, 689]. Socio-cognitive [1792]. Socio-cognitively [671]. socio-spatio-temporal [689]. sodium [1861]. Soft [216, 1751, 1110, 1415, 1786, 1145, 68, 796, 1217, 217]. Software [57, 1672, 1256, 75, 53, 1758, 54, 69, 1322, 55, 696, 79, 1445, 26, 473, 740, 248, 95, 46, 1314, 1330, 195, 810, 214, 783]. software-defined [696, 740]. soils [1415]. solar [1695, 1282, 513]. solid [1760, 1784, 555, 782, 411]. solidification [631, 1253]. Solubilities [498]. solubility [403]. Solution [38, 803, 280, 43, 34, 1703, 75, 1503, 1727, 135, 1702, 1469, 148, 139, 1047, 1698, 1609, 1718, 1464, 203, 653, 655, 1641, 430, 1413, 1573, 691, 1580, 1401, 1586, 1357]. Solutions [30, 127, 1152, 1874, 123, 1319, 1518, 168, 162, 161, 167, 1134, 350]. solve [577, 1434, 1512, 1299, 1887, 1033]. solved [1369]. solver [529, 1192, 1514, 1513, 923, 1302, 313, 1553, 1543, 1094, 1575, 1900, 1630, 149, 548, 9, 896, 136, 1262, 326, 1860, 919]. solvers [152, 1558, 1263, 1197, 154, 575, 1403, 592, 433, 1310, 438, 1866, 1304, 1839, 436, 1896]. solves [826]. Solving [202, 54, 510, 1628, 1774, 653, 655, 1578, 1465, 1367, 993, 145, 850, 1308, 391, 1864, 808, 1676, 187, 777, 1799, 715, 1233, 1749, 600, 1576, 1930, 1205, 612, 776, 1142, 770, 1492, 1073, 1802, 1886, 1307, 1840, 1811, 1092, 1196, 1551, 630, 1835]. Some [950, 427, 1916, 73, 26, 1413]. Sonar [1082]. SOR [959]. sort [1785]. sorting [714, 1625, 243]. sound [1670]. source [1051, 1256, 1648, 1520, 1823, 426, 1151, 1797, 279, 1049, 20, 1811, 1854, 1551, 1888, 614]. sources [311, 414]. sourcing [257]. soybean [839]. space [1674, 1080, 1936, 1149, 729, 1747, 176, 1437, 514, 1866, 1562, 1888, 1822]. space-dependent [1888]. space-splitting [1149]. spaces [1354]. spammers [724]. spanning [826]. spark [1121, 583]. Sparse [896, 1522, 1183, 577, 1404, 1302, 1482, 1931, 1502, 578, 518, 846, 1517, 1033]. sparsified [1860]. Sparsistent [1652]. Spatial [1028, 968, 24, 1826, 591, 1712, 739, 1547, 648, 388, 943, 814, 1771, 1724]. Spatially [1925, 1343, 1204, 204]. Spatio [191, 689, 1158, 1655]. Spatio-temporal [191, 1158, 1655]. SPE10 [373]. Special [230, 239, 265, 536, 1763, 286, 549, 1026, 753, 1324]. species [1372, 35]. species- [35]. Specific [586, 745, 404, 1710, 1106, 1310, 1793, 783, 130, 35]. specification [625]. specificity [260]. spectral [1098, 811, 1566, 168, 1231, 814, 1619, 1568, 630]. spectrin [1693]. spectrometry [133]. spectroscopic [409]. Spectroscopy [1773, 155, 839]. spectrum [1398]. spectrum-based [1398]. speed [1575, 378, 340, 877]. Speeding [1013]. Speedup [1663]. SPH [522, 927]. sphere [50, 88, 304]. spherical [1771]. Sphericity [983]. spike [1662]. spin [1436, 76, 120, 1541]. spine [1460]. spins [7]. spiral [1020, 1249, 50, 1540]. spiral-groove [1249]. spirometry [1567]. spline [434, 1759, 1616]. splines [141]. split [1445]. splitter [1671]. splitting [1585, 1149, 427, 335]. spraying [1025]. spread [24, 337, 737, 138, 77, 940]. spreaders [1114]. spreading [677]. spring [712]. Spy [112]. square [1660, 424]. squares [577, 1094, 1749, 1142, 483, 245, 1522]. SrH [1278]. SrHfO [78]. SRL [1803]. SRL-assisted [1803]. SRTI [1397]. ss [496]. ss-DNA [496]. SSA [1472]. St [209]. St. [146]. Stability [1372, 1906, 1854, 1198, 1004, 350, 1505, 1466]. stabilization [1487, 1630]. stabilized [1864, 1546]. stabilizing [182]. Stable [1897, 890, 1651, 1187, 1162, 193, 548, 1414]. stack [1470]. stack-type [1470]. stacked [834, 1461]. Stackelberg [1717]. stacker [399]. stacking [1059, 1455]. stage [242, 1443, 577, 354, 1780, 1475, 798]. staggered [1077]. stagnation [1854]. stakeholders [1680]. stall [1856]. star [1785]. StarSs [215]. start [965]. starting [1391]. STAT [404]. State [1243, 1242, 1363, 541, 554, 1303, 1720, 283, 908]. states [729, 1751, 1618]. static [1775, 1384, 367]. stationary [360, 824, 700, 989]. stations [1238]. Statistical [238, 1029, 163, 337, 472]. statistics [1408]. Steady [1363, 176, 1720]. Steady-state [1363, 1720]. Steam [1926]. steganalysis [648]. Steinbergs [66]. stem [1198]. stencil [1916, 223, 58]. stencil-based [223]. stenosis [1382, 1014]. stent [42, 749]. stents [1106]. step [1932, 1513, 1103, 231, 271, 987, 1384]. step-control [987]. step-size [271]. stepping [958, 1666]. Stereo [912]. stick [1477]. stiff [203]. stimulation [1605]. STM [585]. Stochastic [1093, 541, 632, 318, 447, 423, 201, 781, 1017, 1936, 1047, 83, 1685, 1381, 1541, 692, 629, 764, 474, 1899, 685, 1042, 134, 1307, 1105, 1091, 1569, 1060, 766, 540, 754, 1732, 365, 880, 1159]. stochastic-flow [692]. stock [41, 1215, 601, 460, 636, 582, 944]. Stokes [1421, 1932, 529, 1558, 434, 1263, 575, 1667, 1276, 572, 1583]. stone [1627]. storage [1714, 1187, 1080, 1280, 1039, 300, 1408, 862, 858]. store [1870]. storm [1820]. storms [375, 1882]. STQCA [1470]. STQCA-FFT [1470]. strain [154]. strategic [957]. strategies [1917, 1318, 1049, 1565, 1556, 1475, 981, 1471, 214, 58, 1709]. Strategy [547, 1104, 832, 701, 404, 1010, 1605, 1608, 1640, 1205, 1748, 870, 1587, 730, 174, 1822, 1050, 1472]. stratified [994]. straw [839]. stream [668, 1279]. streamflow [1566]. streaming [294, 1016, 1664]. streams [1484, 360, 948]. street [1284, 920]. strengths [290]. stress [522, 851, 154, 385, 1923, 1283, 61, 1083]. stretch [1120]. stretchable [1560]. stretching [1861, 1892]. striated [1117]. strict [1180]. Striving [1015]. structural [609, 462, 757, 267, 990, 405, 1088]. structurally [531]. Structure [401, 1860, 1278, 1823, 1529, 1318, 1301, 801, 599, 1547, 1606, 1850, 988, 1126, 1447, 981, 571, 1369, 1272, 1253, 1291]. Structure-exploiting [1860]. structure-preserving [1823]. structured [565, 574, 273, 1077]. Structures [1644, 462, 719, 1169, 1261, 231, 1139, 597, 593, 342, 1894, 1383]. Student [725]. Studies [557, 401, 127, 409, 815, 352, 287, 623]. Study [641, 1537, 968, 1581, 978, 850, 625, 72, 1183, 560, 1040, 737, 977, 306, 1376, 1911, 575, 221, 938, 49, 739, 1239, 1370, 1892, 403, 1175, 615, 405, 1627, 889, 240, 1607, 1337, 357, 1330, 487, 894, 1671, 1384, 432, 1340, 1839, 1511, 1820, 866, 1616, 1236, 791, 109, 877, 367, 569]. studying [1368]. Sub [223, 1195, 1354, 430]. sub-domains [430]. Sub-exponential [223]. Sub-Kolmogorov [1195]. sub-spaces [1354]. subdiffusion [1337]. subgroups [531]. Subject [881, 1775, 1788, 783]. Subject-independent [881]. subject-specific [783]. subjected [520]. submersible [495]. subsea [495]. subsequences [1130]. subsequent [1392]. subsonic [642]. subspaces [1749]. Substrate [834, 1365]. substructure [1398]. Subsystem [181]. subtasks [1807]. successful [1714]. succession [636]. suction [1854]. suggests [748, 1143, 638]. suitable [743, 1569]. suite [56]. sum [1718]. Summarization [1153]. super [1655]. super-resolution [1655]. supercomputers [53, 1870, 1485]. supernodal [896]. superseding [1020]. supersonic [1239, 919, 1409]. supervised [1826, 1022, 1131, 1214, 1803, 1766, 1497]. supplier [1132, 659, 333]. supply [1762, 914, 775, 274, 608, 540, 1762]. Support [394, 905, 125, 1862, 470, 1758, 1767, 627, 743, 146, 209, 333, 527, 413, 483, 124, 245, 617, 424, 1386]. supported [305]. supporting [441, 1072, 876]. Supra [1504]. Supra-Laplacian [1504]. surface [978, 1861, 385, 1365, 454, 555, 1489, 1036, 1498]. surfaces [565, 1308, 550, 1741]. surge [146, 209, 1820]. surgery [1460]. surgical [1460]. Surrogate [1375, 1451, 1723, 84, 1474, 1761, 17, 390, 883, 461, 1475, 848, 208]. Surrogate-assisted [1375, 1451, 1723]. Surrogate-based [84, 208]. surrogate-model-assisted [461]. surrogates [1145, 224, 517]. surveillance [1493, 812]. survey [872, 815, 652, 695]. survival [1461, 1370, 511, 1497]. suspect [605]. suspension [364]. Sustainability [396]. Sustainable [608, 1758, 659, 1745, 909]. sustained [419]. SVD [845]. SVM [799, 1083]. SVM-based [1083]. Swarm [27, 1110, 818, 809, 942, 1252, 974, 1364, 1137, 926, 1082, 96, 474, 651, 224, 1902, 869, 251, 933, 1062, 1822]. swarming [966]. swarms [1010, 1053, 1005, 1082]. swelling [1349]. SWIFT [928]. swim [136]. Swin [1813]. Swin-FlowNet [1813]. Swin-Transformer [1813]. Switchable [273]. switched [740]. Switching [1753, 538, 1060]. SWMM [1330]. symbiotic [866]. symbolic [314, 1821]. symmetric [75, 88, 1687]. symmetrical [976]. Symmetry [1077, 304]. Symmetry-preserving [1077]. Synapse [901]. synchronization [1149, 1218, 1383]. synchronous [593, 1736]. Synergy [1173]. syntactic [1672]. Synthesis [409, 1221, 1184, 1208, 475, 62]. Synthesizing [855]. Synthetic [375, 901, 1329]. System [1237, 272, 1186, 1066, 788, 668, 666, 669, 1308, 958, 466, 1779, 1670, 520, 434, 1089, 1313, 1872, 216, 472, 706, 1897, 1863, 1698, 921, 1605, 1065, 248, 36, 803, 1356, 333, 1005, 1458, 925, 829, 1012, 124, 1740, 1921, 920, 419, 1782, 246, 1283, 1280, 541, 1851, 1059, 1141, 1831, 1918, 617, 1737, 1369, 699, 1340, 694, 540, 1651, 830, 911, 933, 878, 868, 952, 954, 35, 1505, 666, 337, 394, 1238, 1073, 513]. system-on-chip [878]. Systematic [261, 424]. systems [65, 145, 965, 276, 151, 703, 1474, 719, 391, 1176, 396, 169, 1341, 1131, 727, 577, 1240, 702, 1676, 1404, 1298, 87, 1333, 707, 898, 601, 154, 1344, 271, 739, 693, 1053, 1094, 348, 855, 1309, 162, 203, 1645, 267, 538, 371, 519, 270, 593, 1834, 386, 524, 149, 240, 273, 1015, 730, 1539, 1716, 1658, 122, 1281, 1526, 249, 1569, 316, 344, 1894, 1595, 1511, 766, 582, 1170, 1026, 1185, 713, 856, 1732, 630, 853, 465, 858, 852, 1054, 670].

T [401, 231]. table [924]. tabular [1598]. tabulation [1725]. Tailored [1588, 1549]. Taiwan [780]. take [143]. talent [633]. Tamoxifen [1497]. Tamoxifen-treated [1497]. tandem [1671]. TAO [1231]. Target [68, 1203, 649, 966, 805, 401, 1748]. targets [552, 1712]. task [958, 1067, 1038, 1435, 974, 1364, 1825, 1599, 1608, 343, 606, 436, 853, 1584]. task-based [958]. task-parallelism [1067, 1038]. Tasks [1153, 956, 1005, 1631]. taxes [614]. taxonomic [810]. taxonomy [906]. TCP [740]. tea [1643]. Teaching [1565, 485]. Technical [1464]. technique [1436, 663, 977, 1483, 1149, 803, 1576, 607, 612, 1923, 1748, 335, 253, 810, 878]. Techniques [1066, 680, 1375, 416, 1567, 1075, 139, 1404, 1319, 1181, 925, 793, 1759, 1142, 1903, 1871, 686, 281, 359, 762, 573, 1720]. technologies [1320, 654, 876]. technology [1068, 849]. temperature [454, 1158, 1440, 308, 530]. templates [814]. Temporal [1443, 1566, 1146, 689, 943, 1158, 1899, 191, 1655, 1724]. tendon [1449]. tensile [61]. Tensor [1715, 1618, 1215, 1353, 516, 1600, 1039]. tensor-based [1215]. TensorFloat [1715]. TensorFloat-32 [1715]. terabyte [1094]. terabyte-sized [1094]. Term [1713, 1423, 885, 829, 1907, 1551, 1888]. terminal [911, 118]. terminals [1059]. terms [1041, 117]. terrains [1681]. test [1696, 187, 294, 176, 1390]. test-scenario [187]. testbeds [695]. testing [1398, 851, 419, 1314, 465]. tests [46, 1330]. tethered [1309]. tethered-wing [1309]. tetragonal [78]. tetrahedron [1819]. Text [1464, 242, 503, 650, 661, 943, 1076]. texture [967]. textured [555]. their [201, 831, 1294, 1392, 622]. theoretic [957, 992, 1800, 1627, 1168, 1732, 533]. theoretical [1511]. Theories [239]. Theory [1048, 841, 1826, 1022, 1131, 333, 277, 534, 586, 1584, 867, 1395]. therapeutic [552, 746]. therapy [1198, 327, 243, 1570]. Thermal [1788, 606, 1051, 903, 885, 647, 1861, 1201, 1509, 1350, 1821, 1854, 1639]. Thermal-aware [606]. thermal-lattice [1201]. thermalized [1569]. thermochemistry [927]. thermodiffusive [311]. thermodynamic [291]. thermophoresis [1788]. thermophysical [493]. thermoplastic [493]. thin [1229, 1573]. things [136, 1538, 1610, 1602, 1281, 960]. third [352, 1580]. third-order [352, 1580]. thousand [101]. thousand-digit [101]. Thread [1832]. Thread-safe [1832]. Threaded [372, 787]. threat [699]. Three [565, 522, 136, 558, 902, 242, 54, 1235, 1372, 738, 1711, 1751, 657, 550, 438, 1618]. three-body [1751, 1618]. Three-dimensional [565, 522, 558, 902, 54, 738, 1711]. three-pass [657]. three-stage [242]. threes [136]. threshold [1342, 805]. threshold-based [1342]. thresholding [250]. throughput [1164]. thyroid [1875]. Thyrotoxic [1288]. Ti6Al4Vartificial [1471]. ties [1742]. tiled [1254]. tilt [1165]. tilt-rotor [1165]. Time [1713, 175, 1577, 1397, 1582, 397, 514, 1692, 1703, 721, 1818, 642, 431, 1320, 1090, 152, 1192, 958, 1864, 956, 1621, 1881, 1738, 1753, 1513, 1764, 116, 1628, 1103, 1638, 1712, 1119, 56, 1830, 1363, 1047, 1795, 133, 417, 1698, 1331, 1359, 1566, 1005, 354, 63, 1178, 946, 1873, 1016, 1349, 386, 524, 1343, 1189, 937, 232, 904, 1384, 1866, 308, 1783, 709, 1374, 1595, 1907, 732, 1568, 1433, 1408, 1651, 350, 911, 1531, 1732, 1444, 1611, 878, 1666, 1729]. Time-based [397]. time-constrained [1005]. time-critical [946]. time-delayed [946]. time-dependent [1818, 642, 1192, 116, 1119, 63, 1343, 1568]. Time-efficient [1582]. time-fractional [1864, 1628, 1830]. time-harmonic [152]. time-marching [1866]. time-parallel [232]. time-propagation [63]. time-resolved [1408]. time-sensitive [937]. time-series [1444]. time-stepping [1666]. time-varying [1753, 350]. timely [1801]. Times [81, 116, 1047]. timetabling [315].

Timing [1506, 855]. tiny [844]. tinyMD [1373]. tips [1442]. tissue [1503, 1355, 1020, 746, 1786, 1126, 1358]. tolerance [482]. tolerant [219, 1246, 879, 1595]. tomograms [1659]. tomography [1422]. tonal [1700]. tone [750]. tool [616, 426, 707, 1072, 927, 547, 726, 1154, 908, 342]. toolbox [1648]. toolchain [269]. toolkit [1339, 199]. tools [279, 68, 215]. Top [215]. topic [841, 871, 952]. topographies [1573]. topological [1352, 1726]. Topologically [949, 565]. Topology [1437, 182, 1675, 1732]. Topology-based [1437]. TOPSIS [235]. toroidal [1911]. torsion [1702]. touch [392]. tournament [1030]. tournament-based [1030]. toxicity [260]. Toy [1728]. Track [1393, 334, 1371]. track-containing [334]. Tracker [1393]. tracking [466, 986, 1603]. tracks [971]. trading [6]. Traffic [696, 1029, 515, 1065, 763, 1161, 829, 920, 1488, 1894, 1134, 365, 1224, 1491]. traffic-aware [763]. trail [563]. Training [1068, 633, 485, 484, 1398, 68, 395, 509, 245, 275]. trajectories [560]. Trajectory [1911, 1454, 986, 1006]. trajectory-based [1454]. trans [462]. trans-level [462]. transaction [1766]. transactions [640]. transcriptional [259]. transcriptomes [1225]. transcriptomic [1462]. transductive [1778]. Transfer [1502, 1624, 1585, 1352, 1469, 522, 1712, 154, 1400, 984, 828, 1794, 253, 899, 323, 1893, 1517]. TransFlowNet [1655]. transform [667, 1855, 1452, 617, 1783, 1470]. Transformation [1396, 802]. transformations [1310, 713]. Transformer [1655, 1813]. transformers [1443]. Transforming [1316]. transient [903, 1818, 1685, 1205, 572]. transition [1161]. transitioning [1838]. transitions [1808]. Translating [1318, 1321, 1319, 991]. Translation [1324, 472]. Translational [1313, 1322, 1315, 1323, 1317, 1320, 1321]. Transmedia [275]. transmission [1693, 1900, 651]. transonic [162, 17]. transplantation [511, 1548]. Transport [1238, 993, 834, 931, 1250, 1366, 707, 1363, 117, 83, 1849, 176, 1927, 1894, 622]. Transportation [1546, 1576, 371, 519, 1798, 1204]. transporting [1477]. Transprecision [1577]. TraTSA [1577]. travel [116, 1107]. traveler [780]. Travelers [932]. travelling [1396, 1829]. traversal [1829]. traversal-dependent [1829]. trawl [390]. trawl-doors [390]. treated [1497]. treatment [1765, 87, 1365, 124, 130]. treatments [448]. tree [1254, 313, 108, 826, 1804, 59, 892]. tree-based [313]. trees [1785]. Trefftz [776]. trend [527]. trending [336]. Trends [1110, 115, 670]. Tri [1773, 1717, 1680, 1750]. Tri-Factorization [1773]. tri-level [1717]. tri-nanoadditives [1750]. tri-stakeholders [1680]. trials [1314]. triangle [1819]. tridiagonal [145, 1002]. triggered [911]. triggerless [418]. trigonometric [1586]. TripAdvisor [932]. Triplet [418, 1371]. triplex [131]. trivial [1281]. Trojans [906]. true [23]. Trust [933, 1378, 848, 830]. trustworthiness [867]. truthful [587]. TSD [244]. TSIRM [577]. tsunami [1615]. tube [54]. tumor [809, 129, 967, 996, 1770]. tumours [1370]. tunable [617]. tunable- [617]. TunaOil [1587]. tuning [416, 1093, 1587, 1852, 1270, 1266, 481]. tunnel [340, 1901]. tunnels [612]. turbulence [14, 5, 1195, 638]. turbulent [1194, 1216, 970, 1408]. turns [971].

Tutorial [1339]. tweets [952]. Twin [413]. Twitter [137, 41, 503, 941, 336, 636, 1076]. Two [42, 454, 568, 165, 1687, 1491, 1920, 1528, 463, 125, 1662, 642, 1560, 577, 1132, 278, 1363, 1486, 162, 354, 297, 1578, 430, 1468, 103, 1884, 1737, 1927, 578, 555, 1025, 1383, 1551, 303, 1896, 1611, 1591]. two-compartmental [1662]. Two-dimensional [42, 642, 162, 1927, 1551]. two-electron [278]. two-gas [1896]. Two-lane [1491]. Two-layer [454]. two-level [297]. Two-phase [568, 125, 1578, 1468, 555]. Two-point [1687]. two-relaxation-time [1363]. two-scale [430]. two-stage [577, 354]. two-time [1611]. two-way [1486]. Two-Way-Random [1591]. twofold [969]. type [1841, 1853, 1470, 205, 1886, 103, 1580, 1840, 159]. types [382].

U [205]. U-type [205]. U.S [739]. UAV [708]. UAVs [1865, 1099, 1690]. UFS [1493]. UFS-Net [1493]. Ukraine [1769]. ultra [857]. ultra-dense [857]. ultrametric [184]. ultrasonic [327, 1780]. ultrasonic-assisted [1780]. ultrasound [1589, 652]. UML [293]. unbalanced [1280]. uncertain [1753, 1383]. Uncertainty [1474, 1014, 1339, 1409, 1152, 233, 1328, 426, 519, 1926, 1274, 1064, 1223, 1230, 521, 1523, 480, 1844, 1070]. uncertainty-quantification [1523]. Unconditionally [1162]. unconstrained [325]. under-resolved [1194]. undergraduate [1495]. understand [349]. Understanding [1293, 645, 947, 1221, 1184, 1208, 1206, 1084, 1210, 1070, 1134, 1790]. undirected [1517]. unfolding [4]. Unified [1017, 1752, 1493]. uniform [1775, 642, 220, 88]. uniformly [50, 304, 1773]. unimodal [241]. unimolecular [288]. Unit [1330, 855, 803, 50, 88, 304, 1541, 880]. unit-based [1541]. Unite [478]. units [202, 1250, 1627, 382, 781]. units-of-measure [382]. unity [1024]. Universal [1211, 648, 112]. University [315]. unknown [1049, 1595, 766]. Unleashing [1196]. unmanned [1435, 1054]. unrelated [1499]. unsharp [669]. Unstable [5]. unsteady [1513, 602, 1231]. unstructured [574, 1197, 1261, 924, 1803, 1033]. unsupervised [242]. Unveil [1215]. unwrapping [1553]. Update [1472]. updating [192, 1638, 1723, 530]. Upgrades [1772]. Uphill [4]. upscaling [1821]. UQpy [1223, 1230]. Urban [968, 1155, 117, 739, 1547, 1065, 1882, 732, 366, 1919, 458, 1934, 1224]. urgent [626]. Usability [728]. usage [969, 1542, 1922, 1459]. use [1460, 1149, 1526, 836, 118]. used [1141, 1562]. useful [1423]. usefulness [601]. user [666, 466, 833, 1334, 792, 871]. user-oriented [666]. users [689, 1694]. Using [628, 76, 781, 1763, 309, 1533, 259, 1050, 1611, 424, 844, 1174, 818, 841, 616, 1538, 819, 1528, 1695, 1121, 835, 903, 609, 202, 250, 666, 875, 470, 1812, 1567, 315, 1308, 1826, 152, 1761, 510, 1598, 1443, 1462, 1588, 396, 163, 682, 760, 1670, 805, 737, 1022, 983, 522, 687, 1513, 295, 141, 761, 667, 1561, 591, 845, 429, 1915, 1122, 1764, 602, 1460, 1911, 806, 1398, 1787, 1746, 472, 627, 79, 1421, 1825, 385, 595, 1797, 1470, 1667, 1292, 221, 1415, 1493, 1365, 1786, 74, 1843, 1139, 402, 661, 811, 799, 812, 1049, 348, 779, 763]. using [803, 1356, 333, 455, 926, 1225, 1306, 1849, 297, 422, 1200, 1574, 1640, 1120, 17, 390, 243, 1423, 294, 1534, 1649, 1672, 1600, 1373, 1231, 1447, 386, 524, 961, 1590, 1923, 1195, 771, 618, 653, 655, 1578, 387, 1502, 1343, 1748, 1465, 1922, 1540, 932, 430, 1360, 314, 896, 1393, 1058, 691, 84, 208, 1437, 756, 511, 1146, 1856, 613, 517, 1441, 1411, 586, 967, 1188, 1802, 1477, 102, 996, 459, 300, 705, 425, 1886, 617, 513, 972, 894, 765, 1671, 807, 492, 252, 652, 1902, 1440, 1105, 15, 1927, 518, 1339, 792, 1180, 282, 1631, 619, 782, 491, 1427, 1644]. using [766, 326, 1380, 798, 881, 1810, 1185, 1009, 436, 933, 109, 1497, 1444, 853, 1386, 861, 838]. utility [534, 508]. Utilizing [181]. UWB [1502].

vacancies [1080]. vaccination [1102, 1506]. Vacuum [278, 403]. Validating [189]. Validation [994, 97, 1339, 1433, 707, 750, 320, 1871, 644]. validity [1809]. Valley [451]. valuation [1024, 928]. value [1174, 1863, 1755, 1002]. valued [1740]. values [564, 1714, 838]. valve [745, 1301]. VANET [1237]. VaNetLayer [441]. VAPER [1593]. vapor [499, 1929]. variability [598, 673, 377]. Variable [789, 25, 520, 1628, 1103, 389, 1525, 1380, 1395, 1519]. variable-fidelity [389]. Variance [629, 427, 1518, 1037]. Variance-reduced [629]. variant [732]. variants [850, 1506]. variation [1509, 1849, 709]. variational [1585, 1341, 1667, 1609, 1309, 460, 1535]. Variations [1724, 674, 1079]. varied [1669]. variety [550]. various [930, 174]. varying [1753, 1095, 926, 454, 350]. vascular [969, 535, 349, 381, 168, 750, 1532, 1126]. vasculature [368, 619]. Vascún [451]. VCCC [400, 549, 286]. VCCC-2013 [286]. VCCC-2014 [400]. VCCC-2015 [549]. VECMA [1339]. vector [905, 1862, 627, 1482, 527, 413, 483, 245, 103, 1697, 424, 1459]. Vehicle [1047, 777, 1251, 615, 653, 655]. vehicles [641, 1435, 1488, 458, 1224, 1707, 1526, 1224]. vehicular [441]. velocities [1235, 1161, 1260]. velocity [1128, 1235, 440, 961, 989]. ventilator [49]. ventilator-dependent [49]. Verification [1398, 1638, 1339, 1575, 1882]. version [283]. versus [615, 1590, 797, 505]. Vertex [1804, 1428, 770]. vertical [1568]. vertices [1785, 983]. Very [464, 1104, 1207]. vessel [667]. VHDL [1072]. VHDL-based [1072]. via [1152, 463, 930, 1215, 1353, 1594, 1895, 201, 667, 1460, 1412, 1372, 1400, 728, 936, 855, 800, 1175, 1768, 1663, 890, 1354, 541, 335, 1015, 1821, 323, 1918, 1783, 1489, 1639, 1619, 1694, 804, 989]. viability [1485]. vibration [445, 1395]. vibrational [405]. vibrissae [563]. video [666, 591, 1493]. video-based [666]. Vietoris [1920]. View [808, 66, 33]. viewpoint [743]. vine [760]. Viral [77, 172, 1507, 258, 1827]. Virtual [286, 641, 832, 1017, 1115, 763, 387, 1314, 1154, 1118, 1071, 742, 400, 549]. virtualization [441]. virtualized [956]. virus [1849, 1900]. virus-laden [1849]. viscoelastic [970]. viscosity [1583]. viscous [306]. Visitors [1332]. vista [888]. Visual [19, 823, 669, 1833, 545, 960]. visualisation [7]. Visualization [1072, 19, 327, 675, 1061]. visualizing [1719, 801]. visually [1720]. vitro [172]. VMs [587]. volatility [632, 1119, 1306, 1105, 1060, 754]. voltammetry [449]. volume [1436, 1116, 1558, 958, 1440, 1842]. volumes [1104]. Voronoi [1275, 132]. VOVS [513]. VPH [752, 783]. VPH-HF [783]. VPIC [1459]. VRP [1798, 1066]. vs [905]. vs. [978]. Vulnerability [751, 698, 810, 35].

wake [1671]. WaLBerla [55]. walk [1594, 949, 1292, 309, 597]. walks [1889, 455, 425]. walls [1560, 1301, 1745]. waning [1102]. warehouse [1684, 1859]. warning [1906, 386, 524]. Wasatch [586]. Wasserstein [1649]. waste [1798]. WatchMan [197]. Water [505, 1146, 1929, 1528, 276, 175, 1036]. watermarking [607]. Waters [419]. Watson [849]. wave [978, 850, 1828, 175, 1328, 1020, 1061, 1817, 162, 1129, 1540, 150, 430, 1092]. wave-packet [150]. waveguide [231]. Wavelet [25, 1307, 667, 1628, 617, 1783]. wavelets [135, 691, 1886, 1367]. waves [1882]. Way [1591, 418, 1801, 1486, 1432, 1768]. Weak [1150, 5, 1884, 167]. weakly [1830, 1312]. weaning [49]. weather [1274]. web [703, 441, 705, 791, 185, 1181, 1069, 700]. webgraph [48]. WECOI [1484]. weight [1881, 1114, 1467]. Weighted [1279, 1909, 1277, 360, 1534, 1760, 1732, 1484]. weighting [1132]. weights [1386]. welding [522]. wet [616]. wetted [555]. we’ve [1207]. Whale [1004]. wheelchairs [779]. which [534]. while [594]. whirl [877]. whole [126]. Wide [491, 1311, 1197, 697]. wide-SIMD [1197]. WIFIRE [1317]. Wikipedia [599, 66]. Wild [954]. wildfire [448, 677, 1489]. Wildland [38, 1221, 1184, 1208]. Wind [929, 14, 340, 1280, 480, 580, 1489]. windows [1047]. wing [1309]. wireless [299, 442, 182]. wires [1025]. wise [1399, 1390]. within [1366, 441, 69, 1126, 1871, 1923, 1511]. within-project [1871]. without [174]. Wolf [354, 898, 1492, 1396, 1081]. word [1058, 948]. work [1600]. Workflow [490, 69, 1313, 1244, 1017, 1301, 262, 493, 397]. Workflows [1163, 8, 955, 854, 1608, 623, 783]. workload [791]. workloads [771, 1587]. workplace [1283, 317]. World [1234, 682, 1746, 7, 66, 762, 1899, 465]. worm [644]. wrong [759]. WSC [1786]. WSC-integrated [1786]. WSN [295].

X [1563]. X-ray [1563]. XAI [1433]. Xeon [581]. XU4 [523].

years [1219, 1338]. yields [13]. Young [143]. YUKI [1394]. Yuva [646].

Zakharov [1698, 1651]. zeolite [1245]. zero [1718]. zero-sum [1718]. ZIF [889]. ZIF-8 [889].

References

Seidel:2010:P

[1] Edward Seidel and Jeannette M. Wing. Preface. Journal of Computational Science, 1(1):1–2, May 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000189.

Sloot:2010:P

[2] Peter Sloot, Peter Coveney, and Jack Dongarra. Preface. Journal of Computational Science, 1(1):3–4, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000177.

Anonymous:2010:PN

[3] Anonymous. Publisher’s note. Journal of Computational Science, 1(1):5, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750310000050.

Albrecht:2010:UUN

[4] A. A. Albrecht, L. Kapsokalivas, and K. Steinhöfel. Uphill unfolding of native protein conformations in cubic lattices. Journal of Computational Science, 1(1):6–12, May 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000025.

Fazendeiro:2010:UPO

[5] L. Fazendeiro, B. M. Boghosian, P. V. Coveney, and J. Lätt. Unstable periodic orbits in weak turbulence. Journal of Computational Science, 1(1):13–23, May 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000062.

Pellicer-Lostao:2010:CGL

[6] C. Pellicer-Lostao and R. López-Ruiz. A chaotic gas-like model for trading markets. Journal of Computational Science, 1(1):24–32, May 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000074.

Leist:2010:IVS

[7] A. Leist, D. P. Playne, and K. A. Hawick. Interactive visualisation of spins and clusters in regular and smallworld Ising models with CUDA on GPUs. Journal of Computational Science, 1(1):33–40, May 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000153.

Abramson:2010:EOC

[8] David Abramson, Blair Bethwaite, Colin Enticott, Slavisa Garic, Tom Peachey, Anushka Michailova, and Saleh Amirriazi. Embedding optimization in computational science workflows. Journal of Computational Science, 1(1):41–47, May 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000165.

Paszynski:2010:PMF

[9] Maciej Paszyński, David Pardo, and Anna Paszyńska. Parallel multi-frontal solver for p adaptive finite element modeling of multi-physics computational problems. Journal of Computational Science, 1(1):48–54, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000049.

Manson:2010:DRA

[10] J. Russell Manson and Robert J. Olsen. Diagnostics and rubrics for assessing learning across the computational science curriculum. Journal of Computational Science, 1(1):55–61, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750310000141.

Anonymous:2010:EBa

[11] Anonymous. Editorial Board. Journal of Computational Science, 1 (1):CO2, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750310000244.

Anonymous:2010:PM

[12] Anonymous. Pages 1–62 (May 2010). Journal of Computational Science, 1 (1):??, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Fenn:2010:CCE

[13] Daniel J. Fenn, Zhenyuan Zhao, Pak Ming Hui, and Neil F. Johnson. Competitive carbon emission yields the possibility of global selfcontrol. Journal of Computational Science, 1(2):63–74, June 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750310000037.

DeTommasi:2010:DWF

[14] Luciano De Tommasi, Madeleine Gibescu, and Arno J. Brand. A dynamic wind farm aggregate model for the simulation of power fluctuations due to wind turbulence. Journal of Computational Science, 1(2):75–81, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031000013X.

Southern:2010:SCE

[15] J. Southern, G. J. Gorman, M. D. Piggott, P. E. Farrell, M. O. Bernabeu, and J. Pitt-Francis. Simulating cardiac electrophysiology using anisotropic mesh adaptivity. Journal of Computational Science, 1(2):82–88, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000128.

Buntinas:2010:MDC

[16] Darius Buntinas, Alexis J. Malozemoff, and Jean Utke. Multithreaded derivative computation with generated libraries. Journal of Computational Science, 1(2):89–97, June 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000116.

Leifsson:2010:MFD

[17] Leifur Leifsson and Slawomir Koziel. Multi-fidelity design optimization of transonic airfoils using physicsbased surrogate modeling and shapepreserving response prediction. Journal of Computational Science, 1 (2):98–106, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000098.

Aberer:2010:PPP

[18] Andre J. Aberer, Nicholas D. Pattengale, and Alexandros Stamatakis. Parallelized phylogenetic post-analysis on multi-core architectures. Journal of Computational Science, 1 (2):107–114, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750310000086.

Cai:2010:VLB

[19] Yang Cai, Rafael de M. Franco, and Manuel García-Herranz. Visual latency-based interactive visualization for digital forensics. Journal of Computational Science, 1 (2):115–120, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750310000207.

Lawniczak:2010:ISL

[20] Anna T. Lawniczak and Shengkun Xie. Impact of source load and routing on QoS of packets delivery. Journal of Computational Science, 1(2):121–129, June 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750310000190.

Anonymous:2010:EBb

[21] Anonymous. Editorial Board. Journal of Computational Science, 1 (2):CO2, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750310000335.

Anonymous:2010:PJ

[22] Anonymous. Pages 63–130 (June 2010). Journal of Computational Science, 1(2):??, June 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Sloot:2010:CDR

[23] Peter M. A. Sloot. The crossdisciplinary road to true computational science. Journal of Computational Science, 1(3):131, August 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000451.

Balcan:2010:MSS

[24] Duygu Balcan, Bruno Gonçalves, Hao Hu, José J. Ramasco, Vittoria Colizza, and Alessandro Vespignani. Modeling the spatial spread of infectious diseases: the GLobal Epidemic and Mobility computational model. Journal of Computational Science, 1(3): 132–145, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000438.

Chen:2010:WMC

[25] Yiming Chen, Yongbing Wu, Yuhuan Cui, Zhuangzhuang Wang, and Dongmei Jin. Wavelet method for a class of fractional convection-diffusion equation with variable coefficients. Journal of Computational Science, 1(3): 146–149, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000426.

Glavelis:2010:CES

[26] Themistoklis Glavelis, Nikolaos Ploskas, and Nikolaos Samaras. A computational evaluation of some free mathematical software for scientific computing. Journal of Computational Science, 1(3):150–158, August 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000414.

Beji:2010:HPS

[27] Noura Beji, Bassem Jarboui, Mansour Eddaly, and Habib Chabchoub. A hybrid particle swarm optimization algorithm for the redundancy allocation problem. Journal of Computational Science, 1(3):159–167, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000402.

Chorley:2010:PAH

[28] Martin J. Chorley and David W. Walker. Performance analysis of a hybrid MPI/OpenMP application on multi-core clusters. Journal of Computational Science, 1(3): 168–174, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000396.

Adamatzky:2010:ES

[29] Andrew Adamatzky. On excitable βskeletons. Journal of Computational Science, 1(3):175–186, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031000044X.

Anonymous:2010:CPA

[30] Anonymous. Call for paper: Advanced computing solutions for health care and medicine. Journal of Computational Science, 1(3):187, August 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000566.

Anonymous:2010:EBc

[31] Anonymous. Editorial Board. Journal of Computational Science, 1 (3):CO2, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000505.

Anonymous:2010:PA

[32] Anonymous. Pages 131–188 (August 2010). Journal of Computational Science, 1(3):??, August 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Sloot:2010:CSK

[33] Peter M. A. Sloot. Computational science: a kaleidoscopic view into science. Journal of Computational Science, 1(4):189, December 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000694.

Abbasi:2010:NIL

[34] A. Abbasi and B. Vosoughi Vahdat. A non-iterative linear inverse solution for the block approach in EIT. Journal of Computational Science, 1 (4):190–196, December 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031000061X.

Zheng:2010:ESS

[35] Huiru Zheng, Haiying Wang, and Francisco Azuaje. eNelator: a simulation system for large-scale vulnerability analysis of species-, diseaseand process-specific protein networks. Journal of Computational Science, 1 (4):197–205, December 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000608.

Jiang:2010:PSS

[36] Nanyan Jiang and Manish Parashar. A programming system for sensorbased scientific applications. Journal of Computational Science, 1(4): 206–220, December 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000487.

Tofani:2010:OAS

[37] Alberto Tofani, Elisa Castorini, Paolo Palazzari, Andrij Usov, Cesaire Beyel, Erich Rome, and Paolo Servillo. An ontological approach to simulate critical infrastructures. Journal of Computational Science, 1(4): 221–228, December 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000475.

Bianchini:2010:WFG

[38] Germán Bianchini, Mónica Denham, Ana Cortés, Tomàs Margalef, and Emilio Luque. Wildland fire growth prediction method based on multiple overlapping solution. Journal of Computational Science, 1(4): 229–237, December 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000463.

Anonymous:2010:EBd

[39] Anonymous. Editorial Board. Journal of Computational Science, 1(4): CO2, December 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000633.

Anonymous:2010:PD

[40] Anonymous. Pages 189–238 (December 2010). Journal of Computational Science, 1(4):??, December 2010. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Bollen:2011:TMP

[41] Johan Bollen, Huina Mao, and Xiaojun Zeng. Twitter mood predicts the stock market. Journal of Computational Science, 2(1):1–8, March 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031100007X.

Caiazzo:2011:CAA

[42] Alfonso Caiazzo, David Evans, Jean-Luc Falcone, Jan Hegewald, Eric Lorenz, Bernd Stahl, Dinan Wang, Jörg Bernsdorf, Bastien Chopard, Julian Gunn, Rod Hose, Manfred Krafczyk, Pat Lawford, Rod Smallwood, Dawn Walker, and Alfons Hoekstra. A complex automata approach for in-stent restenosis: Twodimensional multiscale modelling and simulations. Journal of Computational Science, 2(1):9–17, March 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000056.

Save:2011:SPD

[43] Yogesh Dilip Save, H. Narayanan, and Sachin B. Patkar. Solution of partial differential equations by electrical analogy. Journal of Computational Science, 2(1):18–30, March 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000044.

Easwaramoorthy:2011:IGF

[44] D. Easwaramoorthy and R. Uthayakumar. Improved generalized fractal dimensions in the discrimination between healthy and epileptic EEG signals. Journal of Computational Science, 2(1):31–38, March 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000032.

Iliopoulos:2011:PIA

[45] A. P. Iliopoulos, J. G. Michopoulos, S. G. Lambrakos, and N. Bernstein. Performance of inverse atomistic scale fracture modeling on GPGPU architectures. Journal of Computational Science, 2(1):39–46, March 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000020.

Kelly:2011:ERD

[46] Diane Kelly, Robert Gray, and Yizhen Shao. Examining random and designed tests to detect code mistakes in scientific software. Journal of Computational Science, 2(1):47–56, March 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031000075X.

Alonso-Sanz:2011:BSA

[47] Ramón Alonso-Sanz and Andrew Adamatzky. On beta-skeleton automata with memory. Journal of Computational Science, 2(1): 57–66, March 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000748.

daSilva:2011:MWE

[48] Roberto da Silva, Luciana S. Buriol, Leila Ribeiro, and Fernando L. Dotti. Modeling the webgraph evolution. Journal of Computational Science, 2(1):67–79, March 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000736.

Huang:2011:SWR

[49] Chao-Yi Huang and Jong-Chen Chen. On the study of the weaning results of ventilator-dependent patients with closest reasonable centroids. Journal of Computational Science, 2 (1):80–87, March 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750310000724.

Koay:2011:AES

[50] Cheng Guan Koay. Analytically exact spiral scheme for generating uniformly distributed points on the unit sphere. Journal of Computational Science, 2(1):88–91, March 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750310000761.

Anonymous:2011:EBa

[51] Anonymous. Editorial Board. Journal of Computational Science, 2 (1):CO2, March 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000202.

Anonymous:2011:PM

[52] Anonymous. Pages 1–92 (March 2011). Journal of Computational Science, 2(1):??, March 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Bader:2011:SSS

[53] Michael Bader, Miriam Mehl, Ulrich Rüde, and Gerhard Wellein. Simulation software for supercomputers. Journal of Computational Science, 2(2):93–94, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000342.

Bucker:2011:SPE

[54] H. Martin Bücker, Oliver Fortmeier, and Monika Petera. Solving a parameter estimation problem in a threedimensional conical tube on a parallel and distributed software infrastructure. Journal of Computational Science, 2(2):95–104, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000068.

Feichtinger:2011:WHS

[55] C. Feichtinger, S. Donath, H. Köstler, J. Götz, and U. Rüde. WaL-Berla: HPC software design for computational engineering simulations. Journal of Computational Science, 2 (2):105–112, May 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000111.

Geveler:2011:SSL

[56] Markus Geveler, Dirk Ribbrock, Sven Mallach, and Dominik Göddeke. A simulation suite for lattice-Boltzmann based real-time CFD applications exploiting multi-level parallelism on modern multi- and many-core architectures. Journal of Computational Science, 2(2):113–123, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000159.

Buchholz:2011:SDH

[57] M. Buchholz, H.-J. Bungartz, and J. Vrabec. Software design for a highly parallel molecular dynamics simulation framework in chemical engineering. Journal of Computational Science, 2(2):124–129, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000160.

Treibig:2011:EMA

[58] Jan Treibig, Gerhard Wellein, and Georg Hager. Efficient multicoreaware parallelization strategies for iterative stencil computations. Journal of Computational Science, 2 (2):130–137, May 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000172.

Speck:2011:TPT

[59] R. Speck, L. Arnold, and P. Gibbon. Towards a petascale tree code: Scaling and efficiency of the PEPC library. Journal of Computational Science, 2(2):138–143, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000184.

Capobianco:2011:NEB

[60] Enrico Capobianco. On network entropy and bio-interactome applications. Journal of Computational Science, 2(2):144–152, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000093.

Ouinas:2011:NMI

[61] D. Ouinas, B. Bachir Bouiadjra, N. Benderdouche, B. Ait Saadi, and J. Viña. Numerical modelling of the interaction macro–multimicrocracks in a plate under tensile stress. Journal of Computational Science, 2 (2):153–164, May 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031100010X.

Mattiussi:2011:BGN

[62] Claudio Mattiussi, Peter Dürr, Daniel Marbach, and Dario Floreano. Beyond graphs: a new synthesis. Journal of Computational Science, 2 (2):165–177, May 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000147.

Kormann:2011:GEC

[63] K. Kormann, S. Holmgren, and H. O. Karlsson. Global error control of the time-propagation for the Schrödinger equation with a time-dependent Hamiltonian. Journal of Computational Science, 2(2):178–187, May 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000287.

Anonymous:2011:EBb

[64] Anonymous. Editorial Board. Journal of Computational Science, 2 (2):CO2, May 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000366.

Agarwal:2011:SCS

[65] Nitin Agarwal and Xiaowei Xu. Social computational systems. Journal of Computational Science, 2(3): 189–192, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000688.

Overell:2011:VWA

[66] S. E. Overell and S. Rüger. View of the world according to Wikipedia: Are we all little Steinbergs? Journal of Computational Science, 2(3): 193–197, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000494.

Safar:2011:NRI

[67] Maytham Safar, Khaled Mahdi, and Sadeq Torabi. Network robustness and irreversibility of information diffusion in complex networks. Journal of Computational Science, 2(3): 198–206, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000482.

Kantola:2011:TTS

[68] Jussi Kantola, Antti Piirto, Jarmo Toivonen, Yoon Chang, and Hannu Vanharanta. Target training with soft computing tools. Journal of Computational Science, 2(3): 207–215, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000500.

Cakici:2011:WSD

[69] Baki Cakici and Magnus Boman. A workflow for software development within computational epidemiology. Journal of Computational Science, 2(3):216–222, August 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000470.

Berlingerio:2011:PHA

[70] Michele Berlingerio, Michele Coscia, Fosca Giannotti, Anna Monreale, and Dino Pedreschi. The pursuit of hubbiness: Analysis of hubs in large multidimensional networks. Journal of Computational Science, 2(3): 223–237, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000524.

Dabbaghian:2011:SIH

[71] Vahid Dabbaghian, Valerie Spicer, Suraj K. Singh, Peter Borwein, and Patricia Brantingham. The social impact in a high-risk community: a cellular automata model. Journal of Computational Science, 2(3): 238–246, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000512.

Brahim:2011:CAB

[72] Abdelhamid Salah Brahim, Bénédicte Le Grand, Lionel Tabourier, and Matthieu Latapy. Citations among blogs in a hierarchy of communities: Method and case study. Journal of Computational Science, 2(3): 247–252, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000536.

Leung:2011:ESC

[73] Ricky C. Leung and Kalyan S. Pasupathy. The economics of social computing: Some preliminary findings on healthcare organizations. Journal of Computational Science, 2(3): 253–261, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031100055X.

Huntemann:2011:DDA

[74] Marcus Huntemann, Georg Heygster, and Gang Hong. Discrete dipole approximation simulations on GPUs using OpenCL — application on cloud ice particles. Journal of Computational Science, 2(3):262–271, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000548.

Auckenthaler:2011:DAS

[75] T. Auckenthaler, H.-J. Bungartz, T. Huckle, L. Krämer, B. Lang, and P. Willems. Developing algorithms and software for the parallel solution of the symmetric eigenvalue problem. Journal of Computational Science, 2(3):272–278, August 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000330.

Cortie:2011:UCF

[76] David Cortie and John Pillans. Using a custom-FPGA architecture to extend the scale of atomistic magnetic spin simulations. Journal of Computational Science, 2(3):279–285, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000317.

Youssef:2011:VCQ

[77] Mina Youssef, Robert Kooij, and Caterina Scoglio. Viral conductance: Quantifying the robustness of networks with respect to spread of epidemics. Journal of Computational Science, 2(3):286–298, August 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000299.

Feng:2011:FPC

[78] Li ping Feng, Zheng tang Liu, and Qi jun Liu. First-principles calculations on electronic, chemical bonding and optical properties of tetragonal SrHfO_{3}. Journal of Computational Science, 2(3):299–303, August 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000275.

Friborg:2011:RDS

[79] Rune Møllegaard Friborg and Brian Vinter. Rapid development of scalable scientific software using a process oriented approach. Journal of Computational Science, 2(3): 304–313, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000263.

Anonymous:2011:EBc

[80] Anonymous. Editorial Board. Journal of Computational Science, 2 (3):CO2, August 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000743.

Shublaq:2011:ETC

[81] Nour Shublaq, Stefan Zasada, and Peter Coveney. Exciting times in computational science. Journal of Computational Science, 2 (4):315, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000925.

Tria:2011:FNR

[82] Francesca Tria, Animesh Mukherjee, Andrea Baronchelli, Andrea Puglisi, and Vittorio Loreto. A fast no-rejection algorithm for the Category Game. Journal of Computational Science, 2 (4):316–323, December 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000883.

Helmuth:2011:ESS

[83] Jo A. Helmuth, Sylvain Reboux, and Ivo F. Sbalzarini. Exact stochastic simulations of intra-cellular transport by mechanically coupled molecular motors. Journal of Computational Science, 2(4):324–334, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000846.

Priess:2011:SBO

[84] M. Prieß, S. Koziel, and T. Slawig. Surrogate-based optimization of climate model parameters using response correction. Journal of Computational Science, 2(4):335–344, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031100072X.

Neme:2011:EPM

[85] Antonio Neme, Sergio Hernández, and Omar Neme. An electoral preferences model based on self-organizing maps. Journal of Computational Science, 2 (4):345–352, December 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000718.

Pendharkar:2011:MAM

[86] Parag C. Pendharkar. A multi-agent memetic algorithm approach for distributed object allocation. Journal of Computational Science, 2(4): 353–364, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000706.

Emans:2011:CGT

[87] Maximilian Emans. Coarse-grid treatment in parallel AMG for coupled systems in CFD applications. Journal of Computational Science, 2(4): 365–376, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031100069X.

Koay:2011:SSG

[88] Cheng Guan Koay. A simple scheme for generating nearly uniform distribution of antipodally symmetric points on the unit sphere. Journal of Computational Science, 2(4):377–381, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000615.

Anonymous:2011:EBd

[89] Anonymous. Editorial Board. Journal of Computational Science, 2(4): CO2, December 2011. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000949.

Anonymous:2011:PD

[90] Anonymous. Pages 315–382 (December 2011). Journal of Computational Science, 2(4):??, December 2011. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2012:EBa

[91] Anonymous. Editorial Board. Journal of Computational Science, 3(1– 2):CO2, ???? 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000142.

Anzai:2012:OFD

[92] Hitomi Anzai, Makoto Ohta, Jean-Luc Falcone, and Bastien Chopard. Optimization of flow diverters for cerebral aneurysms. Journal of Computational Science, 3(1–2):1–7, ???? 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000026.

Southern:2012:PAM

[93] J. Southern, G. J. Gorman, M. D. Piggott, and P. E. Farrell. Parallel anisotropic mesh adaptivity with dynamic load balancing for cardiac electrophysiology. Journal of Computational Science, 3(1–2):8–16, ???? 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001025.

Giabbanelli:2012:MIS

[94] Philippe J. Giabbanelli, Azadeh Alimadad, Vahid Dabbaghian, and Diane T. Finegood. Modeling the influence of social networks and environment on energy balance and obesity. Journal of Computational Science, 3(1–2):17–27, ???? 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000063.

Kang:2012:IMA

[95] Pilsung Kang, Naresh K. C. Selvarasu, Naren Ramakrishnan, Calvin J. Ribbens, Danesh K. Tafti, Yang Cao, and Srinidhi Varadarajan. Implementing modular adaptation of scientific software. Journal of Computational Science, 3(1–2):28–45, ???? 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000099.

Lopez:2012:OCG

[96] Luis Fernando de Mingo López, Nuria Gómez Blas, and Alberto Arteta. The optimal combination: Grammatical swarm, particle swarm optimization and neural networks. Journal of Computational Science, 3(1– 2):46–55, ???? 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311001128.

Roadknight:2012:VMP

[97] Chris Roadknight, Uwe Aickelin, and Galina Sherman. Validation of a microsimulation of the Port of Dover. Journal of Computational Science, 3 (1–2):56–66, ???? 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000676.

Uthayakumar:2012:MSC

[98] R. Uthayakumar, G. Arockia Prabakar, and S. Abdul Azis. Multifractal scaling of crack images from pyroligneous acid dried sludge. Journal of Computational Science, 3(1–2):67–73, ???? 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001074.

Anonymous:2012:PJM

[99] Anonymous. Pages 1–74 (January– March 2012). Journal of Computational Science, 3(1–2):??, ???? 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

deDoncker:2012:ACM

[100] Elise de Doncker. Advances in computational methods: a compilation. Journal of Computational Science, 3(3):75–76, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000294.

Bailey:2012:HHC

[101] David H. Bailey and Jonathan M. Borwein. Hand-to-hand combat with thousand-digit integrals. Journal of Computational Science, 3 (3):77–86, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750310000773.

Saito:2012:AGM

[102] Tsubasa Saito, Emiko Ishiwata, and Hidehiko Hasegawa. Analysis of the GCR method with mixed precision arithmetic using QuPAT. Journal of Computational Science, 3 (3):87–91, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750311000329.

Sidi:2012:RTV

[103] Avram Sidi. Review of two vector extrapolation methods of polynomial type with applications to large-scale problems. Journal of Computational Science, 3(3):92–101, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000123.

deDoncker:2012:QCF

[104] Elise de Doncker, Junpei Fujimoto, Nobuyuki Hamaguchi, Tadashi Ishikawa, Yoshimasa Kurihara, Yoshimitsu Shimizu, and Fukuko Yuasa. Quadpack computation of Feynman loop integrals. Journal of Computational Science, 3(3):102–112, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750311000573.

Luo:2012:DMD

[105] Xin long Luo. A dynamical method of DAEs for the smallest eigenvalue problem. Journal of Computational Science, 3(3):113–119, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031200004X.

Karlin:2012:MMP

[106] Ian Karlin, Elizabeth Jessup, and Erik Silkensen. Modeling the memory and performance impacts of loop fusion. Journal of Computational Science, 3(3):120–126, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000305.

Bigham:2012:NPP

[107] Bahram Sadeghi Bigham, Marzieh Eskandari, and Maryam Tahmasbi. Nearpole polar diagram of objects and duality. Journal of Computational Science, 3(3):127–131, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000834.

Nakasato:2012:IPT

[108] Naohito Nakasato. Implementation of a parallel tree method on a GPU. Journal of Computational Science, 3 (3):132–141, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000135.

Yang:2012:CSC

[109] Yang Yang and William W. Liou. Computational study of compressive loading of carbon nanotubes using quasi-continuum method. Journal of Computational Science, 3 (3):142–149, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000081.

Arai:2012:HMP

[110] Kohei Arai, Achmad Basuki, and Tri Harsono. Hot mudflow prediction area model and simulation based on Cellular Automata for LUSI mud plume at Sidoarjo in East Java. Journal of Computational Science, 3 (3):150–158, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000901.

Lorin:2012:EAN

[111] E. Lorin and A. D. Bandrauk. Efficient and accurate numerical modeling of a micro–macro nonlinear optics model for intense and short laser pulses. Journal of Computational Science, 3(3):159–168, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001116.

Mazzeo:2012:SEM

[112] Marco D. Mazzeo. The Spy Element Method — a universal approach to complex computing on manycore processors. Journal of Computational Science, 3(3):169–180, May 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000221.

Shamir:2012:ADP

[113] Lior Shamir. Automatic detection of peculiar galaxies in large datasets of galaxy images. Journal of Computational Science, 3 (3):181–189, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000245.

Anonymous:2012:EBb

[114] Anonymous. Editorial Board. Journal of Computational Science, 3 (3):CO2, May 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031200035X.

Munuzuri:2012:RAF

[115] Jesús Muñuzuri and Pablo Cortés. Recent advances and future trends in city logistics. Journal of Computational Science, 3(4):191–192, July 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000518.

Ehmke:2012:ARC

[116] Jan Fabian Ehmke, André Steinert, and Dirk Christian Mattfeld. Advanced routing for city logistics service providers based on time-dependent travel times. Journal of Computational Science, 3(4):193–205, July 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000087.

Gonzalez-Feliu:2012:SFE

[117] Jesus Gonzalez-Feliu, Christian Ambrosini, Pascal Pluvinet, Florence Toilier, and Jean-Louis Routhier. A simulation framework for evaluating the impacts of urban goods transport in terms of road occupancy. Journal of Computational Science, 3 (4):206–215, July 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750312000312.

vanDuin:2012:TGN

[118] J. H. R. van Duin and R. E. C. M. van der Heijden. Towards governance on noise between municipality and terminal operator by the use of simulation modelling. Journal of Computational Science, 3(4):216–227, July 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031200021X.

Munuzuri:2012:SLM

[119] Jesús Muñuzuri, Pablo Cortés, Rafael Grosso, and José Guadix. Selecting the location of minihubs for freight delivery in congested downtown areas. Journal of Computational Science, 3(4):228–237, July 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311001098.

Jahan:2012:CLG

[120] Majid Vafaei Jahan and Mohammad-R. Akbarzadeh-T. Composing local and global behaviors: Higher performance of spin glass based portfolio selection. Journal of Computational Science, 3(4):238–245, July 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000324.

Anonymous:2012:EBc

[121] Anonymous. Editorial Board. Journal of Computational Science, 3 (4):CO2, July 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000531.

Sloot:2012:IPP

[122] Peter M. A. Sloot and Rick Quax. Information processing as a paradigm to model and simulate complex systems. Journal of Computational Science, 3 (5):247–249, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000725.

Cannataro:2012:ACS

[123] Mario Cannataro, Rodrigo Weber dos Santos, Joakim Sundnes, and Pierangelo Veltri. Advanced computing solutions for health care and medicine. Journal of Computational Science, 3 (5):250–253, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000737.

Mago:2012:CDS

[124] Vijay Kumar Mago, Nitin Bhatia, Ajay Bhatia, and Anjali Mago. Clinical decision support system for dental treatment. Journal of Computational Science, 3(5):254–261, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000117.

Antal:2012:TPD

[125] Bálint Antal, András Hajdu, Zsuzsanna Maros-Szabó, Zsolt Török, Adrienne Csutak, and Tünde Pető. A two-phase decision support framework for the automatic screening of digital fundus images. Journal of Computational Science, 3(5):262–268, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000038.

Samsi:2012:ECF

[126] Siddharth Samsi, Ashok K. Krishnamurthy, and Metin N. Gurcan. An efficient computational framework for the analysis of whole slide images: Application to follicular lymphoma immunohistochemistry. Journal of Computational Science, 3(5): 269–279, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000129.

Dickmann:2012:SBG

[127] Frank Dickmann, Jürgen Falkner, Wilfried Gunia, Jochen Hampe, Michael Hausmann, Alexander Herrmann, Nick Kepper, Tobias A. Knoch, Svenja Lauterbach, Jörg Lippert, Kathrin Peter, Eberhard Schmitt, Ulrich Schwardmann, Juri Solodenko, Dietmar Sommerfeld, Thomas Steinke, Anette Weisbecker, and Ulrich Sax. Solutions for biomedical grid computing — case studies from the D-Grid project Services@MediGRID. Journal of Computational Science, 3(5): 280–297, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000603.

Nozaki:2012:EEF

[128] Kazunori Nozaki, Masanori Nakamura, Haruka Takimoto, and Shigeo Wada. Effect of expiratory flow rate on the acoustic characteristics of sibilant /s/. Journal of Computational Science, 3 (5):298–305, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000585.

Karantasis:2012:ASB

[129] Konstantinos I. Karantasis, Eleftherios D. Polychronopoulos, Konstantinos T. Panourgias, and John A. Ekaterinaris. Accelerating the simulation of brain tumor proliferation with manycore GPUs. Journal of Computational Science, 3(5):306–313, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000597.

Zasada:2012:IIE

[130] Stefan J. Zasada, Tao Wang, Ali Haidar, Enjie Liu, Norbert Graf, Gordon Clapworthy, Steven Manos, and Peter V. Coveney. IMENSE: an e-infrastructure environment for patient specific multiscale data integration, modelling and clinical treatment. Journal of Computational Science, 3 (5):314–327, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000639.

Schmitt:2012:CSS

[131] Eberhard Schmitt, Jenny Wagner, and Michael Hausmann. Combinatorial selection of short triplex forming oligonucleotides for fluorescence in situ hybridisation COMBO-FISH. Journal of Computational Science, 3 (5):328–334, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000871.

Zivanic:2012:VDG

[132] M. Zivanic, O. Daescu, A. Kurdia, and S. R. Goodman. The Voronoi diagram for graphs and its application in the Sickle Cell Disease research. Journal of Computational Science, 3 (5):335–343, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001001.

Gullo:2012:TSA

[133] Francesco Gullo, Giovanni Ponti, Andrea Tagarelli, Giuseppe Tradigo, and Pierangelo Veltri. A time series approach for clustering mass spectrometry data. Journal of Computational Science, 3(5):344–355, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000627.

Schumm:2012:BSG

[134] Phillip Schumm and Caterina Scoglio. Bloom: a stochastic growth-based fast method of community detection in networks. Journal of Computational Science, 3(5):356–366, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000269.

Chen:2012:EAN

[135] Yiming Chen, Mingxu Yi, and Chunxiao Yu. Error analysis for numerical solution of fractional differential equation by Haar wavelets method. Journal of Computational Science, 3 (5):367–373, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000452.

Pickl:2012:AGT

[136] Kristina Pickl, Jan Götz, Klaus Iglberger, Jayant Pande, Klaus Mecke, Ana-Sunčana Smith, and Ulrich Rüde. All good things come in threes — three beads learn to swim with lattice Boltzmann and a rigid body solver. Journal of Computational Science, 3 (5):374–387, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000464.

Bliss:2012:TRR

[137] Catherine A. Bliss, Isabel M. Kloumann, Kameron Decker Harris, Christopher M. Danforth, and Peter Sheridan Dodds. Twitter reciprocal reply networks exhibit assortativity with respect to happiness. Journal of Computational Science, 3(5): 388–397, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031200049X.

Denham:2012:DDD

[138] Mónica Denham, Kerstin Wendt, Germán Bianchini, Ana Cortés, and Tomàs Margalef. Dynamic Data-Driven Genetic Algorithm for forest fire spread prediction. Journal of Computational Science, 3(5): 398–404, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750312000658.

Dhawan:2012:CRT

[139] S. Dhawan, S. Kapoor, S. Kumar, and S. Rawat. Contemporary review of techniques for the solution of nonlinear Burgers equation. Journal of Computational Science, 3(5): 405–419, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031200066X.

Voigt:2012:MMF

[140] A. Voigt and T. Witkowski. A multimesh finite element method for Lagrange elements of arbitrary degree. Journal of Computational Science, 3 (5):420–428, September 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000671.

Dhawan:2012:NMA

[141] S. Dhawan, S. Kapoor, and S. Kumar. Numerical method for advection diffusion equation using FEM and Bsplines. Journal of Computational Science, 3(5):429–437, September 2012.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000695.

Anonymous:2012:EBd

[142] Anonymous. Editorial Board. Journal of Computational Science, 3(5): CO2, September 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000750.

Sloot:2012:YRR

[143] Peter M. A. Sloot and Alexander V. Boukhanovsky. Young Russian researchers take up challenges in the computational sciences. Journal of Computational Science, 3(6): 439–440, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000981.

Fadeev:2012:PNS

[144] Daniil A. Fadeev. To the problem of numerical simulation of gas optical breakdown on GPU. Journal of Computational Science, 3(6): 441–444, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000920.

Akimova:2012:PAS

[145] Elena N. Akimova and Dmitry V. Belousov. Parallel algorithms for solving linear systems with block-tridiagonal matrices on multi-core CPU with GPU. Journal of Computational Science, 3 (6):445–449, November 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000932.

Ivanov:2012:SBC

[146] Sergey V. Ivanov, Sergey S. Kosukhin, Anna V. Kaluzhnaya, and Alexander V. Boukhanovsky. Simulationbased collaborative decision support for surge floods prevention in St. Petersburg. Journal of Computational Science, 3(6):450–455, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000944. See

erratum [209].

Lind:2012:PCD

[147] Yuliya B. Lind. Parallel computations in drilling process. Journal of Computational Science, 3(6): 456–459, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000968.

Demidov:2012:MAM

[148] D. E. Demidov and D. V. Shevchenko. Modification of algebraic multigrid for effective GPGPU-based solution of nonstationary hydrodynamics problems. Journal of Computational Science, 3(6):460–462, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031200097X.

Mitin:2012:PIS

[149] Igor Mitin, Alexander Kalinkin, and Yuri Laevsky. A parallel iterative solver for positive-definite systems with hybrid MPI–OpenMP parallelization for multi-core clusters. Journal of Computational Science, 3(6): 463–468, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000993.

Nikitin:2012:PAW

[150] Viktor V. Nikitin, Anton A. Duchkov, and Fredrik Andersson. Parallel algorithm of 3D wave-packet decomposition of seismic data: Implementation and optimization for GPU. Journal of Computational Science, 3 (6):469–473, November 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312001007.

Bastrakov:2012:PCP

[151] S. Bastrakov, R. Donchenko, A. Gonoskov, E. Efimenko, A. Malyshev, I. Meyerov, and I. Surmin. Particle-in-cell plasma simulation on heterogeneous cluster systems. Journal of Computational Science, 3(6):474–479, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312001019.

Butyugin:2012:EIS

[152] Dmitry Butyugin. Efficient iterative solvers for time-harmonic Maxwell equations using domain decomposition and algebraic multigrid. Journal of Computational Science, 3(6): 480–485, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312001020.

Fesko:2012:PAI

[153] Oles Fesko. A parallel approach to improvement and estimation of the approximate optimal control. Journal of Computational Science, 3(6): 486–491, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312001032.

Getmanskiy:2012:CSM

[154] Victor V. Getmanskiy, Alexander S. Gorobtsov, Efim S. Sergeev, Timur D. Ismailov, and Oleg V. Shapovalov. Concurrent simulation of multibody systems coupled with stress–strain and heat transfer solvers. Journal of Computational Science, 3(6): 492–497, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312001044.

Gorshkov:2012:MCS

[155] Anton V. Gorshkov and Mikhail Yu. Kirillin. Monte Carlo simulation of brain sensing by optical diffuse spectroscopy. Journal of Computational Science, 3(6):498–503, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312001056.

Knyazkov:2012:CSI

[156] Konstantin V. Knyazkov, Sergey V. Kovalchuk, Timofey N. Tchurov, Sergey V. Maryin, and Alexander V. Boukhanovsky. CLAVIRE: e-science infrastructure for data-driven computing. Journal of Computational Science, 3(6):504–510, November 2012. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000956.

Anonymous:2012:EBe

[157] Anonymous. Editorial Board. Journal of Computational Science, 3(6): CO2, November 2012. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312001111.

Tesdall:2013:P

[158] Allen M. Tesdall, Jae-Hun Jung, Ilias Kotsireas, and Roderick Melnik. Preface. Journal of Computational Science, 4(1–2):1–2, January 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312001238.

Steinmoeller:2013:PMB

[159] D. T. Steinmoeller, M. Stastna, and K. G. Lamb. Pseudospectral methods for Boussinesq-type equations in an annular domain with applications to midsized lakes. Journal of Computational Science, 4(1–2):3–11, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000075.

Wang:2013:CSM

[160] Ying Wang and Chiu-Yen Kao. Central schemes for the modified Buckley–Leverett equation. Journal of Computational Science, 4(1– 2):12–23, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750312000105.

Qin:2013:DGM

[161] Ruibin Qin and Lilia Krivodonova. A discontinuous Galerkin method for solutions of the Euler equations on Cartesian grids with embedded geometries. Journal of Computational Science, 4(1–2):24–35, January 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000282.

Kim:2013:NSS

[162] Eun Heui Kim and Chung-Min Lee. Numerical solutions to shock reflection and shock interaction problems for the self-similar transonic twodimensional nonlinear wave systems. Journal of Computational Science, 4 (1–2):36–45, January 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000270.

Chakraborty:2013:EDC

[163] Debananda Chakraborty and Jae-Hun Jung. Efficient determination of the critical parameters and the statistical quantities for Klein–Gordon and sine-Gordon equations with a singular potential using generalized polynomial chaos methods. Journal of Computational Science, 4(1– 2):46–61, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000300.

Berzins:2013:DRB

[164] Martin Berzins. Data and rangebounded polynomials in ENO methods. Journal of Computational Science, 4(1–2):62–70, January 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000439.

Pitman:2013:TPS

[165] E. Bruce Pitman, Abani K. Patra, Dinesh Kumar, Kouichi Nishimura, and Jiro Komori. Two phase simulations of glacier lake outburst flows. Journal of Computational Science, 4 (1–2):71–79, January 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000440.

Yakovlev:2013:LDF

[166] Sergey Yakovlev, Liwei Xu, and Fengyan Li. Locally divergence-free central discontinuous Galerkin methods for ideal MHD equations. Journal of Computational Science, 4(1– 2):80–91, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000506.

Tesdall:2013:SSS

[167] Allen M. Tesdall and John K. Hunter. Self-similar solutions for the diffraction of weak shocks. Journal of Computational Science, 4(1– 2):92–100, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750312000610.

Jung:2013:RIM

[168] Jae-Hun Jung, Joseph Lee, Kenneth R. Hoffmann, Todd Dorazio, and E. Bruce Pitman. A rapid interpolation method of finding vascular CFD solutions with spectral collocation methods. Journal of Computational Science, 4(1–2):101–110, January 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000646.

Chalmers:2013:NAO

[169] N. Chalmers and E. Lorin. On the numerical approximation of onedimensional nonconservative hyperbolic systems. Journal of Computational Science, 4(1–2):111–124, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750312000919.

Anonymous:2013:EBa

[170] Anonymous. Editorial Board. Journal of Computational Science, 4(1– 2):CO2, January 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000033.

Paszynski:2013:P

[171] Maciej Paszyński, Krzysztof Cetnarowicz, Robert Schaefer, David Pardo, and Victor Calo. Preface. Journal of Computational Science, 4 (3):125–126, May 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031300029X.

Bankhead:2013:SFI

[172] Armand Bankhead, Emiliano Mancini, Amy C. Sims, Ralph S. Baric, Shannon McWeeney, and Peter M. A. Sloot. A simulation framework to investigate in vitro viral infection dynamics. Journal of Computational Science, 4(3):127–134, May 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000858.

Wang:2013:CMM

[173] Gong Wang, T. N. Wong, and Chunxia Yu. A computational model for multi-agent e-commerce negotiations with adaptive negotiation behaviors. Journal of Computational Science, 4 (3):135–143, May 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311000895.

Sniezynski:2013:CSL

[174] Bartłomiej Śnieżyński and Jacek Dajda. Comparison of strategy learning methods in Farmer–Pest problem for various complexity environments without delays. Journal of Computational Science, 4(3):144–151, May 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000233.

Collier:2013:TAD

[175] Nathan Collier, Hany Radwan, Lisandro Dalcin, and Victor M. Calo. Time adaptivity in the diffusive wave approximation to the shallow water equations. Journal of Computational Science, 4(3):152–156, May 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000664.

Niemi:2013:DPG

[176] Antti H. Niemi, Nathaniel O. Collier, and Victor M. Calo. Discontinuous Petrov–Galerkin method based on the optimal test space norm for steady transport problems in one space dimension. Journal of Computational Science, 4(3):157–163, May 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000652.

Gurgul:2013:AMA

[177] Piotr Gurgul, Marcin Sieniek, Krzysztof Magiera, and Marcin Skotniczny. Application of multi-agent paradigm to hp-adaptive projection-based interpolation operator. Journal of Computational Science, 4(3):164–169, May 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750311000640.

Szymczak:2013:PDD

[178] Arkadiusz Szymczak, Anna Paszyńska, Maciej Paszyński, and David Pardo. Preventing deadlock during anisotropic 2D mesh adaptation in hp-adaptive FEM. Journal of Computational Science, 4(3):170–179, May 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031100086X.

Anonymous:2013:EBb

[179] Anonymous. Editorial Board. Journal of Computational Science, 4 (3):CO2, May 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000264.

Fouchal:2013:P

[180] Hacène Fouchal and M. Sechi Moretti Annoni Notare. Preface. Journal of Computational Science, 4 (4):181–182, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000501.

Hasswa:2013:UIM

[181] Ahmed Hasswa and Hossam Hassanein. Utilizing the IP multimedia subsystem to create an extensible service-oriented architecture. Journal of Computational Science, 4(4):183–198, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000208.

Ben-Othman:2013:SSA

[182] Jalel Ben-Othman, Karim Bessaoud, Alain Bui, and Laurence Pilard. Selfstabilizing algorithm for efficient topology control in Wireless Sensor Networks. Journal of Computational Science, 4(4):199–208, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000051.

Klasen:2013:ECD

[183] Bernd Klasen. Efficient content distribution in social-aware hybrid networks. Journal of Computational Science, 4(4):209–218, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001104.

Fouchal:2013:CCA

[184] S. Fouchal, M. Ahat, S. Ben Amor, I. Lavallée, and M. Bui. Competitive clustering algorithms based on ultrametric properties. Journal of Computational Science, 4(4):219–231, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001062.

Haddad:2013:BMF

[185] Serge Haddad, Lynda Mokdad, and Samir Youcef. Bounding models families for performance evaluation in composite Web services. Journal of Computational Science, 4 (4):232–241, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750311001050.

Sager:2013:SCC

[186] Sebastian Sager, Katja Mombaur, and Joachim Funke. Scientific computing for the cognitive sciences. Journal of Computational Science, 4 (4):242–244, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031200124X.

Engelhart:2013:DAN

[187] Michael Engelhart, Joachim Funke, and Sebastian Sager. A decomposition approach for a new testscenario in complex problem solving. Journal of Computational Science, 4(4):245–254, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000683.

Felis:2013:MIE

[188] Martin L. Felis, Katja Mombaur, Hideki Kadone, and Alain Berthoz. Modeling and identification of emotional aspects of locomotion. Journal of Computational Science, 4 (4):255–261, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312001226.

Gonzalez:2013:VIB

[189] Cleotilde Gonzalez, Varun Dutt, and Christian Lebiere. Validating instancebased learning mechanisms outside of ACT-R. Journal of Computational Science, 4(4):262–268, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311001086.

Taix:2013:GHL

[190] Michel Taïx, Minh Tuan Tran, Philippe Souères, and Emmanuel Guigon. Generating human-like reaching movements with a humanoid robot: a computational approach. Journal of Computational Science, 4 (4):269–284, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000907.

Volchenkov:2013:STA

[191] Dimitri Volchenkov and Bettina Bläsing. Spatio-temporal analysis of kinematic signals in classical ballet. Journal of Computational Science, 4(4):285–292, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000713.

Augustin:2013:SUR

[192] Thomas Augustin, Cord Hockemeyer, Michael D. Kickmeier-Rust, Patrick Podbregar, Reinhard Suck, and Dietrich Albert. The simplified updating rule in the formalization of digital educational games. Journal of Computational Science, 4(4):293–303, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312001093.

Mukovskiy:2013:DSC

[193] Albert Mukovskiy, Jean-Jacques E. Slotine, and Martin A. Giese. Dynamically stable control of articulated crowds. Journal of Computational Science, 4(4):304–310, July 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312001081.

Anonymous:2013:EBc

[194] Anonymous. Editorial Board. Journal of Computational Science, 4 (4):CO2, July 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000446.

Pradal:2013:PSS

[195] Christophe Pradal, Gaël Varoquaux, and Hans Peter Langtangen. Publishing scientific software matters. Journal of Computational Science, 4(5): 311–312, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000938.

Fiers:2013:IDC

[196] Martin Fiers, Emmanuel Lambert, Shibnath Pathak, Bjorn Maes, Peter Bienstman, Wim Bogaerts, and Pieter Dumon. Improving the design cycle for nanophotonic components. Journal of Computational Science, 4 (5):313–324, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000732.

Bianchi:2013:WPP

[197] Riccardo Maria Bianchi and Renaud Brunelière. WatchMan project — a Python CASE framework for High Energy Physics data analysis in the LHC era. Journal of Computational Science, 4(5):325–333, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312000336.

Walter:2013:ADP

[198] Sebastian F. Walter and Lutz Lehmann. Algorithmic differentiation in Python with AlgoPy. Journal of Computational Science, 4(5):334–344, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311001013.

Wilbert:2013:BEF

[199] Niko Wilbert, Tiziano Zito, Rike-Benjamin Schuppner, Zbigniew Jȩdrzejewski-Szmek, Laurenz Wiskott, and Pietro Berkes. Building extensible frameworks for data processing: the case of MDP, Modular toolkit for Data Processing. Journal of Computational Science, 4 (5):345–351, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750311000913.

Rossant:2013:PPL

[200] Cyrille Rossant, Bertrand Fontaine, and Dan F. M. Goodman. Playdoh: a lightweight Python library for distributed computing and optimisation. Journal of Computational Science, 4 (5):352–359, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750311000561.

Das:2013:FSI

[201] Barun Das and Manoranjan Maiti. Fuzzy stochastic inequality and equality possibility constraints and their application in a production-inventory model via optimal control method. Journal of Computational Science, 4 (5):360–369, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750312000257.

Amorim:2013:SCB

[202] Ronan Mendonça Amorim and Rodrigo Weber dos Santos. Solving the cardiac bidomain equations using graphics processing units. Journal of Computational Science, 4(5): 370–376, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312000701.

Kroshko:2013:ESS

[203] Andrew Kroshko and Raymond J. Spiteri. Efficient SIMD solution of multiple systems of stiff IVPs. Journal of Computational Science, 4(5): 377–385, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750312001068.

Wang:2013:CFS

[204] D. Wang, S. Kang, J. Nichols, W. Post, S. Liu, and Z. Zhao. A computational framework for spatially explicit agroecosystem modeling: Application to regional simulation. Journal of Computational Science, 4(5): 386–392, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031200107X.

Nourmohammadi:2013:ICA

[205] A. Nourmohammadi, M. Zandieh, and R. Tavakkoli-Moghaddam. An imperialist competitive algorithm for multiobjective u-type assembly line design. Journal of Computational Science, 4 (5):393–400, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750312001160.

Leiter:2013:AMP

[206] Kenneth W. Leiter, Joshua C. Crone, and Jaroslaw Knap. An algorithm for massively parallel dislocation dynamics simulations of small scale plasticity. Journal of Computational Science, 4(5):401–411, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000227.

Groen:2013:AMP

[207] Derek Groen, James Hetherington, Hywel B. Carver, Rupert W. Nash, Miguel O. Bernabeu, and Peter V. Coveney. Analysing and modelling the performance of the HemeLB lattice-Boltzmann simulation environment. Journal of Computational Science, 4 (5):412–422, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000240.

Priess:2013:MEM

[208] M. Prieß, S. Koziel, and T. Slawig. Marine ecosystem model calibration with real data using enhanced surrogate-based optimization. Journal of Computational Science, 4(5): 423–437, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000409.

Ivanov:2013:ESB

[209] Sergey V. Ivanov, Sergey S. Kosukhin, Anna V. Kaluzhnaya, and Alexander V. Boukhanovsky. Erratum to “Simulation-based collaborative decision support for surge floods prevention in St. Petersburg” [J. Comput. Sci. 3 (2012) 450–455]. Journal of Computational Science, 4 (5):438, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000136. See

[146].

Anonymous:2013:EBd

[210] Anonymous. Editorial Board. Journal of Computational Science, 4(5): CO2, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000860.

Anonymous:2013:PS

[211] Anonymous. Pages 311–438 (September 2013). Journal of Computational Science, 4(5):??, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Alexandrov:2013:TSM

[212] Vassil Alexandrov. Towards scalable mathematics and scalable algorithms for extreme scale computing. Journal of Computational Science, 4(6):iii–v, November 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313001208.

Rajovic:2013:LPA

[213] Nikola Rajovic, Lluis Vilanova, Carlos Villavieja, Nikola Puzovic, and Alex Ramirez. The low power architecture approach towards exascale computing. Journal of Computational Science, 4 (6):439–443, November 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000148.

Trefethen:2013:EAS

[214] Anne E. Trefethen and Jeyarajan Thiyagalingam. Energy-aware software: Challenges, opportunities and strategies. Journal of Computational Science, 4(6):444–449, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000173.

Subotic:2013:PPE

[215] Vladimir Subotić, Steffen Brinkmann, Vladimir Marjanović, Rosa M. Badia, Jose Gracia, Christoph Niethammer, Eduard Ayguade, Jesus Labarta, and Mateo Valero. Programmability and portability for exascale: Top down programming methodology and tools with StarSs. Journal of Computational Science, 4(6):450–456, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000203.

Du:2013:SER

[216] Peng Du, Piotr Luszczek, Stan Tomov, and Jack Dongarra. Soft error resilient QR factorization for hybrid system with GPGPU. Journal of Computational Science, 4(6): 457–464, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000161.

Wu:2013:LSE

[217] Panruo Wu, Chong Ding, Longxiang Chen, Teresa Davies, Christer Karlsson, and Zizhong Chen. On-line soft error correction in matrix–matrix multiplication. Journal of Computational Science, 4(6):465–472, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000641.

Strassburg:2013:FAM

[218] Janko Straßburg and Vassil N. Alexandrov. Facilitating analysis of Monte Carlo dense matrix inversion algorithm scaling behaviour through simulation. Journal of Computational Science, 4 (6):473–479, November 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031300015X.

Gansterer:2013:SFT

[219] Wilfried N. Gansterer, Gerhard Niederbrucker, Hana Straková, and Stefan Schulze Grotthoff. Scalable and fault tolerant orthogonalization based on randomized distributed data aggregation. Journal of Computational Science, 4(6):480–488, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000185.

Goodall:2013:NUD

[220] Tabitha Goodall, David Pettinger, and Giuseppe Di Fatta. Non-uniform data distribution for communicationefficient parallel clustering. Journal of Computational Science, 4(6): 489–495, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000197.

He:2013:LAS

[221] Jun He, Jim Kowalkowski, Marc Paterno, Don Holmgren, James Simone, and Xian-He Sun. Layout-aware scientific computing: a case study using the MILC code. Journal of Computational Science, 4(6):496–506, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000690.

Anonymous:2013:EBe

[222] Anonymous. Editorial Board. Journal of Computational Science, 4(6): CO2, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001142.

Lulfesmann:2014:SEG

[223] Michael Lülfesmann and Ken ichi Kawarabayashi. Sub-exponential graph coloring algorithm for stencilbased Jacobian computations. Journal of Computational Science, 5 (1):1–11, January 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000781.

Regis:2014:PSR

[224] Rommel G. Regis. Particle swarm with radial basis function surrogates for expensive black-box optimization. Journal of Computational Science, 5 (1):12–23, January 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000847.

Oxman:2014:CMC

[225] Gadi Oxman, Shlomo Weiss, and Yair Be’ery. Computational methods for Conway’s Game of Life cellular automaton. Journal of Computational Science, 5(1):24–31, January 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031300094X.

Huang:2014:SEM

[226] Guang qiu Huang. SIS epidemic model-based optimization. Journal of Computational Science, 5(1): 32–50, January 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313001117.

Zasada:2014:FCE

[227] Stefan J. Zasada, David C. W. Chang, Ali N. Haidar, and Peter V. Coveney. Flexible composition and execution of large scale applications on distributed e-infrastructures. Journal of Computational Science, 5 (1):51–62, January 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313001269.

Anonymous:2014:EBa

[228] Anonymous. Editorial Board. Journal of Computational Science, 5 (1):CO2, January 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031300135X.

Anonymous:2014:PJa

[229] Anonymous. Pages 1–62 (January 2014). Journal of Computational Science, 5(1):??, January 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ali:2014:SII

[230] Hesham Ali, Yong Shi, and Deepak Khazanchi. Special issue on ICCS 2012 [Preface]. Journal of Computational Science, 5(2):63–64, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031400012X. Held

in Omaha, NE, June 4–6, 2012.

Gomez-Revuelto:2014:HAF

[231] I. Gomez-Revuelto, L. E. Garcia-Castillo, S. Llorente-Romano, and D. Pardo. 3D hp-adaptive finite element simulations of bend, step, and magic-T electromagnetic waveguide structures. Journal of Computational Science, 5(2):65–75, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000689.

Rao:2014:ABS

[232] Vishwas Rao and Adrian Sandu. An adjoint-based scalable algorithm for time-parallel integration. Journal of Computational Science, 5 (2):76–84, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000380.

Archibald:2014:CEE

[233] R. Archibald, M. Chakoumakos, and T. Zhuang. Characterizing the elements of Earth’s radiative budget: Applying uncertainty quantification to the CESM. Journal of Computational Science, 5(2):85–89, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000239.

Ciznicki:2014:BJI

[234] Miłosz Ciżnicki, Michał Kierzynka, Piotr Kopta, Krzysztof Kurowski, and Paweł Gepner. Benchmarking JPEG 2000 implementations on modern CPU and GPU architectures. Journal of Computational Science, 5(2):90–98, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000410.

Zhu:2014:TMQ

[235] Xiaoqian Zhu, Fei Wang, Haiyan Wang, Changzhi Liang, Run Tang, Xiaolei Sun, and Jianping Li. TOPSIS method for quality credit evaluation: a case of air-conditioning market in China. Journal of Computational Science, 5(2):99–105, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000215.

Nunez:2014:FFA

[236] Alberto Núñez and Mercedes G. Merayo. A formal framework to analyze cost and performance in Map– Reduce based applications. Journal of Computational Science, 5(2): 106–118, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000422.

Yang:2014:MMP

[237] Xin-She Yang, Mehmet Karamanoglu, Tao Luan, and Slawomir Koziel. Mathematical modelling and parameter optimization of pulsating heat pipes. Journal of Computational Science, 5 (2):119–125, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001427.

Dinh:2014:SAM

[238] Minh Ngoc Dinh, David Abramson, and Chao Jin. Statistical assertion: a more powerful method for debugging scientific applications. Journal of Computational Science, 5 (2):126–134, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001415.

Bansal:2014:SIB

[239] Jagdish Chand Bansal, Pramod Kumar Singh, Kusum Deep, Atulya Nagar, and Millie Pant. Special issue on bioinspired computing: Theories and applications: Journal of Computational Science. Journal of Computational Science, 5(2):135–136, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000258.

Muniyandi:2014:MFM

[240] Ravie Chandren Muniyandi and Abdullah Mohd Zin. Modeling framework for membrane computing in biological systems: Evaluation with a case study. Journal of Computational Science, 5(2):137–143, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001439.

Kumar:2014:PAH

[241] Vijay Kumar, Jitender Kumar Chhabra, and Dinesh Kumar. Parameter adaptive harmony search algorithm for unimodal and multimodal optimization problems. Journal of Computational Science, 5(2):144–155, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001403.

Bharti:2014:TSU

[242] Kusum Kumari Bharti and P. K. Singh. A three-stage unsupervised dimension reduction method for text clustering. Journal of Computational Science, 5(2):156–169, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001336.

Lin:2014:IND

[243] Chih-Hao Lin and Pei-Ling Lin. Improving the non-dominated sorting genetic algorithm using a gene-therapy method for multi-objective optimization. Journal of Computational Science, 5(2):170–183, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001324.

Pandey:2014:TBF

[244] Subhash Chandra Pandey and Gora Chand Nandi. TSD based framework for mining the induction rules. Journal of Computational Science, 5(2): 184–195, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001312.

Mustaffa:2014:EAB

[245] Zuriani Mustaffa, Yuhanis Yusof, and Siti Sakira Kamaruddin. Enhanced artificial bee colony for training least squares support vector machines in commodity price forecasting. Journal of Computational Science, 5(2):196–205, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001300.

Mozaffari:2014:EMS

[246] Ahmad Mozaffari, Mohammadreza Azimi, and Mofid Gorji-Bandpy. Ensemble mutable smart bee algorithm and a robust neural identifier for optimal design of a large scale power system. Journal of Computational Science, 5(2):206–223, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001245.

Gandomi:2014:CBA

[247] Amir H. Gandomi and Xin-She Yang. Chaotic bat algorithm. Journal of Computational Science, 5(2): 224–232, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001099.

Jha:2014:OCS

[248] P. C. Jha, Vikram Bali, Sonam Narula, and Mala Kalra. Optimal component selection based on cohesion & coupling for component based software system under build-or-buy scheme. Journal of Computational Science, 5(2):233–242, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000835.

Subramanian:2014:ASP

[249] K. G. Subramanian, Pradeep Isawasan, Ibrahim Venkat, Linqiang Pan, and Atulya Nagar. Array P systems with permitting features. Journal of Computational Science, 5(2): 243–250, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000823.

Bakhshali:2014:SCL

[250] Mohamad Amin Bakhshali and Mousa Shamsi. Segmentation of color lip images by optimal thresholding using bacterial foraging optimization (BFO). Journal of Computational Science, 5(2):251–257, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000811.

Yadav:2014:ECS

[251] Anupam Yadav and Kusum Deep. An efficient co-swarm particle swarm optimization for non-linear constrained optimization. Journal of Computational Science, 5(2):258–268, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000768.

Singh:2014:OCD

[252] Manohar Singh, B. K. Panigrahi, A. R. Abhyankar, and Swagatam Das. Optimal coordination of directional over-current relays using informative differential evolution algorithm. Journal of Computational Science, 5(2):269–276, March 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000756.

Paul:2014:NPS

[253] P. Victer Paul, A. Ramalingam, R. Baskaran, P. Dhavachelvan, K. Vivekanandan, and R. Subramanian. A new population seeding technique for permutation-coded Genetic Algorithm: Service transfer approach. Journal of Computational Science, 5 (2):277–297, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000744.

Thakur:2014:NGA

[254] Manoj Thakur. A new genetic algorithm for global optimization of multimodal continuous functions. Journal of Computational Science, 5(2):

298–311, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000677.

Sharma:2014:SBD

[255] Harish Sharma, Jagdish Chand Bansal, and K. V. Arya. Self balanced differential evolution. Journal of Computational Science, 5(2): 312–323, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750312001251.

Anonymous:2014:EBb

[256] Anonymous. Editorial Board. Journal of Computational Science, 5 (2):CO2, March 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031400009X.

Baev:2014:NGS

[257] Vesselin Baev, Gaurav Sablok, and Ivan Minkov. Next generation sequencing crowd sourcing at BIOCOMP: What promises it holds for us in future? Journal of Computational Science, 5(3):325–326, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000404.

Plakunmonthon:2014:CPH

[258] Sasitorn Plakunmonthon, Nattanan Panjaworayan T-Thienprasert, Kritsada Khongnomnan, Yong Poovorawan, and Sunchai Payungporn. Computational prediction of hybridization patterns between hepatitis C viral genome and human microR-NAs. Journal of Computational Science, 5(3):327–331, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001440.

Politano:2014:UBN

[259] Gianfranco Politano, Alessandro Savino, Alfredo Benso, Stefano Di Carlo, Hafeez Ur Rehman, and Alessandro Vasciaveo. Using Boolean networks to model post-transcriptional regulation in gene regulatory networks. Journal of Computational Science, 5(3):332–344, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001129.

Bolboaca:2014:SSA

[260] Sorana D. Bolboacă and Lorentz Jäntschi. Sensitivity, specificity, and accuracy of predictive models on phenols toxicity. Journal of Computational Science, 5(3):345–350, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001105.

Mitiushkina:2014:BDG

[261] Natalia V. Mitiushkina, Aglaya G. Iyevleva, Ekatherina Sh. Kuligina, Alexandr V. Togo, Yoshio Miki, and Evgeny N. Imyanitov. Biased detection of guanine-rich microRNAs by array profiling: Systematic error or biological phenomenon? Journal of Computational Science, 5 (3):351–356, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313001075.

Krachunov:2014:AMD

[262] Milko Krachunov and Dimitar Vassilev. An approach to a metagenomic data processing workflow. Journal of Computational Science, 5 (3):357–362, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000999.

Jantschi:2014:RNC

[263] Lorentz Jäntschi and Sorana D. Bolboacă. Rarefaction on natural compound extracts diversity among genus. Journal of Computational Science, 5 (3):363–367, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000987.

Boeva:2014:CAD

[264] Veselka Boeva. Clustering approaches for dealing with multiple DNA microarray datasets. Journal of Computational Science, 5(3):368–376, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313000653.

Bruzzone:2014:SIM

[265] Agostino G. Bruzzone. Special issue: Multidisciplinary M&S. Journal of Computational Science, 5 (3):377–379, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000532.

Sokolov:2014:MMI

[266] Boris V. Sokolov, Vyacheslav A. Zelentsov, Rafael M. Yusupov, and Yuri A. Merkuryev. Multiple models of information fusion processes: Quality definition and estimation. Journal of Computational Science, 5 (3):380–386, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000027.

Latorre-Biel:2014:SDD

[267] Juan-Ignacio Latorre-Biel, Emilio Jiménez-Macías, and Mercedes Pérez-Parte. Sequence of decisions on discrete event systems modeled by Petri nets with structural alternative configurations. Journal of Computational Science, 5(3):387–394, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001063.

Massei:2014:DPB

[268] Marina Massei, Simonluca Poggi, Matteo Agresta, and Angelo Ferrando. Development planning based on interoperable agent driven simulation. Journal of Computational Science, 5 (3):395–407, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314000052.

Rogovchenko-Buffoni:2014:ITM

[269] Lena Rogovchenko-Buffoni, Andrea Tundis, Muhammed Zoheb Hossain, Mattias Nyberg, and Peter Fritzson. An integrated toolchain for model based functional safety analysis. Journal of Computational Science, 5(3):408–414, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313001051.

Louca:2014:MAR

[270] Loucas S. Louca. Modal analysis reduction of multi-body systems with generic damping. Journal of Computational Science, 5(3):415–426, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031300104X.

Hafner:2014:ICS

[271] Irene Hafner, Matthias Rößler, Bernhard Heinzl, Andreas Körner, Michael Landsiedl, and Felix Breitenecker. Investigating communication and stepsize behaviour for co-simulation of hybrid physical systems. Journal of Computational Science, 5 (3):427–438, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313001038.

Garro:2014:RAA

[272] Alfredo Garro, Johannes Groß, Marius Riestenpatt gen. Richter, and Andrea Tundis. Reliability analysis of an Attitude Determination and Control System (ADCS) through the RAMSAS method. Journal of Computational Science, 5(3):439–449, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313000793.

Nacusse:2014:SSB

[273] Matías A. Nacusse and Sergio J. Junco. Switchable structured bond: a bond graph device for modeling power coupling/decoupling of physical systems. Journal of Computational Science, 5(3):450–462, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031300080X.

Hennies:2014:MSC

[274] Til Hennies, Tobias Reggelin, Juri Tolujew, and Pierre-Alain Piccut. Mesoscopic supply chain simulation. Journal of Computational Science, 5 (3):463–470, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001002.

Raybourn:2014:NPS

[275] Elaine M. Raybourn. A new paradigm for serious games: Transmedia learning for more effective training and education. Journal of Computational Science, 5(3):471–481, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001014.

Balino:2014:MSS

[276] J. L. Baliño. Modeling and simulation of severe slugging in airwater systems including inertial effects. Journal of Computational Science, 5(3):482–495, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750313001026.

Kotsireas:2014:HPC

[277] Ilias S. Kotsireas, Lilia Krivodonova, Sabine McConnell, and Erik Schnetter. High performance computing theory and applications — proceedings of SHARCNET Research Day 2012 (Guelph, Ontario). Journal of Computational Science, 5 (3):497–498, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314000581.

Frolov:2014:VPL

[278] Alexei M. Frolov and David M. Wardlaw. Vacuum polarization in light two-electron atoms and ions. Journal of Computational Science, 5 (3):499–506, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750313000392.

Green:2014:OST

[279] Kevin R. Green and Lennaert van Veen. Open-source tools for dynamical analysis of Liley’s mean-field cortex model. Journal of Computational Science, 5(3):507–516, May 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031300077X.

Kasinathan:2014:SLG

[280] Dhanaraja Kasinathan, Kirsten Morris, and Steven Yang. Solution of large generalized H_{∞} algebraic Riccati equations. Journal of Computational Science, 5(3):517–526, 2014. ISSN 1877-7503 (print), 1877-7511 (electronic).

Nika:2014:EAC

[281] Varvara Nika, Paul Babyn, and Hongmei Zhu. EigenBlock algorithm for change detection — an application of adaptive dictionary learning techniques. Journal of Computational Science, 5(3):527–535, 2014. ISSN 1877-7503 (print), 1877-7511 (electronic).

Sykes:2014:IPI

[282] Edward R. Sykes and Wesley Skoczen. An improved parallel implementation of RainbowCrack using MPI. Journal of Computational Science, 5(3): 536–541, May 2014. ISSN 1877-7503 (print), 1877-7511 (electronic).

Yuen:2014:GSE

[283] Wai Kong Yuen, Egor Ospadov, and Stuart M. Rothstein. Ground-state electronic properties of LiH calculated from the “bounce” version of quantum Monte Carlo. Journal of Computational Science, 5(3):542–548, 2014. ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2014:EBc

[284] Anonymous. Editorial Board. Journal of Computational Science, 5 (3):CO2, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000374.

Anonymous:2014:PM

[285] Anonymous. Pages 325–556 (May 2014). Journal of Computational Science, 5(3):??, May 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ramasami:2014:PSI

[286] Ponnadurai Ramasami. Preface for the special issue of the Journal of Computational Science (Proceedings of the Virtual Conference on Computational Chemistry, VCCC-2013). Journal of Computational Science, 5(4):557, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000842.

Kakkar:2014:IPD

[287] Rita Kakkar, Richa Arora, Pragya Gahlot, and Deepti Gupta. An insight into pyruvate dehydrogenase kinase (PDHK) inhibition through pharmacophore modeling and QSAR studies. Journal of Computational Science, 5(4):558–567, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000489.

Almatarneh:2014:HLI

[288] Mansour H. Almatarneh, Mohammednoor Altarawneh, Raymond A. Poirier, and Ibrahim A. Saraireh. High level ab initio, DFT, and RRKM calculations for the unimolecular decomposition reaction of ethylamine. Journal of Computational Science, 5 (4):568–575, July 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031400026X.

Shen:2014:SIM

[289] Ning Shen, Ye Fan, and Sudhakar Pamidighantam. E-science infrastructures for molecular modeling and parametrization. Journal of Computational Science, 5(4):576–589, July 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000064.

Murray:2014:FAS

[290] Jane S. Murray, Laura Macaveiu, and Peter Politzer. Factors affecting the strengths of σ-hole electrostatic potentials. Journal of Computational Science, 5(4):590–596, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000039.

Jantschi:2014:FOD

[291] Lorentz Jäntschi and Sorana D. Bolboacă. First order derivatives of thermodynamic functions under assumption of no chemical changes revisited. Journal of Computational Science, 5(4):597–602, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001452.

Debnath:2014:PP

[292] Narayan Debnath and Hacène Fouchal. Preface — PEDISWESA’13. Journal of Computational Science, 5 (4):603–605, July 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314000908.

Goto:2014:GUP

[293] Takaaki Goto, Tadaaki Kirishima, Tetsuro Nishino, Takeo Yaku, and Kensei Tsuchida. Generation of UML package diagrams based on an attribute graph grammar. Journal of Computational Science, 5(4):606–615, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000283.

Linck:2014:AMS

[294] S. Linck, E. Mory, J. Bourgeois, E. Dedu, and F. Spies. Adaptive multimedia streaming using a simulation test bed. Journal of Computational Science, 5(4):616–623, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000271.

Dessart:2014:DCU

[295] Nathalie Dessart and Philippe Hunel. Data collection using WSN for counting individuals and habitat characterization. Journal of Computational Science, 5(4):624–632, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000301.

Gharbi:2014:IBD

[296] Nawel Gharbi, Besma Nemmouchi, Lynda Mokdad, and Jalel Ben-Othman. The impact of breakdowns disciplines and repeated attempts on performances of Small Cell Networks. Journal of Computational Science, 5(4):633–644, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000349.

Koneru:2014:HER

[297] Sindoora Koneru, Bidyut Gupta, Shahram Rahimi, Ziping Liu, and Narayan Debnath. A highly efficient RPF-based broadcast protocol using a new two-level pruning mechanism. Journal of Computational Science, 5(4):645–652, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000337.

Haraty:2014:ACM

[298] Ramzi A. Haraty, Mirna F. Naous, and Azzam Mourad. Assuring consistency in mixed models. Journal of Computational Science, 5(4):653–663, July 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000325.

Alghamdi:2014:MWB

[299] Bandar Alghamdi and Hacène Fouchal. A mobile wireless body area network platform. Journal of Computational Science, 5(4):664–674, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000313.

Sen:2014:DDQ

[300] Soumya Sen, Santanu Roy, Anirban Sarkar, Nabendu Chaki, and Narayan C. Debnath. Dynamic discovery of query path on the lattice of cuboids using hierarchical data granularity and storage hierarchy. Journal of Computational Science, 5(4):675–683, July 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000295.

Anonymous:2014:EBd

[301] Anonymous. Editorial Board. Journal of Computational Science, 5(4):ifc, July 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750314000787.

Anonymous:2014:PJb

[302] Anonymous. Pages 557–684 (July 2014). Journal of Computational Science, 5(4):??, July 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Xie:2014:DAP

[303] Youxiang Xie, Zhaohui Yuan, and Linjun Wang. Dynamic analysis of pest control model with population dispersal in two patches and impulsive effect. Journal of Computational Science, 5(5):685–695, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000957.

Koay:2014:DPU

[304] Cheng Guan Koay. Distributing points uniformly on the unit sphere under a mirror reflection symmetry constraint. Journal of Computational Science, 5 (5):696–700, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000763.

Trevisan:2014:LCM

[305] Diogo Fernando Trevisan, João Paulo Gois, and Harlen Costa Batagelo. A low-cost-memory CUDA implementation of the conjugate gradient method applied to globally supported radial basis functions implicits. Journal of Computational Science, 5(5):701–708, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000751.

Dadvand:2014:IBL

[306] Abdolrahman Dadvand, Masoud Baghalnezhad, Iraj Mirzaee, Boo Cheong Khoo, and Soheila Ghoreishi. An immersed boundary–lattice Boltzmann approach to study the dynamics of elastic membranes in viscous shear flows. Journal of Computational Science, 5 (5):709–718, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031400074X.

Borgdorff:2014:DMC

[307] J. Borgdorff, M. Mamonski, B. Bosak, K. Kurowski, M. Ben Belgacem, B. Chopard, D. Groen, P. V. Coveney, and A. G. Hoekstra. Distributed multiscale computing with MUSCLE 2, the Multiscale Coupling Library and Environment. Journal of Computational Science, 5(5):719–731, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000465.

Sturmer:2014:RTS

[308] Markus Stürmer, Johannes Dagner, Paul Manstetten, and Harald Köstler. Real-time simulation of temperature in hot rolling rolls. Journal of Computational Science, 5(5): 732–742, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314000453.

Kherici:2014:UPW

[309] Nada Kherici and Yamina Mohamed Ben Ali. Using PSO for a walk of a biped robot. Journal of Computational Science, 5(5):743–749, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000076.

Bliss:2014:EAA

[310] Catherine A. Bliss, Morgan R. Frank, Christopher M. Danforth, and Peter Sheridan Dodds. An evolutionary algorithm approach to link prediction in dynamic social networks.

Journal of Computational Science, 5 (5):750–764, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000040.

Srinivasan:2014:CEM

[311] Seshasai Srinivasan and M. Ziad Saghir. Computational evaluation of micro-scale and macro-scale error sources in a thermodiffusive cell. Journal of Computational Science, 5 (5):765–776, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001294.

Behnamian:2014:PCC

[312] J. Behnamian. A parallel competitive colonial algorithm for JIT flowshop scheduling. Journal of Computational Science, 5(5):777–783, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750313001282.

Hasert:2014:CFS

[313] Manuel Hasert, Kannan Masilamani, Simon Zimny, Harald Klimach, Jiaxing Qi, Jörg Bernsdorf, and Sabine Roller. Complex fluid simulations with the parallel tree-based Lattice Boltzmann solver Musubi. Journal of Computational Science, 5(5): 784–794, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313001270.

Pang:2014:QMN

[314] Wei Pang and George Macleod Coghill. QML-Morven: a novel framework for learning qualitative differential equation models using both symbolic and evolutionary approaches. Journal of Computational Science, 5 (5):795–808, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000702.

Bolaji:2014:UCT

[315] Asaju La’aro Bolaji, Ahamad Tajudin Khader, Mohammed Azmi Al-Betar, and Mohammed A. Awadallah. University course timetabling using hybridized artificial bee colony with hill climbing optimizer. Journal of Computational Science, 5(5):809– 818, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314000441.

Topa:2014:CCS

[316] Paweł Topa and Jarosław Wa̧s. Complex collective systems. Journal of Computational Science, 5(5):819– 820, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314001094. See

corrigendum [344].

Saravakos:2014:MEB

[317] Petros Saravakos and Georgios Ch. Sirakoulis. Modeling employees behavior in workplace dynamics. Journal of Computational Science, 5(5):

821–833, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000672.

Bezbradica:2014:CAA

[318] Marija Bezbradica, Heather J. Ruskin, and Martin Crane. Comparative analysis of asynchronous cellular automata in stochastic pharmaceutical modelling. Journal of Computational Science, 5(5):834–840, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000659.

Dietrich:2014:BGC

[319] Felix Dietrich, Gerta Köster, Michael Seitz, and Isabella von Sivers. Bridging the gap: From cellular automata to differential equation models for pedestrian dynamics. Journal of Computational Science, 5(5): 841–846, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000738.

Kirik:2014:VSP

[320] Ekaterina Kirik and Andrey Malyshev. On validation of SigmaEva pedestrian evacuation computer simulation module with bottleneck flow. Journal of Computational Science, 5 (5):847–850, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000684.

Anonymous:2014:EBe

[321] Anonymous. Editorial Board. Journal of Computational Science, 5 (5):ifc, September 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314001021.

Anonymous:2014:PS

[322] Anonymous. Pages 685–850 (September 2014). Journal of Computational Science, 5(5):??, September 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ramli:2014:PSI

[323] Muhamad Azfar Ramli and Christopher Pineda Monterola. Probing and shaping the information transfer of noise-perturbed complex networks via Markov chain analysis. Journal of Computational Science, 5 (6):851–860, November 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314001185.

Maroosi:2014:EMC

[324] Ali Maroosi and Ravie Chandren Muniyandi. Enhancement of membrane computing model implementation on GPU by introducing matrix representation for balancing occupancy and reducing inter-block communications. Journal of Computational Science, 5 (6):861–871, November 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000994.

urRehman:2014:EKB

[325] Samee ur Rehman, Matthijs Langelaar, and Fred van Keulen. Efficient kriging-based robust optimization of unconstrained problems. Journal of Computational Science, 5(6): 872–881, November 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000477.

VanSchyndel:2014:DCF

[326] Michael Van Schyndel, Gabriel A. Wainer, Rhys Goldstein, Jeremy P. M. Mogk, and Azam Khan. On the definition of a computational fluid dynamic solver using cellular discrete-event simulation. Journal of Computational Science, 5(6):882–890, November 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000696.

Clees:2014:FUT

[327] T. Clees, N. Hornung, I. Nikitin, L. Nikitina, D. Steffes-lai, and S. Klimenko. Focused ultrasonic therapy planning: Metamodeling, optimization, visualization. Journal of Computational Science, 5(6):891– 897, November 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314000714.

Anonymous:2014:EBf

[328] Anonymous. Editorial Board. Journal of Computational Science, 5 (6):ifc, November 2014. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314001409.

Anonymous:2014:PN

[329] Anonymous. Pages 851–898 (November 2014). Journal of Computational Science, 5(6):??, November 2014. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2015:PJ

[330] Anonymous. Pages 1–66 (January 2015). Journal of Computational Science, 6:1–66, January 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Kampis:2015:ASM

[331] George Kampis, Jan W. Kantelhardt, Kamil Kloch, and Paul Lukowicz. Analytical and simulation models for collaborative localization. Journal of Computational Science, 6:1–10, January 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314001197.

Zhong:2015:DBE

[332] Jinghui Zhong, Nan Hu, Wentong Cai, Michael Lees, and Linbo Luo. Density-based evolutionary framework for crowd model calibration. Journal of Computational Science, 6: 11–22, January 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314001203.

Kar:2015:HGD

[333] Arpan Kumar Kar. A hybrid group decision support system for supplier selection using analytic hierarchy process, fuzzy set theory and neural network. Journal of Computational Science, 6:23–33, January 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001586.

Mykytenko:2015:CMI

[334] N. Mykytenko, D. Fink, and A. Kiv. Computer modeling of ion current pulsations in track-containing foils. Journal of Computational Science, 6:34–39, January 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750314001550.

Pastel:2015:EDL

[335] Rudy Pastel, Jérôme Morio, and François Le Gland. Extreme density level set estimation for input–output functions via the adaptive splitting technique. Journal of Computational Science, 6:40–46, January 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001574.

Kaleel:2015:CDT

[336] Shakira Banu Kaleel and Abdolreza Abhari. Cluster-discovery of Twitter messages for event detection and trending. Journal of Computational Science, 6:47–57, January 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314001604.

Bianchini:2015:ESS

[337] G. Bianchini, P. Caymes-Scutari, and M. Méndez-Garabetti. Evolutionary-Statistical System: a parallel method for improving forest fire spread prediction. Journal of Computational Science, 6:58–66, January 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314001628.

Anonymous:2015:EBa

[338] Anonymous. Editorial Board. Journal of Computational Science, 6:ifc, January 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750314001689.

Anonymous:2015:PMa

[339] Anonymous. Pages 1–48 (March 2015). Journal of Computational Science, 7:1– 48, March 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Leifsson:2015:SDD

[340] Leifur Leifsson and Slawomir Koziel. Simulation-driven design of low-speed wind tunnel contraction. Journal of Computational Science, 7:1–12, March 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750314001653.

Mohseniahouei:2015:ASC

[341] Yasaman Mohseniahouei, Kenzu Abdella, and Marco Pollanen. The application of the Sinc-Collocation approach based on derivative interpolation in numerical oceanography. Journal of Computational Science, 7:13–25, March 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001665.

Onu:2015:LTC

[342] K. Onu, F. Huhn, and G. Haller. LCS Tool: a computational platform for Lagrangian coherent structures. Journal of Computational Science, 7:26–36, March 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031400163X.

Pendharkar:2015:ACO

[343] Parag C. Pendharkar. An ant colony optimization heuristic for constrained task allocation problem. Journal of Computational Science, 7:37–47, March 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000058.

Topa:2015:CEC

[344] Paweł Topa and Jarosław Was. Corrigendum to the editorial of “Complex collective systems” [J. Comput. Sci. 5 (2014) 819–820]. Journal of Computational Science, 7:48, March 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000083. See

[316].

Anonymous:2015:EBb

[345] Anonymous. Editorial Board. Journal of Computational Science, 7:ifc, March 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750315000186.

Anonymous:2015:PMb

[346] Anonymous. Pages 1–120 (May 2015). Journal of Computational Science, 8:1– 120, May 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Bartuschat:2015:PMS

[347] Dominik Bartuschat and Ulrich Rüde. Parallel multiphysics simulations of charged particles in microfluidic flows. Journal of Computational Science, 8:1–19, May 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000162.

Jiang:2015:PIC

[348] Qiaoyong Jiang, Lei Wang, and Xinhong Hei. Parameter identification of chaotic systems using artificial raindrop algorithm. Journal of Computational Science, 8:20–31, May 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000149.

Garbey:2015:MCF

[349] M. Garbey, M. Rahman, and S. Berceli. A multiscale computational framework to understand vascular adaptation. Journal of Computational Science, 8:32–47, May 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000125.

Xu:2015:EGE

[350] Changjin Xu and Qiming Zhang. Existence and global exponential stability of anti-periodic solutions of highorder bidirectional associative memory (BAM) networks with time-varying delays on time scales. Journal of Computational Science, 8:48–61, May 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000265.

Huang:2015:FOA

[351] Guang qiu Huang. Function optimization algorithm based on SIRQV epidemic dynamic model. Journal of Computational Science, 8:62–92, May 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000289.

Gonzalez-Dominguez:2015:GAE

[352] Jorge González-Domínguez and Bertil Schmidt. GPU-accelerated exhaustive search for third-order epistatic interactions in case–control studies. Journal of Computational Science, 8:93–100, May 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000393.

Wang:2015:ECG

[353] Linjun Wang, Huiping Cao, Xu Han, Jie Liu, and Youxiang Xie. An efficient conjugate gradient method and application to dynamic force reconstruction. Journal of Computational Science, 8:101–108, May 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000356.

Komaki:2015:GWO

[354] G. M. Komaki and Vahid Kayvanfar. Grey Wolf Optimizer algorithm for the two-stage assembly flow shop scheduling problem with release time. Journal of Computational Science, 8:109–120, May 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000381.

Anonymous:2015:EBc

[355] Anonymous. Editorial Board. Journal of Computational Science, 8:ifc, May 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750315000708.

Koziel:2015:E

[356] Slawomir Koziel and Leifur Leifsson. Editorial. Journal of Computational Science, 9:iii, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315001039.

Norman:2015:CSC

[357] Matthew Norman, Jeffrey Larkin, Aaron Vose, and Katherine Evans. A case study of CUDA FORTRAN and OpenACC for an atmospheric climate kernel. Journal of Computational Science, 9:1–6, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000605.

Cannataro:2015:GRP

[358] Mario Cannataro, Pietro H. Guzzi, and Marianna Milano. GoD: an rpackage based on ontologies for prioritization of genes with respect to diseases. Journal of Computational Science, 9:7–13, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000551.

Papa:2015:MSD

[359] João P. Papa, Gustavo H. Rosa, Aparecido N. Marana, Walter Scheirer, and David D. Cox. Model selection for Discriminative Restricted Boltzmann Machines through meta-heuristic techniques. Journal of Computational Science, 9:14–18, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000526.

Krawczyk:2015:IWO

[360] Bartosz Krawczyk and Michał Woźniak. Incremental weighted one-class classifier for mining stationary data streams. Journal of Computational Science, 9:19–25, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000629.

Zhong:2015:DES

[361] Jinghui Zhong and Wentong Cai. Differential evolution with sensitivity analysis and the Powell’s method for crowd model calibration. Journal of Computational Science, 9:26–32, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000514.

Ozog:2015:FEC

[362] David Ozog, Jay McCarty, Grant Gossett, Allen D. Malony, and Marina Guenza. Fast equilibration of coarse-grained polymeric liquids. Journal of Computational Science, 9:33–38, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000538.

Paredes:2015:CPI

[363] Simão Paredes, Teresa Rocha, Paulo de Carvalho, Jorge Henriques, Diana Mendes, Ricardo Cabete, Ramona Cabiddu, Anna Bianchi, and João Morais. The CardioRisk project: Improvement of cardiovascular risk assessment. Journal of Computational Science, 9:39–44, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000630.

Mountrakis:2015:PPI

[364] Lampros Mountrakis, Eric Lorenz, Orestis Malaspinas, Saad Alowayyed, Bastien Chopard, and Alfons G. Hoekstra. Parallel performance of an IB-LBM suspension simulation framework. Journal of Computational Science, 9:45–50, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000447.

Zamith:2015:NSC

[365] Marcelo Zamith, Regina Célia P. Leal-Toledo, Esteban Clua, Elson M. Toledo, and Guilherme V. P. de Magalhães. A new stochastic cellular automata model for traffic flow simulation with drivers’ behavior prediction. Journal of Computational Science, 9:51–56, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000435.

Vallejo:2015:GAE

[366] Marta Vallejo, Verena Rieser, and David W. Corne. Genetic algorithm evaluation of green search allocation policies in multilevel complex urban scenarios. Journal of Computational Science, 9:57–63, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000423.

Zhang:2015:PNM

[367] Tao Zhang, Amgad Salama, Shuyu Sun, and Mohamed F. El-Amin. Pore network modeling of drainage process in patterned porous media: a quasistatic study. Journal of Computational Science, 9:64–69, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000484.

Randles:2015:MPS

[368] Amanda Randles, Erik W. Draeger, and Peter E. Bailey. Massively parallel simulations of hemodynamics in the primary large arteries of the human vasculature. Journal of Computational Science, 9:70–75, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000411.

Hell:2015:SBF

[369] Sascha Hell and Wilfried Becker. The scaled boundary finite element method for the analysis of 3D crack interaction. Journal of Computational Science, 9:76–81, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000459.

Gardner:2015:NCI

[370] William B. Gardner, Gary Grewal, Deborah Stacey, David A. Calvert, Stefan C. Kremer, and Fangju Wang. A new Canadian interdisciplinary Ph.D. in computational sciences. Journal of Computational Science, 9:82–87, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000666.

Litescu:2015:IIT

[371] Sorina Litescu, Vaisagh Viswanathan, Michael Lees, Alois Knoll, and Heiko Aydt. Information impact on transportation systems. Journal of Computational Science, 9:88–93, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000575.

Richie:2015:TMP

[372] David Richie, James Ross, Song Park, and Dale Shires. Threaded MPI programming model for the Epiphany RISC array processor. Journal of Computational Science, 9:94–100, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000617.

Meister:2015:APP

[373] Oliver Meister and Michael Bader. 2D adaptivity for 3D problems: Parallel SPE10 reservoir simulation on dynamically adaptive prism grids. Journal of Computational Science, 9:101–106, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031500054X.

DeVos:2015:SLG

[374] Dirk De Vos, Emil De Borger, Jan Broeckhove, and Gerrit T. S. Beemster. Simulating leaf growth dynamics through Metropolis–Monte Carlo based energy minimization. Journal of Computational Science, 9:107–111, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315000642.

Kalyuzhnaya:2015:SSR

[375] Anna V. Kalyuzhnaya, Alexander A. Visheratin, Alexey Dudko, Denis Nasonov, and Alexander V. Boukhanovsky. Synthetic storms reconstruction for coastal floods risks assessment. Journal of Computational Science, 9:112–117, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000678.

Olsen:2015:CEP

[376] Megan M. Olsen and Rachel Fraczkowski. Co-evolution in predator prey through reinforcement learning. Journal of Computational Science, 9:118–124, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750315000496.

Zhong:2015:MVM

[377] Chen Zhong, Ed Manley, Stefan Müller Arisona, Michael Batty, and Gerhard Schmitt. Measuring variability of mobility patterns from multiday smartcard data. Journal of Computational Science, 9:125–130, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000599.

Kassens:2015:HSE

[378] Jan Christian Kässens, Lars Wienbrandt, Jorge González-Domínguez, Bertil Schmidt, and Manfred Schimmler. High-speed exhaustive 3-locus interaction epistasis analysis on FP-GAs. Journal of Computational Science, 9:131–136, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031500068X.

Meng:2015:IFC

[379] Haiyan Meng, Rupa Kommineni, Quan Pham, Robert Gardner, Tanu Malik, and Douglas Thain. An invariant framework for conducting reproducible computational science. Journal of Computational Science, 9:137–142, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000502.

Gomez-Revuelto:2015:HAA

[380] Ignacio Gomez-Revuelto, Luis E. Garcia-Castillo, and David Pardo. High-accuracy adaptive modeling of the energy distribution of a meniscusshaped cell culture in a Petri dish. Journal of Computational Science, 9:143–149, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315000654.

Itani:2015:AME

[381] Mohamed A. Itani, Ulf D. Schiller, Sebastian Schmieschek, James Hetherington, Miguel O. Bernabeu, Hoskote Chandrashekar, Fergus Robertson, Peter V. Coveney, and Derek Groen. An automated multiscale ensemble simulation approach for vascular blood flow. Journal of Computational Science, 9:150–155, July 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000460.

Orchard:2015:EFT

[382] Dominic Orchard, Andrew Rice, and Oleg Oshmyan. Evolving Fortran types with inferred units-of-measure. Journal of Computational Science, 9:156–162, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315000563.

Anonymous:2015:EBd

[383] Anonymous. Editorial Board. Journal of Computational Science, 9:ifc, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750315000903.

Anonymous:2015:PS

[384] Anonymous. Pages 1–382 (September 2015). Journal of Computational Science, 10:1–382, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ghaffari:2015:ISS

[385] R. Ghaffari, K. Alipour, S. Solgi, S. Irani, and H. Haddadpour. Investigation of surface stress effect in 3D complex nano parts using FEM and modified boundary Cauchy–Born method. Journal of Computational Science, 10:1–12, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000782.

Melnikova:2015:EUF

[386] N. B. Melnikova, D. Jordan, and V. V. Krzhizhanovskaya. Experience of using FEM for real-time flood early warning systems: Monitoring and modeling Boston levee instability. Journal of Computational Science, 10:13–25, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000800. See

corrigendum [524].

Monfared:2015:RDP

[387] V. Monfared, S. Daneshmand, and J. N. Reddy. Rate dependent plastic deformation analysis of creeping short fiber composites using the virtual fiber method in the non-reinforced regions. Journal of Computational Science, 10:26–35, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000873.

Jiang:2015:MPF

[388] Yanqun Jiang, Shuguang Zhou, and Fang-Bao Tian. Macroscopic pedestrian flow model with degrading spatial information. Journal of Computational Science, 10:36–44, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300016.

Leifsson:2015:ASO

[389] Leifur Leifsson and Slawomir Koziel. Aerodynamic shape optimization by variable-fidelity computational fluid dynamics models: a review of recent progress. Journal of Computational Science, 10:45–54, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000071.

Leifsson:2015:OSD

[390] Leifur Leifsson, Elvar Hermannsson, and Slawomir Koziel. Optimal shape design of multi-element trawl-doors using local surrogate models. Journal of Computational Science, 10: 55–62, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000101.

Bruzzone:2015:PMA

[391] Agostino G. Bruzzone. Perspectives of modeling & applied simulation: “Modeling, Algorithms and Simulations: advances and novel researches for problem-solving and decision-making in complex, multiscale and multi-domain systems”. Journal of Computational Science, 10: 63–65, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315001027.

Dittmar:2015:NAT

[392] Tim Dittmar, Claudia Krull, and Graham Horton. A new approach for touch gesture recognition: Conversive hidden non-Markovian models. Journal of Computational Science, 10:66–76, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000290.

Merkuryeva:2015:ARF

[393] Galina Merkuryeva, Yuri Merkuryev, Boris V. Sokolov, Semyon Potryasaev, Viacheslav A. Zelentsov, and Arnis Lektauers. Advanced river flood monitoring, modelling and forecasting. Journal of Computational Science, 10:77–85, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001240.

Fanti:2015:SBD

[394] M. P. Fanti, G. Iacobellis, W. Ukovich, V. Boschian, G. Georgoulas, and C. Stylios. A simulation based Decision Support System for logistics management. Journal of Computational Science, 10:86–96, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001239.

Longo:2015:OPC

[395] Francesco Longo, Alessandro Chiurco, Roberto Musmanno, and Letizia Nicoletti. Operative and procedural cooperative training in marine ports. Journal of Computational Science, 10:97–107, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314001227.

Castro:2015:SAC

[396] Rodrigo D. Castro, François E. Cellier, and Andreas Fischlin. Sustainability analysis of complex dynamic systems using embodied energy flows: The ecobond graphs modeling and simulation framework. Journal of Computational Science, 10:108–125, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314001215.

Ribault:2015:TBO

[397] Judicaël Ribault and Gregory Zacharewicz. Time-based orchestration of workflow, interoperability with G-Devs/Hla. Journal of Computational Science, 10:126–136, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750314000982.

Seiger:2015:MCF

[398] Ronny Seiger, Christine Keller, Florian Niebling, and Thomas Schlegel. Modelling complex and flexible processes for smart cyber-physical environments. Journal of Computational Science, 10:137–148, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000970.

vanVianen:2015:SBR

[399] Teus van Vianen, Jaap Ottjes, and Gabriël Lodewijks. Simulation-based rescheduling of the stacker–reclaimer operation. Journal of Computational Science, 10:149–154, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314000726.

Ramasami:2015:PVC

[400] Ponnadurai Ramasami. Perspectives of Virtual Conference on Computational Chemistry (VCCC-2014). Journal of Computational Science, 10:155, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315001118.

Choudhary:2015:SBD

[401] Deepika Choudhary, Girish Kumar Gupta, Sukhbir Lal Khokra, and Nisha. Structure based designing and ADME-T studies of butenolide derivatives as potential agents against receptor ICAM-1: a drug target for cerebral malaria. Journal of Computational Science, 10:156–165, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000848.

Issar:2015:AMB

[402] Upasana Issar, Tripti Kumari, and Rita Kakkar. Assessment of molecular binding of Hoechst 33258 analogues into DNA using docking and MM/GBSA approach. Journal of Computational Science, 10:166–177, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000836.

Kumar:2015:CVM

[403] Dinesh Kumar, Rajesh Thipparaboina, and Nalini R. Shastri. Can vacuum morphologies predict solubility and intrinsic dissolution rate? a case study with felodipine polymorph form IV. Journal of Computational Science, 10:178–185, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000368.

Czerwoniec:2015:CNS

[404] Anna Czerwoniec, Malgorzata Szelag, Kajetan Juszczak, Joanna Wesoly, and Hans A. R. Bluyssen. CAVS — novel in silico selection strategy of specific STAT inhibitory compounds. Journal of Computational Science, 10:186–194, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000277.

Mishra:2015:DSS

[405] Abhishek Kumar Mishra. DFT study of structural, vibrational and electronic properties of polyaniline pernigraniline model compounds. Journal of Computational Science, 10: 195–208, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000137.

Murray:2015:ICN

[406] Jane S. Murray, Zenaida Peralta-Inga Shields, Paul G. Seybold, and Peter Politzer. Intuitive and counterintuitive noncovalent interactions of aromatic π regions with the hydrogen and the nitrogen of HCN. Journal of Computational Science, 10:209–216, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000113.

Kharkar:2015:SNA

[407] P. S. Kharkar, S. Borhade, A. Dangi, and S. Warrier. In search of novel anti-inflammatory agents: Computational repositioning of approved drugs. Journal of Computational Science, 10:217–224, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031500006X.

Sharma:2015:ASM

[408] Neha Sharma and Rita Kakkar. Adsorption of sarin on MgO nanotubes: Role of doped and defect sites. Journal of Computational Science, 10: 225–236, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750314001641.

Diwaker:2015:SSC

[409] Diwaker, C. S. Chidan Kumar, Ashish Kumar, Siddegowda Chandraju, Ching Kheng Quah, and Hoong-Kun Fun. Synthesis, spectroscopic characterization, electronic and optical studies of (2 Z )-5,6-dimethyl-2-[(4-nitrophenyl)methylidene]-2,3-dihydro-1-benzofuran-3-one. Journal of Computational Science, 10:237–246, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S1877750314001616.

Abramson:2015:PIC

[410] David Abramson, Valeria V. Krzhizhanovskaya, and Michael Lees. Perspectives of the International Conference of Computational Science 2014. Journal of Computational Science, 10: 247–248, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315001180.

Valero-Lara:2015:AFS

[411] Pedro Valero-Lara, Francisco D.

Igual, Manuel Prieto-Matías, Alfredo Pinelli, and Julien Favier. Accelerating fluid–solid simulations (Lattice-Boltzmann & Immersed-Boundary) on heterogeneous architectures. Journal of Computational Science, 10: 249–261, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315300028.

Dixit:2015:FGH

[412] Mayank Kumar Dixit, Amar Srivastava, and B. L. Tembe. Fourier Grid Hamiltonian Method for calculating the Einstein coefficients, Franck– Condon factors, r-Centroids, average internuclear separations and radiative lifetimes for N_{2} and CO molecules. Journal of Computational Science, 10:262–269, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000952.

Li:2015:TSV

[413] Dewei Li and Yingjie Tian. Twin support vector machine in linear programs. Journal of Computational Science, 10:270–277, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031500085X.

Zang:2015:LMS

[414] Wenyu Zang, Peng Zhang, Chuan Zhou, and Li Guo. Locating multiple sources in social networks under the SIR model: a divide-and-conquer approach. Journal of Computational Science, 10:278–287, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000824.

Krzhizhanovskaya:2015:RDD

[415] Valeria V. Krzhizhanovskaya, Alexey V. Dukhanov, Anna Bilyatdinova, Alexander V. Boukhanovsky, and Peter M. A. Sloot. Russian-Dutch double-degree Master’s programme in computational science in the age of global education. Journal of Computational Science, 10:288–298, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000812.

Bernabe:2015:ATT

[416] Gregorio Bernabé, Javier Cuenca, Luis-Pedro Garc´ ıa, and Domingo Giménez. Auto-tuning techniques for linear algebra routines on hybrid platforms. Journal of Computational Science, 10:299–310, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000794.

Hanasaki:2015:RTS

[417] Itsuo Hanasaki, Satoshi Uehara, and Satoyuki Kawano. Role of time scales for the non-Gaussianity of the Brownian motion combined with intermittent adsorption. Journal of Computational Science, 10:311–316, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031500040X.

Adinetz:2015:TFW

[418] Andrew Adinetz, Andreas Herten, Jiri Kraus, Marius C. Mertens, Dirk Pleiter, Tobias Stockmanns, and Peter Wintz. Triplet finder: On the way to triggerless online reconstruction with GPUs for the P̄ANDA experiment. Journal of Computational Science, 10:

317–326, September 2015. ISSN 1877-7503 (print), 1877-7511 (electronic).

Mendes:2015:DTS

[419] Celso L. Mendes, Brett Bode, Gregory H. Bauer, Jeremy Enos, Cristina Beldica, and William T. Kramer. Deployment and testing of the sustained petascale Blue Waters system. Journal of Computational Science, 10:327–337, September 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000344.

Anonymous:2015:EBe

[420] Anonymous. Editorial Board. Journal of Computational Science, 10: ifc, September 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315001222.

Anonymous:2015:PN

[421] Anonymous. Pages 1–330 (November 2015). Journal of Computational Science, 11:1–330, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Koziel:2015:EKB

[422] Slawomir Koziel and Leifur Leifsson. Efficient knowledge-based optimization of expensive computational models using adaptive response correction. Journal of Computational Science, 11:1–11, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300089.

Brumm:2015:SHD

[423] Johannes Brumm, Dmitry Mikushin, Simon Scheidegger, and Olaf Schenk. Scalable high-dimensional dynamic stochastic economic modeling. Journal of Computational Science, 11: 12–25, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315300053.

Yuan:2015:ULS

[424] Fong-Ching Yuan and Chao-Hui Lee. Using least square support vector regression with genetic algorithm to forecast beta systematic risk. Journal of Computational Science, 11: 26–33, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315300090.

Sevilla:2015:NSU

[425] Andrés Sevilla, Alberto Mozo, and Antonio Fernández Anta. Node sampling using Random Centrifugal Walks. Journal of Computational Science, 11:34–45, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300156.

Feinberg:2015:COS

[426] Jonathan Feinberg and Hans Petter Langtangen. Chaospy: an open source tool for designing methods of uncertainty quantification. Journal of Computational Science, 11: 46–57, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315300119.

Jacquemart:2015:SIB

[427] Damien Jacquemart, Jérôme Morio, and François Le Gland. Some ideas for bias and variance reduction in the splitting algorithm for diffusion processes. Journal of Computational Science, 11:58–68, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300193.

Elsebakhi:2015:LSM

[428] Emad Elsebakhi, Frank Lee, Eric Schendel, Anwar Haque, Nagarajan Kathireason, Tushar Pathare, Najeeb Syed, and Rashid Al-Ali. Largescale machine learning based on functional networks for biomedical big data with high performance computing platforms. Journal of Computational Science, 11:69–81, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300223.

Dongale:2015:MNM

[429] T. D. Dongale, K. P. Patil, S. R. Vanjare, A. R. Chavan, P. K. Gaikwad, and R. K. Kamat. Modelling of nanostructured memristor device characteristics using Artificial Neural Network (ANN). Journal of Computational Science, 11:82–90, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300302.

Noack:2015:TSM

[430] Marcus Noack. A two-scale method using a list of active sub-domains for a fully parallelized solution of wave equations. Journal of Computational Science, 11:91–101, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300296.

Berardi:2015:POD

[431] V. Berardi, R. Carretero-González, N. E. Klepeis, A. Palacios, J. Bellettiere, S. Hughes, S. Obayashi, and M. F. Hovell. Proper orthogonal decomposition methods for the analysis of real-time data: Exploring peak clustering in a secondhand smoke exposure intervention. Journal of Computational Science, 11:102–111, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300284.

Tavakoli:2015:SSS

[432] Hossein Tavakoli and Jahan B. Ghasemi. An in silico screening study and design of potent cognition agents. Journal of Computational Science, 11:112–120, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300211.

Paszynski:2015:ABS

[433] Maciej Paszyński. Agent-based simulations, adaptive algorithms and solvers.

Journal of Computational Science, 11:121–122, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315001295.

Cortes:2015:PEB

[434] A. M. A. Côrtes, A. L. G. A. Coutinho, L. Dalcin, and V. M. Calo. Performance evaluation of block-diagonal preconditioners for the divergence-conforming B-spline discretization of the Stokes system. Journal of Computational Science, 11: 123–136, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000095.

Alvarez-Aramberri:2015:SFB

[435] J. Alvarez-Aramberri, D. Pardo, and H. Barucq. A secondary field based hp-finite element method for the simulation of magnetotelluric measurements. Journal of Computational Science, 11:137–144, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315000150.

Wozniak:2015:FGI

[436] Maciej Woźniak. Fast GPU integration algorithm for isogeometric finite element method solvers using task dependency graphs. Journal of Computational Science, 11: 145–152, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315000253.

Krzywicki:2015:MCA

[437] D. Krzywicki, W. Turek, A. Byrski, and M. Kisiel-Dorohinicki. Massively concurrent agent-based evolutionary computing. Journal of Computational Science, 11:153–162, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300041.

Schaefer:2015:QLC

[438] Robert Schaefer, Marcin Łoś, Marcin Sieniek, Leszek Demkowicz, and Maciej Paszyński. Quasi-linear computational cost adaptive solvers for three dimensional modeling of heating of a human head induced by cell phone. Journal of Computational Science, 11:163–174, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300235.

Hsu:2015:MSI

[439] Ching-Hsien Hsu and Shangguang Wang. Modeling and simulation for intelligent mobility. Journal of Computational Science, 11:175– 176, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315001301.

Jabbarpour:2015:AIV

[440] Mohammad Reza Jabbarpour, Ali Jalooli, Erfan Shaghaghi, Alireza Marefat, Rafidah Md Noor, and Jason J. Jung. Analyzing the impacts of velocity and density on intelligent position-based routing protocols. Journal of Computational Science, 11:177–184, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001598.

Bravo-Torres:2015:VVL

[441] Jack F. Bravo-Torres, Martín López-Nores, Yolanda Blanco-Fernández, José J. Pazos-Arias, and Esteban F. Ordóñez-Morales. VaNetLayer: a virtualization layer supporting access to web contents from within vehicular networks. Journal of Computational Science, 11:185–195, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001008.

Li:2015:MEA

[442] Xuran Li, Jianlong Xu, Hong-Ning Dai, Qinglin Zhao, Chak Fong Cheang, and Qiu Wang. On modeling eavesdropping attacks in wireless networks. Journal of Computational Science, 11:196–204, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750314001562.

Haryanto:2015:GRK

[443] Anasthasia Agnes Haryanto, David Taniar, and Kiki Maulana Adhinugraha. Group Reverse kNN Query optimisation. Journal of Computational Science, 11:205–221, November 2015.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031530020X.

Was:2015:CAA

[444] Jarosław Wa̧s and Georgios Ch. Sirakoulis. Cellular automata applications for research and industry. Journal of Computational Science, 11: 223–225, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315001313.

Yamagishi:2015:FVM

[445] Seiya Yamagishi and Shin Morishita. Friction vibration modeling with detachment and adhesion by cellular automata. Journal of Computational Science, 11:226–232, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300326.

Sen:2015:THD

[446] Bibhash Sen, Anirban Nag, Asmit De, and Biplab K. Sikdar. Towards the hierarchical design of multilayer QCA logic circuit. Journal of Computational Science, 11:233–244, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300247.

Bolt:2015:DSC

[447] Witold Bołt, Jan M. Baetens, and Bernard De Baets. On the decomposition of stochastic cellular automata. Journal of Computational Science, 11:245–257, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300181.

Arca:2015:GAM

[448] Bachisio Arca, Tiziano Ghisu, and Giuseppe A. Trunfio. GPU-accelerated multi-objective optimization of fuel treatments for mitigating wildfire hazard. Journal of Computational Science, 11:258–268, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300120.

Perez-Brokate:2015:CVS

[449] Cristian Felipe Pérez-Brokate, Dung di Caprio, Éric Mahé, Damien Féron, and Jacques de Lamare. Cyclic voltammetry simulations with cellular automata. Journal of Computational Science, 11:269–278, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300107.

Mozafari:2015:CLM

[450] Milad Mozafari, Mohammad Ebrahim Shiri, and Hamid Beigy. A cooperative learning method based on cellular learning automata and its application in optimization problems. Journal of Computational Science, 11:279–288, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300077.

Machado:2015:CMS

[451] G. Machado, V. Lupiano, M. V. Avolio, F. Gullace, and S. Di Gregorio. A cellular model for secondary lahars and simulation of cases in the Vascún Valley, Ecuador. Journal of Computational Science, 11:289–299, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300065.

Dzwinel:2015:IEC

[452] Witold Dzwinel and Krzysztof Magiera. Irreducible elementary cellular automata found. Journal of Computational Science, 11:300–308, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S187775031530003X.

Bartosik:2015:CAM

[453] Łukasz Bartosik, Janusz Stafiej, and Dung Di Caprio. Cellular automata model of anodization. Journal of Computational Science, 11: 309–316, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315000964.

Kireeva:2015:TLC

[454] Anastasiia Kireeva. Two-layer CA for simulation of catalytic reaction at dynamically varying surface temperature. Journal of Computational Science, 11:317–325, November 2015. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315000885.

Karafyllidis:2015:QWG

[455] Ioannis G. Karafyllidis. Quantum walks on graphene nanoribbons using quantum gates as coins. Journal of Computational Science, 11: 326–330, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315000861.

Anonymous:2015:EBf

[456] Anonymous. Editorial Board. Journal of Computational Science, 11: ifc, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315001337.

Anonymous:2016:PJ

[457] Anonymous. Pages 1–94 (January 2016). Journal of Computational Science, 12:1–94, January 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Viswanathan:2016:SAE

[458] Vaisagh Viswanathan, Daniel Zehe, Jordan Ivanchev, Dominik Pelzer, Alois Knoll, and Heiko Aydt. Simulationassisted exploration of charging infrastructure requirements for electric vehicles in urban environments. Journal of Computational Science, 12:1–10, January 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031530034X.

Sandhu:2016:SMC

[459] Rajinder Sandhu, Harsuminder K. Gill, and Sandeep K. Sood. Smart monitoring and controlling of Pandemic Influenza A (H1N1) using Social Network Analysis and cloud computing. Journal of Computational Science, 12:11–22, January 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300351.

Lahmiri:2016:ISP

[460] Salim Lahmiri. Intraday stock price forecasting based on variational mode decomposition. Journal of Computational Science, 12:23–27, January 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300466.

Liu:2016:MFS

[461] Bo Liu, Slawomir Koziel, and Qingfu Zhang. A multi-fidelity surrogatemodel-assisted evolutionary algorithm for computationally expensive optimization problems. Journal of Computational Science, 12: 28–37, January 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315300387.

Bin:2016:MSM

[462] Sun Bin and Zhaoxia Li. Multiscale modeling and trans-level simulation from material meso-damage to structural failure of reinforced concrete frame structures under seismic loading. Journal of Computational Science, 12:38–50, January 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300375.

Alizadeh:2016:SMT

[463] As’ad Alizadeh and Abdolrahman Dadvand. Simulation of the motion of two elastic membranes in Poiseuille shear flow via a combined immersed boundary-lattice Boltzmann method. Journal of Computational Science, 12:51–61, January 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300429.

Mattila:2016:PCP

[464] Keijo Mattila, Tuomas Puurtinen, Jari Hyväluoma, Rodrigo Surmas, Markko Myllys, Tuomas Turpeinen, Fredrik Robertsén, Jan Westerholm, and Jussi Timonen. A prospect for computing in porous materials research: Very large fluid flow simulations. Journal of Computational Science, 12:62–76, January 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300478.

Zhao:2016:ECT

[465] Guannan Zhao, Zhenyuan Zhao, and Neil F. Johnson. Efficient computational testing of scale-free behavior in real-world systems. Journal of Computational Science, 12: 77–82, January 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315300430.

Celdran:2016:DRS

[466] Alberto Huertas Celdrán, Manuel Gil Pérez, Félix J. García Clemente, and Gregorio Martínez Pérez. Design of a recommender system based on users’ behavior and collaborative location and tracking. Journal of Computational Science, 12:83–94, January 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300454.

Anonymous:2016:EBa

[467] Anonymous. Editorial Board. Journal of Computational Science, 12:ifc, January 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S187775031600003X.

Anonymous:2016:PM

[468] Anonymous. Pages 1–96 (March 2016). Journal of Computational Science, 13: 1–96, March 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Martinez:2016:ADP

[469] V´ ıctor Mart´ ınez, Fernando Berzal, and Juan-Carlos Cubero. Adaptive degree penalization for link prediction. Journal of Computational Science, 13:1–9, March 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031530051X.

Bashir:2016:HMD

[470] Saba Bashir, Usman Qamar, Farhan Hassan Khan, and Lubna Naseem. HMV: a medical decision support framework using multi-layer classifiers for disease prediction. Journal of Computational Science, 13:10–25, March 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300011.

Celik:2016:DMD

[471] Fatih Çelik. DEVS-m: a discrete event simulation framework for MANETs. Journal of Computational Science, 13:26–36, March 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300442.

Fernandez:2016:BPS

[472] M. Fernández, Juan C. Pichel, José C. Cabaleiro, and Tomás F. Pena. Boosting performance of a Statistical Machine Translation system using dynamic parallelism. Journal of Computational Science, 13:37–48, March 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300023.

Hosseini:2016:MPM

[473] Soodeh Hosseini, Mohammad Abdollahi Azgomi, and Adel Torkaman Rahmani. Malware propagation modeling considering software diversity and immunization. Journal of Computational Science, 13:49–67, March 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300035.

Ngo:2016:CPS

[474] Thi Thuy Ngo, Ali Sadollah, and Joong Hoon Kim. A cooperative particle swarm optimizer with stochastic movements for computationally expensive numerical optimization problems. Journal of Computational Science, 13:68–82, March 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300047.

He:2016:DSC

[475] Hongsheng He, Dali Wang, Yang Xu, and Jindong Tan. Data synthesis in the Community Land Model for ecosystem simulation. Journal of Computational Science, 13:83–95, March 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300059.

Anonymous:2016:EBb

[476] Anonymous. Editorial Board. Journal of Computational Science, 13:ifc, March 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750316300187.

Alexandrov:2016:REN

[477] Vassil Alexandrov. Route to exascale: Novel mathematical methods, scalable algorithms and Computational Science skills. Journal of Computational Science, 14:1–4, May 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300527.

Emad:2016:UCA

[478] Nahid Emad and Serge Petiton. Unite and conquer approach for high scale numerical computing. Journal of Computational Science, 14:5–14, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300084.

Vazquez:2016:AME

[479] Mariano Vázquez, Guillaume Houzeaux, Seid Koric, Antoni Artigues, Jazmin Aguado-Sierra, Ruth Arís, Daniel Mira, Hadrien Calmet, Fernando Cucchietti, Herbert Owen, Ahmed Taha, Evan Dering Burness, José María Cela, and Mateo Valero. Alya: Multiphysics engineering simulation toward exascale. Journal of Computational Science, 14:15–27, May 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300521.

Sanjuan:2016:DMP

[480] Gemma Sanjuan, Carlos Brun, Tomàs Margalef, and Ana Cortés. Determining map partitioning to minimize wind field uncertainty in forest fire propagation prediction. Journal of Computational Science, 14:28–37, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300060.

Song:2016:CCO

[481] Sukhyun Song and Jeffrey K. Hollingsworth. Computation–communication overlap and parameter auto-tuning for scalable parallel 3-D FFT. Journal of Computational Science, 14:38–50, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031530048X.

Elliott:2016:EDR

[482] J. Elliott, M. Hoemmen, and F. Mueller. Exploiting data representation for fault tolerance. Journal of Computational Science, 14:51–60, May 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300491.

Liu:2016:RLL

[483] Dalian Liu, Yong Shi, Yingjie Tian, and Xiankai Huang. Ramp loss least squares support vector machine. Journal of Computational Science, 14:61–68, May 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300096.

Alexandrov:2016:ETE

[484] Nia Alexandrov. Education and training for exascale. Journal of Computational Science, 14:69–73, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300515.

Sancho:2016:BBP

[485] Maria-Ribera Sancho. BSC Best Practices in Professional Training and Teaching for the HPC Ecosystem. Journal of Computational Science, 14:74–77, May 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300508.

Frank:2016:NPE

[486] Anton Frank, Matti Heikkurinen, Ferdinand Jamitzky, Helmut Satzger, and Dieter Kranzlmüller. In need of partnerships — an essay about the collaboration between computational sciences and IT services. Journal of Computational Science, 14:78–84, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300072.

Schiffers:2016:DCS

[487] Michael Schiffers, Nils gentschen Felde, and Dieter Kranzlmüller. I have a DRIHM: a case study in lifting computational science services up to the scientific mainstream. Journal of Computational Science, 14:85–89, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300533.

Anonymous:2016:EBc

[488] Anonymous. Editorial Board. Journal of Computational Science, 14:ifc, May 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750316300540.

Bouhali:2016:ICS

[489] O. Bouhali, I. G. Economou, and F. El-Mellouhi. The International Computational Science and Engineering Conference. Journal of Computational Science, 15:1–2, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316000089.

Al-Ali:2016:WOP

[490] Rashid Al-Ali, Nagarajan Kathiresan, Mohammed El Anbari, Eric R. Schendel, and Tariq Abu Zaid. Workflow optimization of performance and quality of service for bioinformatics application in high performance computing. Journal of Computational Science, 15:3–10, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300278.

Thareja:2016:GBM

[491] Gaurav Thareja, Ziad Kronfol, Karsten Suhre, and Pankaj Kumar. A graph based method for depicting population characteristics using Genome Wide Data. Journal of Computational Science, 15:11–17, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300545.

Silani:2016:NDM

[492] Mohammad Silani, Hossein Talebi, Abdel Magid Hamouda, and Timon Rabczuk. Nonlocal damage modelling in clay/epoxy nanocomposites using a multiscale approach.

Journal of Computational Science, 15:18–23, July 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315300417.

Laurini:2016:PMN

[493] Erik Laurini, Paola Posocco, Maurizio Fermeglia, and Sabrina Pricl. MoDeNa Nanotools: an integrated multiscale simulation workflow to predict thermophysical properties of thermoplastic polyurethanes. Journal of Computational Science, 15:24–33, July 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300405.

Baklouti:2016:IRK

[494] Raoudha Baklouti, Majdi Mansouri, Mohamed Nounou, Hazem Nounou, and Ahmed Ben Hamida. Iterated Robust kernel Fuzzy Principal Component Analysis and application to fault detection. Journal of Computational Science, 15:34–49, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300399.

Mohammadzaheri:2016:IAO

[495] Morteza Mohammadzaheri, Reza Tafreshi, Zurwa Khan, Matthew Franchek, and Karolos Grigoriadis. An intelligent approach to optimize multiphase subsea oil fields lifted by electrical submersible pumps. Journal of Computational Science, 15:50–59, July 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750315300314.

Lee:2016:HDS

[496] One-Sun Lee. How does ss-DNA recognize the chirality of carbon nanotubes? Journal of Computational Science, 15:60–64, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300260.

Luthi:2016:QCS

[497] Hans P. Lüthi, Stefan Heinen, Gisbert Schneider, Andreas Glöss, Martin P. Brändle, Rollin A. King, Edward Pyzer-Knapp, Fahhad H. Alharbi, and Sabre Kais. The quantum chemical search for novel materials and the issue of data processing: the InfoMol project. Journal of Computational Science, 15:65–73, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300272.

Ramdin:2016:SCC

[498] Mahinder Ramdin, Qu Chen, Sayee Prasaad Balaji, José Manuel Vicent-Luna, Ariana Torres-Knoop, David Dubbeldam, Sofía Calero, Theo W. de Loos, and Thijs J. H. Vlugt. Solubilities of CO_{2}, CH_{4}, C_{2}H_{6}, and SO_{2} in ionic liquids and Selexol from Monte Carlo simulations. Journal of Computational Science, 15:74–80, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750315300168.

Cheimarios:2016:MPM

[499] N. Cheimarios, G. Kokkoris, and A. G. Boudouvis. A multi-parallel multiscale computational framework for chemical vapor deposition processes. Journal of Computational Science, 15:81–85, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300132.

Becker:2016:IPE

[500] Tim M. Becker, David Dubbeldam, Li-Chiang Lin, and Thijs J. H. Vlugt. Investigating polarization effects of CO _{2} adsorption in MgMOF-

74. Journal of Computational Science, 15:86–94, July 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300144.

Anonymous:2016:EBd

[501] Anonymous. Editorial Board. Journal of Computational Science, 15:ifc, July 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S187775031630093X.

Anonymous:2016:PS

[502] Anonymous. Pages 1–226 (September 2016). Journal of Computational Science, 16:1–226, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Clark:2016:SRO

[503] Eric M. Clark, Jake Ryland Williams, Chris A. Jones, Richard A. Galbraith, Christopher M. Danforth, and Peter Sheridan Dodds. Sifting robotic from organic text: a natural language approach for detecting automation on Twitter. Journal of Computational Science, 16:1–7, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300363.

Ahmad:2016:TSL

[504] Firdous Ahmad, Ghulam Mohiuddin Bhat, Hossein Khademolhosseini, Saeid Azimi, Shaahin Angizi, and Keivan Navi. Towards single layer quantumdot cellular automata adders based on explicit interaction of cells. Journal of Computational Science, 16: 8–15, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316300126.

Prestininzi:2016:CBM

[505] P. Prestininzi, V. Lombardi, and M. La Rocca. Curved boundaries in multi-layer Shallow Water Lattice Boltzmann Methods: bounce back versus immersed boundary. Journal of Computational Science, 16: 16–28, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316300151.

Korecki:2016:ASF

[506] Tomasz Korecki, Dariusz Pałka, and Jarosław Wa̧s. Adaptation of Social Force Model for simulation of downhill skiing. Journal of Computational Science, 16:29–42, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300138.

Liu:2016:LPA

[507] Jie Liu, Baomin Xu, Xiang Xu, and Tinglin Xin. A link prediction algorithm based on label propagation. Journal of Computational Science, 16:43–50, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300382.

Luo:2016:DML

[508] Peng Luo, Yongli Li, Chong Wu, and Kun Chen. Detecting the missing links in social networks based on utility analysis. Journal of Computational Science, 16:51–58, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300746.

Lopez:2016:MNA

[509] Francisco Javier Mart´ ınez López, José Antonio Torres Arriaza, Sergio Mart´ ınez Puertas, and Mar´ ıa Mercedes Peralta López. Multilevel neuronal architecture to resolve classification problems with large training sets: Parallelization of the training process. Journal of Computational Science, 16:59–64, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300400.

Canaval:2016:SOP

[510] Sandra Gómez Canaval, José Ramón Sánchez Couso, and Meritxell Vinyals. Solving optimization problems by using networks of evolutionary processors with quantitative filtering. Journal of Computational Science, 16: 65–71, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316300850.

Raji:2016:GSP

[511] C. G. Raji and S. S. Vinod Chandra. Graft survival prediction in liver transplantation using artificial neural network models. Journal of Computational Science, 16: 72–78, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316300825.

Loffler:2016:NPS

[512] Frank Löffler, Zhoujian Cao, Steven R. Brandt, and Zhihui Du. A new parallelization scheme for adaptive mesh refinement. Journal of Computational Science, 16:79–88, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300795.

Sharp:2016:GEB

[513] P. W. Sharp and W. I. Newman. GPUenabled N-body simulations of the Solar System using a VOVS Adams integrator. Journal of Computational Science, 16:89–97, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300412.

Sitko:2016:TLS

[514] Mateusz Sitko, Maciej Pietrzyk, and Lukasz Madej. Time and length scale issues in numerical modelling of dynamic recrystallization based on the multi space cellular automata method. Journal of Computational Science, 16:98–113, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300849.

Heeroo:2016:LCA

[515] Kelvin N. S. Heeroo, Oomesh Gukhool, and Dristesh Hoorpah. A Ludo Cellular Automata model for microscopic traffic flow. Journal of Computational Science, 16:114–127, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300734.

Harrison:2016:NTF

[516] Adam P. Harrison and Dileepan Joseph. Numeric tensor framework: Exploiting and extending Einstein notation. Journal of Computational Science, 16:128–139, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300813.

Regis:2016:MOC

[517] Rommel G. Regis. Multi-objective constrained black-box optimization using radial basis function surrogates.

Journal of Computational Science, 16:140–155, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300904.

Streletz:2016:ISS

[518] Gregory J. Streletz, Geoffrey Gebbie, Oliver Kreylos, Bernd Hamann, Louise H. Kellogg, and Howard J. Spero. Interpolating sparse scattered data using flow information. Journal of Computational Science, 16:156–169, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300394.

Litescu:2016:EIU

[519] Sorina Costache Litescu, Vaisagh Viswanathan, Heiko Aydt, and Alois Knoll. The effect of information uncertainty in road transportation systems. Journal of Computational Science, 16:170–176, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300801.

Cordero-Sanchez:2016:PNS

[520] S. Cordero-Sánchez, F. Rojas-González, G. Román-Alonso, M. A. Castro-García, M. Aguilar-Cornejo, and J. Matadamas-Hernández. Pore networks subjected to variable connectivity and geometrical restrictions: a simulation employing a multicore system. Journal of Computational Science, 16:177–189, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301041.

Racca:2016:ROF

[521] P. Racca, R. Casarin, F. Squazzoni, and P. Dondio. Resilience of an online financial community to market uncertainty shocks during the recent financial crisis. Journal of Computational Science, 16:190–199, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301156.

Das:2016:TDM

[522] R. Das and P. W. Cleary. Threedimensional modelling of coupled flow dynamics, heat transfer and residual stress generation in arc welding processes using the mesh-free SPH method. Journal of Computational Science, 16:200–216, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300266.

Fernandes:2016:OXB

[523] Steven Lawrence Fernandes and G. Josemin Bala. ODROID XU4 based implementation of decision level fusion approach for matching computer generated sketches. Journal of Computational Science, 16: 217–224, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301211.

Melnikova:2016:CEU

[524] N. B. Melnikova, D. Jordan, and V. V. Krzhizhanovskaya. Corrigendum to “Experience of using FEM for real-time flood early warning systems: Monitoring and modelling Boston levee instability”. Journal of Computational Science, 16:225–226, September 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630014X. See

[386].

Anonymous:2016:EBe

[525] Anonymous. Editorial Board. Journal of Computational Science, 16: ifc, September 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316301302.

Anonymous:2016:PN

[526] Anonymous. Pages 1–306 (November 2016). Journal of Computational Science, 17 (part 1):1–306, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Kumar:2016:PSV

[527] Deepak Kumar, Suraj S. Meghwani, and Manoj Thakur. Proximal support vector machine based hybrid prediction models for trend forecasting in financial markets. Journal of Computational Science, 17 (part 1):1–13, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316301144.

Silva:2016:BDM

[528] Orlando Silva. Black Death — model and simulation. Journal of Computational Science, 17 (part 1): 14–34, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301247.

Bruneau:2016:HPC

[529] Charles-Henri Bruneau and Khodor Khadra. Highly parallel computing of a multigrid solver for 3D Navier–Stokes equations. Journal of Computational Science, 17 (part 1): 35–46, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301478.

Xiong:2016:SUM

[530] Heng Xiong and Rogemar Mamon. A self-updating model driven by a higher-order hidden Markov chain for temperature dynamics. Journal of Computational Science, 17 (part 1): 47–61, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301466.

Sinkovits:2016:FDS

[531] Robert S. Sinkovits, James Moody, B. Tolga Oztan, and Douglas R. White. Fast determination of structurally cohesive subgroups in large networks. Journal of Computational Science, 17 (part 1):62–72, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316302058.

Shakibian:2016:MAL

[532] Hadi Shakibian, Nasrollah Moghadam Charkari, and Saeed Jalili. A multilayered approach for link prediction in heterogeneous complex networks. Journal of Computational Science, 17 (part 1):73–82, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031630179X.

Zarei:2016:IGT

[533] Mohammad Zarei and Abolfazl Salami. Improving the game theoretic analysis of electricity auctions applied in medium markets. Journal of Computational Science, 17 (part 1): 83–96, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316302307.

LimadeCastro:2016:EUP

[534] Paulo André Lima de Castro, Anderson Rodrigo Barreto Teodoro, Luciano Irineu de Castro, and Simon Parsons. Expected utility or prospect theory: which better fits agent-based modeling of markets? Journal of Computational Science, 17 (part 1): 97–102, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301806.

Dogra:2016:CGS

[535] Ayush Dogra, Sunil Agrawal, Bhawna Goyal, Niranjan Khandelwal, and Chirag Kamal Ahuja. Color and grey scale fusion of osseous and vascular information. Journal of Computational Science, 17 (part 1): 103–114, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301454.

Dohi:2016:SSC

[536] Tadashi Dohi and Hisashi Yamamoto. Special section on computational reliability in engineering and science. Journal of Computational Science, 17 (part 1):115–116, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031630357X.

Bae:2016:BDM

[537] Suk Joo Bae, Tao Yuan, and Seong joon Kim. Bayesian degradation modeling for reliability prediction of organic light-emitting diodes. Journal of Computational Science, 17 (part 1):117–125, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301387.

Li:2016:RNR

[538] Yan Li, Lirong Cui, and He Yi. Reliability of non-repairable systems with cyclic-mission switching and multimode failure components. Journal of Computational Science, 17 (part 1):126–138, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300916.

Chen:2016:SMF

[539] Shin-Guang Chen and Yi-Kuei Lin. Searching for d-MPs with fast enumeration. Journal of Computational Science, 17 (part 1):139– 147, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316300886.

Wen:2016:SNP

[540] Yanqing Wen, Lirong Cui, Shubin Si, and Baoliang Liu. Several new performance measures for Markov system with stochastic supply patterns and stochastic demand patterns. Journal of Computational Science, 17 (part 1):148–155, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300862.

Ohi:2016:SEM

[541] Fumio Ohi. Stochastic evaluation methods of a multi-state system via a modular decomposition. Journal of Computational Science, 17 (part 1):156–169, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630045X.

Cotta:2016:CM

[542] Carlos Cotta and Robert Schaefer. Complex metaheuristics. Journal of Computational Science, 17 (part 1):171–173, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301028.

Li:2016:DBC

[543] Hongye Li, Lei Wang, and Xinhong Hei. Decomposition-based chemical reaction optimization (CRO) and an extended CRO algorithms for multiobjective optimization. Journal of Computational Science, 17 (part 1):174–204, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031530017X.

Harrison:2016:MAC

[544] Kyle Robert Harrison, Mario Ventresca, and Beatrice M. Ombuki-Berman. A meta-analysis of centrality measures for comparing and generating complex network models. Journal of Computational Science, 17 (part 1):205–215, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750315300259.

Hernandez:2016:EMO

[545] Daniel E. Hernández, Eddie Clemente, Gustavo Olague, and José L. Briseño. Evolutionary multi-objective visual cortex for object classification in natural images. Journal of Computational Science, 17 (part 1):216– 233, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315300338.

Turek:2016:HSE

[546] Wojciech Turek, Jan Stypka, Daniel Krzywicki, Piotr Anielski, Kamil Pietak, Aleksander Byrski, and Marek Kisiel-Dorohinicki. Highly scalable Erlang framework for agent-based metaheuristic computing. Journal of Computational Science, 17 (part 1):234–248, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300242.

Lazarz:2016:HGS

[547] Radosław Łazarz, Michał Idzik, Konrad Ga̧dek, and Ewa Gajda-Zagórska. Hierarchic genetic strategy with maturing as a generic tool for multiobjective optimization. Journal of Computational Science, 17 (part 1): 249–260, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316300254.

Obuchowicz:2016:ASM

[548] Adam K. Obuchowicz and Maciej Smołka. Application of α-stable mutation in a hierarchic evolutionary inverse solver. Journal of Computational Science, 17 (part 1):261– 269, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750315300557.

Ramasami:2016:PVC

[549] Ponnadurai Ramasami. Proceedings of the Virtual Conference on Computational Science (VCCC-2015): Special Issue of the Journal of Computational Science. Journal of Computational Science, 17 (part 1):271– 272, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316303581.

Riley:2016:CAE

[550] Kevin E. Riley, Khanh-An Tran, Pat Lane, Jane S. Murray, and Peter Politzer. Comparative analysis of electrostatic potential maxima and minima on molecular surfaces, as determined by three methods and a variety of basis sets. Journal of Computational Science, 17 (part 1): 273–284, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031630031X.

Choong:2016:GOO

[551] Yee Siew Choong and Roy Lee Yung-Hung. A general overview on outer membrane protein (Omp) simulations. Journal of Computational Science, 17 (part 1):285–291, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300291.

Ahmad:2016:ICT

[552] K. Ahmad, M. H. Baig, Girish Kumar Gupta, M. A. Kamal, Neelam Pathak, and I. Choi. Identification of common therapeutic targets for selected neurodegenerative disorders: an in silico approach. Journal of Computational Science, 17 (part 1):292–306, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630028X.

Henrich:2016:ICD

[553] Oliver Henrich, Timm Krüger, Rupert W. Nash, Dhiraj V. Patil, and Kevin Stratford. The 24th International Conference on Discrete Simulation of Fluid Dynamics in Edinburgh, Scotland. Journal of Computational Science, 17 (part 2):307– 308, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316302800. Held

July 13–17, 2015.

Reijers:2016:AML

[554] S. A. Reijers, H. Gelderblom, and F. Toschi. Axisymmetric multiphase lattice Boltzmann method for generic equations of state. Journal of Computational Science, 17 (part 2):309–314, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300898.

Takada:2016:PFM

[555] Naoki Takada, Junichi Matsumoto, Sohei Matsumoto, and Kazuma Kurihara. Phase-field model-based simulation of two-phase fluid motion on partially wetted and textured solid surface. Journal of Computational Science, 17 (part 2):315–324, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300874.

Lauricella:2016:DMR

[556] Marco Lauricella, Giuseppe Pontrelli, Dario Pisignano, and Sauro Succi. Dynamic mesh refinement for discrete models of jet electro-hydrodynamics. Journal of Computational Science, 17 (part 2):325–333, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300783.

Otomo:2016:SAM

[557] Hiroshi Otomo, Hongli Fan, Yong Li, Marco Dressler, Ilya Staroselsky, Raoyang Zhang, and Hudong Chen. Studies of accurate multicomponent lattice Boltzmann models on benchmark cases required for engineering applications. Journal of Computational Science, 17 (part 2):334–339, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300771.

Wang:2016:TDP

[558] Ningning Wang, Haihu Liu, and Chuhua Zhang. Three-dimensional phase-field lattice Boltzmann model for incompressible multiphase flows. Journal of Computational Science, 17 (part 2):340–356, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300503.

DiStaso:2016:DLM

[559] G. Di Staso, H. J. H. Clercx, S. Succi, and F. Toschi. DSMC– LBM mapping scheme for rarefied and non-rarefied gas flows. Journal of Computational Science, 17 (part 2):357–369, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300497.

Busik:2016:SSR

[560] Martin Bušík, Iveta Jančigová, Renáta Tóthová, and Ivan Cimrák. Simulation study of rare cell trajectories and capture rate in periodic obstacle arrays. Journal of Computational Science, 17 (part 2):370–376, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300473.

Montessori:2016:EKD

[561] A. Montessori, P. Prestininzi, M. La Rocca, G. Falcucci, S. Succi, and E. Kaxiras. Effects of Knudsen diffusivity on the effective reactivity of nanoporous catalyst media. Journal of Computational Science, 17 (part 2):377–383, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300448.

Kullmer:2016:NOP

[562] Knut Küllmer, Andreas Krämer, Dirk Reith, Wolfgang Joppich, and Holger Foysi. Numerical optimisation of the pseudopotential-based lattice Boltzmann method. Journal of Computational Science, 17 (part 2):384–393, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300436.

Morrison:2016:SFT

[563] Helen E. Morrison, Martin Brede, Guido Dehnhardt, and Alfred Leder. Simulating the flow and trail following capabilities of harbour seal vibrissae with the Lattice Boltzmann Method. Journal of Computational Science, 17 (part 2):394–402, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300424.

Ambrus:2016:AMQ

[564] Victor E. Ambruş and Victor Sofonea. Application of mixed quadrature lattice Boltzmann models for the simulation of Poiseuille flow at non-negligible values of the Knudsen number. Journal of Computational Science, 17 (part 2):403–417, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300370.

Ba:2016:TDL

[565] Yan Ba, Qinjun Kang, Haihu Liu, Jinju Sun, and Chao Wang. Threedimensional lattice Boltzmann simulations of microdroplets including contact angle hysteresis on topologically structured surfaces. Journal of Computational Science, 17 (part 2):418–430, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300369.

Mink:2016:LBM

[566] Albert Mink, Gudrun Thäter, Hermann Nirschl, and Mathias J. Krause. A 3D Lattice Boltzmann method for light simulation in participating media. Journal of Computational Science, 17 (part 2):431–437, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300357.

Trunk:2016:IDP

[567] Robin Trunk, Thomas Henn, Willy Dörfler, Hermann Nirschl, and Mathias J. Krause. Inertial dilute particulate fluid flow simulations with an Euler–Euler lattice Boltzmann method. Journal of Computational Science, 17 (part 2):438–445, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300345.

Korotkin:2016:TPF

[568] Ivan Korotkin, Dmitry Nerukh, Elvira Tarasova, Vladimir Farafonov, and Sergey Karabasov. Two-phase flow analogy as an effective boundary condition for modelling liquids at atomistic resolution. Journal of Computational Science, 17 (part 2):446– 456, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316300333.

Zhang:2016:BRD

[569] Yi Ran Zhang, Xi Zhuo Jiang, and Kai Hong Luo. Bounce regime of droplet collisions: a molecular dynamics study. Journal of Computational Science, 17 (part 2):457– 462, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316300321.

Ioannou:2016:DDC

[570] N. Ioannou, H. Liu, and Y. H. Zhang. Droplet dynamics in confinement. Journal of Computational Science, 17 (part 2):463–474, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300308.

Shan:2016:MSL

[571] Xiaowen Shan. The mathematical structure of the lattices of the lattice Boltzmann method. Journal of Computational Science, 17 (part 2):475–481, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316300163.

Neumann:2016:THL

[572] Philipp Neumann. On transient hybrid Lattice Boltzmann–Navier– Stokes flow simulations. Journal of Computational Science, 17 (part 2):482–490, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300114.

Strazdins:2016:RAP

[573] Peter Strazdins, Raphaël Couturier, and Laurence T. Yang. Recent advances in parallel techniques for scientific computing [preface]. Journal of Computational Science, 17 (part 3):491–493, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316303714.

Eghbal:2016:AME

[574] A. Eghbal, A. G. Gerber, and E. Aubanel. Algebraic multigrid employing mixed structured–unstructured data on manycore hardware. Journal of Computational Science, 17 (part 3):494–508, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316301120.

Gmeiner:2016:QPS

[575] Björn Gmeiner, Markus Huber, Lorenz John, Ulrich Rüde, and Barbara Wohlmuth. A quantitative performance study for Stokes solvers at the extreme scale. Journal of Computational Science, 17 (part 3):

509–521, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301077.

Fasi:2016:BFR

[576] Massimiliano Fasi, Julien Langou, Yves Robert, and Bora Uçar. A backward/forward recovery approach for the preconditioned conjugate gradient method. Journal of Computational Science, 17 (part 3):522– 534, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316300461.

Couturier:2016:TTS

[577] Raphaël Couturier, Lilia Ziane Khodja, and Christophe Guyeux. TSIRM: a two-stage iteration with leastsquares residual minimization algorithm to solve large sparse linear and nonlinear systems. Journal of Computational Science, 17 (part 3): 535–546, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316302721.

Strazdins:2016:DAT

[578] Peter E. Strazdins, Md. Mohsin Ali, and Brendan Harding. Design and analysis of two highly scalable sparse grid combination algorithms. Journal of Computational Science, 17 (part 3):547–561, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316301053.

Fanfakh:2016:OEC

[579] Ahmed Fanfakh, Jean-Claude Charr, Raphaël Couturier, and Arnaud Giersch. Optimizing the energy consumption of message passing applications with iterations executed over grids. Journal of Computational Science, 17 (part 3):562–575, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630120X.

Sanjuan:2016:HAA

[580] G. Sanjuan, T. Margalef, and A. Cortés. Hybrid application to accelerate wind field calculation. Journal of Computational Science, 17 (part 3):576–590, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316301193.

Zhu:2016:PNL

[581] Huming Zhu, Yanfei Wu, Pei Li, Duo Wang, Wei Shi, Peng Zhang, and Licheng Jiao. A parallel non-local means denoising algorithm implementation with OpenMP and OpenCL on Intel Xeon Phi Coprocessor. Journal of Computational Science, 17 (part 3):591–598, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316301090.

Vazquez:2016:NPS

[582] Sergio Vázquez, María J. Martín, Basilio B. Fraguela, Andrés Gómez, Aurelio Rodríguez, and Bjarki Thór Elvarsson. Novel parallelization of simulated annealing and Hooke & Jeeves search algorithms for multicore systems with application to complex fisheries stock assessment models. Journal of Computational Science, 17 (part 3):599–608, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301119.

Eiras-Franco:2016:MSP

[583] Carlos Eiras-Franco, Verónica Bolón-Canedo, Sabela Ramos, Jorge González-Domínguez, Amparo Alonso-Betanzos, and Juan Touriño. Multithreaded and Spark parallelization of feature selection filters. Journal of Computational Science, 17 (part 3): 609–619, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301107.

Fonseca:2016:CGA

[584] Alcides Fonseca and Bruno Cabral. Controlling the granularity of automatic parallel programs. Journal of Computational Science, 17 (part 3):620–629, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301065.

Pedrero:2016:CAS

[585] Manuel Pedrero, Eladio Gutierrez, Sergio Romero, and Oscar Plata. A comparative analysis of STM approaches to reduction operations in irregular applications. Journal of Computational Science, 17 (part 3): 630–638, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316301272.

Saad:2016:WAP

[586] Tony Saad and James C. Sutherland. Wasatch: an architectureproof multiphysics development environment using a Domain Specific Language and graph theory. Journal of Computational Science, 17 (part 3):639–646, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316300485.

Gu:2016:PTV

[587] Yonggen Gu, Jie Tao, Guoqiang Li, Daniel W. Sun, Xiaohong Wu, Prem Prakash Jayaraman, and Rajiv Ranjan. A preemptive truthful VMs allocation online mechanism in private cloud. Journal of Computational Science, 17 (part 3): 647–653, November 2016. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316300837.

Anonymous:2016:EBf

[588] Anonymous. Editorial Board. Journal of Computational Science, 17 (part 1):ifc, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303404.

Anonymous:2016:EBg

[589] Anonymous. Editorial Board. Journal of Computational Science, 17 (part 2):ifc, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303611.

Anonymous:2016:EBh

[590] Anonymous. Editorial Board. Journal of Computational Science, 17 (part 3):ifc, November 2016. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303763.

Dolly:2017:PES

[591] D. Raveena Judie Dolly, G. Josemin Bala, and J. Dinesh Peter. Performance enhanced spatial video compression using global affine frame reconstruction. Journal of Computational Science, 18: 1–11, January 2017. ISSN 1877-7503 (print), 1877-7511 (electronic).

Jacobs:2017:OFA

[592] Christian T. Jacobs, Satya P. Jammy, and Neil D. Sandham. OpenSBLI: A framework for the automated derivation and parallel execution of finite difference solvers on a range of computer architectures. Journal of Computational Science, 18:12–23, January 2017. ISSN 1877-7503 (print), 1877-7511 (electronic).

Magdevska:2017:CDS

[593] Lidija Magdevska, Žiga Pušnik, Miha Mraz, Nikolaj Zimic, and Miha Moškon. Computational design of synchronous sequential structures in biological systems. Journal of Computational Science, 18:24–31, 2017. ISSN 1877-7503 (print), 1877-7511 (electronic).

Westera:2017:HPL

[594] Wim Westera. How people learn while playing serious games: a computational modelling approach. Journal of Computational Science, 18:32–45, 2017. ISSN 1877-7503 (print), 1877-7511 (electronic).

Ghidouche:2017:EHD

[595] Kahina Ghidouche, Abderrahmane Sider, Raphaël Couturier, and Christophe Guyeux. Efficient high degree polynomial root finding using GPU. Journal of Computational Science, 18:46– 56, 2017. ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2017:PMa

[596] Anonymous. Pages 1–164 (March 2017). Journal of Computational Science, 19:1–164, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Lang:2017:EPA

[597] Jens Lang and Janett Prehl. An embarrassingly parallel algorithm for random walk simulations on random fractal structures. Journal of Computational Science, 19:1–10, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304057.

Garcia-Magarino:2017:AMA

[598] Iván García-Magariño and Inmaculada Plaza. ABS-MindHeart: an agent based simulator of the influence of mindfulness programs on heart rate variability. Journal of Computational Science, 19:11–20, March 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316305348.

Ibrahim:2017:CEB

[599] Mark Ibrahim, Christopher M. Danforth, and Peter Sheridan Dodds. Connecting every bit of knowledge: the structure of Wikipedia’s First Link Network. Journal of Computational Science, 19:21–30, March 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304471.

Kaboli:2017:RFO

[600] S. Hr. Aghay Kaboli, J. Selvaraj, and N. A. Rahim. Rain-fall optimization algorithm: a population based algorithm for solving constrained optimization problems. Journal of Computational Science, 19:31–42, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316305336.

Galvez:2017:AUO

[601] Ramiro H. Gálvez and Agustín Gravano. Assessing the usefulness of online message board mining in automatic stock prediction systems. Journal of Computational Science, 19:43–56, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317300091.

Eghbal:2017:AUH

[602] Araz Eghbal, Andrew G. Gerber, and Eric Aubanel. Acceleration of unsteady hydrodynamic simulations using the parareal algorithm. Journal of Computational Science, 19:57–76, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304987.

Mirsepahi:2017:CIM

[603] Ali Mirsepahi, Arash Mehdizadeh, Lei Chen, Brian O’Neill, and Morteza Mohammadzaheri. Comparison of inverse modelling and optimizationbased methods in the heat flux estimation problem of an irradiative dryer/furnace. Journal of Computational Science, 19:77–85, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300881.

Mordhorst:2017:PDR

[604] M. Mordhorst, T. Strecker, D. Wirtz, T. Heidlauf, and O. Röhrle. POD-DEIM reduction of computational EMG models. Journal of Computational Science, 19:86–96, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300960.

Hillerman:2017:ACA

[605] Tiago Hillerman, João Carlos F. Souza, Ana Carla B. Reis, and Rommel N. Carvalho. Applying clustering and AHP methods for evaluating suspect healthcare claims. Journal of Computational Science, 19:97–111, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302041.

Tyagi:2017:TAP

[606] Sumarga Kumar Sah Tyagi, Deepak Kumar Jain, Steven Lawrence Fernandes, and Pranab K. Muhuri. Thermalaware power-efficient deadline based task allocation in multi-core processor. Journal of Computational Science, 19:112–120, March 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031630388X.

Kumar:2017:RSB

[607] Shishir Kumar, Neha Jain, and Steven Lawrence Fernandes. Rough set based effective technique of image watermarking. Journal of Computational Science, 19:121–137, March 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303854.

Jauhar:2017:SES

[608] Sunil Kumar Jauhar, Millie Pant, and Atulya K. Nagar. Sustainable educational supply chain performance measurement through DEA and differential evolution: a case on Indian HEI. Journal of Computational Science, 19:138–152, March 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630206X.

Amin:2017:MDC

[609] Javeria Amin, Muhammad Sharif, Mussarat Yasmin, Hussam Ali, and Steven Lawrence Fernandes. A method for the detection and classification of diabetic retinopathy using structural predictors of bright lesions. Journal of Computational Science, 19:153–164, March 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317300546.

Anonymous:2017:EBa

[610] Anonymous. Editorial Board. Journal of Computational Science, 19:ifc, March 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S187775031730265X.

Anonymous:2017:PMb

[611] Anonymous. Pages 1–214 (May 2017). Journal of Computational Science, 20: 1–214, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ligeiro:2017:LMQ

[612] Rui Ligeiro. Linked Markovian quantum tunnels: an approximation technique for solving the bin packing problem. Journal of Computational Science, 20:1–7, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730248X.

Redlarski:2017:GOP

[613] Grzegorz Redlarski, Mariusz Dabkowski, and Aleksander Palkowski. Generating optimal paths in dynamic environments using River Formation Dynamics algorithm. Journal of Computational Science, 20:8–16, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302417.

Yuan:2017:OCG

[614] Yujing Yuan, Dong Liang, and Hongmei Zhu. Optimal control of groundwater pollution combined with source abatement costs and taxes. Journal of Computational Science, 20:17–29, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303095.

Li:2017:SVJ

[615] Bai Li, Youmin Zhang, Zhijiang Shao, and Ning Jia. Simultaneous versus joint computing: a case study of multi-vehicle parking motion planning. Journal of Computational Science, 20:30–40, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031630254X.

Acharya:2017:AST

[616] U. Rajendra Acharya, Yuki Hagiwara, Joel E. W. Koh, Jen Hong Tan, Sulatha V. Bhandary, A. Krishna Rao, and U. Raghavendra. Automated screening tool for dry and wet agerelated macular degeneration (ARMD) using pyramid of histogram of oriented gradients (PHOG) and nonlinear features. Journal of Computational Science, 20:41–51, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302533.

Sharma:2017:DSS

[617] Rajeev Sharma, Mohit Kumar, Ram Bilas Pachori, and U. Rajendra Acharya. Decision support system for focal EEG signals using tunable- Q wavelet transform. Journal of Computational Science, 20:52–60, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303435.

Mohammed:2017:ASA

[618] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, Raed Ibraheem Hamed, Mohamad Khir Abdullah, and Dheyaa Ahmed Ibrahim. Automatic segmentation and automatic seed point selection of nasopharyngeal carcinoma from microscopy images using region growing based approach. Journal of Computational Science, 20:61–69, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302892.

Tan:2017:SOD

[619] Jen Hong Tan, U. Rajendra Acharya, Sulatha V. Bhandary, Kuang Chua Chua, and Sobha Sivaprasad. Segmentation of optic disc, fovea and retinal vasculature using a single convolutional neural network. Journal of Computational Science, 20:70–79, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302028.

Kovalchuk:2017:DTC

[620] Sergey V. Kovalchuk, Tesfamariam M. Abuhay, Ilkay Altintas, Michael L. Norman, Michael H. Lees, Valeria V. Krzhizhanovskaya, Jack Dongarra, and Peter M. A. Sloot. Data through the Computational Lens. Journal of Computational Science, 20:81–84, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317305008.

Abdelfattah:2017:FCF

[621] Ahmad Abdelfattah, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Fast Cholesky factorization on GPUs for batch and native modes in MAGMA. Journal of Computational Science, 20:85–93, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316305154.

Wang:2017:CEL

[622] Yunsong Wang, Emeric Brun, Fausto Malvagi, and Christophe Calvin. Competing energy lookup algorithms in Monte Carlo neutron transport calculations and their optimization on CPU and Intel MIC architectures. Journal of Computational Science, 20:94–102, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317300777.

Owsiak:2017:RSK

[623] Michal Owsiak, Marcin Plociennik, Bartek Palak, Tomasz Zok, Cedric Reux, Luc Di Gallo, Denis Kalupin, Thomas Johnson, and Mireille Schneider. Running simultaneous Kepler sessions for the parallelization of parametric scans and optimization studies applied to complex workflows. Journal of Computational Science, 20:103–111, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304884.

Korczynski:2017:BLS

[624] Wojciech Korczynski, Aleksander Byrski, and Marek Kisiel-Dorohinicki. Buffered local search for efficient memetic agent-based continuous optimization. Journal of Computational Science, 20:112–117, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301345.

Bi:2017:ICB

[625] Ran Bi, Jiajian Xiao, Vaisagh Viswanathan, and Alois Knoll. Influence of charging behaviour given charging infrastructure specification: a case study of Singapore. Journal of Computational Science, 20: 118–128, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031730306X.

Visheratin:2017:MLM

[626] Alexander A. Visheratin, Tamara B. Trofimenko, Ksenia D. Mukhina, Denis Nasonov, and Alexander V. Boukhanovsky. A multi-layer model for diffusion of urgent information in mobile networks. Journal of Computational Science, 20:129–142, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301333.

Fisher:2017:ADE

[627] Wendy D. Fisher, Tracy K. Camp, and Valeria V. Krzhizhanovskaya. Anomaly detection in earth dam and levee passive seismic data using support vector machines and automatic feature selection. Journal of Computational Science, 20:143–153, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304185.

Artetxe:2017:UAH

[628] Arkaitz Artetxe, Borja Ayerdi, Manuel Graña, and Sebastian Rios. Using Anticipative Hybrid Extreme Rotation Forest to predict emergency service readmission risk. Journal of Computational Science, 20:154–161, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316305130.

Melis:2017:VRM

[629] Ward Melis and Giovanni Samaey. Variance-reduced multiscale simulation of slow–fast stochastic differential equations. Journal of Computational Science, 20:162–176, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630432X.

Yeung:2017:SPP

[630] Man-Chung Yeung, Craig C. Douglas, and Long Lee. A spectral projection preconditioner for solving ill conditioned linear systems. Journal of Computational Science, 20:177–186, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317300571.

Fu:2017:BMP

[631] Yao Fu, John G. Michopoulos, and Jeong-Hoon Song. Bridging the multi phase-field and molecular dynamics models for the solidification of nanocrystals. Journal of Computational Science, 20:187–197, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316302526.

Balajewicz:2017:ROM

[632] Maciej Balajewicz and Jari Toivanen. Reduced order models for pricing European and American options under stochastic volatility and jump-diffusion models. Journal of Computational Science, 20:198–204, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730056X.

Purawat:2017:BBD

[633] Shweta Purawat, Charles Cowart, Rommie E. Amaro, and Ilkay Altintas. Biomedical Big Data Training Collaborative (BBDTC): an effort to bridge the talent gap in biomedical science and research. Journal of Computational Science, 20:205–214, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302909.

Anonymous:2017:EBb

[634] Anonymous. Editorial Board. Journal of Computational Science, 20:ifc, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750317305021.

Anonymous:2017:PJ

[635] Anonymous. Pages 1–494 (July 2017). Journal of Computational Science, 21: 1–494, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Leitch:2017:TMC

[636] Darren Leitch and Mohamed Sherif. Twitter mood, CEO succession announcements and stock returns. Journal of Computational Science, 21:1–10, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316302113.

Liu:2017:ABM

[637] Zhengchun Liu, Dolores Rexachs, Francisco Epelde, and Emilio Luque. An agent-based model for quantitatively analyzing and predicting the complex behavior of emergency departments. Journal of Computational Science, 21:11–23, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031730011X.

Pechenick:2017:LEG

[638] Eitan Adam Pechenick, Christopher M. Danforth, and Peter Sheridan Dodds. Is language evolution grinding to a halt? The scaling of lexical turbulence in English fiction suggests it is not. Journal of Computational Science, 21:24–37, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317304799.

Gao:2017:DAC

[639] X. Gao, Y. Wang, N. Overton, M. Zupanski, and X. Tu. Data-assimilated computational fluid dynamics modeling of convection-diffusion-reaction problems. Journal of Computational Science, 21:38–59, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316303222.

Garcia-Magarino:2017:ABS

[640] Iván García-Magariño and Raquel Lacuesta. Agent-based simulation of real-estate transactions. Journal of Computational Science, 21:60–76, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750317301011.

Chen:2017:SAV

[641] Qi Chen, Jianqiang Chen, Yufei Xie, and Xianxu Yuan. Study and application of virtual flight simulation for rolling control of vehicles. Journal of Computational Science, 21:77–85, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303210.

Barhoumi:2017:ITD

[642] Bassem Barhoumi and Jamel Bessrour. An improved time-dependent Boundary Element Method for twodimensional acoustic problems in a subsonic uniform flow. Journal of Computational Science, 21:86–112, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301515.

Wainer:2017:MMA

[643] Gabriel Wainer and Sixuan Wang. MAMS: Mashup architecture with modeling and simulation as a service. Journal of Computational Science, 21:113–131, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317305975.

Xiao:2017:MVS

[644] Xi Xiao, Peng Fu, Qing Li, Guangwu Hu, and Yong Jiang. Modeling and validation of SMS worm propagation over social networks. Journal of Computational Science, 21:132–139, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317305501.

Gao:2017:UDE

[645] Shangce Gao, Yirui Wang, Jiahai Wang, and JiuJun Cheng. Understanding differential evolution: a Poisson law derived from population interaction network. Journal of Computational Science, 21:140–149, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317300467.

Srivastava:2017:YHM

[646] Shilpa Srivastava, Millie Pant, and Atulya Nagar. Yuva: an e-health model for dealing with psychological issues of adolescents. Journal of Computational Science, 21:150–163, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317306609.

Do:2017:RMC

[647] T. V. Do, T. Q. Bui, T. T. Yu, D. T. Pham, and C. T. Nguyen. Role of material combination and new results of mechanical behavior for FG sandwich plates in thermal environment. Journal of Computational Science, 21:164–181, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303964.

J:2017:BIO

[648] Anita Christaline J., Ramesh R., Gomathy C., and Vaishali D. Bio inspired optimization for universal spatial image steganalysis. Journal of Computational Science, 21:182–188, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302818.

Yu:2017:TIM

[649] Xiang Yu and Xuening Fei. Target image matching algorithm based on pyramid model and higher moments. Journal of Computational Science, 21:189–194, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730724X.

Du:2017:ATC

[650] Jian hai Du. Automatic text classification algorithm based on Gauss improved convolutional neural network. Journal of Computational Science, 21:195–200, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317307238.

P:2017:ASB

[651] Malathy P. and Shunmugalatha A. Application of swarm based intelligent computing algorithms for dynamic evaluation of maximum loadability of transmission network. Journal of Computational Science, 21: 201–222, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317304635.

Singh:2017:SPC

[652] Raman Preet Singh, Savita Gupta, and U. Rajendra Acharya. Segmentation of prostate contours for automated diagnosis using ultrasound images: a survey. Journal of Computational Science, 21:223–231, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304611.

Mohammed:2017:SVRa

[653] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, Raed Ibraheem Hamed, Salama A. Mostafa, Dheyaa Ahmed Ibrahim, Humam Khaled Jameel, and Ahmed Hamed Alallah. Solving vehicle routing problem by using improved k-nearest neighbor algorithm for best solution. Journal of Computational Science, 21:232–240, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730426X.

Mohammed:2017:AEM

[654] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, Raed Ibraheem Hamed, and Dheyaa Ahmed Ibrahim. Analysis of an electronic methods for nasopharyngeal carcinoma: Prevalence, diagnosis, challenges and technologies. Journal of Computational Science, 21:241–254, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304003.

Mohammed:2017:SVRb

[655] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, Raed Ibraheem Hamed, Salama A. Mostafa, Mohd Sharifuddin Ahmad, and Dheyaa Ahmed Ibrahim. Solving vehicle routing problem by using improved genetic algorithm for optimal solution. Journal of Computational Science, 21:255–262, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303848.

Mohammed:2017:ANN

[656] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, Raed Ibraheem Hamed, Dheyaa Ahmed Ibrahim, and Mohamad Khir Abdullah. Artificial neural networks for automatic segmentation and identification of nasopharyngeal carcinoma. Journal of Computational Science, 21:263–274, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317303617.

Nithya:2017:GAB

[657] S. Nithya and K. Meena alias Jeyanthi. Genetic algorithm based bacterial foraging optimization with threepass protocol concept for heterogeneous network security enhancement. Journal of Computational Science, 21:275–282, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317303447.

Mohammed:2017:RNC

[658] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, Raed Ibraheem Hamed, and Dheyaa Ahmed Ibrahim. Review on Nasopharyngeal Carcinoma: Concepts, methods of analysis, segmentation, classification, prediction and impact: a review of the research literature. Journal of Computational Science, 21:283–298, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317303423.

Jauhar:2017:IDM

[659] Sunil Kumar Jauhar and Millie Pant. Integrating DEA with DE and MODE for sustainable supplier selection. Journal of Computational Science, 21:299–306, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317302090.

Sharma:2017:OBL

[660] Tarun Kumar Sharma and Millie Pant. Opposition based learning ingrained shuffled frog-leaping algorithm. Journal of Computational Science, 21:307–315, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730203X.

Jain:2017:EEM

[661] Vinay Kumar Jain, Shishir Kumar, and Steven Lawrence Fernandes. Extraction of emotions from multilingual text using intelligent text processing and computational linguistics. Journal of Computational Science, 21:316–326, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301035.

Basha:2017:DIL

[662] S. Mahaboob Basha and M. Kannan. Design and implementation of low-power motion estimation based on modified full-search block motion estimation. Journal of Computational Science, 21:327–332, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304215.

Boukamcha:2017:RTS

[663] Hamdi Boukamcha, Abdelkader Ben Amara, Fethi Smach, and Mohamed Atri. Robust technique for 3D shape reconstruction. Journal of Computational Science, 21:333–339, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304203.

Boukamcha:2017:ALD

[664] Hamdi Boukamcha, Mohamed Hallek, Fethi Smach, and Mohamed Atri. Automatic landmark detection and 3D Face data extraction. Journal of Computational Science, 21: 340–348, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316304197.

Abd:2017:EAM

[665] Sura Khalil Abd, S. A. R. Al-Haddad, Fazirulhisyam Hashim, Azizol B. H. J. Abdullah, and Salman Yussof. An effective approach for managing power consumption in cloud computing infrastructure. Journal of Computational Science, 21:349–360, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316303830.

Bao:2017:OFS

[666] Le Nguyen Bao, Dac-Nhuong Le, Gia Nhu Nguyen, Vikrant Bhateja, and Suresh Chandra Satapathy. Optimizing feature selection in videobased recognition using Max–Min Ant System for the online video contextual advertisement user-oriented system. Journal of Computational Science, 21:361–370, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031630271X.

Dogra:2017:BVI

[667] Ayush Dogra, Bhawna Goyal, and Sunil Agrawal. Bone vessel image fusion via generalized reisz wavelet transform using averaging fusion rule. Journal of Computational Science, 21:371–378, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316302101.

B:2017:GML

[668] Sudeepa K. B. and Ganesh Aithal. Generation of maximum length nonbinary key sequence and its application for stream cipher based on residue number system. Journal of Computational Science, 21:379–386, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316302071.

Bhateja:2017:HVS

[669] Vikrant Bhateja, Mukul Misra, and Shabana Urooj. Human visual system based unsharp masking for enhancement of mammographic images. Journal of Computational Science, 21:387–393, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301235.

Topa:2017:NTC

[670] Paweł Topa and Jarosław Wa̧s. New trends in Complex Collective Systems. Journal of Computational Science, 21:395–396, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306002.

Byrski:2017:SCI

[671] Aleksander Byrski, Ewelina Świderska, Jakub Łasisz, Marek Kisiel-Dorohinicki, Tom Lenaerts, Dana Samson, Bipin Indurkhya, and Ann Nowé. Sociocognitively inspired ant colony optimization. Journal of Computational Science, 21:397–406, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316302198.

Komosinski:2017:ASM

[672] Maciej Komosinski. Applications of a similarity measure in the analysis of populations of 3D agents. Journal of Computational Science, 21:407–418, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316302046.

Komosinski:2017:MAS

[673] Maciej Komosinski, Agnieszka Mensfelt, Jarosław Tyszka, and Jan Goleń. Multi-agent simulation of benthic foraminifera response to annual variability of feeding fluxes. Journal of Computational Science, 21: 419–431, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316301715.

Kechaidou:2017:GLV

[674] M. G. Kechaidou and G. Ch. Sirakoulis. Game of Life variations for image scrambling. Journal of Computational Science, 21:432–447, July 2017.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301442.

Dzwinel:2017:IFF

[675] Witold Dzwinel, Rafał Wcisło, and Wojciech Czech. ivga: a fast force-directed method for interactive visualization of complex networks. Journal of Computational Science, 21:448–459, July 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301430.

Gasior:2017:SCA

[676] Jakub Ga̧sior and Franciszek Seredyński. A Sandpile cellular automata-based scheduler and load balancer. Journal of Computational Science, 21: 460–468, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301284.

Ntinas:2017:PFC

[677] Vasileios G. Ntinas, Byron E. Moutafis, Giuseppe A. Trunfio, and Georgios Ch. Sirakoulis. Parallel fuzzy cellular automata for data-driven simulation of wildfire spreading. Journal of Computational Science, 21: 469–485, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316301260.

Hrabak:2017:IAH

[678] Pavel Hrabák and Marek Bukáček. Influence of agents heterogeneity in cellular model of evacuation. Journal of Computational Science, 21: 486–493, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316301259.

Anonymous:2017:IFCa

[679] Anonymous. Inside front cover — Editorial Board page. Journal of Computational Science, 21:ifc, July 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317308256.

Anonymous:2017:PNP

[680] Anonymous. Publisher note: “Parallel multi-projection preconditioned methods based on semi-aggregation techniques” [Moutafis:2017:PMP]. Journal of Computational Science, 22: iii, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317312012.

Anonymous:2017:PS

[681] Anonymous. Pages 1–314 (September 2017). Journal of Computational Science, 22:1–314, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Chalupa:2017:MRS

[682] David Chalupa and Christian Blum. Mining k-reachable sets in real-world networks using domination in shortcut graphs. Journal of Computational Science, 22:1–14, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316302642.

Alowayyed:2017:MCE

[683] Saad Alowayyed, Derek Groen, Peter V. Coveney, and Alfons G. Hoekstra. Multiscale computing in the exascale era. Journal of Computational Science, 22:15–25, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316302988.

Rabanal:2017:ARF

[684] Pablo Rabanal, Ismael Rodr´ ıguez, and Fernando Rubio. Applications of river formation dynamics. Journal of Computational Science, 22: 26–35, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317307184.

Sahneh:2017:GSS

[685] Faryad Darabi Sahneh, Aram Vajdi, Heman Shakeri, Futing Fan, and Caterina Scoglio. GEMFsim: a stochastic simulator for the generalized epidemic modeling framework. Journal of Computational Science, 22: 36–44, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317305227.

Moutafis:2017:PMP

[686] Byron E. Moutafis, Christos K. Filelis-Papadopoulos, and George A. Gravvanis. Parallel multi-projection preconditioned methods based on semi-aggregation techniques. Journal of Computational Science, 22: 45–54, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317304647.

De:2017:DNC

[687] Debashis De and Jadav Chandra Das. Design of novel carry save adder using quantum dot-cellular automata. Journal of Computational Science, 22:54–68, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317301722.

Oyebamiji:2017:GPE

[688] O. K. Oyebamiji, D. J. Wilkinson, P. G. Jayathilake, T. P. Curtis, S. P. Rushton, B. Li, and P. Gupta. Gaussian process emulation of an individual-based model simulation of microbial communities. Journal of Computational Science, 22:69–84, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317300443.

Celik:2017:DSS

[689] Mete Celik and Ahmet Sakir Dokuz. Discovering socio-spatio-temporal important locations of social media users. Journal of Computational Science, 22:85–98, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308165.

Verbosio:2017:ESS

[690] Fabio Verbosio, Arne De Coninck, Drosos Kourounis, and Olaf Schenk. Enhancing the scalability of selected inversion factorization algorithms in genomic prediction. Journal of Computational Science, 22: 99–108, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317301473.

Prakash:2017:NSN

[691] Bijil Prakash, Amit Setia, and Deepak Alapatt. Numerical solution of nonlinear fractional SEIR epidemic model by using Haar wavelets. Journal of Computational Science, 22: 109–118, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750316305300.

Lin:2017:MFM

[692] Yi-Kuei Lin and Shin-Guang Chen. A maximal flow method to search for d-MPs in stochastic-flow networks. Journal of Computational Science, 22:119–125, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304581.

Iacono:2017:MEH

[693] Mauro Iacono, Marco Gribaudo, Joanna Kołodziej, and Florin Pop. Modeling and evaluation of highly complex computer systems architectures. Journal of Computational Science, 22:126–130, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317310992.

Tychalas:2017:HPS

[694] Dimitrios Tychalas and Helen Karatza. High performance system based on Cloud and beyond: Jungle Computing. Journal of Computational Science, 22:131–147, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317303113.

Tsai:2017:CFN

[695] Pang-Wei Tsai, Francesco Piccialli, Chun-Wei Tsai, Mon-Yen Luo, and Chu-Sing Yang. Control frameworks in network emulation testbeds: a survey. Journal of Computational Science, 22:148–161, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317302429.

Fiore:2017:TME

[696] Ugo Fiore, Paolo Zanetti, Francesco Palmieri, and Francesca Perla. Traffic matrix estimation with softwaredefined NFV: Challenges and opportunities. Journal of Computational Science, 22:162–170, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302405.

Kim:2017:ADR

[697] Dae-Young Kim, Seokhoon Kim, Houcine Hassan, and Jong Hyuk Park. Adaptive data rate control in low power wide area networks for long range IoT services. Journal of Computational Science, 22:171–178, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304428.

Ficco:2017:SPC

[698] Massimo Ficco, Michał Choraś, and Rafał Kozik. Simulation platform for cyber-security and vulnerability analysis of critical infrastructures. Journal of Computational Science, 22: 179–186, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317303605.

Szynkiewicz:2017:DES

[699] Paweł Szynkiewicz and Adam Kozakiewicz. Design and evaluation of a system for network threat signatures generation. Journal of Computational Science, 22:187–197, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305197.

Suchacka:2017:MNS

[700] Grażyna Suchacka and Daria Wotzka. Modeling a non-stationary bots’ arrival process at an e-commerce Web site. Journal of Computational Science, 22:198–208, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305720.

Czerniak:2017:AAO

[701] Jacek M. Czerniak and Hubert Zarzycki. Artificial Acari Optimization as a new strategy for global optimization of multimodal functions. Journal of Computational Science, 22:209–227, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306610.

Cuomo:2017:PPB

[702] Salvatore Cuomo, Pasquale De Michele, Ardelio Galletti, and Livia Marcellino. A parallel PDE-based numerical algorithm for computing the Optical Flow in hybrid systems. Journal of Computational Science, 22: 228–236, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317303010.

Bensaber:2017:PEC

[703] Boucif Amar Bensaber and Luca Foschini. Performance evaluation of communications in distributed systems and web based service architectures. Journal of Computational Science, 22:237–239, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317311006.

Damasevicius:2017:IMD

[704] Robertas Damaševičius, Christian Napoli, Tatjana Sidekerskienė, and Marcin Woźniak. IMF mode demixing in EMD for jitter analysis. Journal of Computational Science, 22: 240–252, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317304180.

Serrai:2017:TEM

[705] Walid Serrai, Abdelkrim Abdelli, Lynda Mokdad, and Youcef Hammal. Towards an efficient and a more accurate web service selection using MCDM methods. Journal of Computational Science, 22:253–267, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306154.

Ferreira:2017:ALB

[706] Carlos Ferreira, Susana Sargento, and Arnaldo Oliveira. An architecture for a learning-based autonomic decision system. Journal of Computational Science, 22:268–282, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304209.

Fouchal:2017:VTC

[707] Hacène Fouchal, Emilien Bourdy, Geoffrey Wilhelm, and Marwane Ayaida. A validation tool for cooperative intelligent transport systems. Journal of Computational Science, 22: 283–288, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317306300.

Kim:2017:CFR

[708] Hyunbum Kim, Jalel Ben-Othman, and Paolo Bellavista. Collision-free reinforced barriers in UAV networks. Journal of Computational Science, 22:289–300, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317306282.

Tanaka:2017:AMD

[709] Yasuyuki Tanaka, Mitsuaki Akiyama, and Atsuhiro Goto. Analysis of malware download sites by focusing on time series variation of malware. Journal of Computational Science, 22:301–313, September 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317306294.

Anonymous:2017:IFCb

[710] Anonymous. Inside front cover — Editorial Board page. Journal of Computational Science, 22:ifc, September 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031731102X.

Anonymous:2017:PNb

[711] Anonymous. Pages 1–248 (November 2017). Journal of Computational Science, 23:1–248, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

daSilva:2017:DDO

[712] Josildo Pereira da Silva, Gilson A. Giraldi, and Antonio L. Apolinário, Jr. Data-driven optimization approach for mass-spring models parametrization based on isogeometric analysis. Journal of Computational Science, 23:1–19, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310311.

Wojnicki:2017:ADG

[713] Igor Wojnicki, Leszek Kotulski, Adam Sȩdziwy, and Sebastian Ernst. Application of distributed graph transformations to automated generation of control patterns for intelligent lighting systems. Journal of Computational Science, 23:20–30, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303228.

Bao:2017:NND

[714] Chunteng Bao, Lihong Xu, Erik D. Goodman, and Leilei Cao. A novel non-dominated sorting algorithm for evolutionary multi-objective optimization. Journal of Computational Science, 23:31–43, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310530.

Horng:2017:MCS

[715] Shih-Cheng Horng and Shieh-Shing Lin. Merging crow search into ordinal optimization for solving equality constrained simulation optimization problems. Journal of Computational Science, 23:44–57, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304441.

Boulila:2017:SAA

[716] Wadii Boulila, Zouhayra Ayadi, and Imed Riadh Farah. Sensitivity analysis approach to model epistemic and aleatory imperfection: Application to Land Cover Change prediction model. Journal of Computational Science, 23:58–70, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303368.

Dinkar:2017:OBL

[717] Shail Kumar Dinkar and Kusum Deep. Opposition based Laplacian Ant Lion Optimizer. Journal of Computational Science, 23:71–90, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317307093.

Kaur:2017:RBC

[718] Mankirat Kaur and Sarbjeet Singh. Ranking based comparative analysis of graph centrality measures to detect negative nodes in online social networks. Journal of Computational Science, 23:91–108, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303945.

Brown:2017:RDS

[719] Joshua M. Brown, Terry Bossomaier, and Lionel Barnett. Review of data structures for computationally efficient nearest-neighbour entropy estimators for large systems with periodic boundary conditions. Journal of Computational Science, 23: 109–117, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317300807.

Veiga:2017:RAC

[720] Luís Veiga, Didier El Baz, and João M. P. Cardoso. Recent advances in computational science and engineering research. Journal of Computational Science, 23:118–119, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310359.

Arabnejad:2017:MCR

[721] Hamid Arabnejad and Jorge G. Barbosa. Maximizing the completion rate of concurrent scientific applications under time and budget constraints. Journal of Computational Science, 23:120–129, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316302447.

Banati:2017:MMB

[722] Hema Banati and Reshu Chaudhary. Multi-Modal Bat Algorithm with Improved Search (MMBAIS). Journal of Computational Science, 23:

130–144, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031630463X.

Branitskiy:2017:HCI

[723] A. Branitskiy and I. Kotenko. Hybridization of computational intelligence methods for attack detection in computer networks. Journal of Computational Science, 23: 145–156, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316301181.

Azad:2017:EIS

[724] Muhammad Ajmal Azad and Ricardo Morla. Early identification of spammers through identity linking, social network and call features. Journal of Computational Science, 23: 157–172, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750316302824.

Ashtari:2017:SPC

[725] Sadaf Ashtari and Ali Eydgahi. Student perceptions of cloud applications effectiveness in higher education. Journal of Computational Science, 23:173–180, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316304975.

Saadaoui:2017:FFB

[726] Amina Saâdaoui, Nihel Ben Youssef Ben Souayeh, and Adel Bouhoula. FARE: FDD-based firewall anomalies resolution tool. Journal of Computational Science, 23:181–191, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750317309894.

Cotta:2017:BAC

[727] Carlos Cotta and Robert Schaefer. Bioinspired algorithms and complex systems. Journal of Computational Science, 23:192–194, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312681.

Gupta:2017:UFS

[728] Deepak Gupta and Anil K. Ahlawat. Usability feature selection via MB-BAT: a novel approach. Journal of Computational Science, 23: 195–203, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317307081.

Karcz-Duleba:2017:MDS

[729] Iwona Karcz-Duleba and Andrzej Cichon. Modeling dynamics of small populations in a simple phenotypic evolutionary algorithm. A space of population states approach. Journal of Computational Science, 23: 204–215, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317307895.

Poteralski:2017:HAI

[730] Arkadiusz Poteralski. Hybrid artificial immune strategy in identification and optimization of mechanical systems. Journal of Computational Science, 23:216–225, November 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031730460X.

Qi:2017:NMH

[731] Xiangbo Qi, Yunlong Zhu, and Hao Zhang. A new meta-heuristic butterfly-inspired algorithm. Journal of Computational Science, 23: 226–239, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317306853.

Turek:2017:RTM

[732] Wojciech Turek, Leszek Siwik, Marek Kisiel-Dorohinicki, Sebastian Łakomy, Piotr Kala, and Aleksander Byrski. Real-time metaheuristic-based urban crossroad management with multi-variant planning. Journal of Computational Science, 23:240– 248, November 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317304623.

Anonymous:2017:IFCc

[733] Anonymous. Inside front cover — Editorial Board page. Journal of Computational Science, 23:ifc, November 2017.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312802.

Anonymous:2018:EBa

[734] Anonymous. Editorial Board. Journal of Computational Science, 24:ii, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S187775031830053X.

Anonymous:2018:PJa

[735] Anonymous. Pages 1–426 (January 2018). Journal of Computational Science, 24:1–426, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Alowayyed:2018:LBP

[736] S. Alowayyed, G. Závodszky, V. Azizi, and A. G. Hoekstra. Load balancing of parallel cell-based blood flow simulations. Journal of Computational Science, 24:1–7, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302788.

Chen:2018:NSF

[737] T. B. Y. Chen, A. C. Y. Yuen, G. H. Yeoh, V. Timchenko, S. C. P. Cheung, Q. N. Chan, W. Yang, and H. Lu. Numerical study of fire spread using the level-set method with large eddy simulation incorporating detailed chemical kinetics gas-phase combustion model. Journal of Computational Science, 24:8–23, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304088.

Joo:2018:PCT

[738] HyunShig Joo, Haeseong Cho, Seil Kim, SangJoon Shin, and JunYoung Kwak. Parallel computation for threedimensional shell analysis of curved configuration based on domain decomposition method. Journal of Computational Science, 24:24–33, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317307445.

Huynh:2018:CCS

[739] Hoai Nguyen Huynh, Evgeny Makarov, Erika Fille Legara, Christopher Monterola, and Lock Yue Chew. Characterisation and comparison of spatial patterns in urban systems: a case study of U.S. cities. Journal of Computational Science, 24:34–43, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730769X.

Imran:2018:PET

[740] Muhammad Imran, Martin Collier, Pascal Landais, and Kostas Katrinis. Performance evaluation of TCP over software-defined optical burstswitched data centre network. Journal of Computational Science, 24: 44–53, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031731356X.

Arya:2018:FBB

[741] Rubi Arya, Pitam Singh, and Deepak Bhati. A fuzzy based branch and bound approach for multi-objective linear fractional (MOLF) optimization problems. Journal of Computational Science, 24:54–64, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301862.

Hoekstra:2018:VPH

[742] Alfons G. Hoekstra, Frans van de Vosse, and Oliver Röhrle. The Virtual Physiological Human Conference 2016. Journal of Computational Science, 24:65–67, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313467.

Huberts:2018:WNM

[743] Wouter Huberts, Stefan G. H. Heinen, Niek Zonnebeld, Daniel A. F. van den Heuvel, Jean-Paul P. M. de Vries, Jan H. M. Tordoir, D. Rodney Hose, Tammo Delhaas, and Frans N. van de Vosse. What is needed to make cardiovascular models suitable for clinical decision support? a viewpoint paper. Journal of Computational Science, 24:68–84, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317307901.

Finsberg:2018:ECC

[744] Henrik Finsberg, Gabriel Balaban, Stian Ross, Trine F. Håland, Hans Henrik Odland, Joakim Sundnes, and Samuel Wall. Estimating cardiac contraction through high resolution data assimilation of a personalized mechanical model. Journal of Computational Science, 24:85–90, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308190.

Bruening:2018:IPS

[745] Jan Bruening, Florian Hellmeier, Pavlo Yevtushenko, Marcus Kelm, Sarah Nordmeyer, Simon H. Sündermann, Titus Kuehne, and Leonid Goubergrits. Impact of patient-specific LVOT inflow profiles on aortic valve prosthesis and ascending aorta hemodynamics. Journal of Computational Science, 24:91–100, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312310.

Fink:2018:ADM

[746] D. Fink, A. Wagner, and W. Ehlers. Application-driven model reduction for the simulation of therapeutic infusion processes in multi-component brain tissue. Journal of Computational Science, 24:101–115, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310840.

Kenny:2018:MPS

[747] Allanah Kenny, Constantine Zakkaroff, Michael J. Plank, and Tim David. Massively parallel simulations of neurovascular coupling with extracellular diffusion. Journal of Computational Science, 24:116–124, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317307652.

Ingham-Dempster:2018:CBM

[748] Tim Ingham-Dempster, Bernard Corfe, and Dawn Walker. A cellular based model of the colon crypt suggests novel effects for Apc phenotype in colorectal carcinogenesis. Journal of Computational Science, 24: 125–131, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317307342.

Nolan:2018:IDM

[749] David R. Nolan and Caitríona Lally. An investigation of damage mechanisms in mechanobiological models of in-stent restenosis. Journal of Computational Science, 24: 132–142, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317304192.

Keijsers:2018:MRV

[750] J. M. T. Keijsers, C. A. D. Leguy, A. J. Narracott, J. Rittweger, F. N. van de Vosse, and W. Huberts. Modeling regulation of vascular tone following muscle contraction: Model development, validation and global sensitivity analysis. Journal of Computational Science, 24:143–159, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317304106.

Tobar:2018:VRI

[751] Andrés Mena Tobar, José M. Ferrero, Francesco Migliavacca, and José F. Rodríguez Matas. Vulnerability in regionally ischemic human heart. effect of the extracellular potassium concentration. Journal of Computational Science, 24:160–168, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317312668.

Nowakowski:2018:CCI

[752] Piotr Nowakowski, Marian Bubak, Tomasz Bartyński, Tomasz Gubała, Daniel Harȩżlak, Marek Kasztelnik, Maciej Malawski, and Jan Meizner. Cloud computing infrastructure for the VPH community. Journal of Computational Science, 24: 169–179, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317307330.

intHout:2018:SIC

[753] Karel in ’t Hout, Andrey Itkin, Lina von Sydow, and Jari Toivanen. Special issue — computational and algorithmic finance. Journal of Computational Science, 24:180–181, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317314138.

Wyns:2018:AME

[754] Maarten Wyns and Karel J. in ’t Hout. An adjoint method for the exact calibration of stochastic local volatility models. Journal of Computational Science, 24:182–194, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301709.

Itkin:2018:FGS

[755] Andrey Itkin and Alexander Lipton. Filling the gaps smoothly. Journal of Computational Science, 24: 195–208, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317301539.

Rad:2018:FDP

[756] Jamal Amani Rad, Josef Höök, Elisabeth Larsson, and Lina von Sydow. Forward deterministic pricing of options using Gaussian radial basis functions. Journal of Computational Science, 24:209–217, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305537.

Kaushansky:2018:NAE

[757] Vadim Kaushansky, Alexander Lipton, and Christoph Reisinger. Numerical analysis of an extended structural default model with mutual liabilities and jump risk. Journal of Computational Science, 24:218–231, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317305513.

Fernandez:2018:GPI

[758] José L. Fernández, Ana M. Ferreiro-Ferreiro, José A. Garc´ ıa-Rodr´ ıguez, and Carlos Vázquez. GPU parallel implementation for asset-liability management in insurance companies. Journal of Computational Science, 24: 232–254, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317305495.

Lapshin:2018:IBM

[759] Victor Lapshin. Inconsistencies in bond market quotes: Is it the wrong model or the wrong data? Journal of Computational Science, 24: 255–265, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317305525.

Chatrabgoun:2018:ANG

[760] Omid Chatrabgoun, Amin Hosseinian-Far, Victor Chang, Nigel G. Stocks, and Alireza Daneshkhah. Approximating non-Gaussian Bayesian networks using minimum information vine model with applications in financial modelling. Journal of Computational Science, 24:266–276, January 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317309882.

Dixon:2018:SCL

[761] Matthew Dixon. Sequence classification of the limit order book using recurrent neural networks. Journal of Computational Science, 24: 277–286, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317309675.

Rezvanian:2018:NAL

[762] Alireza Rezvanian, S. Mehdi Vahidipour, and Mehdi Esnaashari. New applications of learning automata-based techniques in real-world environments. Journal of Computational Science, 24: 287–289, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317312000.

Jobava:2018:AIT

[763] Akaki Jobava, Anis Yazidi, B. John Oommen, and Kyrre Begnum. On achieving intelligent traffic-aware consolidation of virtual machines in a data center using Learning Automata. Journal of Computational Science, 24: 290–312, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317300054.

Moradabadi:2018:LPS

[764] Behnaz Moradabadi and Mohammad Reza Meybodi. Link prediction in stochastic social networks: Learning automata approach. Journal of Computational Science, 24:

313–328, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317300534.

Shirvani:2018:EOM

[765] Abdolreza Shirvani and B. John Oommen. On enhancing the object migration automaton using the Pursuit paradigm. Journal of Computational Science, 24:329–342, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317302259.

Vahidipour:2018:PAQ

[766] S. Mehdi Vahidipour and Mehdi Esnaashari. Priority assignment in queuing systems with unknown characteristics using learning automata and adaptive stochastic Petri nets. Journal of Computational Science, 24: 343–357, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317302326.

Esnaashari:2018:DIC

[767] Mehdi Esnaashari and Mohammad Reza Meybodi. Dynamic irregular cellular learning automata. Journal of Computational Science, 24: 358–370, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317302983.

Saghiri:2018:ECA

[768] Ali Mohammad Saghiri and Mohammad Reza Meybodi. On expediency of Closed Asynchronous Dynamic Cellular Learning Automata. Journal of Computational Science, 24: 371–378, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317302235.

Hasanzadeh-Mofrad:2018:LAC

[769] Mohammad Hasanzadeh-Mofrad and Alireza Rezvanian. Learning Automata Clustering. Journal of Computational Science, 24:379–388, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302247.

Mirsaleh:2018:MMA

[770] Mehdi Rezapoor Mirsaleh and Mohammad Reza Meybodi. A Michigan memetic algorithm for solving the vertex coloring problem. Journal of Computational Science, 24: 389–401, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317301680.

Moghadam:2018:MRD

[771] Mahshid Helali Moghadam and Seyed Morteza Babamir. Makespan reduction for dynamic workloads in clusterbased data grids using reinforcementlearning based scheduling. Journal of Computational Science, 24: 402–412, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317302314.

Khomami:2018:NCL

[772] Mohammad Mehdi Daliri Khomami, Alireza Rezvanian, and Mohammad Reza Meybodi. A new cellular learning automata-based algorithm for community detection in complex social networks. Journal of Computational Science, 24:413–426, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317302119.

Anonymous:2018:EBb

[773] Anonymous. Editorial Board. Journal of Computational Science, 25:ii, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750318303466.

Anonymous:2018:PMa

[774] Anonymous. Pages 1–474 (March 2018). Journal of Computational Science, 25:1–474, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

He:2018:ABM

[775] Zhou He, Jichang Dong, and Lean Yu. An agent-based model for investigating the impact of distorted supply– demand information on China’s resale housing market. Journal of Computational Science, 25:1–15, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308141.

Lin:2018:NTM

[776] Ji Lin, Chein-Shan Liu, Wen Chen, and Linlin Sun. A novel Trefftz method for solving the multidimensional direct and Cauchy problems of Laplace equation in an arbitrary domain. Journal of Computational Science, 25:16–27, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305367.

Goel:2018:HAC

[777] Rajeev Goel and Raman Maini. A hybrid of ant colony and firefly algorithms (HAFA) for solving vehicle routing problems. Journal of Computational Science, 25:28–37, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730813X.

Cao:2018:GBG

[778] Zijian Cao, Lei Wang, and Xinhong Hei. A global-best guided phase based optimization algorithm for scalable optimization problems and its application. Journal of Computational Science, 25:38–49, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304039.

Jiao:2018:PPM

[779] Zhuqing Jiao, Kai Ma, Yiling Rong, Peng Wang, Hongkai Zhang, and Shuihua Wang. A path planning method using adaptive polymorphic ant colony algorithm for smart wheelchairs. Journal of Computational Science, 25:50–57, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308840.

Cheng:2018:LTM

[780] Ching-Hsue Cheng, You-Shyang Chen, Arun Kumar Sangaiah, Jieh-Ren Chang, and Ting-Yu Wang. Lessoning travelers’ motivating behavior for a growing phenomenon in Taiwan by an advanced hybrid object-oriented hierarchical model. Journal of Computational Science, 25:58–75, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317309833.

Denham:2018:UEP

[781] Mónica Denham and Karina Laneri. Using efficient parallelization in Graphic Processing Units to parameterize stochastic fire propagation models. Journal of Computational Science, 25:76–88, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308773.

Tan:2018:PFS

[782] Jifu Tan, Talid R. Sinno, and Scott L. Diamond. A parallel fluid–solid coupling model using lammps and Palabos based on the immersed boundary method. Journal of Computational Science, 25:89–100, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730981X.

Viceconti:2018:VHS

[783] Marco Viceconti, Simone Bnà, Daniele Tartarini, Stelios Sfakianakis, James Grogan, Dawn Walker, Samuel Gamble, and Debora Testi. VPH-HF: a software framework for the execution of complex subject-specific physiology modelling workflows. Journal of Computational Science, 25: 101–114, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317305732.

Jung:2018:EFB

[784] Pawel S. Jung, Katarzyna Rutkowska, and Miroslaw A. Karpierz. Evanescent field boundary conditions for modelling of light propagation. Journal of Computational Science, 25: 115–121, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317312577.

Sattari:2018:CID

[785] Mohammad Sattari and Kamran Zamanifar. A cascade information diffusion based label propagation algorithm for community detection in dynamic social networks. Journal of Computational Science, 25: 122–133, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317303009.

Gonzalez-Dominguez:2018:PFP

[786] Jorge González-Domínguez, Christian Hundt, and Bertil Schmidt. parSRA: a framework for the parallel execution of short read aligners on compute clusters. Journal of Computational Science, 25:134–139, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317300893.

Catalan:2018:MTD

[787] Sandra Catalán, José R. Herrero, Francisco D. Igual, Rafael Rodríguez-Sánchez, Enrique S. Quintana-Ortí, and Chris Adeniyi-Jones. Multithreaded dense linear algebra libraries for low-power asymmetric multicore processors. Journal of Computational Science, 25:140–151, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316302812.

Aljawarneh:2018:ABI

[788] Shadi Aljawarneh, Monther Aldwairi, and Muneer Bani Yassein. Anomaly-based intrusion detection system through feature selection analysis and building hybrid efficient model. Journal of Computational Science, 25:152–160, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316305099.

Zhai:2018:VDK

[789] Yanhui Zhai, Deyu Li, and Jing Zhang. Variable decision knowledge representation: a logical description. Journal of Computational Science, 25:161–169, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304616.

Min:2018:ACO

[790] Fan Min, Zhi-Heng Zhang, and Ji Dong. Ant colony optimization with partial-complete searching for attribute reduction. Journal of Computational Science, 25:170–182, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316304604.

Yan:2018:ECS

[791] Yongquan Yan, Ping Guo, Bin Cheng, and Zhigao Zheng. An experimental case study on the relationship between workload and resource consumption in a commercial web server. Journal of Computational Science, 25:183–192, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301382.

Sun:2018:DUA

[792] Xiao Sun, Chen Zhang, Guoqiang Li, Daniel Sun, Fuji Ren, Albert Zomaya, and Rajiv Ranjan. Detecting users’ anomalous emotion using social media for business intelligence. Journal of Computational Science, 25:193–200, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317302879.

Kumar:2018:ICT

[793] Shishir Kumar, Prabhat Mahanti, and Su-Jing Wang. Intelligent computational techniques. Journal of Computational Science, 25: 201–203, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318301467.

Kant:2018:NBC

[794] Surya Kant and Tripti Mahara. Nearest biclusters collaborative filtering framework with fusion. Journal of Computational Science, 25: 204–212, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775031730323X.

Gupta:2018:IGB

[795] Nidhi Gupta, Pushpraj Bhatele, and Pritee Khanna. Identification of gliomas from brain MRI through adaptive segmentation and run length of centralized patterns. Journal of Computational Science, 25:213–220, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317302077.

Smiti:2018:SSC

[796] Abir Smiti and Zied Elouedi. SCBM: soft case base maintenance method based on competence model. Journal of Computational Science, 25: 221–227, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317310414.

Nguyen:2018:AVS

[797] Thin Nguyen, Svetha Venkatesh, and Dinh Phung. Academia versus social media: a psycho-linguistic analysis. Journal of Computational Science, 25:228–237, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317309122.

Verma:2018:SCC

[798] Om Prakash Verma, Gaurav Manik, and Vinay Kumar Jain. Simulation and control of a complex nonlinear dynamic behavior of multi-stage evaporator using PID and Fuzzy-PID controllers. Journal of Computational Science, 25:238–251, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730368X.

Jain:2018:AHI

[799] Deepak Kumar Jain, Surendra Bilouhan Dubey, Rishin Kumar Choubey, Amit Sinhal, Siddharth Kumar Arjaria, Amar Jain, and Haoxiang Wang. An approach for hyperspectral image classification by optimizing SVM using self organizing map. Journal of Computational Science, 25: 252–259, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317308530.

Kokkinos:2018:SPS

[800] Yiannis Kokkinos and Konstantinos G. Margaritis. Simulating parallel scalable probabilistic neural networks via exemplar selection and EM in a ring pipeline. Journal of Computational Science, 25:260–279, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308050.

Hamid:2018:FHD

[801] Isma Hamid, Yu Wu, Qamar Nawaz, and Runqian Zhao. A fast heuristic detection algorithm for visualizing structure of large community. Journal of Computational Science, 25:280–288, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317307780.

Jia:2018:MMT

[802] Xitong Jia, Xianye Ben, Hui Yuan, Kidiyo Kpalma, and Weixiao Meng. Macro-to-micro transformation model for micro-expression recognition. Journal of Computational Science, 25: 289–297, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317303198.

K:2018:SUC

[803] Srikanth Reddy K., Lokesh Kumar Panwar, B. K. Panigrahi, and Rajesh Kumar. Solution to unit commitment in power system operation planning using binary coded modified moth flame optimization algorithm (BMMFOA): a flame selection based computational technique. Journal of Computational Science, 25:298–317, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304210.

Yan:2018:MDM

[804] Wen-Jing Yan and Yu-Hsin Chen. Measuring dynamic micro-expressions via feature extraction methods. Journal of Computational Science, 25: 318–326, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317302089.

Chen:2018:DFU

[805] Yee Ming Chen, Chi-Shun Hsueh, Chu-Kai Wang, and Tai-Yi Wu. Decision fusion using fuzzy threshold scheme for target detection in sensor networks. Journal of Computational Science, 25:327–338, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317309584.

Elhoseny:2018:BCB

[806] Mohamed Elhoseny, Alaa Tharwat, and Aboul Ella Hassanien. Bézier curve based path planning in a dynamic field using modified genetic algorithm. Journal of Computational Science, 25:339–350, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308906.

Sikha:2018:SRD

[807] O. K. Sikha, S. Sachin Kumar, and K. P. Soman. Salient region detection and object segmentation in color images using dynamic mode decomposition. Journal of Computational Science, 25:351–366, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308049.

Das:2018:PMS

[808] Sapan Kumar Das, S. A. Edalatpanah, and T. Mandal. A proposed model for solving fuzzy linear fractional programming problem: Numerical point of view. Journal of Computational Science, 25:367–375, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313571.

Anter:2018:CIO

[809] Ahmed M. Anter and Aboul Ella Hassenian. Computational intelligence optimization approach based on particle swarm optimizer and neutrosophic set for abdominal CT liver tumor segmentation. Journal of Computational Science, 25:376–387, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318300218.

Srivastava:2018:ECT

[810] Amit Kumar Srivastava and Shishir Kumar. An effective computational technique for taxonomic position of security vulnerability in software development. Journal of Computational Science, 25:388–396, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308918.

Jain:2018:ABA

[811] Deepak Kumar Jain, Neha Jain, Shishir Kumar, Amit Kumar, Raj Kumar, and Haoxiang Wang. An approach for behavior analysis using correlation spectral embedding method. Journal of Computational Science, 25:397–405, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317308128.

Jain:2018:ESP

[812] Vinay Kumar Jain and Shishir Kumar. Effective surveillance and predictive mapping of mosquito-borne diseases using social media. Journal of Computational Science, 25: 406–415, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317307834.

Weber:2018:CSE

[813] Gerhard-Wilhelm Weber, Pandian Vasant, and Jose Antonio Marmolejo Saucedo. Computer science and engineering. Journal of Computational Science, 25:416–418, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303429.

Matveev:2018:IDF

[814] Ivan Matveev, Vladimir Novik, and Igor Litvinchev. Influence of degrading factors on the optimal spatial and spectral features of biometric templates. Journal of Computational Science, 25:419–424, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301916.

Fuchigami:2018:SCS

[815] Helio Yochihiro Fuchigami and Socorro Rangel. A survey of case studies in production scheduling: Analysis and perspectives. Journal of Computational Science, 25:425–436, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300741.

Zelinka:2018:NAE

[816] Ivan Zelinka, Lukas Tomaszek, Pandian Vasant, Tran Trong Dao, and Duy Vo Hoang. A novel approach on evolutionary dynamics analysis — a progress report. Journal of Computational Science, 25: 437–445, March 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317301126.

Beaudou:2018:EPO

[817] Laurent Beaudou, Kaoutar Ghazi, Giacomo Kahn, Olivier Raynaud, and Eric Thierry. Encoding partial orders through modular decomposition. Journal of Computational Science, 25:446–455, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300704.

Abualigah:2018:NFS

[818] Laith Mohammad Abualigah, Ahamad Tajudin Khader, and Essam Said Hanandeh. A new feature selection method to improve the document clustering using particle swarm optimization algorithm. Journal of Computational Science, 25:456–466, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316305002.

Ahmed:2018:SSA

[819] Abdulghani Ali Ahmed and Mohammed Falah Mohammed. SAIRF: a similarity approach for attack intention recognition using fuzzy min-max neural network. Journal of Computational Science, 25:467–473, March 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300716.

Anonymous:2018:EBc

[820] Anonymous. Editorial Board. Journal of Computational Science, 26:ii, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750318305088.

Anonymous:2018:PMb

[821] Anonymous. Pages 1–532 (May 2018). Journal of Computational Science, 26: 1–532, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Hotzer:2018:PMP

[822] J. Hötzer, A. Reiter, H. Hierl, P. Steinmetz, M. Selzer, and Britta Nestler. The parallel multi-physics phase-field framework Pace3D. Journal of Computational Science, 26:1–12, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317310116.

Antonov:2018:IVA

[823] Anatoliy Antonov and Lars Linsen. Interactive visual analysis and classification of hyperspectral imaging data. Journal of Computational Science, 26:13–21, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317310220.

Lovetskiy:2018:RCO

[824] K. P. Lovetskiy, L. A. Sevastianov, D. S. Kulyabov, and N. E. Nikolaev. Regularized computation of oscillatory integrals with stationary points. Journal of Computational Science, 26:22–27, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317314151.

Laurini:2018:MAA

[825] Erik Laurini, Domenico Marson, Maurizio Fermeglia, and Sabrina Pricl. Multimodel approach for accurate determination of industry-driven properties for Polymer Nanocomposite Materials. Journal of Computational Science, 26:28–38, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312292.

Roy:2018:CAS

[826] Souvik Roy, Arunima Ray, and Sukanta Das. A cellular automaton that solves distributed spanning tree problem. Journal of Computational Science, 26:39–54, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317309006.

Rathore:2018:NFI

[827] Shailendra Rathore, Arun Kumar Sangaiah, and Jong Hyuk Park. A novel framework for internet of knowledge protection in social networking services. Journal of Computational Science, 26:55–65, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310724.

Madej:2018:DED

[828] Lukasz Madej, Mateusz Sitko, Adam Legwand, Konrad Perzynski, and Kazimierz Michalik. Development and evaluation of data transfer protocols in the fully coupled random cellular automata finite element model of dynamic recrystallization. Journal of Computational Science, 26:66–77, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313224.

Kong:2018:CIT

[829] Fanhui Kong, Jian Li, and Zhihan Lv. Construction of intelligent traffic information recommendation system based on long short-term memory. Journal of Computational Science, 26:78–86, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317314084.

Xu:2018:PPM

[830] Kun Xu, Weidong Zhang, and Zheng Yan. A privacy-preserving mobile application recommender system based on trust evaluation. Journal of Computational Science, 26:87–107, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031731428X.

Devasia:2018:IDC

[831] Jeethu V. Devasia and Priya Chandran. Inferring disease causing genes and their pathways: GpRr method. Journal of Computational Science, 26:108–117, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317309821.

Couturier:2018:BES

[832] Raphaël Couturier, Arnaud Giersch, and Mourad Hakem. Best effort strategy and virtual load for asynchronous iterative load balancing. Journal of Computational Science, 26:118–127, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313893.

Li:2018:MRB

[833] Jing Li, Wentao Xu, Wenbo Wan, and Jiande Sun. Movie recommendation based on bridging movie feature and user interest. Journal of Computational Science, 26:128–134, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318300012.

Abdullah:2018:SET

[834] Hasan M. Abdullah and H. Bahlouli. Substrate effects on transport properties of a biased AA-stacked bilayer graphene. Journal of Computational Science, 26:135–140, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317309341.

Alostad:2018:DSB

[835] Jasem M. Alostad. Dimensionality scale back in massive datasets using PDLPP. Journal of Computational Science, 26:141–146, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317314126.

Taou:2018:IUB

[836] Nadia S. Taou, David W. Corne, and Michael A. Lones. Investigating the use of Boolean networks for the control of gene regulatory networks. Journal of Computational Science, 26:147–156, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317307196.

Lusher:2018:HPC

[837] J. Lusher, J. Ji, and J. Orr. High-Performance Correlation and Mapping Engine for rapid generating brain connectivity networks from big fMRI data. Journal of Computational Science, 26:157–164, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317310633.

Zhu:2018:ARV

[838] B. Zhu, F. Ortega, J. Bobadilla, and A. Gutiérrez. Assigning reliability values to recommendations using matrix factorization. Journal of Computational Science, 26:165–177, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317307524.

Wang:2018:BDD

[839] Yan Wang, Qingfen Liu, Han dan Hou, Seungmin Rho, Brij Gupta, Ying xin Mu, and Wei zheng Shen. Big data driven outlier detection for soybean straw near infrared spectroscopy. Journal of Computational Science, 26:178–189, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303708.

Kovalchuk:2018:ACS

[840] Sergey V. Kovalchuk, Valeria V. Krzhizhanovskaya, Petros Koumoutsakos, Eleni Chatzi, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. The art of computational science: Bridging gaps — forming alloys. Journal of Computational Science, 26:190–192, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318304095.

Abuhay:2018:APA

[841] Tesfamariam M. Abuhay, Sergey V. Kovalchuk, Klavdiya Bochenina, Gali-Ketema Mbogo, Alexander A. Visheratin, George Kampis, Valeria V. Krzhizhanovskaya, and Michael H. Lees. Analysis of publication activity of computational science society in 2001–2017 using topic modelling and graph theory. Journal of Computational Science, 26:193–204, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318302461.

Nathan:2018:NAC

[842] Eisha Nathan, Geoffrey Sanders, Van Emden Henson, and David A. Bader. Numerically approximating centrality for graph ranking guarantees. Journal of Computational Science, 26:205–216, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317311304.

Wermelinger:2018:PSC

[843] F. Wermelinger, U. Rasthofer, P. E. Hadjidoukas, and P. Koumoutsakos. Petascale simulations of compressible flows with interfaces. Journal of Computational Science, 26: 217–225, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317312152.

Abdelfattah:2018:BOS

[844] Ahmad Abdelfattah, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Batched one-sided factorizations of tiny matrices using GPUs: Challenges and countermeasures. Journal of Computational Science, 26:226–236, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317311456.

Dong:2018:ASB

[845] Tingxing Dong, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Accelerating the SVD bi-diagonalization of a batch of small matrices using GPUs. Journal of Computational Science, 26:237–245, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031731150X.

Tang:2018:OSC

[846] Meng Tang, Mohamed Gadou, Steven Rennich, Timothy A. Davis, and Sanjay Ranka. Optimized sparse Cholesky factorization on hybrid multicore architectures. Journal of Computational Science, 26:246–253, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312164.

Silva-Palacios:2018:PCH

[847] Daniel Silva-Palacios, Cèsar Ferri, and María José Ramírez-Quintana. Probabilistic class hierarchies for multiclass classification. Journal of Computational Science, 26:254–263, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312188.

Nino-Ruiz:2018:ISM

[848] Elias D. Nino-Ruiz. Implicit surrogate models for trust region based methods. Journal of Computational Science, 26:264–274, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317311912.

Memeti:2018:PPP

[849] Suejb Memeti and Sabri Pllana. PAPA: a parallel programming assistant powered by IBM Watson cognitive computing technology. Journal of Computational Science, 26:275–284, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317311493.

Baumann:2018:CCS

[850] Manuel Baumann and Martin B. van Gijzen. Convergence and complexity study of GMRES variants for solving multi-frequency elastic wave propagation problems. Journal of Computational Science, 26:285–293, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317310852.

Gao:2018:CDS

[851] Gelin Gao, Bud Mishra, and Daniele Ramazzotti. Causal data science for financial stress testing. Journal of Computational Science, 26: 294–304, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317311377.

Zhu:2018:ODI

[852] Xiaomin Zhu, Laurence T. Yang, Hai Jiang, Parimala Thulasiraman, and Beniamino Di Martino. Optimization in distributed information systems. Journal of Computational Science, 26:305–306, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318304435.

Yu:2018:DTS

[853] Siyang Yu, Kenli Li, and Yuming Xu. A DAG task scheduling scheme on heterogeneous cluster systems using discrete IWO algorithm. Journal of Computational Science, 26:307–317, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301582.

Casas:2018:GEE

[854] Israel Casas, Javid Taheri, Rajiv Ranjan, Lizhe Wang, and Albert Y. Zomaya. GA-ETI: an enhanced genetic algorithm for the scheduling of scientific workflows in cloud environments. Journal of Computational Science, 26:318–331, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301399.

Jiang:2018:SDP

[855] Weiwen Jiang, Edwin H.-M. Sha, Xianzhang Chen, Lin Wu, and Qingfeng Zhuge. Synthesizing distributed pipelining systems with timing constraints via optimal functional unit assignment and communication selection. Journal of Computational Science, 26:332–343, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303265.

Xiao:2018:MRE

[856] Xiongren Xiao, Guoqi Xie, Cheng Xu, Chunnian Fan, Renfa Li, and Keqin Li. Maximizing reliability of energy constrained parallel applications on heterogeneous distributed systems. Journal of Computational Science, 26:344–353, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304933.

Gu:2018:HGM

[857] Xin Gu, Xiaoyong Zhang, Yijun Cheng, Zhuofu Zhou, and Jun Peng. A hybrid game method for interference management with energy constraint in 5G ultra-dense HetNets. Journal of Computational Science, 26:354–362, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317310049.

Zhou:2018:OCG

[858] Jingya Zhou, Jianxi Fan, Juncheng Jia, Baolei Cheng, and Zhao Liu. Optimizing cost for geo-distributed storage systems in online social networks. Journal of Computational Science, 26:363–374, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317308724.

Zeng:2018:CES

[859] Xuezhi Zeng, Saurabh Kumar Garg, Zhenyu Wen, Peter Strazdins, Albert Y. Zomaya, and Rajiv Ranjan. Cost efficient scheduling of MapReduce applications on public clouds. Journal of Computational Science, 26:375–388, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317308542.

Shi:2018:FMR

[860] Jiyuan Shi, Junzhou Luo, Fang Dong, Jiahui Jin, and Jun Shen. Fast multi-resource allocation with patterns in large scale cloud data center. Journal of Computational Science, 26:389–401, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305185.

Zhao:2018:CAO

[861] Hui Zhao, Meikang Qiu, Min Chen, and Keke Gai. Cost-aware optimal data allocations for multiple dimensional heterogeneous memories using dynamic programming in big data. Journal of Computational Science, 26:402–408, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031630103X.

Zhou:2018:IBD

[862] Wei Zhou, Dan Feng, Zhipeng Tan, and Yingfei Zheng. Improving big data storage performance in hybrid environment. Journal of Computational Science, 26:409–418, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300558.

Wang:2018:AHA

[863] Yuxiang Wang, Yixing Xia, Qiming Fang, and Xiaoliang Xu. AQP++: a hybrid approximate query processing framework for generalized aggregation queries. Journal of Computational Science, 26:419–431, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304921.

Sahal:2018:ECG

[864] Radhya Sahal, Mohamed H. Khafagy, and Fatma A. Omara. Exploiting coarse-grained reused-based opportunities in Big Data multi-query optimization. Journal of Computational Science, 26:432–452, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317306142.

Zhang:2018:CNC

[865] Zhiyong Zhang, Kim-Kwang Raymond Choo, and Brij B. Gupta. The convergence of new computing paradigms and big data analytics methodologies for online social networks. Journal of Computational Science, 26:453–455, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318303570.

Wang:2018:PGM

[866] Derek Wang, Tingmin Wu, Sheng Wen, Donghai Liu, Yang Xiang, Wanlei Zhou, Houcine Hassan, and Abdulhameed Alelaiwi. Pokémon GO in Melbourne CBD: a case study of the cyber-physical symbiotic social networks. Journal of Computational Science, 26:456–467, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317303277.

Zhang:2018:NCE

[867] Zhiyong Zhang, Jing Wen, Xiaoxue Wang, and Changwei Zhao. A novel crowd evaluation method for security and trustworthiness of online social networks platforms based on signaling theory. Journal of Computational Science, 26:468–477, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730337X.

Zhang:2018:HDB

[868] Yangyang Zhang, Jianxin Li, Chenggen Sun, Md Zakirul Alam Bhuiyan, Weiren Yu, and Richong Zhang. HotML: a DSM-based machine learning system for social networks. Journal of Computational Science, 26: 478–487, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317303459.

Xia:2018:HOB

[869] Xuewen Xia, Ling Gui, Guoliang He, Chengwang Xie, Bo Wei, Ying Xing, Ruifeng Wu, and Yichao Tang. A hybrid optimizer based on firefly algorithm and particle swarm optimization algorithm. Journal of Computational Science, 26:488–500, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303538.

Peng:2018:EDE

[870] Hu Peng, Zhaolu Guo, Changshou Deng, and Zhijian Wu. Enhancing differential evolution with random neighbors based strategy. Journal of Computational Science, 26: 501–511, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317303587.

Zhang:2018:MBT

[871] Shunxiang Zhang, Wenjuan Liu, XiaoLu Deng, Zheng Xu, and Kim-Kwang Raymond Choo. Microblog topic recommendation based on knowledge flow and user selection. Journal of Computational Science, 26:512–521, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317303629.

Al-Ayyoub:2018:DLA

[872] Mahmoud Al-Ayyoub, Aya Nuseir, Kholoud Alsmearat, Yaser Jararweh, and Brij Gupta. Deep learning for Arabic NLP: a survey. Journal of Computational Science, 26: 522–531, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317303757.

Anonymous:2018:EBd

[873] Anonymous. Editorial Board. Journal of Computational Science, 27:ii, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750318308196.

Anonymous:2018:PJb

[874] Anonymous. Pages 1–476 (July 2018). Journal of Computational Science, 27: 1–476, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Basha:2018:MFD

[875] Nour Basha, Mohamed Nounou, and Hazem Nounou. Multivariate fault detection and classification using interval principal component analysis. Journal of Computational Science, 27:1–9, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317312450.

Stamatakis:2018:MSS

[876] Emmanuel Stamatakis, Andreas Yiotis, Stella Giannissi, Ilias Tolias, and Athanassios Stubos. Modeling and simulation supporting the application of fuel cell & hydrogen technologies. Journal of Computational Science, 27:10–20, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031731339X.

Yuen:2018:NSD

[877] A. C. Y. Yuen, G. H. Yeoh, S. C. P. Cheung, Q. N. Chan, T. B. Y. Chen, W. Yang, and H. Lu. Numerical study of the development and angular speed of a small-scale fire whirl. Journal of Computational Science, 27:21–34, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310141.

Zhai:2018:RTA

[878] Xiaojun Zhai, Mohammad Eslami, Ealaf Sayed Hussein, Maroua Salem Filali, Salma Tarek Shalaby, Abbes Amira, Faycal Bensaali, Sarada Dakua, Julien Abinahed, Abdulla Al-Ansari, and Ayman Z. Ahmed. Real-time automated image segmentation technique for cerebral aneurysm on reconfigurable system-on-chip. Journal of Computational Science, 27:35–45, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313005.

Toye:2018:FTH

[879] Habib Toye, Samuel Kortas, Peng Zhan, and Ibrahim Hoteit. A faulttolerant HPC scheduler extension for large and operational ensemble data assimilation: Application to the Red Sea. Journal of Computational Science, 27:46–56, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312905.

Zhang:2018:ABI

[880] Yu-Dong Zhang, Chichun Pan, Xianqing Chen, and Fubin Wang. Abnormal breast identification by nine-layer convolutional neural network with parametric rectified linear unit and rank-based stochastic pooling. Journal of Computational Science, 27:57–68, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318302655.

Wahid:2018:SIH

[881] Md Ferdous Wahid, Reza Tafreshi, Mubarak Al-Sowaidi, and Reza Langari. Subject-independent hand gesture recognition using normalization and machine learning algorithms. Journal of Computational Science, 27:69–76, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312632.

Fabregat-Traver:2018:AMD

[882] Diego Fabregat-Traver, Ahmed E. Ismail, and Paolo Bientinesi. Accelerating molecular dynamics codes by performance and accuracy modeling. Journal of Computational Science, 27:77–90, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305549.

Leiter:2018:ASB

[883] Kenneth W. Leiter, Brian C. Barnes, Richard Becker, and Jaroslaw Knap. Accelerated scale-bridging through adaptive surrogate model evaluation. Journal of Computational Science, 27:91–106, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313807.

Gavin:2018:FEN

[884] Brendan Gavin, Agnieszka Miȩdlar, and Eric Polizzi. FEAST eigensolver for nonlinear eigenvalue problems. Journal of Computational Science, 27:107–117, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318302096.

Barth:2018:ITE

[885] Nicolas Barth, Zaid S. Al Otaibi, and Saïd Ahzi. Irradiance, thermal and electrical coupled modeling of photovoltaic panels with long-term simulation periods under service in harsh desert conditions. Journal of Computational Science, 27:118–129, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318300632.

Zhuge:2018:ABM

[886] Chengxiang Zhuge and Chunfu Shao. Agent-based modelling of purchasing, renting and investing behaviour in dynamic housing markets. Journal of Computational Science, 27: 130–146, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317313546.

Bartuschat:2018:SMA

[887] Dominik Bartuschat and Ulrich Rüde. A scalable multiphysics algorithm for massively parallel direct numerical simulations of electrophoretic motion. Journal of Computational Science, 27:147–167, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317311298.

Roy:2018:ENV

[888] Arup Roy, Soumya Banerjee, Manash Sarkar, Ashraf Darwish, Mohamed Elhoseny, and Aboul Ella Hassanien. Exploring new vista of intelligent collaborative filtering: a restaurant recommendation paradigm. Journal of Computational Science, 27:168–182, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318301650.

Mohamed:2018:CSM

[889] Amro Mohamed, Panagiotis Krokidas, and Ioannis G. Economou. CO 2 selective metal organic framework ZIF-8 modified through ionic liquid encapsulation: a computational study. Journal of Computational Science, 27:183–191, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313364.

Marquez-Vera:2018:SFC

[890] M. A. Márquez-Vera, L. E. Ramos-Velasco, and B. D. Balderrama-Hernández. Stable fuzzy control and observer via LMIs in a fermentation process. Journal of Computational Science, 27:192–198, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303301.

Zhang:2018:FEB

[891] Nan Zhang, Yi Chen, Maolong Xi, Fangqin Wang, and Yanwen Qu. Feature extraction based on low-rank affinity matrix for biological recognition. Journal of Computational Science, 27:199–205, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318303338.

Wu:2018:GTA

[892] Jiehua Wu. A generalized tree augmented naive Bayes link prediction model. Journal of Computational Science, 27:206–217, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317301400.

Elejalde:2018:QED

[893] Erick Elejalde, Leo Ferres, Eelco Herder, and Johan Bollen. Quantifying the ecological diversity and health of online news. Journal of Computational Science, 27:218–226, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317311870.

Sheriff:2018:PMU

[894] M. Ziyan Sheriff, M. Nazmul Karim, Hazem N. Nounou, and Mohamed N. Nounou. Process monitoring using PCA-based GLR methods: a comparative study. Journal of Computational Science, 27:227–246, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317312590.

Akila:2018:CSR

[895] S. Akila and U. Srinivasulu Reddy. Cost-sensitive Risk Induced Bayesian Inference Bagging (RIBIB) for credit card fraud detection. Journal of Computational Science, 27:247–254, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750317311729.

Pichon:2018:SSS

[896] Grégoire Pichon, Eric Darve, Mathieu Faverge, Pierre Ramet, and Jean Roman. Sparse supernodal solver using block low-rank compression: Design, performance and analysis. Journal of Computational Science, 27:255–270, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317314497.

Chen:2018:IEC

[897] Shu sheng Chen, Chao Yan, Shuai Lou, and Bo xi Lin. An improved entropyconsistent Euler flux in low Mach number. Journal of Computational Science, 27:271–283, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318301509.

Gaidhane:2018:HGW

[898] Prashant J. Gaidhane and Madhav J. Nigam. A hybrid grey wolf optimizer and artificial bee colony algorithm for enhancing the performance of complex systems. Journal of Computational Science, 27:284–302, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318301844.

Peterson:2018:DGC

[899] Brad Peterson, Alan Humphrey, John Holmen, Todd Harman, Martin Berzins, Dan Sunderland, and H. Carter Edwards. Demonstrating GPU code portability and scalability for radiative heat transfer computations. Journal of Computational Science, 27:303–319, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317314485.

Jiang:2018:IRI

[900] Yimin Jiang, Kuangrong Hao, Xin Cai, and Yongsheng Ding. An improved reinforcement-immune algorithm for agricultural resource allocation optimization. Journal of Computational Science, 27:320–328, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318302333.

Merzky:2018:SSA

[901] Andre Merzky, Ming Tai Ha, Matteo Turilli, and Shantenu Jha. Synapse: Synthetic application profiler and emulator. Journal of Computational Science, 27:329–344, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318300966.

Barsamian:2018:EDL

[902] Yann Barsamian, Sever A. Hirstoaga, and Éric Violard. Efficient data layouts for a three-dimensional electrostatic Particle-in-Cell code. Journal of Computational Science, 27: 345–356, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317313856.

Aly:2018:FTN

[903] Shahzada Pamir Aly, Nicolas Barth, Benjamin W. Figgis, and Said Ahzi. A fully transient novel thermal model for in-field photovoltaic modules using developed explicit and implicit finite difference schemes. Journal of Computational Science, 27:357–369, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317311481.

Singh:2018:HFT

[904] Pritpal Singh and Gaurav Dhiman. A hybrid fuzzy time series forecasting model based on granular computing and bio-inspired optimization approaches. Journal of Computational Science, 27:370–385, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300923.

Al-Smadi:2018:DRN

[905] Mohammad Al-Smadi, Omar Qawasmeh, Mahmoud Al-Ayyoub, Yaser Jararweh, and Brij Gupta. Deep Recurrent neural network vs. support vector machine for aspect-based sentiment analysis of Arabic hotels’ reviews. Journal of Computational Science, 27:386–393, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305252.

Kiwia:2018:CKC

[906] Dennis Kiwia, Ali Dehghantanha, Kim-Kwang Raymond Choo, and Jim Slaughter. A cyber kill chain based taxonomy of banking Trojans for evolutionary computational intelligence. Journal of Computational Science, 27:394–409, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304957.

Moreira:2018:ERB

[907] Mário W. L. Moreira, Joel J. P. C. Rodrigues, Neeraj Kumar, Jalal Al-Muhtadi, and Valeriy Korotaev. Evolutionary radial basis function network for gestational diabetes data analytics. Journal of Computational Science, 27:410–417, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317304726.

Zhou:2018:EBR

[908] Yu Zhou and Tingling Wang. ENNbased recognition method for tool cutting state. Journal of Computational Science, 27:418–427, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317301989.

Kumar:2018:BDD

[909] Ajay Kumar, Ravi Shankar, and Lakshman S. Thakur. A big data driven sustainable manufacturing framework for condition-based maintenance prediction. Journal of Computational Science, 27:428–439, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316305129.

Arunkumar:2018:EFF

[910] N. Arunkumar, K. Ram Kumar, and V. Venkataraman. Entropy features for focal EEG and non focal EEG. Journal of Computational Science, 27:440–444, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318301364.

Xu:2018:RSM

[911] Yuanyuan Xu, Hongjuan Ge, and Jingzhong Yang. Research on the scheduling method of AFDX terminal system based on time triggered and event triggered. Journal of Computational Science, 27:445–453, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313996.

Jianhui:2018:SMM

[912] Nie Jianhui. Stereo matching method for non-coding circular reference points based on motion consistency. Journal of Computational Science, 27:454–461, July 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317313984.

Yongdong:2018:BLP

[913] Pan Yongdong. Bi-level programming optimization method for cloud manufacturing service composition based on harmony search. Journal of Computational Science, 27:462–468, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317313583.

Feng:2018:EBG

[914] Wang Feng. Enterprise benefit game model of collaborative supply chain in logistics industry park. Journal of Computational Science, 27:469–475, July 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317313558.

Anonymous:2018:EBe

[915] Anonymous. Editorial Board. Journal of Computational Science, 28: ii, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318310810.

Anonymous:2018:PS

[916] Anonymous. Pages 1–482 (September 2018). Journal of Computational Science, 28:1–482, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Zhang:2018:MSI

[917] Yu-Dong Zhang, Chichun Pan, Junding Sun, and Chaosheng Tang. Multiple sclerosis identification by convolutional neural network with dropout and parametric ReLU. Journal of Computational Science, 28:1–10, 2018. ISSN 1877-7503 (print), 1877-7511 (electronic).

Mohammed:2018:CLE

[918] Lawal Mohammed, Muhammad A. Saeed, Qinfang Zhang, and Auwalu Musa. Core-level excitation in polymorph of AS_{2}S_{3} and β-In_{2}S_{3}. Journal of Computational Science, 28:11– 17, September 2018. ISSN 1877-7503 (print), 1877-7511 (electronic).

Zang:2018:AOS

[919] B. Zang, Vevek US, H. D. Lim, X. Wei, and T. H. New. An assessment of OpenFOAM solver on RANS simulations of round supersonic free jets. Journal of Computational Science, 28:18–31, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303958.

Malecki:2018:CST

[920] Krzysztof Małecki. A computer simulation of traffic flow with on-street parking and drivers’ behaviour based on cellular automata and a multi-agent system. Journal of Computational Science, 28:32–42, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312279.

Inan:2018:MGP

[921] Emrah Inan, Fatih Tekbacak, and Cemalettin Ozturk. Moreopt: a goal programming based movie recommender system. Journal of Computational Science, 28:43–50, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317314540.

Tsompanas:2018:FGS

[922] Michail-Antisthenis Tsompanas and Andrew Adamatzky. Fluidic gates simulated with lattice Boltzmann method under different Reynolds numbers. Journal of Computational Science, 28:51–58, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318302412.

Gao:2018:DPD

[923] Xiang Gao, Chuanfu Xu, Yidao Dong, Min Xiong, Dali Li, Zhenghua Wang, and Xiaogang Deng. Developing a parallel density-based implicit solver with mesh deformation in Open-FOAM. Journal of Computational Science, 28:59–69, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318301613.

Rubino:2018:LTM

[924] A. Rubino, M. Pini, M. Kosec, S. Vitale, and P. Colonna. A look-up table method based on unstructured grids and its application to non-ideal compressible fluid dynamic simulations. Journal of Computational Science, 28:70–77, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318305891.

Kilani:2018:GAB

[925] Yousef Kilani, Ahmed Fawzi Otoom, Ayoub Alsarhan, and Manal Almaayah. A genetic algorithms-based hybrid recommender system of matrix factorization and neighborhood-based techniques. Journal of Computational Science, 28:78–93, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318301704.

Khan:2018:FRU

[926] Sajid Ali Khan, Muhammad Ishtiaq, Muhammad Nazir, and Muhammad Shaheen. Face recognition under varying expressions and illumination using particle swarm optimization. Journal of Computational Science, 28:94–100, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317312255.

Lanzafame:2018:GTL

[927] G. Lanzafame. A general tool for LTE thermochemistry for adiabatic non-diffusive reactive fluid dynamics: Applications to 2D planar discontinuity flows in SPH. Journal of Computational Science, 28: 101–119, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318301595.

Leitao:2018:SVD

[928] Álvaro Leitao, Luis Ortiz-Gracia, and Emma I. Wagner. SWIFT valuation of discretely monitored arithmetic Asian options. Journal of Computational Science, 28:120–139, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031830228X.

Chow:2018:WAN

[929] W. K. Chow and J. Li. Wind action on natural smoke exhaust in atria. Journal of Computational Science, 28:140–147, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318302485.

Ayyildiz:2018:MVB

[930] Ezgi Ayyıldız and Vilda Purutçuoğlu. Modeling of various biological networks via LCMARS. Journal of Computational Science, 28:148– 154, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317304489.

Amikiya:2018:MSR

[931] Adoliwine E. Amikiya and Mapundi K. Banda. Modelling and simulation of reactive transport phenomena. Journal of Computational Science, 28: 155–167, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318302631.

Nilashi:2018:TDM

[932] Mehrbakhsh Nilashi, Othman Ibrahim, Elaheh Yadegaridehkordi, Sarminah Samad, Elnaz Akbari, and Azar Alizadeh. Travelers decision making using online review in social network sites: a case on TripAdvisor. Journal of Computational Science, 28:168–179, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303363.

Yadav:2018:TAR

[933] Sambhav Yadav, Vikesh Kumar, Shreyam Sinha, and Sushama Nagpal. Trust aware recommender system using swarm intelligence. Journal of Computational Science, 28: 180–192, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317314588.

Kemmer:2018:NJE

[934] Thomas Kemmer, Sergej Rjasanow, and Andreas Hildebrandt. NESSie.jl — efficient and intuitive finite element and boundary element methods for nonlocal protein electrostatics in the Julia language. Journal of Computational Science, 28: 193–203, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775031730738X.

Hu:2018:IDO

[935] Ying Hu, Marco Aiello, and Changjun Hu. Information diffusion in online social networks: a compilation. Journal of Computational Science, 28:204–205, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318309566.

He:2018:NPI

[936] Jing He, Xin Li, and Lejian Liao. Next point-of-interest recommendation via a category-aware Listwise Bayesian Personalized Ranking. Journal of Computational Science, 28: 206–216, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317306476.

Quan:2018:RPI

[937] Yong Quan, Yan Jia, Bin Zhou, Weihong Han, and Shudong Li. Repost prediction incorporating timesensitive mutual influence in social networks. Journal of Computational Science, 28:217–227, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317306543.

Hu:2018:ESS

[938] Jingyuan Hu, Yingwei Luo, and Jian Yu. An empirical study on selectiviey of retweeting behaviors under multiple exposures in social networks. Journal of Computational Science, 28:228–235, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306439.

Li:2018:CSB

[939] Lei Li, Chuan Zhou, Jianping He, Jiamiao Wang, Xin Li, and Xindong Wu. Collective semantic behavior extraction in social networks. Journal of Computational Science, 28:236–244, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305793.

Zhu:2018:IEM

[940] Xiang Zhu, Zhefeng Wang, Yu Yang, Bin Zhou, and Yan Jia. Influence efficiency maximization: How can we spread information efficiently? Journal of Computational Science, 28: 245–256, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317305288.

Hoang:2018:PID

[941] Thi Bich Ngoc Hoang and Josiane Mothe. Predicting information diffusion on Twitter — analysis of predictive features. Journal of Computational Science, 28:257–264, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305860.

Atay:2018:SIB

[942] Yilmaz Atay, Murat Aslan, and Halife Kodaz. A swarm intelligence-based hybrid approach for identifying network modules. Journal of Computational Science, 28:265–280, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306208.

Kou:2018:SMM

[943] Feifei Kou, Junping Du, Zijian Lin, Meiyu Liang, Haisheng Li, Lei Shi, and Congxian Yang. A semantic modeling method for social network short text based on spatial and temporal characteristics. Journal of Computational Science, 28:281–293, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305318.

Zhang:2018:EIS

[944] Xi Zhang, Jiawei Shi, Di Wang, and Binxing Fang. Exploiting investors social network for stock prediction in China’s market. Journal of Computational Science, 28: 294–303, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317306361.

Zhou:2018:EDD

[945] Chuan Zhou, Wei-Xue Lu, Jingzun Zhang, Lei Li, Yue Hu, and Li Guo. Early detection of dynamic harmful cascades in large-scale networks. Journal of Computational Science, 28:304–317, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305434.

Li:2018:DCI

[946] Huijuan Li, Li Pan, and Peng Wu. Dominated competitive influence maximization with time-critical and timedelayed diffusion in social networks. Journal of Computational Science, 28:318–327, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306117.

Xu:2018:UPP

[947] Wenwen Xu, Peng Shi, Jianyi Huang, and Feng Liu. Understanding and predicting the peak popularity of bursting hashtags. Journal of Computational Science, 28:328–335, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317305914.

Zhou:2018:EJN

[948] Pengpeng Zhou, Zhen Cao, Bin Wu, Chunzi Wu, and Shuqi Yu. EDM-JBW: a novel event detection model based on JS-ID’F_{order} and Bikmeans with word embedding for news streams. Journal of Computational Science, 28:336–342, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317306646.

Ding:2018:TBR

[949] Cangfeng Ding and Kan Li. Topologically biased random walk for diffusions on multiplex networks. Journal of Computational Science, 28: 343–356, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317305604.

Bouhali:2018:SRD

[950] O. Bouhali, I. G. Economou, and F. El-Mellouhi. Some recent developments on computational science and engineering. Journal of Computational Science, 28: 357–359, September 2018. ISSN 1877-7503 (print), 1877-7511 (electronic).

Li:2018:KSA

[951] Junke Li, Bing Guo, Yan Shen, Deguang Li, and Yanhui Huang. Kernel scheduling approach for reducing GPU energy consumption. Journal of Computational Science, 28: 360–368, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317313376.

Zhang:2018:PBT

[952] Lu Zhang, Zhiang Wu, Zhan Bu, Ye Jiang, and Jie Cao. A patternbased topic detection and analysis system on Chinese tweets. Journal of Computational Science, 28: 369–381, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317309511.

Liu:2018:RPS

[953] Yizhi Liu, Jianxun Liu, Zhuhua Liao, Mingdong Tang, and Jinjun Chen. Recommending a personalized sequence of pick-up points. Journal of Computational Science, 28:

382–388, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317305173.

Zhang:2018:WPD

[954] Yuzuo Zhang and Boyi Li. Wild plant data collection system based on distributed location. Journal of Computational Science, 28:389– 397, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317304313.

Benoit:2018:MLC

[955] Anne Benoit, Aurélien Cavelan, Yves Robert, and Hongyang Sun. Multilevel checkpointing and silent error detection for linear workflows. Journal of Computational Science, 28: 398–415, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750317303599.

Chen:2018:EEE

[956] Huangke Chen, Guipeng Liu, Shu Yin, Xiaocheng Liu, and Dishan Qiu. ERECT: Energy-efficient reactive scheduling for real-time tasks in heterogeneous virtualized clouds. Journal of Computational Science, 28:416–425, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317303204.

Esposito:2018:DSL

[957] Christian Esposito, Aniello Castiglione, Francesco Palmieri, and Florin Pop. Distributed strategic learning and game theoretic formulation of network embedded coding. Journal of Computational Science, 28: 426–438, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317303058.

Carpaye:2018:DAT

[958] Jean Marie Couteyen Carpaye, Jean Roman, and Pierre Brenner. Design and analysis of a task-based parallelization over a runtime system of an explicit finite-volume CFD code with adaptive time stepping. Journal of Computational Science, 28: 439–454, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317302880.

Guo:2018:SOS

[959] Kehua Guo, Zhonghe Liang, Yayuan Tang, and Tao Chi. SOR: an optimized semantic ontology retrieval algorithm for heterogeneous multimedia big data. Journal of Computational Science, 28:455–465, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031730176X.

Zhong:2018:HVF

[960] Fangming Zhong, Zhikui Chen, Zhaolong Ning, Geyong Min, and Yueming Hu. Heterogeneous visual features integration for image recognition optimization in Internet of Things. Journal of Computational Science, 28:466–475, September 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303076.

Meng:2018:ANP

[961] Jianping Meng, Xiao-Jun Gu, and David R. Emerson. Analysis of nonphysical slip velocity in lattice Boltzmann simulations using the bounceback scheme. Journal of Computational Science, 28:476–482, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317300583.

Anonymous:2018:EBf

[962] Anonymous. Editorial Board. Journal of Computational Science, 29: ii, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S1877750318312614.

Anonymous:2018:PN

[963] Anonymous. Pages 1–168 (November 2018). Journal of Computational Science, 29:1–168, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

R:2018:FSM

[964] Simi V. R., Damodar Reddy Edla, and Justin Joseph. A fuzzy sharpness metric for magnetic resonance images. Journal of Computational Science, 29:1–8, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318307543.

Anwaar:2018:HCH

[965] Fahad Anwaar, Naima Iltaf, Hammad Afzal, and Raheel Nawaz. HRS-CE: a hybrid framework to integrate content embeddings in recommender systems for cold start items. Journal of Computational Science, 29:9–18, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318300036.

Alfeo:2018:DSE

[966] Antonio L. Alfeo, Mario G. C. A. Cimino, Nicoletta De Francesco, Massimiliano Lega, and Gigliola Vaglini. Design and simulation of the emergent behavior of small drones swarming for distributed target localization. Journal of Computational Science, 29:19–33, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318302898.

Sadad:2018:FCM

[967] Tariq Sadad, Asim Munir, Tanzila Saba, and Ayyaz Hussain. Fuzzy C-means and region growing based classification of tumor from mammograms using hybrid texture feature. Journal of Computational Science, 29:34–45, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317310062.

Zhuge:2018:ABS

[968] Chengxiang Zhuge, Chunfu Shao, and Binru Wei. An agent-based spatial urban social network generator: a case study of Beijing, China. Journal of Computational Science, 29: 46–58, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318300103.

Casarin:2018:TUA

[969] Stefano Casarin, Scott A. Berceli, and Marc Garbey. A twofold usage of an agent-based model of vascular adaptation to design clinical experiments. Journal of Computational Science, 29:59–69, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318308731.

Li:2018:PFP

[970] Jingfa Li, Bo Yu, Shuyu Sun, Dongliang Sun, and Yasuo Kawaguchi. An N-parallel FENE-P constitutive model and its application in largeeddy simulation of viscoelastic turbulent drag-reducing flow. Journal of Computational Science, 29: 70–80, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318309049.

Ezzat:2018:BRS

[971] Mahmoud Ezzat, Mahmoud Sakr, Rania Elgohary, and Mohammed Essam Khalifa. Building road segments and detecting turns from GPS tracks. Journal of Computational Science, 29:81–93, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318302813.

Shende:2018:CHR

[972] Suraj B. Shende, Ashish B. Deoghare, and Krishna M. Pandey. Characterization of harmonic response of human middle ear using finite element approach. Journal of Computational Science, 29:94–98, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031830557X.

Robnik-Sikonja:2018:MIF

[973] Marko Robnik-Šikonja, Miloš Radović, Smiljana – Dorović, Bojana An– delković-Ćirković, and Nenad Filipović. Modeling ischemia with finite elements and automated machine learning. Journal of Computational Science, 29: 99–106, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318304526.

Garcia-Aunon:2018:COA

[974] Pablo Garcia-Aunon and Antonio Barrientos Cruz. Control optimization of an aerial robotic swarm in a search task and its adaptation to different scenarios. Journal of Computational Science, 29:107–118, November 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318308482. See

corrigendum [1364].

Dinkar:2018:EOB

[975] Shail Kumar Dinkar and Kusum Deep. An efficient opposition based Lévy Flight Antlion optimizer for optimization problems. Journal of Computational Science, 29:119– 141, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318306525.

Zhang:2018:CFB

[976] Tao Zhang, Lei Li, Lu Shi-hong, Jia bin Zhang, and Hai Gong. Comparisons of flow behavior characteristics and microstructure between asymmetrical shear rolling and symmetrical rolling by macro / micro coupling simulation. Journal of Computational Science, 29:142–152, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318304770.

Chen:2018:NSS

[977] Qi Chen, Xianxu Yuan, Xinguang Wang, Jianqiang Chen, and Yufei Xie. Numerical study of saddlenode bifurcation for longitudinal flight with CFD/RBD technique. Journal of Computational Science, 29: 153–162, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317311651.

Adhikari:2018:BPP

[978] Kapil Adhikari, Aurab Chakrabarty, Othmane Bouhali, Normand Mousseau, Charlotte S. Becquart, and Fedwa El-Mellouhi. Benchmarking the performance of plane-wave vs. localized orbital basis set methods in DFT modeling of metal surface: a case study for Fe-(110). Journal of Computational Science, 29:163–167, November 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317313261.

Anonymous:2019:EBa

[979] Anonymous. Editorial Board. Journal of Computational Science, 30:ii, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750319300420.

Anonymous:2019:PJa

[980] Anonymous. Pages 1–222 (January 2019). Journal of Computational Science, 30:1–222, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ramazzotti:2019:ECS

[981] Daniele Ramazzotti, Marco S. Nobile, Marco Antoniotti, and Alex Graudenzi. Efficient computational strategies to learn the structure of probabilistic graphical models of cumulative phenomena. Journal of Computational Science, 30:1–10, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318309967.

Yang:2019:AEL

[982] Jingyuan Yang, Qiaoyong Jiang, Lei Wang, Shuai Liu, Yu-Dong Zhang, Wei Li, and Bin Wang. An adaptive encoding learning for artificial bee colony algorithms. Journal of Computational Science, 30:11–27, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310123.

Cruz-Matias:2019:SRC

[983] Irving Cruz-Matías, Dolors Ayala, Daniel Hiller, Sebastian Gutsch, Margit Zacharias, Sònia Estradé, and Francesca Peiró. Sphericity and roundness computation for particles using the extreme vertices model. Journal of Computational Science, 30:28–40, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318304757.

Lu:2019:PBD

[984] Siyuan Lu, Zhihai Lu, and Yu-Dong Zhang. Pathological brain detection based on AlexNet and transfer learning. Journal of Computational Science, 30:41–47, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318309116.

Zang:2019:MCR

[985] Sheng Zang, Chao Wang, and Jian Dong. A multilinear collaborative representation preserving projections method for feature extraction. Journal of Computational Science, 30:48–54, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318306860.

Filho:2019:CST

[986] Pedro Pedrosa Rebouças Filho, Suane Pires P. da Silva, Victor N. Praxedes, Jude Hemanth, and Victor Hugo C. de Albuquerque. Control of singularity trajectory tracking for robotic manipulator by genetic algorithms. Journal of Computational Science, 30:55–64, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318307634.

Lyu:2019:ISA

[987] Shilei Lyu, Zhen Li, Yonglin Huang, Jianhua Wang, and Jie Hu. Improved self-adaptive bat algorithm with step-control and mutation mechanisms. Journal of Computational Science, 30:65–78, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317314102.

Leng:2019:NSI

[988] Wei Leng, Chen-Song Zhang, Pengtao Sun, Bin Gao, and Jinchao Xu. Numerical simulation of an immersed rotating structure in fluid for hemodynamic applications. Journal of Computational Science, 30:79–89, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318309979.

Yang:2019:EDM

[989] Xianfeng Yang and Jian Yang. Efficient diffeomorphic metric image registration via stationary velocity. Journal of Computational Science, 30:90–97, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318308755.

Mallek:2019:ELP

[990] Sabrine Mallek, Imen Boukhris, Zied Elouedi, and Eric Lefèvre. Evidential link prediction in social networks based on structural and social information. Journal of Computational Science, 30:98–107, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303053.

Zhu:2019:NTG

[991] Qiannan Zhu, Xiaofei Zhou, Peng Zhang, and Yong Shi. A neural translating general hyperplane for knowledge graph embedding. Journal of Computational Science, 30: 108–117, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318310172.

Guo:2019:GTA

[992] Yunchuan Guo, Han Zhang, Lingcui Zhang, Liang Fang, and Fenghua Li. A game theoretic approach to cooperative intrusion detection. Journal of Computational Science, 30: 118–126, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318310202.

Abbassi:2019:ROI

[993] Abderrahman Abbassi, Ahmed El hilali Alaoui, and Jaouad Boukachour. Robust optimisation of the intermodal freight transport problem: Modeling and solving with an efficient hybrid approach. Journal of Computational Science, 30:127–142, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318306653.

Garcia:2019:VNG

[994] Mariangel Garcia, Paul F. Choboter, Ryan K. Walter, and Jose E. Castillo. Validation of the nonhydrostatic General Curvilinear Coastal Ocean Model (GCCOM) for stratified flows. Journal of Computational Science, 30: 143–156, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318307919.

Puzyrev:2019:PCF

[995] Vladimir Puzyrev, Mehdi Ghommem, and Shiv Meka. pyROM: a computational framework for reduced order modeling. Journal of Computational Science, 30:157–173, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318307518.

Sajjad:2019:MGB

[996] Muhammad Sajjad, Salman Khan, Khan Muhammad, Wanqing Wu, Amin Ullah, and Sung Wook Baik. Multigrade brain tumor classification using deep CNN with extensive data augmentation. Journal of Computational Science, 30:174–182, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318307385.

Wienbrandt:2019:FTP

[997] Lars Wienbrandt, Jan Christian Kässens, Matthias Hübenthal, and David Ellinghaus. 1000× faster than PLINK: Combined FPGA and GPU accelerators for logistic regressionbased detection of epistasis. Journal of Computational Science, 30: 183–193, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318310184.

Oyebamiji:2019:BEC

[998] O. K. Oyebamiji, D. J. Wilkinson, B. Li, P. G. Jayathilake, P. Zuliani, and T. P. Curtis. Bayesian emulation and calibration of an individualbased model of microbial communities. Journal of Computational Science, 30:194–208, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318306094.

Spadon:2019:DMS

[999] Gabriel Spadon, Bruno Brandoli, Danilo M. Eler, and Jose F. Rodrigues-Jr. Detecting multi-scale distancebased inconsistencies in cities through complex-networks. Journal of Computational Science, 30:209–222, January 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750318309955.

Anonymous:2019:EBb

[1000] Anonymous. Editorial Board. Journal of Computational Science, 31:ii, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750319302686.

Anonymous:2019:PF

[1001] Anonymous. Pages 1–184 (February 2019). Journal of Computational Science, 31:1–184, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Tanasescu:2019:FSV

[1002] Andrei Tănăsescu and Pantelimon George Popescu. A fast singular value decomposition algorithm of general k-tridiagonal matrices. Journal of Computational Science, 31: 1–5, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318311931.

Abouaissa:2019:CRM

[1003] Hassane Aboua¨ ıssa and Samira Chouraqui. On the control of robot manipulator: a model-free approach. Journal of Computational Science, 31:6–16, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318305805.

Elhosseini:2019:BRS

[1004] Mostafa A. Elhosseini, Amira Y.

Haikal, Mahmoud Badawy, and Nour Khashan. Biped robot stability based on an A–C parametric Whale Optimization Algorithm. Journal of Computational Science, 31:17–32, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318308524.

Khaluf:2019:LAS

[1005] Yara Khaluf, Seppe Vanhee, and Pieter Simoens. Local ant system for allocating robot swarms to time-constrained tasks. Journal of Computational Science, 31:33–44, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310470.

Liu:2019:SCP

[1006] Shigang Liu, Yali Peng, Zengguo Sun, and Xili Wang. Self-calibration of projective camera based on trajectory basis. Journal of Computational Science, 31:45–53, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318306628.

Turkyilmazoglu:2019:ACA

[1007] Mustafa Turkyilmazoglu. Accelerating the convergence of Adomian decomposition method (ADM). Journal of Computational Science, 31: 54–59, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318310068.

Bauer:2019:LSS

[1008] S. Bauer, M. Huber, S. Ghelichkhan, M. Mohr, U. Rüde, and B. Wohlmuth. Large-scale simulation of mantle convection based on a new matrix-free approach. Journal of Computational Science, 31:60–76, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318309840.

Weskida:2019:FIS

[1009] Michał Weskida and Radosław Michalski. Finding influentials in social networks using evolutionary algorithm. Journal of Computational Science, 31:77–85, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031830070X.

Inacio:2019:PBS

[1010] Fabrício R. Inácio, Douglas G. Macharet, and Luiz Chaimowicz. PSObased strategy for the segregation of heterogeneous robotic swarms. Journal of Computational Science, 31: 86–94, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318308500.

Khaluf:2019:CSE

[1011] Yara Khaluf and Pieter Simoens. Collective sampling of environmental features under limited sampling budget. Journal of Computational Science, 31:95–110, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310275.

Mabrouk:2019:ISP

[1012] Emad Mabrouk, Ahmed Ayman, Yara Raslan, and Abdel-Rahman Hedar. Immune system programming for medical image segmentation. Journal of Computational Science, 31: 111–125, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318311268.

Barabasz:2019:SMO

[1013] Barbara Barabasz, Stephen Barrett, Leszek Siwik, Marcin Łoś, Krzysztof Podsiadło, and Maciej Woźniak. Speeding up multi-objective optimization of liquid fossil fuel reserve exploitation with parallel hybrid memory integration. Journal of Computational Science, 31:126–136, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318308305.

Sturdy:2019:UQC

[1014] Jacob Sturdy, Johannes Kløve Kjernlie, Hallvard Moian Nydal, Vinzenz G. Eck, and Leif R. Hellevik. Uncertainty quantification of computational coronary stenosis assessment and model based mitigation of image resolution limitations. Journal of Computational Science, 31:137–150, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031830351X.

Pietak:2019:SPD

[1015] Kamil Piȩtak, Dominik Żurek, Marcin Pietroń, Andrzej Dymara, and Marek Kisiel-Dorohinicki. Striving for performance of discrete optimisation via memetic agent-based systems in a hybrid CPU/GPU environment. Journal of Computational Science, 31: 151–162, February 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318307774.

Marrinan:2019:PSB

[1016] Thomas Marrinan, Greg Eisenhauer, Matthew Wolf, Joseph A. Insley, Silvio Rizzi, and Michael E. Papka. Parallel streaming between heterogeneous HPC resources for real-time analysis. Journal of Computational Science, 31:163–171, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313686.

Feinauer:2019:MUW

[1017] Julian Feinauer, Simon Hein, Stephan Rave, Sebastian Schmidt, Daniel Westhoff, Jochen Zausch, Oleg Iliev, Arnulf Latz, Mario Ohlberger, and Volker Schmidt. MULTIBAT: Unified workflow for fast electrochemical 3D simulations of lithium-ion cells combining virtual stochastic microstructures, electrochemical degradation models and model order reduction. Journal of Computational Science, 31:172–184, February 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317307627.

Anonymous:2019:EBc

[1018] Anonymous. Editorial Board. Journal of Computational Science, 32:ii, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750319303825.

Anonymous:2019:PMa

[1019] Anonymous. Pages 1–148 (March 2019). Journal of Computational Science, 32:1–148, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Epanchintsev:2019:SSW

[1020] Timofei Epanchintsev, Sergei Pravdin, and Alexander Panfilov. Simulation of spiral wave superseding in the Luo–Rudy anisotropic model of cardiac tissue with circular-shaped fibres. Journal of Computational Science, 32:1–11, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310755.

Yabe:2019:IIH

[1021] Takahiro Yabe and Satish V. Ukkusuri. Integrating information from heterogeneous networks on social media to predict post-disaster returning behavior. Journal of Computational Science, 32:12–20, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318305532.

Chew:2019:FEU

[1022] Alvin Wei Ze Chew and Adrian Wing-Keung Law. Feature engineering using homogenization theory with multiscale perturbation analysis for supervised model-based learning of physical clogging condition in seepage filters. Journal of Computational Science, 32:21–35, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310251.

Byeon:2019:AID

[1023] Wonmin Byeon, Manuel Domínguez-Rodrigo, Georgios Arampatzis, Enrique Baquedano, José Yravedra, Miguel Angel Maté-González, and Petros Koumoutsakos. Automated identification and deep classification of cut marks on bones and its paleoanthropological implications. Journal of Computational Science, 32:36–43, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310226.

Fereshtian:2019:RAP

[1024] Ali Fereshtian, Reza Mollapourasl, and Florin Avram. RBF approximation by partition of unity for valuation of options under exponential Lévy processes. Journal of Computational Science, 32:44–55, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310883.

Tamaki:2019:EPM

[1025] Ryoji Tamaki and Masashi Yamakawa.

Effect of plate mounted between two wires in electric arc spraying. Journal of Computational Science, 32:56–67, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318310160.

Was:2019:SIC

[1026] Jarosław Wa̧s and Paweł Topa. Special issue on complex collective systems. Journal of Computational Science, 32:68–69, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319302649.

Bujas:2019:HPC

[1027] Jakub Bujas, Dawid Dworak, Wojciech Turek, and Aleksander Byrski. High-performance computing framework with desynchronized information propagation for large-scale simulations. Journal of Computational Science, 32:70–86, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303776.

Mitsopoulou:2019:SGM

[1028] Martha Mitsopoulou, Nikolaos I. Dourvas, Georgios Ch. Sirakoulis, and Katsuhiro Nishinari. Spatial games and memory effects on crowd evacuation behavior with Cellular Automata. Journal of Computational Science, 32:87–98, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303272.

Krbalek:2019:TFM

[1029] Milan Krbálek, Jiří Apeltauer, and František Šeba. Traffic flow merging — statistical and numerical modeling of microstructure. Journal of Computational Science, 32:99–105, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303259.

Komosinski:2019:CTB

[1030] Maciej Komosinski and Konrad Miazga. Comparison of the tournamentbased convection selection with the island model in evolutionary algorithms. Journal of Computational Science, 32:106–114, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318304216.

DeRango:2019:FMG

[1031] Alessio De Rango, Davide Spataro, William Spataro, and Donato D’Ambrosio. A first multi-GPU/multi-node implementation of the open computing abstraction layer. Journal of Computational Science, 32: 115–124, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318303922.

Gerakakis:2019:AFC

[1032] Ioannis Gerakakis, Prodromos Gavriilidis, Nikolaos I. Dourvas, Ioakeim G. Georgoudas, Giuseppe A. Trunfio, and Georgios Ch. Sirakoulis. Accelerating fuzzy cellular automata for modeling crowd dynamics. Journal of Computational Science, 32: 125–140, March 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318304022.

Zonnchen:2019:PGS

[1033] Benedikt Zönnchen and Gerta Köster.

A parallel generator for sparse unstructured meshes to solve the eikonal equation. Journal of Computational Science, 32:141–147, March 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318303193.

Anonymous:2019:EBd

[1034] Anonymous. Editorial Board. Journal of Computational Science, 33:ii, April 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750319305381.

Anonymous:2019:PA

[1035] Anonymous. Pages 1–112 (April 2019). Journal of Computational Science, 33: 1–112, April 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Thorimbert:2019:CLB

[1036] Yann Thorimbert, Jonas Lätt, and Bastien Chopard. Coupling of lattice Boltzmann shallow water model with lattice Boltzmann free-surface model. Journal of Computational Science, 33:1–10, April 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317314096.

Peng:2019:LMV

[1037] Yali Peng, Shigang Liu, Yongqian Qiang, XiaoJun Wu, and Ling Hong. A local mean and variance active contour model for biomedical image segmentation. Journal of Computational Science, 33:11–19, April 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031830663X.

Carratala-Saez:2019:ENT

[1038] Roc´ ıo Carratalá-Sáez, Sven Christophersen, José I. Aliaga, Vicenç Beltran, Steffen Börm, and Enrique S. Quintana-Ort´ ı. Exploiting nested task-parallelism in the H-LU factorization. Journal of Computational Science, 33:20–33, April 2019. ISSN 1877-7503 (print), 1877-7511 (electronic).

Pawlowski:2019:MDM

[1039] Filip Pawłowski, Bora Uçar, and Albert-Jan Yzelman. A multidimensional Morton-ordered block storage for mode-oblivious tensor computations. Journal of Computational Science, 33:34–44, April 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031831130X.

Cao:2019:MLP

[1040] Cheng Cao, Hong-Quan Chen, Jia-Le Zhang, and Sheng-Guan Xu. A multi-layered point reordering study of GPU-based meshless method for compressible flow simulations. Journal of Computational Science, 33:45–60, April 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318312961.

Forghani-elahabad:2019:AAR

[1041] Majid Forghani-elahabad and Nelson Kagan. An approximate approach for reliability evaluation of a multistate flow network in terms of minimal cuts. Journal of Computational Science, 33:61–67, April 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318306422.

Scheidegger:2019:MLH

[1042] Simon Scheidegger and Ilias Bilionis.

Machine learning for high-dimensional dynamic stochastic economies. Journal of Computational Science, 33:68–82, April 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750318306161.

Malinowski:2019:MAL

[1043] Artur Malinowski and Paweł Czarnul. Multi-agent large-scale parallel crowd simulation with NVRAM-based distributed cache. Journal of Computational Science, 33:83–94, April 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318307221.

Jain:2019:RAF

[1044] Shashi Jain, Álvaro Leitao, and Cornelis W. Oosterlee. Rolling Adjoints: Fast Greeks along Monte Carlo scenarios for early-exercise options. Journal of Computational Science, 33:95–112, April 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318312547.

Anonymous:2019:EBe

[1045] Anonymous. Editorial Board. Journal of Computational Science, 34:ii, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750319306088.

Anonymous:2019:PMb

[1046] Anonymous. Pages 1–120 (May 2019). Journal of Computational Science, 34: 1–120, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Goel:2019:VRP

[1047] Rajeev Goel, Raman Maini, and Sandhya Bansal. Vehicle routing problem with time windows having stochastic customers demands and stochastic service times: Modelling and solution. Journal of Computational Science, 34:1–10, May 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318308810.

Delkhosh:2019:GPM

[1048] Mehdi Delkhosh and Kourosh Parand.

Generalized pseudospectral method: Theory and applications. Journal of Computational Science, 34:11–32, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750318304678.

Jain:2019:MOS

[1049] Upma Jain, Ritu Tiwari, and W. Wilfred Godfrey. Multiple odor source localization using diverse-PSO and group-based strategies in an unknown environment. Journal of Computational Science, 34:33–47, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318308457.

Sawicki:2019:UCM

[1050] Jakub Sawicki, Marcin Łoś, Maciej Smołka, and Julen Alvarez-Aramberri. Using Covariance Matrix Adaptation Evolutionary Strategy to boost the search accuracy in hierarchic memetic computations. Journal of Computational Science, 34:48–54, May 2019.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318307233.

Ahmmed:2019:OSC

[1051] Mohammad Shakil Ahmmed and Nazmul Huda. An open-source CFD model for computing thermal effect in the context of laser-induced semiconductor processing in photovoltaic applications. Journal of Computational Science, 34:55–65, May 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318311190.

Tan:2019:DDP

[1052] Sing Kuang Tan, Nan Hu, and Wentong Cai. A data-driven path planning model for crowd capacity analysis. Journal of Computational Science, 34:66–79, May 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318309712.

Innocente:2019:SOS

[1053] Mauro S. Innocente and Paolo Grasso. Self-organising swarms of firefighting drones: Harnessing the power of collective intelligence in decentralised multirobot systems. Journal of Computational Science, 34:80–101, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318310238.

Bowley:2019:AAC

[1054] Connor Bowley, Marshall Mattingly, Andrew Barnas, Susan Ellis-Felege, and Travis Desell. An analysis of altitude, citizen science and a convolutional neural network feedback loop on object detection in Unmanned Aerial Systems. Journal of Computational Science, 34:102–116, May 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310287.

Kovalchuk:2019:SID

[1055] Sergey V. Kovalchuk, Valeria V.

Krzhizhanovskaya, Yong Shi, Haohuan Fua, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. Science at the intersection of data, modelling, and computation. Journal of Computational Science, 34:117–119, May 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319305253.

Anonymous:2019:EBf

[1056] Anonymous. Editorial Board. Journal of Computational Science, 35:ii, July 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750319307847.

Anonymous:2019:PJb

[1057] Anonymous. Pages 1–112 (July 2019).

Journal of Computational Science, 35: 1–112, July 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Prajapati:2019:RRE

[1058] Gend Lal Prajapati and Rekha Saha. REEDS: relevance and enhanced entropy based Dempster Shafer approach for next word prediction using language model. Journal of Computational Science, 35:1–11, July 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318307749.

Rekik:2019:MAS

[1059] Ines Rekik and Sabeur Elkosantini. A multi agent system for the online container stacking in seaport terminals. Journal of Computational Science, 35: 12–24, July 2019. ISSN 1877-7503 (print), 1877-7511 (electronic).

Tour:2019:HOR

[1060] Geraldine Tour, Nawdha Thakoor, Désiré Yannick Tangman, and Muddun Bhuruth. A high-order RBF-FD method for option pricing under regime-switching stochastic volatility models with jumps. Journal of Computational Science, 35:25–43, 2019. ISSN 1877-7503 (print), 1877-7511 (electronic).

Figueiras:2019:QPB

[1061] Edgar Figueiras, David Olivieri, Angel Paredes, and Humberto Michinel. QMBlender: Particle-based visualization of 3D quantum wave function dynamics. Journal of Computational Science, 35:44–56, July 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031831278X.

Yang:2019:ABS

[1062] Han Yang, Tao Chen, and Nan jing Huang. An adaptive bird swarm algorithm with irregular random flight and its application. Journal of Computational Science, 35:57–65, July 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318308056.

Zhao:2019:EMC

[1063] Yuqing Zhao, Guangmei Liu, and Zhongguo Zhou. An efficient mass-conserved domain decomposition scheme for the nonlinear parabolic problem. Journal of Computational Science, 35:66–79, 2019. ISSN 1877-7503 (print), 1877-7511 (electronic).

Nikishova:2019:SIU

[1064] Anna Nikishova and Alfons G. Hoekstra. Semi-intrusive uncertainty propagation for multiscale models. Journal of Computational Science, 35:80–90, July 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318313711.

Janczykowski:2019:LSU

[1065] Michal Janczykowski, Wojciech Turek, Maciej Malawski, and Aleksander Byrski. Large-scale urban traffic simulation with Scala and high-performance computing system. Journal of Computational Science, 35:91–101, July 2019. ISSN 1877-7503 (print), 1877-7511 (electronic).

Zapata:2019:AVT

[1066] Angel Augusto Agudelo Zapata, Eduardo Giraldo Suarez, and Jairo Alberto Villegas Florez. Application of VRP techniques to the allocation of resources in an electric power distribution system. Journal of Computational Science, 35:102–109, July 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031830259X.

Carratala-Saez:2019:EEN

[1067] Roc´ ıo Carratalá-Sáez, Sven Christophersen, José I. Aliaga, Vicenç Beltran, Steffen Börm, and Enrique S. Quintana-Ort´ ı. Erratum to “Exploiting nested task-parallelism in the H-LU factorization” [J. Comput. Sci. 33 (2019) 20–33]. Journal of Computational Science, 35:110, 2019. ISSN 1877-7503 (print), 1877-7511 (electronic).

Calhoun:2019:LSR

[1068] S. Patrick Calhoun, David Akin, Brett Zimmerman, and Henry Neeman. Large scale research data archiving: Training for an inconvenient technology. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316301132.

Pileggi:2019:WS

[1069] Salvatore F. Pileggi. Web of similarity. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303842.

Tchernykh:2019:TUU

[1070] Andrei Tchernykh, Uwe Schwiegelsohn, El ghazali Talbi, and Mikhail Babenko. Towards understanding uncertainty in cloud computing with risks of confidentiality, integrity, and availability. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750316303878.

Aluja-Banet:2019:MMV

[1071] Tomàs Aluja-Banet, Maria-Ribera Sancho, and Ivan Vukic. Measuring motivation from the Virtual Learning Environment in secondary education. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317302727.

Garay:2019:VVB

[1072] Godofredo R. Garay, Andrei Tchernykh, Alexander Yu. Drozdov, Sergey N. Garichev, Sergio Nesmachnow, and Moisés Torres-Martinez. Visualization of VHDL-based simulations as a pedagogical tool for supporting computer science education. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031730385X.

Prims:2019:FAS

[1073] Oriol Tintó Prims, Miguel Castrillo, Mario C. Acosta, Oriol Mula-Valls, Alicia Sanchez Lorente, Kim Serradell, Ana Cortés, and Francisco J. Doblas-Reyes. Finding, analysing and solving MPI communication bottlenecks in Earth System models. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318304150.

Karbovskii:2019:IDO

[1074] Vladislav Karbovskii, Oksana Severiukhina, Ivan Derevitskii, Daniil Voloshin, Alva Presbitero, and Michael Lees. The impact of different obstacles on crowd dynamics. Journal of Computational Science, 36: ??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317312140.

Couturier:2019:RAP

[1075] Raphaël Couturier, Peter Strazdins, and Laurence T. Yang. Recent advances in parallel techniques for scientific computing. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319304326.

Sailunaz:2019:ESA

[1076] Kashfia Sailunaz and Reda Alhajj. Emotion and sentiment analysis from Twitter text. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318311037.

vantHof:2019:SPF

[1077] Bas van ’t Hof and Mathea J. Vuik.

Symmetry-preserving finite-difference discretizations of arbitrary order on structured curvilinear staggered grids. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319300791.

Garcia-Magarino:2019:AMA

[1078] Iván García-Magariño, Inmaculada Plaza, and Filippo Neri. ABS-MindBurnout: an agent-based simulator of the effects of mindfulnessbased interventions on job burnout. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318308883.

Raman:2019:CVN

[1079] Sundaresan Raman, Shantanu Singh, Thierry Pécot, Enrico Caserta, Kun Huang, Jens Rittscher, Gustavo Leone, and Raghu Machiraju. Capturing variations in nuclear phenotypes. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318307580.

Defo:2019:PBI

[1080] Rodrick Kuate Defo, Richard Wang, and M. Manjunathaiah. Parallel BFS implementing optimized decomposition of space and kMC simulations for diffusion of vacancies for quantum storage. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318304320.

Vijay:2019:QGW

[1081] Rahul Kumar Vijay and Satyasai Jagannath Nanda. A Quantum Grey Wolf optimizer based declustering model for analysis of earthquake catalogs in an ergodic framework. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318312663.

Leblond:2019:SIR

[1082] Isabelle Leblond, Sébastien Tauvry, and Marc Pinto. Sonar image registration for swarm AUVs navigation: Results from SWARMs project. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750318307786.

Shrivastava:2019:SBC

[1083] Abhijeet Shrivastava, Ajaya Kumar Tripathy, and Pronob Kumar Dalal. A SVM-based classification approach for obsessive compulsive disorder by oxidative stress biomarkers. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318309669.

Li:2019:AUO

[1084] Yu-Sheng Li and Ming-Liang Qi. An approach for understanding offender modus operandi to detect serial robbery crimes. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303412.

Anonymous:2019:EBg

[1085] Anonymous. Editorial Board. Journal of Computational Science, 36: ??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319310075.

Anonymous:2019:S

[1086] Anonymous. September 2019. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Xu:2019:DCL

[1087] Wenqing (William) Xu and Mark Stalzer. Deriving compact laws based on algebraic formulation of a data set. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318307609.

Safaei:2019:EBH

[1088] F. Safaei, S. Tabrizchi, A. H. Hadian Rasanan, and M. Zare. An energybased heterogeneity measure for quantifying structural irregularity in complex networks. Journal of Computational Science, 36:??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750318310469.

Dang:2019:NFD

[1089] Quang A. Dang and Manh Tuan Hoang. Nonstandard finite difference schemes for a general predator–prey system. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319304090.

Borovykh:2019:GFC

[1090] Anastasia Borovykh, Cornelis W. Oosterlee, and Sander M. Bohté. Generalization in fully-connected neural networks for time series forecasting. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319301838.

Teng:2019:AIA

[1091] L. Teng and A. Clevenhaus. Accelerated implementation of the ADI schemes for the Heston model with stochastic correlation. Journal of Computational Science, 36: ??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319304752.

Thakoor:2019:NLR

[1092] Nawdha Thakoor and Muddun Bhuruth. New local radial point interpolation-FD methods for solving fractional diffusion and damped-wave problems. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303503.

Graillat:2019:ATF

[1093] Stef Graillat, Fabienne Jézéquel, Romain Picot, François Févotte, and Bruno Lathuilière. Auto-tuning for floating-point precision with Discrete Stochastic Arithmetic. Journal of Computational Science, 36: ??, September 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318309475.

Iyer:2019:RLS

[1094] Chander Iyer, Haim Avron, Georgios Kollias, Yves Ineichen, Christopher Carothers, and Petros Drineas. A randomized least squares solver for terabyte-sized dense overdetermined systems. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316301508.

Jammy:2019:PEE

[1095] Satya P. Jammy, Christian T. Jacobs, and Neil D. Sandham. Performance evaluation of explicit finite difference algorithms with varying amounts of computational and memory intensity. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316302708.

Anzt:2019:FGB

[1096] Hartwig Anzt, Jack Dongarra, and Enrique S. Quintana-Ort´ ı. Fine-grained bit-flip protection for relaxation methods. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316303891.

Nino-Ruiz:2019:PIE

[1097] Elias D. Nino-Ruiz, Adrian Sandu, and Xinwei Deng. A parallel implementation of the ensemble Kalman filter based on modified Cholesky decomposition. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031730399X.

Deng:2019:ISA

[1098] Quanling Deng, Vladimir Puzyrev, and Victor Calo. Isogeometric spectral approximation for elliptic differential operators. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318301327.

Jain:2019:MBP

[1099] Gatij Jain, Gaurav Yadav, Dhruv Prakash, Anupam Shukla, and Ritu Tiwari. MVO-based path planning scheme with coordination of UAVs in 3-D environment. Journal of Computational Science, 37:??, October 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318310263.

Anonymous:2019:EBh

[1100] Anonymous. Editorial Board. Journal of Computational Science, 37:??, October 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750319311147.

Anonymous:2019:O

[1101] Anonymous. October 2019. Journal of Computational Science, 37:??, October 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ehrhardt:2019:SBM

[1102] Matthias Ehrhardt, Ján Gašper, and Soňa Kilianová. SIR-based mathematical modeling of infectious diseases with vaccination and waning immunity. Journal of Computational Science, 37:??, October 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303163.

Falgout:2019:PTA

[1103] Robert D. Falgout, Matthieu Lecouvez, and Carol S. Woodward. A parallel-in-time algorithm for variable step multistep methods. Journal of Computational Science, 37:??, October 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319304351.

Coudiere:2019:DDS

[1104] Yves Coudière and Rodolphe Turpault. A domain decomposition strategy for a very high-order finite volumes scheme applied to cardiac electrophysiology. Journal of Computational Science, 37:??, October 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303862.

Soleymani:2019:PFE

[1105] Fazlollah Soleymani and Andrey Itkin. Pricing foreign exchange options under stochastic volatility and interest rates using an RBF-FD method. Journal of Computational Science, 37:??, October 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319304508.

Li:2019:CMF

[1106] Sha Li, Bastien Chopard, and Jonas Latt. Continuum model for flow diverting stents in 3D patient-specific simulation of intracranial aneurysms. Journal of Computational Science, 38:??, November 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319310269.

Zhuge:2019:AAB

[1107] Chengxiang Zhuge, Chunfu Shao, and Xiong Yang. Agent- and activitybased large-scale simulation of enroute travel, enroute refuelling and parking behaviours in Beijing, China. Journal of Computational Science, 38:??, November 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318309785.

Anonymous:2019:EBi

[1108] Anonymous. Editorial Board. Journal of Computational Science, 38: ??, November 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319312050.

Anonymous:2019:N

[1109] Anonymous. November 2019. Journal of Computational Science, 38:??, November 2019. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Osaba:2020:SCS

[1110] Eneko Osaba, Javier Del Ser, Andres Iglesias, and Xin-She Yang. Soft computing for swarm robotics: New trends and applications. Journal of Computational Science, 39:??, January 2020.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031931172X.

Anonymous:2020:EBa

[1111] Anonymous. Editorial Board. Journal of Computational Science, 39:??, January 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750320300685.

Anonymous:2020:Ja

[1112] Anonymous. January 2020. Journal of Computational Science, 39:??, January 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Kaba:2020:OKF

[1113] Aziz Kaba and Emre Kıyak. Optimizing a Kalman filter with an evolutionary algorithm for nonlinear quadrotor attitude dynamics. Journal of Computational Science, 39:??, January 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319302376.

Maji:2020:ISI

[1114] Giridhar Maji. Influential spreaders identification in complex networks with potential edge weight based k-shell degree neighborhood method. Journal of Computational Science, 39:??, January 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319304181.

Filelis-Papadopoulos:2020:TSO

[1115] Christos K. Filelis-Papadopoulos, Patricia Takako Endo, Malika Bendechache, Sergej Svorobej, Konstantinos M. Giannoutakis, George A. Gravvanis, Dimitrios Tzovaras, James Byrne, and Theo Lynn. Towards simulation and optimization of cache placement on large virtual content distribution networks. Journal of Computational Science, 39:??, January 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319302406.

Benamour:2020:VPL

[1116] M. Benamour, E. Liberge, and C. Béghein. A volume penalization lattice Boltzmann method for simulating flows in the presence of obstacles. Journal of Computational Science, 39:??, January 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303771.

Stojanovic:2020:MSS

[1117] Boban Stojanovic, Marina Svicevic, Ana Kaplarevic-Malisic, Richard J. Gilbert, and Srboljub M. Mijailovich. Multi-scale striated muscle contraction model linking sarcomere length-dependent cross-bridge kinetics to macroscopic deformation. Journal of Computational Science, 39:??, January 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319301334.

delaTorre:2020:PVS

[1118] Juan Carlos de la Torre, Renzo Massobrio, Patricia Ruiz, Sergio Nesmachnow, and Bernabé Dorronsoro. Parallel virtual savant for the heterogeneous computing scheduling problem. Journal of Computational Science, 39:??, January 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319303321.

Georgiev:2020:CRT

[1119] Slavi G. Georgiev and Lubin G. Vulkov. Computational recovery of time-dependent volatility from integral observations in option pricing. Journal of Computational Science, 39:??, January 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319303709.

LaTorre:2020:MCU

[1120] Antonio LaTorre, Man Ting Kwong, Julián A. Garc´ ıa-Grajales, Riyi Shi, Antoine Jérusalem, and José-Mar´ ıa Peña. Model calibration using a parallel differential evolution algorithm in computational neuroscience: Simulation of stretch induced nerve deficit. Journal of Computational Science, 39:??, January 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319302212.

Al-Sawwa:2020:PEC

[1121] Jamil Al-Sawwa and Simone A. Ludwig. Performance evaluation of a costsensitive differential evolution classifier using spark — imbalanced binary classification. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319302236.

Dziurzanski:2020:SDE

[1122] Piotr Dziurzanski, Shuai Zhao, Michal Przewozniczek, Marcin Komarnicki, and Leandro Soares Indrusiak. Scalable distributed evolutionary algorithm orchestration using Docker containers. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303333.

Lakouari:2020:MSC

[1123] N. Lakouari, O. Oubram, A. Bassam, Saul E. Pomares Hernandez, R. Marzoug, and H. Ez-Zahraouy. Modeling and simulation of CO _{2} emissions in roundabout intersection. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319304843.

Anonymous:2020:EBb

[1124] Anonymous. Editorial Board. Journal of Computational Science, 40: ??, February 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S1877750320302787.

Anonymous:2020:F

[1125] Anonymous. February 2020. Journal of Computational Science, 40:??, February 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Macnamara:2020:CMS

[1126] Cicely K. Macnamara, Alfonso Caiazzo, Ignacio Ramis-Conde, and Mark A. J. Chaplain. Computational modelling and simulation of cancer growth and migration within a 3D heterogeneous tissue: the effects of fibre and vascular structure. Journal of Computational Science, 40:??, February 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319305691.

Goncalves:2020:PEC

[1127] Francisco Gonçalves, Sergio Santander-Jiménez, Leonel Sousa, José M. Granado-Criado, and Aleksandar Ilic. Parallel evolutionary computation for multiobjective gene interaction analysis. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031930794X.

Aristov:2020:KMS

[1128] V. V. Aristov, O. V. Ilyin, and O. A. Rogozin. Kinetic multiscale scheme based on the discretevelocity and lattice-Boltzmann methods. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318313772.

Li:2020:EHA

[1129] Keran Li and Wenyuan Liao. An efficient and high accuracy finitedifference scheme for the acoustic wave equation in 3D heterogeneous media. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303588.

Ngomade:2020:DPB

[1130] Armel Nkonjoh Ngomade, Jean Frédéric Myoupo, and Vianney Kengne Tchendji. A dominant point-based parallel algorithm that finds all longest common subsequences for a constrained-MLCS problem. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319301942.

Chew:2020:HTM

[1131] Alvin Wei Ze Chew and Adrian Wing-Keung Law. Homogenization theory with multiscale perturbation analysis for supervised learning of complex adsorption-desorption process in porous-media systems. Journal of Computational Science, 40:??, February 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319303515.

Davoudabadi:2020:IWR

[1132] Reza Davoudabadi, S. Meysam Mousavi, and Ebrahim Sharifi. An integrated weighting and ranking model based on entropy, DEA and PCA considering two aggregation approaches for resilient supplier selection problem. Journal of Computational Science, 40:??, February 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319305113.

Verma:2020:MSE

[1133] Kevin Verma, Hui Cao, Prithvi Mandapalli, and Robert Wille. Modeling and simulation of electrophoretic deposition coatings. Journal of Computational Science, 41:??, March 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319308506.

Villagra:2020:BUT

[1134] Andrea Villagra, Enrique Alba, and Gabriel Luque. A better understanding on traffic light scheduling: New cellular GAs and new in-depth analysis of solutions. Journal of Computational Science, 41:??, March 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319302169.

Anonymous:2020:EBc

[1135] Anonymous. Editorial Board. Journal of Computational Science, 41:??, March 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750320304221.

Anonymous:2020:Ma

[1136] Anonymous. March 2020. Journal of Computational Science, 41:??, March 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Isiet:2020:SAC

[1137] Mewael Isiet and Mohamed Gadala.

Sensitivity analysis of control parameters in particle swarm optimization. Journal of Computational Science, 41:??, March 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750319300067.

Rodriguez-Baena:2020:ILB

[1138] Domingo S. Rodriguez-Baena, Francisco A. Gomez-Vela, Miguel Garc´ ıa-Torres, Federico Divina, Carlos D. Barranco, Norberto Daz-Diaz, Manuel Jimenez, and Gema Montalvo. Identifying livestock behavior patterns based on accelerometer dataset. Journal of Computational Science, 41:??, March 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750318311359.

Inostroza-Ponta:2020:EHS

[1139] Mario Inostroza-Ponta, Márcio Dorn, Iván Escobar, Leonardo de Lima Correa, Erika Rosas, Nicolás Hidalgo, and Mauricio Marin. Exploring the high selectivity of 3-D protein structures using distributed memetic algorithms. Journal of Computational Science, 41:??, March 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319302170.

He:2020:NCI

[1140] Daobing He and Xiaoyang Liu. Novel competitive information propagation macro mathematical model in online social network. Journal of Computational Science, 41:??, March 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319308294.

Rosalie:2020:BOS

[1141] Martin Rosalie, Emmanuel Kieffer, Matthias R. Brust, Grégoire Danoy, and Pascal Bouvry. Bayesian optimisation to select Rössler system parameters used in Chaotic Ant Colony Optimisation for Coverage. Journal of Computational Science, 41:??, March 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775031930242X.

Lipitakis:2020:NPA

[1142] Anastasia-Dimitra Lipitakis, Christos K. Filelis-Papadopoulos, George A. Gravvanis, and Dimosthenis Anagnostopoulos. A note on parallel approximate pseudoinverse matrix techniques for solving linear least squares problems. Journal of Computational Science, 41:??, March 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750317313510.

Ingham-Dempster:2020:CMC

[1143] Timothy A. Ingham-Dempster, Ria Rosser, Bernard M. Corfe, and Dawn C. Walker. From cell to multicrypt: agent-based models of the human colon suggests novel processes of Field cancerisation. Journal of Computational Science, 41:??, March 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319306465.

Xiang:2020:ELS

[1144] Wentao Xiang, Ahmad Karfoul, Chunfeng Yang, Huazhong Shu, and Régine Le Bouquin Jeannès. An exact line search scheme to accelerate the EM algorithm: Application to Gaussian mixture models identification. Journal of Computational Science, 41:??, March 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318313929.

Kadupitiya:2020:MLS

[1145] J. C. S. Kadupitiya, Fanbo Sun, Geoffrey Fox, and Vikram Jadhao. Machine learning surrogates for molecular dynamics simulations of soft materials. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319310609.

Randall:2020:ITO

[1146] M. Randall, J. Montgomery, and A. Lewis. An introduction to temporal optimisation using a water management problem. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319310993.

Anonymous:2020:EBd

[1147] Anonymous. Editorial Board. Journal of Computational Science, 42:??, April 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750320304373.

Anonymous:2020:A

[1148] Anonymous. April 2020. Journal of Computational Science, 42:??, April 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Hemami:2020:USS

[1149] Mohammad Hemami, Jamal Amani Rad, and Kourosh Parand. The use of space-splitting RBF-FD technique to simulate the controlled synchronization of neural networks arising from brain activity modeling in epileptic seizures. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319309317.

Dur:2020:WCG

[1150] Tolga Hasan Dur, Rossella Arcucci, Laetitia Mottet, Miguel Molina Solana, Christopher Pain, and Yi-Ke Guo. Weak constraint Gaussian processes for optimal sensor placement. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319310543.

Garcia-Garcia:2020:POS

[1151] Adrian Garcia-Garcia, Juan Carlos Saez, José Luis Risco-Martin, and Manuel Prieto-Matias. PBBCache: an open-source parallel simulator for rapid prototyping and evaluation of cache-partitioning and cache-clustering policies. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319303229.

Alekseev:2020:UQE

[1152] A. K. Alekseev, A. E. Bondarev, and A. E. Kuvshinnikov. On uncertainty quantification via the ensemble of independent numerical solutions. Journal of Computational Science, 42:??, April 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319310695.

Zamuda:2020:ODD

[1153] Aleš Zamuda and Elena Lloret. Optimizing data-driven models for summarization as parallel tasks. Journal of Computational Science, 42:??, April 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750319302194.

Shi:2020:HIV

[1154] Harvey Shi, Jeff Ames, and Amanda Randles. Harvis: an interactive virtual reality tool for hemodynamic modification and simulation. Journal of Computational Science, 43:??, May 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319309214.

Curado:2020:ACC

[1155] Manuel Curado, Leandro Tortosa, Jose F. Vicent, and Gevorg Yeghikyan. Analysis and comparison of centrality measures applied to urban networks with data. Journal of Computational Science, 43:??, May 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304282.

Anonymous:2020:EBe

[1156] Anonymous. Editorial Board. Journal of Computational Science, 43:??, May 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750320304518.

Anonymous:2020:Mb

[1157] Anonymous. May 2020. Journal of Computational Science, 43:??, May 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Lai:2020:SBI

[1158] Yingying Lai, Andrew Golightly, and Richard J. Boys. Sequential Bayesian inference for spatio-temporal models of temperature and humidity data. Journal of Computational Science, 43:??, May 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320304269.

deOliveira:2020:CMS

[1159] Alan Delgado de Oliveira and Tiago Pascoal Filomena. Compacting multistage stochastic programming models through a new implicit extensive framework. Journal of Computational Science, 43:??, May 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304294.

Wang:2020:NLS

[1160] Zhixiao Wang, Chengcheng Sun, Guan Yuan, Xiaobin Rui, and Xiaodong Yang. A neighborhood link sensitive dismantling method for social networks. Journal of Computational Science, 43:??, May 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304300.

Khan:2020:MTM

[1161] Zawar H. Khan and T. Aaron Gulliver.

A macroscopic traffic model based on transition velocities. Journal of Computational Science, 43:??, May 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304324.

Fan:2020:USE

[1162] Xiaolin Fan, Zhonghua Qiao, and Shuyu Sun. Unconditionally stable, efficient and robust numerical simulation of isothermal compositional grading by gravity. Journal of Computational Science, 43:??, May 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031931066X.

daSilva:2020:CMA

[1163] Rafael Ferreira da Silva, Henri Casanova, Anne-Cécile Orgerie, Ryan Tanaka, Ewa Deelman, and Frédéric Suter. Characterizing, modeling, and accurately simulating power and energy consumption of I/O-intensive scientific workflows. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304580.

Begy:2020:FNT

[1164] Volodimir Begy, Martin Barisits, Mario Lassnig, and Erich Schikuta. Forecasting network throughput of remote data access in computing grids. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750320304592.

Takii:2020:SDF

[1165] A. Takii, M. Yamakawa, S. Asao, and K. Tajiri. Six degrees of freedom flight simulation of tilt-rotor aircraft with nacelle conversion. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304658.

Anonymous:2020:EBf

[1166] Anonymous. Editorial Board. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750320304920.

Anonymous:2020:Jb

[1167] Anonymous. July 2020. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Toupance:2020:IMC

[1168] Pierre-Alain Toupance, Laurent Lefèvre, and Bastien Chopard. Influence measurement in a complex dynamical model: an information theoretic approach. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319307872.

Chaudhari:2020:CCD

[1169] Harshal Chaudhari and Martin Crane.

Cross-correlation dynamics and community structures of cryptocurrencies. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320304312.

Wang:2020:DAB

[1170] Hao Wang, Shunhuai Chen, and Liang Luo. A diffusion algorithm based on P systems for continuous global optimization. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031931230X.

Ames:2020:MGI

[1171] Jeff Ames, Daniel F. Puleri, Peter Balogh, John Gounley, Erik W. Draeger, and Amanda Randles. Multi-GPU immersed boundary method hemodynamics simulations. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750320304543.

Assad:2020:OCB

[1172] Luiz Paulo de Freitas Assad, Raquel Toste, Carina Stefoni Böck, Douglas Medeiros Nehme, L´ ıvia Sancho, Anderson Elias Soares, and Luiz Landau. Ocean climatology at Brazilian Equatorial Margin: a numerical approach. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304609.

Melab:2020:SBP

[1173] Nouredine Melab, Jan Gmys, Peter Korosec, and Imen Chakroun. Synergy between parallel computing, optimization and simulation. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304695.

Abdelli:2020:DVC

[1174] Abdelkrim Abdelli, Lynda Mokdad, and Youcef Hammal. Dealing with value constraints in decision making using MCDM methods. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304555.

Leonenko:2020:CPI

[1175] Vasiliy Leonenko, Sviatoslav Arzamastsev, and Georgiy Bobashev. Contact patterns and influenza outbreaks in Russian cities: a proof-of-concept study via agent-based modeling. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320304579.

Capacho:2020:SOF

[1176] John William Vásquez Capacho, Carlos Gustavo Perez Zuñiga, Yecid Alfonso Muñoz Maldonado, and Adalberto Ospino Castro. Simultaneous occurrences and false-positives analysis in discrete event dynamic systems. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320304634.

Banerjee:2020:DDO

[1177] Mahan Raj Banerjee, Rashmi Ramadugu, and Santosh Ansumali. Discrete differential operators on a class of lattices. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304737.

Krivovichev:2020:PSS

[1178] Gerasim V. Krivovichev. Parametric schemes for the simulation of the advection process in finite-difference-based single-relaxation-time lattice Boltzmann methods. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032030452X.

Montessori:2020:CLB

[1179] Andrea Montessori, Adriano Tiribocchi, Marco Lauricella, and Sauro Succi. A coupled lattice Boltzmannmultiparticle collision method for multi-resolution hydrodynamics. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750320304610.

Sweidan:2020:OFA

[1180] Zahraa Sweidan, Rania Islambouli, and Sanaa Sharafeddine. Optimized flow assignment for applications with strict reliability and latency constraints using path diversity. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304646.

Hammal:2020:FTC

[1181] Youcef Hammal, Khadidja Salah Mansour, Abdelkrim Abdelli, and Lynda Mokdad. Formal techniques for consistency checking of orchestrations of semantic Web services. Journal of Computational Science, 44:??, July 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032030466X.

Mills:2020:ISL

[1182] Lara Mills, João Janeiro, Antonio Augusto Sepp Neves, and Flávio Martins. The impact of sea level rise in the Guadiana Estuary. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304701.

Brajard:2020:CDA

[1183] Julien Brajard, Alberto Carrassi, Marc Bocquet, and Laurent Bertino. Combining data assimilation and machine learning to emulate a dynamical model from sparse and noisy observations: a case study with the Lorenz 96 model. Journal of Computational Science, 44:??, July 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304725.

Coen:2020:CMEa

[1184] Janice L. Coen, W. Schroeder, S. Conway, and L. Tarnay. Computational modeling of extreme wildland fire events: a synthesis of scientific understanding with applications to forecasting, land management, and firefighter safety. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304531.

Webster:2020:EEE

[1185] Matt Webster, Michael Breza, Clare Dixon, Michael Fisher, and Julie Mc-Cann. Exploring the effects of environmental conditions and design choices on IoT systems using formal methods. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304841.

Krzysztof:2020:SSA

[1186] Mikuta Krzysztof, Konieczny Leszek, and Roterman Irena. System to simulate the activity of living organism — construction of proteome. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750320304968.

Benkaouha:2020:SSM

[1187] Haroun Benkaouha. A stable storage in MANET: Replication or distributed storage. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032030497X.

Saeed:2020:JPR

[1188] H. Faizan Saeed, Sobia Jangsher, Moayad Aloqaily, Hassaan Khaliq Qureshi, and Jalel Ben Othman. Joint pairing and resource allocation for backhaul of small cells using NOMA. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304981.

Qian:2020:RTM

[1189] Kun Qian and Christian Claudel. Realtime mobile sensor management framework for city-scale environmental monitoring. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305068.

Anonymous:2020:EBg

[1190] Anonymous. Editorial Board. Journal of Computational Science, 45: ??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320305342.

Anonymous:2020:S

[1191] Anonymous. September 2020. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Cardesa:2020:ACA

[1192] J. I. Cardesa, L. Hascoët, and C. Airiau. Adjoint computations by algorithmic differentiation of a parallel solver for time-dependent PDEs. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750320304567.

Lima:2020:CAP

[1193] Leandro Lima and Said Sadique Adi. The chain alignment problem. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304622.

Chavez-Modena:2020:OFP

[1194] M. Chávez-Modena, A. Mart´ ınez-Cava, G. Rubio, and E. Ferrer. Optimizing free parameters in the D3Q19 multiple-relaxation lattice Boltzmann methods to simulate under-resolved turbulent flows. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304713.

Milan:2020:SKD

[1195] Felix Milan, Luca Biferale, Mauro Sbragaglia, and Federico Toschi. Sub-Kolmogorov droplet dynamics in isotropic turbulence using a multiscale lattice Boltzmann scheme. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320304798.

Valencia-Parra:2020:UCO

[1196] Álvaro Valencia-Parra, Ángel Jesús Varela-Vaca, Luisa Parody, and Mar´ ıa Teresa Gómez-López. Unleashing constraint optimisation problem solving in Big Data environments. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304816.

Fujita:2020:DEE

[1197] Kohei Fujita, Masashi Horikoshi, Tsuyoshi Ichimura, Larry Meadows, Kengo Nakajima, Muneo Hori, and Lalith Maddegedara. Development of element-by-element kernel algorithms in unstructured finite-element solvers for many-core wide-SIMD CPUs: Application to earthquake simulation. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320304750.

Alqudah:2020:GSN

[1198] Manar A. Alqudah and Noufe H. Aljahdaly. Global stability and numerical simulation of a mathematical model of stem cells therapy of HIV-1 infection. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304774.

Barraza:2020:TGL

[1199] Juan Antonio Reyes Barraza and Ralf Deiterding. Towards a generalised lattice Boltzmann method for aerodynamic simulations. Journal of Computational Science, 45: ??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775032030483X.

Kumar:2020:EDC

[1200] Dheeraj Kumar and Satish V. Ukkusuri. Enhancing demographic coverage of hurricane evacuation behavior modeling using social media. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304853.

Gawas:2020:ATL

[1201] Amitkumar S. Gawas and Dhiraj V. Patil. Axisymmetric thermal-lattice Boltzmann method for Rayleigh– Bénard convection with anisotropic thermal diffusion. Journal of Computational Science, 45:??, September 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304865.

Kolluru:2020:EBM

[1202] Praveen Kumar Kolluru, Mohammad Atif, and Santosh Ansumali. Extended BGK model for diatomic gases. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304804.

Wang:2020:TRN

[1203] Ying Wang, Baichun Hu, Shasha Feng, Jian Wang, and Fengjiao Zhang. Target recognition and network pharmacology for revealing antidiabetes mechanisms of natural product. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304877.

Su:2020:PMS

[1204] Boyi Su, Philipp Andelfinger, Jaeyoung Kwak, David Eckhoff, Henriette Cornet, Goran Marinkovic, Wentong Cai, and Alois Knoll. A passenger model for simulating boarding and alighting in spatially confined transportation scenarios. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304749.

Liang:2020:NAI

[1205] Yu Liang, Xiao-Wei Gao, Bing-Bing Xu, Qiang-Hua Zhu, and Ze-Yan Wu. A new alternating iteration strategy based on the proper orthogonal decomposition for solving large-scaled transient nonlinear heat conduction problems. Journal of Computational Science, 45:??, September 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032030507X.

JohnsonRestrepo:2020:CSA

[1206] D. Dylan Johnson Restrepo, Michael Spagat, Stijn van Weezel, Minzhang Zheng, and Neil F. Johnson. A computational science approach to understanding human conflict. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750319313456.

Wolfram:2020:WWB

[1207] Stephen Wolfram. What we’ve built is a computational language (and that’s very important!). Journal of Computational Science, 46:??, October 2020. ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304336.

Coen:2020:CMEb

[1208] Janice L. Coen, W. Schroeder, S. Conway, and L. Tarnay. Computational modeling of extreme wildland fire events: a synthesis of scientific understanding with applications to forecasting, land management, and firefighter safety. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305299.

Anonymous:2020:EBh

[1209] Anonymous. Editorial Board. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750320305482.

Restrepo:2020:CSA

[1210] D. Dylan Johnson Restrepo, Michael Spagat, Stijn van Weezel, Minzhang Zheng, and Neil F. Johnson. A computational science approach to understanding human conflict. Journal of Computational Science, 46, 2020. ISSN 1877-7503 (print), 1877-7511 (electronic).

Coveney:2020:DHA

[1211] Peter V. Coveney and Roger R. Highfield. From digital hype to analogue reality: Universal simulation beyond the quantum and exascale eras. Journal of Computational Science, 46:??, October 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320300545.

Kuai:2020:EDB

[1212] Hongzhi Kuai and Ning Zhong. The extensible data-brain model: Architecture, applications and directions. Journal of Computational Science, 46:??, October 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320300752.

Yang:2020:NIO

[1213] Xin-She Yang. Nature-inspired optimization algorithms: Challenges and open problems. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320300144.

Lu:2020:SSM

[1214] Ruqian Lu and Shengluan Hou. On semi-supervised multiple representation behavior learning. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750319313389.

Brandi:2020:USC

[1215] Giuseppe Brandi, Ruggero Gramatica, and Tiziana Di Matteo. Unveil stock correlation via a new tensor-based decomposition method. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750319312724.

Glimm:2020:CTP

[1216] James Glimm. The computation of turbulent phenomena. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750319313997.

Succi:2020:TED

[1217] Sauro Succi, Giorgio Amati, Fabio Bonaccorso, Marco Lauricella, M. Bernaschi, Andrea Montessori, and Adriano Tiribocchi. Toward exascale design of soft mesoscale materials. Journal of Computational Science, 46:??, October 2020. ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304762.

Jalan:2020:ESM

[1218] Sarika Jalan, Ajay Deep Kachhvah, and Hawoong Jeong. Explosive synchronization in multilayer dynamically dissimilar networks. Journal of Computational Science, 46:??, October 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304786.

Kovalchuk:2020:YCS

[1219] Sergey V. Kovalchuk, Valeria V.

Krzhizhanovskaya, and Peter M. A. Sloot. 20 years of computational science. Journal of Computational Science, 46:??, ???? 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304889.

Bekisz:2020:CMM

[1220] Sophie Bekisz and Liesbet Geris. Cancer modeling: From mechanistic to data-driven approaches, and from fundamental insights to clinical applications. Journal of Computational Science, 46:??, October 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320304993.

Coen:2020:CME

[1221] Janice L. Coen, W. Schroeder, S. Conway, and L. Tarnay. Computational modeling of extreme wildland fire events: a synthesis of scientific understanding with applications to forecasting, land management, and firefighter safety. Journal of Computational Science, 46, 2020. ISSN 1877-7503 (print), 1877-7511 (electronic).

Vallee:2020:NOR

[1222] S. Vallee, A. Oulamara, and W. Ramdane Cherif-Khettaf. New online reinsertion approaches for a dynamic Diala-Ride Problem. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305007.

Olivier:2020:PUG

[1223] Audrey Olivier, Dimitris G. Giovanis, B. S. Aakash, Mohit Chauhan, Lohit Vandanapu, and Michael D. Shields. UQpy: a general purpose Python package and development environment for uncertainty quantification. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305056.

Zhao:2020:CAM

[1224] Han-Tao Zhao, Xin Zhao, Lu Jiancheng, and Liu yan Xin. Cellular automata model for urban road traffic flow considering Internet of Vehicles and emergency vehicles. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305202.

Khan:2020:DNM

[1225] Asad Khan, Hafeez Ur Rehman, Usman Habib, and Umer Ijaz. Detecting N6-methyladenosine sites from RNA transcriptomes using random forest. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032030538X.

Salkovic:2020:POB

[1226] Edin Salkovic, Mostafa M. Abbas, Samir Brahim Belhaouari, Khaoula Errafii, and Halima Bensmail. OutPyR: Bayesian inference for RNA-Seq outlier detection. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305433.

Anonymous:2020:EBi

[1227] Anonymous. Editorial Board. Journal of Computational Science, 47: ??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032030569X.

Anonymous:2020:N

[1228] Anonymous. November 2020. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Dorywalski:2020:HGG

[1229] Krzysztof Dorywalski, Rüdiger Schmidt-Gründ, and Marius Grundmann. Hybrid GA-gradient method for thin films ellipsometric data evaluation. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750320305020.

Olivier:2020:UGP

[1230] Audrey Olivier, Dimitrios Giovanis, B. S. Aakash, Mohit Chauhan, Lohit Vandanapu, and Michael D. Shields. UQpy: a general purpose Python package and development environment for uncertainty quantification. Journal of Computational Science, 47, 2020. ISSN 1877-7503 (print), 1877-7511 (electronic).

Marin:2020:SMF

[1231] Oana Marin, Emil Constantinescu, and Barry Smith. A scalable matrix-free spectral element approach for unsteady PDE constrained optimization using PETSc/TAO. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305081.

Li:2020:BBB

[1232] Ning Li and Lei Wang. Bare-Bones Based Sine Cosine Algorithm for global optimization. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305196.

Jannelli:2020:NAP

[1233] Alessandra Jannelli. A novel adaptive procedure for solving fractional differential equations. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305184.

Cardoso:2020:CSI

[1234] Pedro J. S. Cardoso, João M. F.

Rodrigues, Jânio Monteiro, Roberto Lam, Valeria V. Krzhizhanovskaya, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. Computational science in the interconnected world: Selected papers from 2019 International Conference on Computational Science [Editorial]. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305214. Held

in Faro, June 12–14, 2019.

Dubois:2020:NNN

[1235] François Dubois, Benjamin Graille, and S. V. Raghurama Rao. A notion of non-negativity preserving relaxation for a mono-dimensional three velocities scheme with relative velocity. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320304828.

Xue:2020:LBS

[1236] Xiao Xue, Luca Biferale, Mauro Sbragaglia, and Federico Toschi. A lattice Boltzmann study on Brownian diffusion and friction of a particle in a confined multicomponent fluid. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319311342.

Bensaber:2020:DMA

[1237] Boucif Amar Bensaber, Caroly Gabriela Pereira Diaz, and Youssef Lahrouni. Design and modeling an Adaptive Neuro-Fuzzy Inference System (ANFIS) for the prediction of a security index in VANET. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305378.

Fouchal:2020:SPB

[1238] Hacène Fouchal. Sharing pseudonyms between Intelligent Transport System stations. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305391.

Kannaiyan:2020:CSE

[1239] Kumaran Kannaiyan. Computational study of the effect of cavity geometry on the supersonic mixing and combustion of ethylene. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032030541X.

Cuevas:2020:NMA

[1240] Erik Cuevas, Jorge Gálvez, Karla Avila, Miguel Toski, and Vahid Rafe. A new metaheuristic approach based on agent systems principles. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305421.

Salkovic:2020:OBI

[1241] Edin Salkovic, Mostafa M. Abbas, Samir Brahim Belhaouari, Khaoula Errafii, and Halima Bensmail. OutPyR: Bayesian inference for RNA-Seq outlier detection. Journal of Computational Science, 47, 2020. ISSN 1877-7503 (print), 1877-7511 (electronic).

Doan:2020:PIE

[1242] N. A. K. Doan, W. Polifke, and L. Magri. Physics-informed echo state networks. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305408.

Yan:2021:AIM

[1243] Ran Yan, Shuaian Wang, and Chuansheng Peng. An artificial intelligence model considering data imbalance for ship selection in port state control based on detention probabilities. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032030555X.

Do:2021:LME

[1244] Tu Mai Anh Do, Loïc Pottier, Silvina Caíno-Lores, Rafael Ferreira da Silva, Michel A. Cuendet, Harel Weinstein, Trilce Estrada, Michela Taufer, and Ewa Deelman. A lightweight method for evaluating in situ workflow efficiency. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305573.

Josephson:2021:AFH

[1245] Tyler R. Josephson, Paul J. Dauenhauer, Michael Tsapatsis, and J. Ilja Siepmann. Adsorption of furan, hexanoic acid, and alkanes in a hierarchical zeolite at reaction conditions: Insights from molecular simulations. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305585.

Sheikh:2021:FTH

[1246] Sophiya Sheikh, A. Nagaraju, and Mohammad Shahid. A faulttolerant hybrid resource allocation model for dynamic computational grid. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305640.

Anonymous:2021:EBa

[1247] Anonymous. Editorial Board. Journal of Computational Science, 48:??, January 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S187775032030586X.

Anonymous:2021:Ja

[1248] Anonymous. January 2021. Journal of Computational Science, 48:??, January 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Iseli:2021:PRF

[1249] Elia Iseli and Jürg Schiffmann. Prediction of the reaction forces of spiralgroove gas journal bearings by artificial neural network regression models. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305548.

Berdiyorov:2021:ETT

[1250] G. R. Berdiyorov and H. Hamoudi.

Electronic transport through molecules containing pyrimidine units: Firstprinciples calculations. Journal of Computational Science, 48:??, January 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750320305597.

Jovanovic:2021:MAP

[1251] Raka Jovanovic, Sertac Bayhan, and Islam Safak Bayram. A multiobjective analysis of the potential of scheduling electrical vehicle charging for flattening the duck curve. Journal of Computational Science, 48:??, January 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305603.

Bazurto-Gomez:2021:MSP

[1252] Nychol Bazurto-Gómez, Carlos Alberto Mart´ ınez-Morales, and Helbert Eduardo Espitia-Cuchango. Multiple swarm particles simulation algorithm applied to coffee berry borer proliferation. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305615.

Wang:2021:FMB

[1253] Jing Jing Wang, Hong Ji Meng, Jian Yang, and Zhi Xie. A fast method based on GPU for solidification structure simulation of continuous casting billets. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305639.

Bielecki:2021:PTC

[1254] Wlodzimierz Bielecki, Piotr Blaszynski, and Maciej Poliwoda. 3D parallel tiled code implementing a modified Knuth’s optimal binary search tree algorithm. Journal of Computational Science, 48:??, January 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305445.

Bauer:2021:PLA

[1255] Martin Bauer, Harald Köstler, and Ulrich Rüde. lbmpy: Automatic code generation for efficient parallel lattice Boltzmann methods. Journal of Computational Science, 49:??, February 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320305652.

Anzt:2021:CPR

[1256] Hartwig Anzt, Eileen Kuehn, and Goran Flegar. Crediting pull requests to open source research software as an academic contribution. Journal of Computational Science, 49:??, February 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320305743.

Anonymous:2021:EBb

[1257] Anonymous. Editorial Board. Journal of Computational Science, 49: ??, February 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S1877750321000120.

Anonymous:2021:F

[1258] Anonymous. February 2021. Journal of Computational Science, 49:??, February 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Bauer:2021:LAC

[1259] Martin Bauer, Harald Köstler, and Ulrich Rüde. lbmpy: automatic code generation for efficient parallel lattice Boltzmann methods. Journal of Computational Science, 49, 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Nabapure:2021:DIR

[1260] Deepak Nabapure and Ram Chandra Murthy K. DSMC investigation of rarefied gas flow in a four-sided lid driven cavity: Effect of rarefaction and lid velocities. Journal of Computational Science, 49:??, February 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032030572X.

Fujita:2021:HFN

[1261] Kohei Fujita, Kentaro Koyama, Kazuo Minami, Hikaru Inoue, Seiya Nishizawa, Miwako Tsuji, Tatsuo Nishiki, Tsuyoshi Ichimura, Muneo Hori, and Lalith Maddegedara. Highfidelity nonlinear low-order unstructured implicit finite-element seismic simulation of important structures by accelerated element-by-element method. Journal of Computational Science, 49:??, February 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305731.

Sahasrabudhe:2021:OHS

[1262] Damodar Sahasrabudhe, Rohit Zambre, Aparna Chandramowlishwaran, and Martin Berzins. Optimizing the Hypre solver for manycore and GPU architectures. Journal of Computational Science, 49:??, February 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305755.

Demidov:2021:ALS

[1263] Denis Demidov, Lin Mu, and Bin Wang. Accelerating linear solvers for Stokes problems with C++ metaprogramming. Journal of Computational Science, 49:??, February 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305809.

Ren:2021:CEH

[1264] Jineng Ren, Jarvis Haupt, and Zehua Guo. Communication-efficient hierarchical distributed optimization for multi-agent policy evaluation. Journal of Computational Science, 49:??, February 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320305767.

Bocewicz:2021:OFN

[1265] Grzegorz Bocewicz, Zbigniew Banaszak, Katarzyna Rudnik, Czeslaw Smutnicki, Marcin Witczak, and Robert Wójcik. An ordered-fuzzy-numbers-driven approach to the milkrun routing and scheduling problem. Journal of Computational Science, 49:??, February 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305822.

Seckler:2021:APL

[1266] Steffen Seckler, Fabio Gratl, Matthias Heinen, Jadran Vrabec, Hans-Joachim Bungartz, and Philipp Neumann. AutoPas in ls1 mardyn: Massively parallel particle simulations with nodelevel auto-tuning. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305901.

Anonymous:2021:EBc

[1267] Anonymous. Editorial Board. Journal of Computational Science, 50:??, March 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750321000338.

Anonymous:2021:M

[1268] Anonymous. March 2021. Journal of Computational Science, 50:??, March 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Benalla:2021:CCD

[1269] Mohammed Benalla, Boujemâa Achchab, and Hamid Hrimech. On the computational complexity of Dempster’s Rule of combination, a parallel computing approach. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305780.

Seckler:2021:ALM

[1270] Steffen Seckler, Fabio Gratl, Matthias Heinen, Jadran Vrabec, Hans-Joachim Bungartz, and Philipp Neumann. AutoPas in ls1 mardyn: massively parallel particle simulations with node-level auto-tuning. Journal of Computational Science, 50, 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Cebrian:2021:MPC

[1271] P. Cebrian and J. C. Moure. Massivelyparallel column-level segmentation of depth images. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000016.

Sun:2021:AAM

[1272] Pengtao Sun, Chen-Song Zhang, Rihui Lan, and Lin Li. An advanced ALE-mixed finite element method for a cardiovascular fluid–structure interaction problem with multiple moving interfaces. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000028.

Lyman:2021:EBF

[1273] Laura Lyman and Gianluca Iaccarino. Extending bluff-and-fix estimates for polynomial chaos expansions. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305810.

Moosavi:2021:MLB

[1274] Azam Moosavi, Vishwas Rao, and Adrian Sandu. Machine learning based algorithms for uncertainty quantification in numerical weather prediction models. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305895.

Zheng:2021:VDM

[1275] A. Z. Zheng, S. J. Bian, E. Chaudhry, J. Chang, H. Haron, L. H. You, and J. J. Zhang. Voronoi diagram and Monte-Carlo simulation based finite element optimization for cost-effective 3D printing. Journal of Computational Science, 50:??, March 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100003X.

Los:2021:DDG

[1276] Marcin Łoś, Sergio Rojas, Maciej Paszyński, Ignacio Muga, and Victor M. Calo. DGIRM: Discontinuous Galerkin based isogeometric residual minimization for the Stokes problem. Journal of Computational Science, 50:??, March 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000077.

Emmendorfer:2021:PIA

[1277] Leonardo Ramos Emmendorfer and Graçaliz Pereira Dimuro. A point interpolation algorithm resulting from weighted linear regression. Journal of Computational Science, 50:??, March 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321000053.

Abuelkher:2021:ESS

[1278] Nariman Abu el kher, Israa Zeid, Nayla El-Kork, and Mahmoud Korek. Electronic structure of the SrH^{+} and BaH^{+} molecules with dipole moment and rovibrational calculations. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750320305627.

Grzyb:2021:HDW

[1279] Joanna Grzyb, Jakub Klikowski, and Michał Woźniak. Hellinger Distance Weighted Ensemble for imbalanced data stream classification. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000156.

Nayak:2021:OAE

[1280] Manas Ranjan Nayak, Diptimayee Behura, and Kumari Kasturi. Optimal allocation of energy storage system and its benefit analysis for unbalanced distribution network with wind generation. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

SousaLago:2021:MNT

[1281] André Sousa Lago, João Pedro Dias, and Hugo Sereno Ferreira. Managing non-trivial Internet-of-Things systems with conversational assistants: a prototype and a feasibility experiment. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Tepner:2021:EAO

[1282] Sebastian Tepner, Linda Ney, Marius Singler, Maximilian Pospischil, Kenji Masuri, and Florian Clement. Evolutionary algorithm optimizes screen design for solar cell metallization. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000235.

Munoz:2021:ABS

[1283] Sergio Muñoz and Carlos A. Iglesias. An agent based simulation system for analyzing stress regulation policies at the workplace. Journal of Computational Science, 51, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Ernst:2021:ARP

[1284] Sebastian Ernst, Leszek Kotulski, Tomasz Lerch, Michał Rad, Adam Sȩdziwy, and Igor Wojnicki. Application of reactive power compensation algorithm for large-scale street lighting. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000363.

Settino:2021:IPM

[1285] Marzia Settino, Mariamena Arbitrio, Francesca Scionti, Daniele Caracciolo, Giuseppe Agapito, Pierfrancesco Tassone, Pierosandro Tagliaferri, Maria Teresa Di Martino, and Mario Cannataro. Identifying prognostic markers for multiple myeloma through integration and analysis of MMRF-CoMMpass data. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000430.

Tepe:2021:ENC

[1286] Cengiz Tepe and Mehmet Can Demir.

The effects of the number of channels and gyroscopic data on the classification performance in EMG data acquired by Myo armband. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Soni:2021:HMH

[1287] Neha Soni, Enakshi Khular Sharma, and Amita Kapoor. Hybrid metaheuristic algorithm based deep neural network for face recognition. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic). See corrigendum [1335].

Derevitskii:2021:HPM

[1288] Ilia V. Derevitskii, Daria A. Savitskaya, Alina Yu. Babenko, and Sergey V. Kovalchuk. Hybrid predictive modelling: Thyrotoxic atrial fibrillation case. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2021:EBd

[1289] Anonymous. Editorial Board. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750321000697.

Anonymous:2021:A

[1290] Anonymous. April 2021. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

elkher:2021:ESM

[1291] Nariman Abu el kher, Israa Zeid, Nayla El-Kork, and Mahmoud Korek. Electronic structure of the SrH^{+} and BaH^{+} molecules with dipole moment and rovibrational calculations. Journal of Computational Science, 51, 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Halim:2021:CGU

[1292] Zahid Halim, Hussain Mahmood Sargana, Aadam, Uzma, and Muhammad Waqas. Clustering of graphs using pseudo-guided random walk. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750320305779.

Curado:2021:UMR

[1293] Manuel Curado, Leandro Tortosa, Jose F. Vicent, and Gevorg Yeghikyan. Understanding mobility in Rome by means of a multiplex network with data. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000065.

Hubert:2021:SAC

[1294] G. Hubert and S. Aubry. Simulation of atmospheric cosmic-rays and their impacts based on pre-calculated databases, physical models and computational methods. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000089.

Zhang:2021:OML

[1295] Kai Zhang, Jiahao Zhu, Yimin Zhang, and Qiujun Huang. Optimization method for linear constraint problems. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000168.

Toye:2021:AEO

[1296] Habib Toye, Peng Zhan, Furrukh Sana, Sivareddy Sanikommu, Naila Raboudi, and Ibrahim Hoteit. Adaptive ensemble optimal interpolation for efficient data assimilation in the Red Sea. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032100017X.

Bazzanini:2021:LBM

[1297] L. Bazzanini, A. Gabbana, D. Simeoni, S. Succi, and R. Tripiccione. A lattice Boltzmann method for relativistic rarefied flows in (2 + 1) dimensions. Journal of Computational Science, 51, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Eleliemy:2021:DCC

[1298] Ahmed Eleliemy and Florina M. Ciorba. A distributed chunk calculation approach for self-scheduling of parallel applications on distributedmemory systems. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305792.

Garcia:2021:CPA

[1299] T. Garcia, P. Spiteri, L. Ziane-Khodja, and R. Couturier. Coupling parallel asynchronous multisplitting methods with Krylov methods to solve pseudolinear evolution 3D problems. Journal of Computational Science, 51, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Nino-Ruiz:2021:DDL

[1300] Elias D. Nino-Ruiz. A data-driven localization method for ensemble based data assimilation. Journal of Computational Science, 51, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Geronzi:2021:HFF

[1301] Leonardo Geronzi, Emanuele Gasparotti, Katia Capellini, Ubaldo Cella, Corrado Groth, Stefano Porziani, Andrea Chiappa, Simona Celi, and Marco Evangelos Biancolini. High fidelity fluid-structure interaction by radial basis functions mesh adaption of moving walls: a workflow applied to an aortic valve. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000259.

Gasparini:2021:HPI

[1302] Leonardo Gasparini, José R. P. Rodrigues, Douglas A. Augusto, Luiz M. Carvalho, Cesar Conopoima, Paulo Goldfeld, Jairo Panetta, João P. Ramirez, Michael Souza, Mateus O. Figueiredo, and Victor M. D. M. Leite. Hybrid parallel iterative sparse linear solver framework for reservoir geomechanical and flow simulation. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000284.

Wu:2021:FDA

[1303] Pin Wu, Xuting Chang, Wenyan Yuan, Junwu Sun, Wenjie Zhang, Rossella Arcucci, and Yike Guo. Fast data assimilation (FDA): Data assimilation by machine learning for faster optimize model state. Journal of Computational Science, 51:??, April 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032100020X.

Totounferoush:2021:DBI

[1304] Amin Totounferoush, Neda Ebrahimi Pour, Juri Schröder, Sabine Roller, and Miriam Mehl. A data-based inter-code load balancing method for partitioned solvers. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Wang:2021:ACD

[1305] Yakun Wang and Xiaodong Han. Attractive community detection in academic social network. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Kim:2021:RLV

[1306] Sangkwon Kim, Hyunsoo Han, Hanbyeol Jang, Darae Jeong, Chaeyoung Lee, Wonjin Lee, and Junseok Kim. Reconstruction of the local volatility function using the Black–Scholes model. Journal of Computational Science, 51:??, April 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000387.

Singh:2021:WCM

[1307] Abhishek Kumar Singh and Mani Mehra. Wavelet collocation method based on Legendre polynomials and its application in solving the stochastic fractional integro-differential equations. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Boltuc:2021:PIE

[1308] Agnieszka Bołtuć and Eugeniusz Zieniuk. Parametric integral equation system (PIES) for solving problems with inclusions and non-homogeneous domains using Bézier surfaces. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Kakavand:2021:NST

[1309] M. Kakavand and A. Nikoobin. Numerical simulation of tethered-wing power systems based on variational integration. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Rauber:2021:MEA

[1310] Thomas Rauber and Gudula Rünger.

Modeling the effect of applicationspecific program transformations on energy and performance improvements of parallel ODE solvers. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Dapelo:2021:LBC

[1311] Davide Dapelo, Stephan Simonis, Mathias J. Krause, and John Bridgeman. Lattice-Boltzmann coupled models for advection — diffusion flow on a wide range of Péclet numbers. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Wilde:2021:CRW

[1312] Dominik Wilde, Andreas Krämer, Mario Bedrunka, Dirk Reith, and Holger Foysi. Cubature rules for weakly and fully compressible off-lattice Boltzmann methods. Journal of Computational Science, 51:??, April 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Deelman:2021:PWM

[1313] Ewa Deelman, Rafael Ferreira da Silva, Karan Vahi, Mats Rynge, Rajiv Mayani, Ryan Tanaka, Wendy Whitcup, and Miron Livny. The Pegasus workflow management system: Translational computer science in practice. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305019.

Passini:2021:VAS

[1314] Elisa Passini, Xin Zhou, Cristian Trovato, Oliver J. Britton, Alfonso Bueno-Orovio, and Blanca Rodriguez. The virtual assay software for human in silico drug trials to augment drug cardiac testing. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305032.

Gropp:2021:TRM

[1315] William Gropp, Rajeev Thakur, and Pavan Balaji. Translational research in the MPICH project. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305044.

Panda:2021:MPT

[1316] Dhabaleswar Kumar Panda, Hari Subramoni, Ching-Hsiang Chu, and Mohammadreza Bayatpour. The MVA-PICH project: Transforming research into high-performance MPI library for HPC community. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305093.

Altintas:2021:BCT

[1317] Ilkay Altintas. Building cyberinfrastructure for translational impact: the WIFIRE example. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305111.

Forsch:2021:CAC

[1318] Nickolas Forsch, Sachin Govil, James C.

Perry, Sanjeet Hegde, Alistair A. Young, Jeffrey H. Omens, and Andrew D. McCulloch. Computational analysis of cardiac structure and function in congenital heart disease: Translating discoveries to clinical strategies. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305123.

Gan:2021:TNH

[1319] Lin Gan, Haohuan Fu, and Guangwen Yang. Translating novel HPC techniques into efficient geoscience solutions. Journal of Computational Science, 52, May 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750320305135.

Bockelman:2021:PTT

[1320] Brian Bockelman, Miron Livny, Brian Lin, and Francesco Prelz. Principles, technologies, and time: the translational journey of the HTCondor-CE. Journal of Computational Science, 52, May 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750320305147.

Foster:2021:TGH

[1321] Ian Foster and Carl Kesselman. Translating the grid: How a translational approach shaped the development of grid computing. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032030510X.

Dongarra:2021:TPM

[1322] Jack Dongarra, Mark Gates, Piotr Luszczek, and Stanimire Tomov. Translational process: Mathematical software perspective. Journal of Computational Science, 52:??, May 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305160.

Johnson:2021:TCS

[1323] Chris Johnson. Translational computer science at the Scientific Computing and Imaging Institute. Journal of Computational Science, 52, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305172.

Abramson:2021:TCS

[1324] David Abramson, Manish Parashar, and Peter Arzberger. Translation computer science — overview of the special issue. Journal of Computational Science, 52:??, May 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305305.

Anonymous:2021:EBe

[1325] Anonymous. Editorial Board. Journal of Computational Science, 52, May 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321000867.

Guiso:2021:IGC

[1326] Simone Guiso, Dung di Caprio, Jacques de Lamare, and Benoît Gwinner. Influence of the grid cell geometry on 3D cellular automata behavior in intergranular corrosion. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000211.

DeLuca:2021:RFB

[1327] P. De Luca, A. Galletti, G. Giunta, and L. Marcellino. Recursive filter based GPU algorithms in a Data Assimilation scenario. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000375.

ElMocayd:2021:NIP

[1328] Nabil El Mocayd, M. Shadi Mohamed, and Mohammed Seaid. Non-intrusive polynomial chaos methods for uncertainty quantification in wave problems at high frequencies. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000417.

Maslyaev:2021:PDE

[1329] Mikhail Maslyaev, Alexander Hvatov, and Anna V. Kalyuzhnaya. Partial differential equations discovery with EPDE framework: Application for real and synthetic data. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000429.

Peng:2021:URT

[1330] Zedong Peng, Xuanyi Lin, Michelle Simon, and Nan Niu. Unit and regression tests of scientific software: a study on SWMM. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000442.

Huber:2021:GBM

[1331] Dominik Huber, Martin Schreiber, and Martin Schulz. Graph-based multicore higher-order time integration of linear autonomous partial differential equations. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000466.

Centorrino:2021:MCM

[1332] P. Centorrino, A. Corbetta, E. Cristiani, and E. Onofri. Managing crowded museums: Visitors flow measurement, analysis, modeling, and optimization. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000521.

Foguelman:2021:EDF

[1333] Daniel Foguelman, Philipp Henning, Adelinde Uhrmacher, and Rodrigo Castro. EB-DEVS: a formal framework for modeling and simulation of emergent behavior in dynamic complex systems. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000752.

Li:2021:EUA

[1334] Tingting Li, Wenqi Niu, and Cun Ji. Edge user allocation by FOA in edge computing environment. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750321000788.

Soni:2021:CHM

[1335] Neha Soni, Enakshi Khular Sharma, and Amita Kapoor. Corrigendum to “Hybrid meta-heuristic algorithm based deep neural network for face recognition” [J. Comput. Sci. 51 (2021) 101352]. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032100079X. See

[1287].

Mielczarek:2021:SAE

[1336] Bożena Mielczarek. A simulation approach to evaluate the effect of demographic changes on projected number of patients across disease categories. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000818.

Nikan:2021:NSN

[1337] O. Nikan, Z. Avazzadeh, and J. A. Tenreiro Machado. Numerical study of the nonlinear anomalous reaction– subdiffusion process arising in the electroanalytical chemistry. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321000806.

Kovalchuk:2021:YCS

[1338] Sergey V. Kovalchuk, Valeria V.

Krzhizhanovskaya, Maciej Paszyński, Gábor Závodszky, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. 20 years of computational science: Selected papers from 2020 International Conference on Computational Science. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100082X.

Suleimenova:2021:TAV

[1339] Diana Suleimenova, Hamid Arabnejad, Wouter N. Edeling, David Coster, Onnie O. Luk, Jalal Lakhlili, Vytautas Jancauskas, Michal Kulczewski, Lourens Veen, Dongwei Ye, Pavel Zun, Valeria Krzhizhanovskaya, Alfons Hoekstra, Daan Crommelin, Peter V. Coveney, and Derek Groen. Tutorial applications for verification, validation and uncertainty quantification using VECMA toolkit. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000892.

Thawani:2021:NMS

[1340] B. Thawani, R. Hazael, and R. Critchley. Numerical modelling study of a modified sandbag system for ballistic protection. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000909.

Cheng:2021:ODC

[1341] Sibo Cheng, Didier Lucor, and Jean-Philippe Argaud. Observation data compression for variational assimilation of dynamical systems. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321000910.

Aubin:2021:HCT

[1342] Mateus Rauback Aubin, Rodrigo da Rosa Righi, Victor Hugo Valiati, Cristiano André da Costa, Rodolfo Stoffel Antunes, and Guilherme Galante. Helastic: On combining thresholdbased and serverless elasticity approaches for optimizing the execution of bioinformatics applications. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321000922.

Mucke:2021:ROM

[1343] Nikolaj T. Mücke, Sander M. Bohté, and Cornelis W. Oosterlee. Reduced order modeling for parameterized time-dependent PDEs using spatially and memory aware deep learning. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000934.

Gomez:2021:ICI

[1344] Henry Duque Gómez, Jose Garcia-Rodriguez, Jorge Azorin-Lopez, David Tomás, Andres Fuster-Guillo, and Higinio Mora-Mora. IA-CPS: Intelligent architecture for cyber-physical systems management. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000946.

Anonymous:2021:EBf

[1345] Anonymous. Editorial Board. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321001010.

Anonymous:2021:Jb

[1346] Anonymous. July 2021. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Luca:2021:RFB

[1347] P. De Luca, A. Galletti, G. Giunta, and L. Marcellino. Recursive filter based GPU algorithms in a data assimilation scenario. Journal of Computational Science, 53, 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Wang:2021:MNL

[1348] Guanghui Wang, Yufei Wang, Jimei Li, and Kaidi Liu. A multidimensional network link prediction algorithm and its application for predicting social relationships. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000533.

Matias:2021:FTT

[1349] André F. V. Matias, Rodrigo C. V. Coelho, José S. Andrade, Jr., and Nuno A. M. Araújo. Flow through time—evolving porous media: swelling and erosion. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000545.

Klass:2021:NEB

[1350] Friedemann Klass, Alessandro Gabbana, and Andreas Bartel. A non-equilibrium bounce-back boundary condition for thermal multispeed LBM. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000569.

Suyanto:2021:ERA

[1351] Suyanto Suyanto, Agung Toto Wibowo, Said Al Faraby, Siti Saadah, and Rita Rismala. Evolutionary Rao algorithm. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000594.

Brown:2021:ITF

[1352] Joshua M. Brown, Terry Bossomaier, and Lionel Barnett. Information transfer in finite flocks with topological interactions. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000600.

Brandi:2021:PMD

[1353] Giuseppe Brandi and T. Di Matteo.

Predicting multidimensional data via tensor learning. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000612.

Nino-Ruiz:2021:LSO

[1354] Elias D. Nino-Ruiz. A line-search optimization method for non-Gaussian data assimilation via random quasiorthogonal sub-spaces. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321000624.

Conradin:2021:PFD

[1355] Raphaël Conradin, Christophe Coreixas, Jonas Latt, and Bastien Chopard. PalaCell2D: a framework for detailed tissue morphogenesis. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000491.

Kapturczak:2021:MBS

[1356] Marta Kapturczak and Eugeniusz Zieniuk. Modeling the boundary shape of the problems described by Navier– Lamé equations using NURBS curves in parametric integral equations system method. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000582.

Zipunova:2021:RPD

[1357] E. Zipunova, A. Perepelkina, A. Zakirov, and S. Khilkov. Regularization and the particles-on-demand method for the solution of the discrete Boltzmann equation. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032100065X.

Pouranbarani:2021:CSC

[1358] Elnaz Pouranbarani, Lucas Arantes Berg, Rafael Sachetto Oliveira, Rodrigo Weber dos Santos, and Anders Nygren. Calibration of singlecell model parameters based on membrane resistance improves the accuracy of cardiac tissue simulations. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321000648.

Hushchyn:2021:GCP

[1359] Mikhail Hushchyn and Andrey Ustyuzhanin. Generalization of change-point detection in time series data based on direct density ratio estimation. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000740.

P:2021:DLC

[1360] Surendar P. and Ponni Bala M. Diagnosis of lung cancer using hybrid deep neural network with adaptive sine cosine crow search algorithm. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000636.

Munoz:2021:EGA

[1361] Alba Muñoz and Fernando Rubio. Evaluating genetic algorithms through the approximability hierarchy. Journal of Computational Science, 53:??, July 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321000764.

Eftekhari:2021:BEP

[1362] Aryan Eftekhari, Dimosthenis Pasadakis, Matthias Bollhöfer, Simon Scheidegger, and Olaf Schenk. Block-enhanced precision matrix estimation for largescale datasets. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000776.

Ginzburg:2021:SST

[1363] Irina Ginzburg. Steady-state tworelaxation-time lattice Boltzmann formulation for transport and flow, closed with the compact multi-reflection boundary and interface-conjugate schemes. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305159.

Garcia-Aunon:2021:CCO

[1364] Pablo Garcia-Aunon and Antonio Barrientos Cruz. Corrigendum to “Control optimization of an aerial robotic swarm in a search task and its adaptation to different scenarios” [J. Comput. Sci. 29 (2018) 107–118]. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305561. See

[974].

Hosseinpour:2021:SIP

[1365] Maryam Hosseinpour, Akbar Zendehnam, Seyedeh Mehri Hamidi Sangdehi, and Hamidreza Ghomi Marzdashti. Simulation of the improvement in the performance of a silver-based surface plasmon resonance biosensor using experimental results of cold plasma treatment of glass substrate. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321000958.

Borthakur:2021:SAA

[1366] Manash Pratim Borthakur, Sauro Succi, Fabio Sterpone, Franck Pérot, Anxhelo Diko, and Simone Melchionna. In-silico analysis of airflow dynamics and particle transport within a human nasal cavity. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100096X.

Xu:2021:SFO

[1367] Xiaoyong Xu, Linchen Xiong, and Fengying Zhou. Solving fractional optimal control problems with inequality constraints by a new kind of Chebyshev wavelets method. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321000971.

Qian:2021:API

[1368] William Qian, Sanjukta Bhowmick, Marty O’Neill, Suhasini Ramisetty-Mikler, and Armin R. Mikler. Applying a Probabilistic Infection Model for studying contagion processes in contact networks. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001046.

Stefanini:2021:CLS

[1369] C. Stefanini, F. Giorgetti, A. Mercuri, A. Facci, and P. Fanelli. Cylinderlamina system fluid–structure interaction problem solved with an original OpenFOAM code. Journal of Computational Science, 54: ??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001058.

Kaura:2021:ECC

[1370] Prerna Kaura, Tanya Mishra, Nishith Verma, Indranil Saha Dalal, and Vivek Sheraton. Effects of combined chemotherapeutic drugs on the growth and survival of cancerous tumours– an in-silico study. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100106X.

Perez:2021:STE

[1371] Daniel Hugo Cámpora Pérez, Niko Neufeld, and Agustín Riscos Núñez. Search by triplet: an efficient local track reconstruction algorithm for parallel architectures. Journal of Computational Science, 54: ??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001071.

Ghosh:2021:SAT

[1372] Uttam Ghosh, Bapin Mondal, Md Sadikur Rahman, and Susmita Sarkar. Stability analysis of a three species food chain model with linear functional response via imprecise and parametric approach. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001083.

Machado:2021:PTM

[1373] Rafael Ravedutti L. Machado, Jonas Schmitt, Sebastian Eibl, Jan Eitzinger, Roland Leißa, Sebastian Hack, Arsène Pérard-Gayot, Richard Membarth, and Harald Köstler. tinyMD: Mapping molecular dynamics simulations to heterogeneous hardware using partial evaluation. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001095.

Treesatayapun:2021:DTD

[1374] Chidentree Treesatayapun and Aldo Jonathan Muñoz-Vázquez. Discrete-time datadriven disturbance-observer control based on fuzzy rules emulating networks. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001101.

Bagheri:2021:SAO

[1375] Samineh Bagheri, Ulf Reinicke, Denis Anders, and Wolfgang Konen. Surrogate-assisted optimization for augmentation of finite element techniques. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001113.

Durukan:2021:ISP

[1376] Ilknur Kars Durukan and Yasemin Oztekin Ciftci. Ab-initio study on physical properties of intermetallic LiPb compound. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001125.

Grulich:2021:ASD

[1377] Lucas Grulich, Ralf Weigel, Andreas Hildebrandt, Michael Wand, and Peter Spichtinger. Automatic shape detection of ice crystals. Journal of Computational Science, 54: ??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321001137.

Hanaforoosh:2021:ABM

[1378] Mahrad Hanaforoosh, Mehrdad Ashtiani, and Mohammad Abdollahi Azgomi. An approach based on multiplex networks for modeling cascading trust failures in social networks. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001149.

Azarang:2021:ADL

[1379] Arian Azarang and Nasser Kehtarnavaz. Application of deep learning models in nonlinear detail map prediction in pansharpening. Journal of Computational Science, 54:

??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001150.

VandenAkker:2021:UVM

[1380] Harry E. A. Van den Akker, Renske Donkers, Githin T. Zachariah, and Orest Shardt. On using variable molecular masses in multicomponent lattice Boltzmann simulations. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001162.

Itkin:2021:FFM

[1381] Andrey Itkin and Fazlollah Soleymani.

Four-factor model of Quanto CDS with jumps-at-default and stochastic recovery. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001174.

Stamou:2021:ESG

[1382] A. C. Stamou, J. Radulovic, and J. M.

Buick. Effect of stenosis growth on blood flow at the bifurcation of the carotid artery. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001186.

Wang:2021:ACS

[1383] Gang Wang, Shiwei Lu, Wenbin Liu, and Runnian Ma. Adaptive complete synchronization of two complex networks with uncertain parameters, structures, and disturbances. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321001198.

Sitko:2021:RCA

[1384] Mateusz Sitko and Lukasz Madej. The role of the cellular automata cell size and time step length in the microstructure evolution model — the static recrystallization case study. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001204.

Dakkak:2021:TAD

[1385] Omar Dakkak, Yousef Fazea, Shahrudin Awang Nor, and Suki Arif. Towards accommodating deadline driven jobs on high performance computing platforms in grid computing environment. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001228.

Zhang:2021:RWU

[1386] Hengshan Zhang, Yimin Zhou, Tianhua Chen, Richard Hill, Zhongmin Wang, and Yanping Chen. Refinement of weights using attribute support for multiple attribute decision making. Journal of Computational Science, 54:??, September 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100123X.

Anonymous:2021:EBg

[1387] Anonymous. Editorial Board. Journal of Computational Science, 54: ??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032100140X.

Anonymous:2021:S

[1388] Anonymous. September 2021. Journal of Computational Science, 54:??, September 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Nian:2021:MIP

[1389] Fuzhong Nian, Yayong Shi, and Jun Cao. Modeling information propagation in high-order networks based on explicit–implicit relationship. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001216.

Paszynski:2021:HOC

[1390] Maciej Paszyński and Marcin Łoś.

Higher order and continuity L^{2} projections with piece-wise constant test functions. Journal of Computational Science, 55:??, October 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Fahim:2021:SMM

[1391] Ahmed Fahim. K and starting means for k-means algorithm. Journal of Computational Science, 55:??, October 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321001277.

Vrba:2021:APA

[1392] Jan Vrba, Charlie Maslen, Jana Maxova, Jan Duras, Ivan Rehor, and Jan Mares. An automated platform for assembling light-powered hydrogel microrobots and their subsequent chemical binding. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001265.

Plazek:2021:TFA

[1393] J. Płażek, J. Smyrski, P. Salabura, W. Przygoda, and K. Korcyl. Track finding algorithm using the track road method for the PANDA Forward Tracker. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001290.

Benaissa:2021:YAP

[1394] Brahim Benaissa, Nourredine Aït Hocine, Samir Khatir, Mohamed Kamel Riahi, and Seyedali Mirjalili. YUKI algorithm and POD-RBF for elastostatic and dynamic crack identification. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001319.

Zhang:2021:AVA

[1395] Y. P. Zhang. Axial vibration analysis of nanorods with variable density based on nonlocal elastic theory and high-order finite difference method. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001307.

Panwar:2021:TOB

[1396] Karuna Panwar and Kusum Deep.

Transformation operators based grey wolf optimizer for travelling salesman problem. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001320.

Lin:2021:SRT

[1397] Szu-Yun Lin, Andrew W. Hlynka, Lichao Xu, Hao Lu, Omar A. Sediek, Sherif El-Tawil, Vineet R. Kamat, Jason McCormick, Carol C. Menassa, Seymour M. J. Spence, Atul Prakash, and Benigno Aguirre. Simple Run-Time Infrastructure (SRTI): an accessible distributed computing platform for interdisciplinary simulation. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001332.

Ellis:2021:VNN

[1398] Joshua D. Ellis, Razib Iqbal, and Keiichi Yoshimatsu. Verification of the neural network training process for spectrum-based chemical substructure prediction using metamorphic testing. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001344.

Berrone:2021:LWR

[1399] Stefano Berrone, Francesco Della Santa, Antonio Mastropietro, Sandra Pieraccini, and Francesco Vaccarino. Layer-wise relevance propagation for backbone identification in discrete fracture networks. Journal of Computational Science, 55:??, October 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001356.

Guo:2021:FEF

[1400] Zhixin Guo, Yifan Xiao, Wenzhi Liao, Peter Veelaert, and Wilfried Philips. FLOPs-efficient filter pruning via transfer scale for neural network acceleration. Journal of Computational Science, 55:??, October 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001368.

Sitnikov:2021:SMC

[1401] S. S. Sitnikov and F. G. Tcheremissine. Simulation of mixture components separation in a rarefied argonneon jet on the basis of direct solution of the Boltzmann kinetic equation. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001447.

Farchi:2021:CCD

[1402] Alban Farchi, Marc Bocquet, Patrick Laloyaux, Massimo Bonavita, and Quentin Malartic. A comparison of combined data assimilation and machine learning methods for offline and online model error correction. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001435.

Hernandez:2021:EMC

[1403] Monica Hernandez. Efficient momentum conservation constrained PDE-LDDMM with Gauss–Newton–Krylov optimization, semi-Lagrangian Runge– Kutta solvers, and the band-limited parameterization. Journal of Computational Science, 55:??, October 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001459.

Du:2021:CTO

[1404] Xin Du, Kife I. Bin Iqbal, M. Monir Uddin, A. Mostakim Fony, Md. Tanzim Hossain, Mian Ilyas Ahmad, and Mohammad Sahadet Hossain. Computational techniques for H_{∈} optimal frequency-limited model order reduction of large-scale sparse linear systems. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001472.

He:2021:ISC

[1405] Wenjuan He, Bing Wang, Ning Li, Xiaojie Gao, Wei Li, and Qiaoyong Jiang. An improved sine–cosine algorithm with dynamic selection pressure. Journal of Computational Science, 55:??, October 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001496.

Anonymous:2021:EBh

[1406] Anonymous. Editorial Board. Journal of Computational Science, 55:??, October 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750321001630.

Anonymous:2021:O

[1407] Anonymous. October 2021. Journal of Computational Science, 55:??, October 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Vela-Martin:2021:LSM

[1408] Alberto Vela-Martín, Miguel P. Encinar, Adrián García-Gutiérrez, and Javier Jiménez. A low-storage method consistent with second-order statistics for time-resolved databases of turbulent channel flow up to Re_{τ} = 5300. Journal of Computational Science, 56:??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

dosSantosCruz:2021:UQR

[1409] Wenna Raissa dos Santos Cruz, Fabio Pereira dos Santos, and Ricardo de Andrade Medronho. Uncertainty quantification of real gas models in CO_{2} supersonic flow. Journal of Computational Science, 56:??, November 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Galvez:2021:NFN

[1410] Akemi Gálvez, Andrés Iglesias, Iztok Fister, Iztok Fister, César Otero, and José A. Díaz. NURBS functional network approach for automatic image segmentation of macroscopic medical images in melanoma detection. Journal of Computational Science, 56:??, November 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001526.

Riego:2021:BDC

[1411] Virginia Riego, Lidia Sánchez-González, Laura Fernández-Robles, Alexis Gutiérrez-Fernández, and Nicola Strisciuglio. Burr detection and classification using RUSTICO and image processing. Journal of Computational Science, 56:??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Farajzadeh:2021:DNN

[1412] Nacer Farajzadeh and Mahdi Hashemzadeh. A deep neural network based framework for restoring the damaged Persian pottery via digital inpainting. Journal of Computational Science, 56:??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750321001551.

Ozdemir:2021:ANS

[1413] Neslihan Ozdemir, Aydin Secer, and Mustafa Bayram. An algorithm for numerical solution of some nonlinear multi-dimensional parabolic partial differential equations. Journal of Computational Science, 56:??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Yang:2021:CSE

[1414] Junxiang Yang, Chaeyoung Lee, Soobin Kwak, Yongho Choi, and Junseok Kim. A conservative and stable explicit finite difference scheme for the diffusion equation. Journal of Computational Science, 56:??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

He:2021:PEB

[1415] Hantao He, Junxing Zheng, Ying Chen, and Yingjie Ning. Physics engine based simulation of shear behavior of granular soils using hard and soft contact models. Journal of Computational Science, 56:??, November 2021. ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2021:EBi

[1416] Anonymous. Editorial Board. Journal of Computational Science, 56: ??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001757.

Anonymous:2021:N

[1417] Anonymous. November 2021. Journal of Computational Science, 56:??, November 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Shahrouzi:2022:EDF

[1418] Mohsen Shahrouzi and Ali Kaveh. An efficient derivative-free optimization algorithm inspired by avian life-saving manoeuvres. Journal of Computational Science, 57:??, January 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001514.

Khouzami:2022:MBA

[1419] Nesrine Khouzami, Friedrich Michel, Pietro Incardona, Jeronimo Castrillon, and Ivo F. Sbalzarini. Model-based autotuning of discretization methods in numerical simulations of partial differential equations. Journal of Computational Science, 57:??, January 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001563.

Kumar:2022:FFB

[1420] Mukesh Kumar, Shivansh Mishra, and Bhaskar Biswas. Features fusion based link prediction in dynamic neworks. Journal of Computational Science, 57:??, January 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001587.

Fuchsberger:2022:IOF

[1421] Jana Fuchsberger, Philipp Aigner, Steven Niederer, Gernot Plank, Heinrich Schima, Gundolf Haase, and Elias Karabelas. On the incorporation of obstacles in a fluid flow problem using a Navier–Stokes–Brinkman penalization approach. Journal of Computational Science, 57:??, January 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001691.

Lei:2022:CII

[1422] Jing Lei, Qibin Liu, and Xueyao Wang. Computational inverse imaging method by machine learning-informed physical model for electrical capacitance tomography. Journal of Computational Science, 57:??, January 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001708.

Lin:2022:RUL

[1423] Ruiguan Lin, Yaowei Yu, Huawei Wang, Changchang Che, and Xiaomei Ni. Remaining useful life prediction in prognostics using multi-scale sequence and Long Short-Term Memory network. Journal of Computational Science, 57:??, January 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100171X.

Anonymous:2022:EBa

[1424] Anonymous. Editorial Board. Journal of Computational Science, 57:??, January 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321001903.

Anonymous:2022:Ja

[1425] Anonymous. January 2022. Journal of Computational Science, 57:??, January 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Breitenbach:2022:PFA

[1426] Tim Breitenbach, Lauritz Rasbach, Chunguang Liang, and Patrick Jahnke. A principal feature analysis. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750321001666.

Torre-Tojal:2022:AGB

[1427] Leyre Torre-Tojal, Aitor Bastarrika, Ana Boyano, Jose Manuel Lopez-Guede, and Manuel Graña. Aboveground biomass estimation from Li-DAR data using random forest algorithms. Journal of Computational Science, 58:??, February 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001800.

Akram:2022:LDA

[1428] Vahid Khalilpour Akram and Onur Ugurlu. A localized distributed algorithm for vertex cover problem. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001794.

Ramazzotti:2022:PLI

[1429] Daniele Ramazzotti, Fabrizio Angaroni, Davide Maspero, Gianluca Ascolani, Isabella Castiglioni, Rocco Piazza, Marco Antoniotti, and Alex Graudenzi. LACE: Inference of cancer evolution models from longitudinal single-cell sequencing data. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001848.

Bai:2022:ESQ

[1430] Mei Bai, Senan Jiang, Xin Zhang, and Xite Wang. An efficient skyline query algorithm in the distributed environment. Journal of Computational Science, 58:??, February 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100185X.

Buizza:2022:DLI

[1431] Caterina Buizza, César Quilodrán Casas, Philip Nadler, Julian Mack, Stefano Marrone, Zainab Titus, Clémence Le Cornec, Evelyn Heylen, Tolga Dur, Luis Baca Ruiz, Claire Heaney, Julio Amador Díaz Lopez, K. S. Sesh Kumar, and Rossella Arcucci. Data learning: Integrating data assimilation and machine learning. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001861.

Lytaev:2022:RIO

[1432] Mikhail S. Lytaev. Rational interpolation of the one-way Helmholtz propagator. Journal of Computational Science, 58:??, February 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001952.

Veerappa:2022:VXE

[1433] Manjunatha Veerappa, Mathias Anneken, Nadia Burkart, and Marco F. Huber. Validation of XAI explanations for multivariate time series classification in the maritime domain. Journal of Computational Science, 58:??, February 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001976.

Durakli:2022:NAB

[1434] Zafer Duraklı and Vasif Nabiyev. A new approach based on Bézier curves to solve path planning problems for mobile robots. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321001988.

Chen:2022:EMO

[1435] Lizhi Chen, Wei-Li Liu, and Jinghui Zhong. An efficient multi-objective ant colony optimization for task allocation of heterogeneous unmanned aerial vehicles. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321002015.

A:2022:DAE

[1436] Shyna A., Ushadevi Amma C., Ansamma John, Kesavadas C., and Bejoy Thomas. Deep-ASL enhancement technique in arterial spin labeling MRI — a novel approach for the error reduction of partial volume correction technique with linear regression algorithm. Journal of Computational Science, 58:??, February 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321002027.

Prokhorov:2022:TBC

[1437] Dmitry Prokhorov, Vadim Lisitsa, Tatyana Khachkova, Yaroslav Bazaikin, and Yongfei Yang. Topology-based characterization of chemically-induced pore space changes using reduction of 3D digital images. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750321002052.

Anonymous:2022:EBb

[1438] Anonymous. Editorial Board. Journal of Computational Science, 58: ??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S1877750322000199.

Anonymous:2022:F

[1439] Anonymous. February 2022. Journal of Computational Science, 58:??, February 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Smejkal:2022:MPC

[1440] Tomáš Smejkal, Jiří Mikyška, and Tissa H. Illangasekare. Multi-phase compositional modeling in porous media using iterative IMPEC scheme and constant volume–temperature flash. Journal of Computational Science, 59:??, March 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321001939.

Renc:2022:TEG

[1441] Paweł Renc, Tomasz Pȩcak, Alessio De Rango, William Spataro, Giuseppe Mendicino, and Jarosław Wa̧s. Towards efficient GPGPU Cellular Automata model implementation using persistent active cells. Journal of Computational Science, 59:??, March 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321001964.

Smolka:2022:AEF

[1442] Krzysztof Smółka, Anna Firych-Nowacka, and Sławomir Wiak. Analysis of the electrostatic field distribution to improve the electrospinning process — practical tips. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100199X.

Casal:2022:TCN

[1443] Ramiro Casal, Leandro E. Di Persia, and Gastón Schlotthauer. Temporal convolutional networks and transformers for classifying the sleep stage in awake or asleep using pulse oximetry signals. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321002003.

Yartu:2022:HFP

[1444] Mercedes Yartu, Carlos Cambra, Milagros Navarro, Carlos Rad, Ángel Arroyo, and Álvaro Herrero. Humidity forecasting in a potato plantation using time-series neural models. Journal of Computational Science, 59:??, March 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750321002039.

Gasque:2022:MMS

[1445] Diógenes Gasque and Pedro Munari. Metaheuristic, models and software for the heterogeneous fleet pickup and delivery problem with split loads. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321002040.

Chen:2022:PSE

[1446] Qipin Chen, Wenrui Hao, and Juncai He. Power series expansion neural network. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321002064.

Maris:2022:AIE

[1447] Pieter Maris, Chao Yang, Dossay Oryspayev, and Brandon Cook. Accelerating an iterative eigensolver for nuclear structure configuration interaction calculations on GPUs using OpenACC. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750321002076.

Nian:2022:SAN

[1448] Fuzhong Nian, Yinuo Qian, and Rendong Liu. Self-adaptive network model based on incentive mechanism. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000035.

Maier:2022:MGM

[1449] Benjamin Maier and Miriam Schulte. Mesh generation and multi-scale simulation of a contracting muscle–tendon complex. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000023.

Ma:2022:LRA

[1450] Linjian Ma and Chao Yang. Low rank approximation in simulations of quantum algorithms. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000047.

Funkner:2022:SAP

[1451] Anastasia A. Funkner, Aleksey N. Yakovlev, and Sergey V. Kovalchuk. Surrogate-assisted performance prediction for data-driven knowledge discovery algorithms: Application to evolutionary modeling of clinical pathways. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000059.

Patikova:2022:ADT

[1452] Zuzana Pátíková and Josef Rebenda. Applications of the differential transform to second-order half-linear Euler equations. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000060.

Mishra:2022:EAM

[1453] Rohit Mishra and Bhagat Singh.

An ensemble approach to maximize metal removal rate for chatter free milling. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000084.

Amigo:2022:SOT

[1454] Daniel Amigo, David Sánchez Pedroche, Jesús García, and José Manuel Molina. Segmentation optimization in trajectory-based ship classification. Journal of Computational Science, 59:??, March 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000096.

Sun:2022:BMA

[1455] Junlong Sun, Shunchuan Wu, Huajin Zhang, Xiaoqiang Zhang, and Tao Wang. Based on multialgorithm hybrid method to predict the slope safety factor–stacking ensemble learning with Bayesian optimization. Journal of Computational Science, 59:??, March 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000242.

Anonymous:2022:EBc

[1456] Anonymous. Editorial Board. Journal of Computational Science, 59:??, March 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322000485.

Anonymous:2022:Ma

[1457] Anonymous. March 2022. Journal of Computational Science, 59:??, March 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Khan:2022:DBA

[1458] Ameer Tamoor Khan, Xinwei Cao, and Shuai Li. Dual Beetle Antennae Search system for optimal planning and robust control of 5-link biped robots. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000011.

Tan:2022:AVP

[1459] Nigel Tan, Robert F. Bird, Guangye Chen, Scott V. Luedtke, Brian J. Albright, and Michela Taufer. Analysis of Vector Particle-In-Cell (VPIC) memory usage optimizations on cuttingedge computer architectures. Journal of Computational Science, 60:??, April 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322000072.

El-Monajjed:2022:HID

[1460] Khaled El-Monajjed and Mark Driscoll. Haptic integration of data-driven forces required to gain access using a probe for minimally invasive spine surgery via cadaveric-based experiments towards use in surgical simulators. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000102.

Kaur:2022:BPB

[1461] Parampreet Kaur, Ashima Singh, and Inderveer Chana. BSense: a parallel Bayesian hyperparameter optimized stacked ensemble model for breast cancer survival prediction. Journal of Computational Science, 60:??, April 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000114.

Cascitti:2022:RSA

[1462] Julian Cascitti, Stefan Niebler, André Müller, and Bertil Schmidt. RNACache: a scalable approach to rapid transcriptomic read mapping using locality sensitive hashing. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000126.

Cardenas:2022:BDH

[1463] Pedro Cárdenas, Ioannis Ivrissimtzis, Boguslaw Obara, Ibad Kureshi, and Georgios Theodoropoulos. Big data for human security: the case of COVID-19. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000138.

Kozik:2022:TSC

[1464] Rafał Kozik, Sebastian Kula, Michał Choraś, and Michał Woźniak. Technical solution to counter potential crime: Text analysis to detect fake news and disinformation. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200014X.

Murua:2022:SMO

[1465] M. Murua, D. Galar, and R. Santana. Solving the multi-objective Hamiltonian cycle problem using a Branch-and-Fix based algorithm. Journal of Computational Science, 60:??, April 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000151.

Clempner:2022:AGA

[1466] Julio B. Clempner. Algorithmicgradient approach for the Price of Anarchy and Stability for incomplete information. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000254.

Meng:2022:NPE

[1467] Lei Meng, Guiqiong Xu, Pingle Yang, and Dengqin Tu. A novel potential edge weight method for identifying influential nodes in complex networks based on neighborhood and position. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000266.

Rocha:2022:MRC

[1468] Franciane F. Rocha, Fabricio S.

Sousa, Roberto F. Ausas, Gustavo C. Buscaglia, and Felipe Pereira. A multiscale Robin-coupled implicit method for two-phase flows in high-contrast formations. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000278.

Cui:2022:NSP

[1469] Miao Cui, Chunyun Zhang, Bowen Zhang, Bingbing Xu, Haifeng Peng, and Xiao wei Gao. Numerical solution of phase change heat transfer problems by effective heat capacity model and element differential method. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200028X.

Gudivada:2022:SFF

[1470] A. Arunkumar Gudivada and Gnanou Florence Sudha. STQCA-FFT: a Fast Fourier Transform architecture using stack-type QCA approach with power and delay reduction. Journal of Computational Science, 60:??, April 2022.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000291.

Sbayti:2022:ESM

[1471] M. Sbayti, A. Ghiotti, R. Bahloul, H. BelhadjSalah, and S. Bruschi. Effective strategies of metamodeling and optimization of hot incremental sheet forming process of Ti6Al4Vartificial hip joint component. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000308.

Yavuz:2022:DPU

[1472] Gürcan Yavuz. Diversified Position Update Equation-Based SSA with Refreshing-Gap Strategy for global optimization. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200031X.

Mitra:2022:PRB

[1473] Suman Mitra and Sriyankar Acharyya. Perturbation and repository based Diversified Cuckoo Search in reconstruction of Gene Regulatory Network: a new Cuckoo Search approach. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000333.

Bhattacharyya:2022:UQD

[1474] Biswarup Bhattacharyya. Uncertainty quantification of dynamical systems by a POD–Kriging surrogate model. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000345.

Lualdi:2022:EOS

[1475] Pietro Lualdi, Ralf Sturm, and Tjark Siefkes. Exploration-oriented sampling strategies for global surrogate modeling: a comparison between one-stage and adaptive methods. Journal of Computational Science, 60:??, April 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000357.

Mishra:2022:MMM

[1476] Shivansh Mishra, Shashank Sheshar Singh, Ajay Kumar, and Bhaskar Biswas. MNERLP-MUL: Merged node and edge relevance based link prediction in multiplex networks. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000369.

Sahu:2022:MRC

[1477] Bandita Sahu, Pradipta Kumar Das, and Manas ranjan Kabat. Multirobot cooperation and path planning for stick transporting using improved Q-learning and democratic robotics PSO. Journal of Computational Science, 60:??, April 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200059X.

Anonymous:2022:EBd

[1478] Anonymous. Editorial Board. Journal of Computational Science, 60:??, April 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322000813.

Anonymous:2022:A

[1479] Anonymous. April 2022. Journal of Computational Science, 60:??, April 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Maryada:2022:RCP

[1480] K. R. Maryada and S. E. Norris. Reduced-communication parallel dynamic mode decomposition. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322000321.

Shevchenko:2022:NLP

[1481] V. Shevchenko and A. Tanashkin. Non-local Potts model on random lattice and chromatic number of a plane. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000370.

Isupov:2022:MPS

[1482] Konstantin Isupov. Multiple-precision sparse matrix–vector multiplication on GPUs. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000382.

Dhou:2022:CCC

[1483] Khaldoon Dhou and Christopher Cruzen. A creative chain coding technique for bi-level image compression inspired by the NetLogo HIV agent-based modeling simulation. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000394.

Czarnowski:2022:WEO

[1484] Ireneusz Czarnowski. Weighted Ensemble with one-class Classification and Over-sampling and Instance selection (WECOI): an approach for learning from imbalanced data streams. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000412.

Sane:2022:DVL

[1485] Sudhanshu Sane, Chris R. Johnson, and Hank Childs. Demonstrating the viability of Lagrangian in situ reduction on supercomputers. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000400.

Jarmatz:2022:MNL

[1486] Piet Jarmatz, Felix Maurer, Helene Wittenberg, and Philipp Neumann. MaMiCo: Non-local means and POD filtering with flexible data-flow for two-way coupled molecular-continuum HPC flow simulation. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000424.

Kusakabe:2022:SLS

[1487] Ryota Kusakabe, Tsuyoshi Ichimura, Kohei Fujita, Muneo Hori, and Lalith Wijerathne. Scalable largescale multi-physics earthquake simulation on multiple GPUs with stabilization. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000436.

Malecki:2022:MCC

[1488] Krzysztof Małecki, Marek Kamiński, and Jarosław Wa̧s. A multi-cell Cellular Automata model of traffic flow with emergency vehicles: Effect of a corridor of life and drivers’ behaviour. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000515.

Tan:2022:EWW

[1489] Li Tan, Raymond A. de Callafon, Jessica Block, Daniel Crawl, Tolga Çağlar, and Ilkay Altıntaş. Estimation of wildfire wind conditions via perimeter and surface area optimization. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000552.

Singh:2022:DCC

[1490] Dhananjay Kumar Singh, Subrata Nandi, Tanmoy Chakraborty, and Prasenjit Choudhury. Disintegrating constant communities in complex networks. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000564.

Zhao:2022:TLM

[1491] Han-Tao Zhao, Hao-Zheng Li, Hao Qin, and Lin-Hui Zheng. Two-lane mixed traffic flow model considering lane changing. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000576.

Nadimi-Shahraki:2022:GGC

[1492] Mohammad H. Nadimi-Shahraki, Shokooh Taghian, Seyedali Mirjalili, Hoda Zamani, and Ardeshir Bahreininejad. GGWO: Gaze cues learning-based grey wolf optimizer and its applications for solving engineering problems. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000588.

Hosseini:2022:UNU

[1493] Ali Hosseini, Mahdi Hashemzadeh, and Nacer Farajzadeh. UFS-net: a unified flame and smoke detection method for early detection of fire in video surveillance applications using CNNs. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000606.

Neves:2022:MDI

[1494] Diogo Telmo Neves, João Alves, Marcel Ganesh Naik, Alberto José Proença, and Fabian Prasser. From missing data imputation to data generation. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000618.

Oprisan:2022:ICC

[1495] Sorinel A. Oprisan. Interdisciplinary curriculum for computational neuroscience at primarily undergraduate institutions. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200062X.

Boulmier:2022:TIP

[1496] Anthony Boulmier, Nabil Abdennadher, and Bastien Chopard. Toward informed partitioning for load balancing: a proof-of-concept. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000631.

Yansari:2022:NSA

[1497] Ramazan Teimouri Yansari, Mitra Mirzarezaee, Mehdi Sadeghi, and Babak Nadjar Araabi. A new survival analysis model in adjuvant Tamoxifentreated breast cancer patients using manifold-based semi-supervised learning. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000643.

Zhu:2022:PPB

[1498] Zaiping Zhu, Anzong Zheng, Andrés Iglesias, Shuangbu Wang, Yu Xia, Ehtzaz Chaudhry, Lihua You, and Jianjun Zhang. PDE patch-based surface reconstruction from point clouds. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000655.

Djurasevic:2022:SDR

[1499] Marko – Durasević and Domagoj Jakobović.

Selection of dispatching rules evolved by genetic programming in dynamic unrelated machines scheduling based on problem characteristics. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000667.

Etminaniesfahani:2022:AHA

[1500] Alireza Etminaniesfahani, Hanyu Gu, and Amir Salehipour. ABFIA: a hybrid algorithm based on artificial bee colony and Fibonacci indicator algorithm. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000679.

Djouzi:2022:NAS

[1501] Kheyreddine Djouzi, Kadda Beghdad-Bey, and Abdenour Amamra. A new adaptive sampling algorithm for big data classification. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000680.

Morawska:2022:TLB

[1502] Barbara Morawska, Piotr Lipinski, Krzysztof Lichy, and Krzysztof Adamkiewicz. Transfer learning-based UWB indoor localization using MHT-MDC and clusterization-based sparse fingerprinting. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000692.

Belhamadia:2022:EPA

[1503] Youssef Belhamadia, Thomas Briffard, and André Fortin. Efficiency of parallel anisotropic mesh adaptation for the solution of the bidomain model in cardiac tissue. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000709.

Lin:2022:IKN

[1504] Wei Lin, Shuming Zhou, Min Li, and Gaolin Chen. Identifying key nodes in interdependent networks based on Supra-Laplacian energy. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322000710.

Zieniuk:2022:RPI

[1505] Eugeniusz Zieniuk and Krzysztof Szerszeń. A regularization of the parametric integral equation system applied to 2D boundary problems for Laplace’s equation with stability evaluation. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000722.

Xavier:2022:TRV

[1506] Carolina Ribeiro Xavier, Rafael Sachetto Oliveira, Vinícius da Fonseca Vieira, Bernardo Martins Rocha, Ruy Freitas Reis, Bárbara de Melo Quintela, Marcelo Lobosco, and Rodrigo Weber dos Santos. Timing the race of vaccination, new variants, and relaxing restrictions during COVID-19 pandemic. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000734.

Fain:2022:GAD

[1507] Baylor G. Fain and Hana M. Dobrovolny. GPU acceleration and data fitting: Agent-based models of viral infections can now be parameterized in hours. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000746.

Macia:2022:AGH

[1508] Sandra Macià, Pedro J. Martínez-Ferrer, Eduard Ayguadé, and Vicenç Beltran. Automated generation of High-Performance Computational Fluid Dynamics Codes. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750322000758.

Jia:2022:MTS

[1509] Wangkun Jia and Ming-C. Cheng. A methodology for thermal simulation of interconnects enabled by model reduction with material property variation. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200076X.

Weron:2022:RCD

[1510] Tomasz Weron and Katarzyna Sznajd-Weron. On reaching the consensus by disagreeing. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000771.

Tsai:2022:TSS

[1511] Pei-Keng Tsai and Mei-Jiau Huang.

A theoretical and simulation study of phonon flow within single-interface systems. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000874.

Fathinavid:2022:MCL

[1512] Amirhossein Fathinavid. Multilayer cellular learning automata: a computational model to solve multilayer infrastructure problems with its application in community detection for multilayer networks. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000916.

Deng:2022:DEL

[1513] Yajun Deng, Lin Zhang, Dongliang Sun, and Bo Yu. Development of an efficient large time step unsteady solver for incompressible flows using the IDEAL algorithm in Open-FOAM. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000977.

Cardesa:2022:OAL

[1514] J. I. Cardesa, R. Fiévet, E. Piot, H. Deniau, and C. Airiau. Optimizing an acoustic liner by automatic differentiation of a compressible flow solver. Journal of Computational Science, 61:??, May 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001041.

Anonymous:2022:EBe

[1515] Anonymous. Editorial Board. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322001235.

Anonymous:2022:Mb

[1516] Anonymous. May 2022. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Wu:2022:TFB

[1517] Jinrong Wu, Jianhua Hu, Yan Song, Guoliang Wei, and Chungen Shen. A transfer framework based on comatrix decomposition for undirected, high-dimensional and sparse networks. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000862.

Javaid:2022:ASP

[1518] Saba Javaid, Asim Aziz, and Taha Aziz. Algebraic solutions for pricing American put options under the constant elasticity of variance (CEV) model: Application of the Lie group approach. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000886.

Soltani-Sarvestani:2022:VCS

[1519] Mohammadamin Soltani-Sarvestani, Zohreh Azimifar, and Alexander Wong. Variable Centrality: a scenario based centrality calculation method. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000898.

Cid:2022:OSF

[1520] Clara Cid, Aitor Baldomir, Miguel Rodríguez-Segade, and Santiago Hernández. An open-source framework for aircraft damage simulation in engine failure events. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000904.

Xie:2022:ONC

[1521] Shengkun Xie, Anna T. Lawniczak, and Chong Gan. Optimal number of clusters in explainable data analysis of agent-based simulation experiments. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322000928.

Sefidgar:2022:SDL

[1522] Seyed Mohammad Hassan Sefidgar, Ali Rahmani Firoozjaee, and Mehdi Dehestani. Sparse discrete least squares meshless method on multicore computers. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200093X.

Rezaeiravesh:2022:UQF

[1523] S. Rezaeiravesh, R. Vinuesa, and P. Schlatter. An uncertaintyquantification framework for assessing accuracy, sensitivity, and robustness in computational fluid dynamics. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322000941.

Jafarimoghaddam:2022:ODM

[1524] Amin Jafarimoghaddam, Manuel Soler, and Abolfazl Simorgh. The optimal decomposition method (ODM) for nonlinear problems. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000965.

Tabatabaei:2022:MAI

[1525] S. Sepehr Tabatabaei and Mohammad Reza Dehghan. Modeling and adaptive identification of arterial behavior; a variable order approach. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000953.

Steclik:2022:AGP

[1526] Tomasz Steclik, Rafal Cupek, and Marek Drewniak. Automatic grouping of production data in Industry 4.0: the use case of internal logistics systems based on Automated Guided Vehicles. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000989.

Maurya:2022:SAN

[1527] Sunil Kumar Maurya, Xin Liu, and Tsuyoshi Murata. Simplifying approach to node classification in Graph Neural Networks. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322000990.

Akram:2022:AEF

[1528] Javaria Akram, Noreen Sher Akbar, and Dharmendra Tripathi. Analysis of electroosmotic flow of silverwater nanofluid regulated by peristalsis using two different approaches for nanofluid. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001004.

Fonzi:2022:ECC

[1529] Nicola Fonzi, Vittorio Cavalieri, Alessandro De Gaspari, and Sergio Ricci. Extended computational capabilities for high-fidelity fluid–structure simulations. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001016.

Zhang:2022:HRL

[1530] Zhihong Zhang, Zhiqiang Li, and Yunke Wu. A hybrid regularized lattice Boltzmann model for convection– diffusion equation. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200103X.

Xu:2022:PBF

[1531] Changjin Xu, Dan Mu, Yuanlu Pan, Chaouki Aouiti, Yicheng Pang, and Lingyun Yao. Probing into bifurcation for fractional-order BAM neural networks concerning multiple time delays. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001028.

Krivovichev:2022:CAO

[1532] Gerasim V. Krivovichev. Computational analysis of one-dimensional models for simulation of blood flow in vascular networks. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001053.

Partee:2022:UML

[1533] Sam Partee, Matthew Ellis, Alessandro Rigazzi, Andrew E. Shao, Scott Bachman, Gustavo Marques, and Benjamin Robbins. Using Machine Learning at scale in numerical simulations with SmartSim: an application to ocean climate modeling. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001065.

Liu:2022:BLU

[1534] Wei Liu, Jiahao Chen, Chuni Liu, Xiaojuan Ban, Boyuan Ma, Hao Wang, Weihua Xue, and Yu Guo. Boundary learning by using weighted propagation in convolution network. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001077.

Matwiejew:2022:QFP

[1535] Edric Matwiejew and Jingbo B. Wang. QuOp MPI: a framework for parallel simulation of quantum variational algorithms. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001089.

Levada:2022:NCF

[1536] Alexandre L. M. Levada. On the numerical computation of Fisher–Rao based distances and KL-divergences between Gaussian random fields. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322001107.

Guo:2022:SIB

[1537] Xiao Guo, Zhenhua Wen, Shuhao Li, and Qingzhen Yang. Study on the influence of blade on electromagnetic scattering characteristics in the openended cavity. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001120.

Ahmad:2022:NPA

[1538] Usman Ahmad. A node pairing approach to secure the Internet of Things using machine learning. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001144.

Sadek:2022:AME

[1539] Lakhlifa Sadek and Hamad Talibi Alaoui. Application of MGA and EGA algorithms on large-scale linear systems of ordinary differential equations. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001156.

Nazimuddin:2022:OWP

[1540] A. K. M. Nazimuddin, M. Humayun Kabir, and M. Osman Gani. Oscillatory wave patterns and spiral breakup in the Brusselator model using numerical bifurcation analysis. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001132.

Liao:2022:NRL

[1541] Zhiqiang Liao, Kaijie Ma, Siyi Tang, Hiroyasu Yamahara, Munetoshi Seki, and Hitoshi Tabata. Nonbistable rectified linear unit-based gain-dissipative Ising spin network with stochastic resonance effect. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200117X.

Namvar:2022:SLB

[1542] Morteza Namvar and Sébastien Leclaire. Simple lattice Boltzmann method algorithm with low memory usage. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322001168.

Huang:2022:CCP

[1543] Yusheng Huang, Dong Chu, Yong Deng, and Kang Hao Cheong. The capacity constraint Physarum solver. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001181.

Kumar:2022:CNC

[1544] Mukesh Kumar, Shivansh Mishra, Rahul Deo Pandey, and Bhaskar Biswas. CFLP: a new cost based feature for link prediction in dynamic networks. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001193.

Ugurlu:2022:CAC

[1545] Onur Ugurlu. Comparative analysis of centrality measures for identifying critical nodes in complex networks. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001314.

Kumar:2022:PMS

[1546] Sandeep Kumar, Pierre Gosselet, Dengpeng Huang, Christian Weißenfels, and Peter Wriggers. Parallel multiphysics simulation for the stabilized Optimal Transportation Meshfree (OTM) method. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001302.

Iserte:2022:AUS

[1547] Sergio Iserte, Aina Macías, Raúl Martínez-Cuenca, Sergio Chiva, Roberto Paredes, and Enrique S. Quintana-Ortí. Accelerating urban scale simulations leveraging local spatial 3D structure. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001326.

Salimian:2022:NSB

[1548] Sina Salimian and Seyed Meysam Mousavi. A new scenario-based robust optimization approach for organ transplantation network design with queue condition and blood compatibility under climate change. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001338.

Shyaman:2022:ATF

[1549] V. P. Shyaman, A. Sreelakshmi, and Ashish Awasthi. An adaptive tailored finite point method for the generalized Burgers’ equations. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200134X.

Kovalchuk:2022:CSB

[1550] Sergey V. Kovalchuk, Valeria V. Krzhizhanovskaya, Maciej Paszyński, Dieter Kranzlmüller, Jack Dongarra, and Peter M. A. Sloot. Computational science for a better future. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322001351.

Wu:2022:HOC

[1551] Yu Wu, Yongbin Ge, and Lin Zhang. A high-order compact LOD difference method for solving the two-dimensional diffusion reaction equation with nonlinear source term. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001375.

Maulik:2022:PSD

[1552] Romit Maulik, Dimitrios K. Fytanidis, Bethany Lusch, Venkatram Vishwanath, and Saumil Patel. Python-FOAM: In-situ data analyses with OpenFOAM and Python. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750322001387.

Heshmat:2022:LCP

[1553] Samia Heshmat, Satoshi Tomioka, Shusuke Nishiyama, and Arata Hirokami. Localized compensator phase unwrapping algorithm based on flux conservable solver. Journal of Computational Science, 62:??, July 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001399.

Anonymous:2022:EBf

[1554] Anonymous. Editorial Board. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750322001703.

Anonymous:2022:Jb

[1555] Anonymous. July 2022. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Lin:2022:DEF

[1556] Jie Lin, Sheng Xin Zhang, Shao Yong Zheng, and Yong Mei Pan. Differential evolution with fusion of local and global search strategies. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001363.

Joseph:2022:SNH

[1557] Sobin Joseph, Lekhapriya Dheeraj Kashyap, and Shashi Jain. Shallow Neural Hawkes: Non-parametric kernel estimation for Hawkes processes. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001405.

Capdeville:2022:GKP

[1558] G. Capdeville. Gas kinetic principles in Navier–Stokes finite-volume solvers. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001417.

Velikajne:2022:PRP

[1559] Nina Velikajne and Miha Moškon. RhythmCount: a Python package to analyse the rhythmicity in count data. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001429.

Biswal:2022:NIN

[1560] Uddhaba Biswal, Snehashish Chakraverty, Bata Krushna Ojha, and Ahmed Kadhim Hussein. Numerical investigation on nanofluid flow between two inclined stretchable walls by Optimal Homotopy Analysis Method. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001430.

Dolezel:2022:CBP

[1561] Petr Dolezel, Pavel Skrabanek, Dominik Stursa, Bruno Baruque Zanon, Hector Cogollos Adrian, and Pavel Kryda. Centroid based person detection using pixelwise prediction of the position. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001442.

Szlachetka:2022:DMU

[1562] Marek Szlachetka, Dariusz Borkowski, and Jarosław Wa̧s. The downselection of measurements used for free space determination in ADAS. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001454.

Akbarimajd:2022:LAI

[1563] Adel Akbarimajd, Nicolas Hoertel, Mohammad Arafat Hussain, Ali Asghar Neshat, Mahmoud Marhamati, Mahdi Bakhtoor, and Mohammad Momeny. Learning-to-augment incorporated noise-robust deep CNN for detection of COVID-19 in noisy X-ray images. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001466.

Hejazi:2022:SBD

[1564] Taha-Hossein Hejazi. A scenario-based desirability function for correlated multi-response optimization problems considering modeling and implementation errors. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001478.

Kommadath:2022:PCS

[1565] Remya Kommadath, Debasis Maharana, Chinta Sivadurgaprasad, and Prakash Kotecha. Parallel computing strategies for Sanitized Teaching Learning Based Optimization. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322001491.

James:2022:TSG

[1566] Nick James and Howard Bondell. Temporal and spectral governing dynamics of Australian hydrological streamflow time series. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200148X.

Bhattacharjee:2022:CON

[1567] Sudipto Bhattacharjee, Banani Saha, Parthasarathi Bhattacharyya, and Sudipto Saha. Classification of obstructive and non-obstructive pulmonary diseases on the basis of spirometry using machine learning techniques. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001508.

Usman:2022:FAJ

[1568] Muhammad Usman, Weaam Alhejaili, Muhammad Hamid, and Nawab Khan. Fractional analysis of Jeffrey fluid over a vertical plate with time-dependent conductivity and diffusivity: a low-cost spectral approach. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200151X.

Tischler:2022:TEM

[1569] Ingo Tischler, Florian Weik, Robert Kaufmann, Michael Kuron, Rudolf Weeber, and Christian Holm. A thermalized electrokinetics model including stochastic reactions suitable for multiscale simulations of reaction– advection–diffusion systems. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001521.

Pavlovskii:2022:HGP

[1570] Vladislav V. Pavlovskii, Ilia V. Derevitskii, and Sergey V. Kovalchuk. Hybrid genetic predictive modeling for finding optimal multipurpose multicomponent therapy. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001545.

Dutta:2022:SFA

[1571] Soumya Dutta, Terece Turton, David Rogers, Jordan M. Musser, James Ahrens, and Ann S. Almgren. In situ feature analysis for large-scale multiphase flow simulations. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322001533.

Yesilkaya:2022:MLM

[1572] Bartu Yesilkaya, Matjaž Perc, and Yalcin Isler. Manifold learning methods for the diagnosis of ovarian cancer. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001557.

Pal:2022:IPM

[1573] S. K. Pal, Y. V. S. S. Sanyasiraju, and R. Usha. Investigation on the performance of meshfree RBF based method for the solution of thin film flows over topographies through depth-averaged Momentum Integral Model. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001569.

Kumar:2022:AGA

[1574] Siddharth Kumar and Jayadeep Pati.

Assessment of groundwater arsenic contamination level in Jharkhand, India using machine learning. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001570.

Jiang:2022:DVH

[1575] Chao Jiang, Jianfeng Pan, Yuejin Zhu, Jianxing Li, and Evans K. Quaye. Development and verification of a highspeed compressible reactive flow solver in OpenFOAM. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001582.

Kacher:2022:FHM

[1576] Yadvendra Kacher and Pitam Singh.

Fuzzy harmonic mean technique for solving fully fuzzy multi-objective transportation problem. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001594.

Fernandez:2022:TTF

[1577] Ivan Fernandez, Ricardo Quislant, Sonia Gonzalez-Navarro, Eladio Gutierrez, and Oscar Plata. TraTSA: a transprecision framework for efficient time series analysis. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001600.

Mollalign:2022:SMO

[1578] Demmelash Mollalign, Allen Mushi, and Berhanu Guta. Solving Multi-Objective Multilevel Programming problems using two-phase Intuitionistic Fuzzy Goal Programming method. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001612.

Yin:2022:DFD

[1579] Kangning Yin, Bin Wu, Rui Zhu, Lin Xiao, Zhuofu Tan, Guofeng He, Zhiguo Wang, and Guangqiang Yin. DLDP-FL: Dynamic local differential privacy federated learning method based on mesh network edge devices. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001648.

Singh:2022:ODM

[1580] Randhir Singh and Mehakpreet Singh. An optimal decomposition method for analytical and numerical solution of third-order Emden–Fowler type equations. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001624.

Abbas:2022:ISN

[1581] Zeesham Abbas, Kisa Fatima, Syed Hassan Abbas Jaffery, Asif Ali, Hafiz Hamid Raza, Shabbir Muhammad, H. Algarni, Sajjad Hussain, and Jongwan Jung. Ab-initio study of Nbbased complex materials: a new class of materials for optoelectronic applications. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200165X.

Qureshi:2022:TER

[1582] Sania Qureshi, Higinio Ramos, Amanullah Soomro, and Evren Hincal. Time-efficient reformulation of the Lobatto III family of order eight. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001636.

Wichrowski:2022:MFM

[1583] Michał Wichrowski and Piotr Krzyżanowski.

A matrix-free multilevel preconditioner for the generalized Stokes problem with discontinuous viscosity. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001661.

Zade:2022:IRF

[1584] B. Mohammad Hasani Zade and N. Mansouri. Improved red fox optimizer with fuzzy theory and game theory for task scheduling in cloud environment. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001776.

Behnoudfar:2022:VSH

[1585] Pouria Behnoudfar, Victor Manuel Calo, Marcin Łoś, Paweł Maczuga, and Maciej Paszyński. A variational splitting of high-order linear multistep methods for heat transfer and advection–diffusion parabolic problems. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001788.

Yang:2022:NSO

[1586] Jianhui Yang and Mingjie Ma. The numerical solution of one-dimensional discrete asset pricing model based on the improved trigonometric extreme learning machine. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200179X.

Portella:2022:TTA

[1587] Felipe Portella, David Buchaca, José Roberto Rodrigues, and Josep Ll. Berral. TunaOil: a tuning algorithm strategy for reservoir simulation workloads. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001806.

Castillo-Reyes:2022:TMP

[1588] Octavio Castillo-Reyes, Adrian Amor-Martin, Arnaud Botella, Pierre Anquez, and Luis Emilio García-Castillo. Tailored meshing for parallel 3D electromagnetic modeling using high-order edge elements. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001818.

Podda:2022:FAD

[1589] Alessandro Sebastian Podda, Riccardo Balia, Silvio Barra, Salvatore Carta, Gianni Fenu, and Leonardo Piano. Fully-automated deep learning pipeline for segmentation and classification of breast ultrasound images. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200182X.

Migallon:2022:PMU

[1590] Violeta Migallón, Francisco J. Navarro-González, Héctor Penadés, José Penadés, and Yolanda Villacampa. A parallel methodology using radial basis functions versus machine learning approaches applied to environmental modelling. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001831.

Curado:2022:CMD

[1591] Manuel Curado, Rocio Rodriguez, Fernando Terroso-Saenz, Leandro Tortosa, and Jose F. Vicent. A centrality model for directed graphs based on the Two-Way-Random Path and associated indices for characterizing the nodes. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001843.

Yahia:2022:PCM

[1592] Eman Yahia and Kannan N. Premnath. Preconditioned central moment lattice Boltzmann method on a rectangular lattice grid for accelerated computations of inhomogeneous flows. Journal of Computational Science, 63:??, September 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001855.

Jung:2022:VDL

[1593] Seungwon Jung, Yoona Noh, Jaeuk Moon, and Eenjun Hwang. VAPER: a deep learning model for explainable probabilistic regression. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001879.

Chen:2022:SAN

[1594] Lei Chen, Yuan Li, Xingye Deng, Zhaohua Liu, Mingyang Lv, and Tingqin He. Semantic-aware network embedding via optimized random walk and paragaraph2vec. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001880.

Treesatayapun:2022:DDF

[1595] Chidentree Treesatayapun. Datadriven fault-tolerant control with fuzzy-rules equivalent model for a class of unknown discrete-time MIMO systems and complex coupling. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001892.

Anonymous:2022:EBg

[1596] Anonymous. Editorial Board. Journal of Computational Science, 63: ??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032200206X.

Anonymous:2022:S

[1597] Anonymous. September 2022. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Carlson:2022:MEE

[1598] Jared Carlson, Sean Couch, Brian W. O’Shea, and Carlo Graziani. Memoryefficient emulation of physical tabular data using quadtree decomposition. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001867.

Hussain:2022:HHA

[1599] Syed Mujtiba Hussain and Gh Rasool Begh. Hybrid heuristic algorithm for cost-efficient QoS aware task scheduling in fog–cloud environment. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001909.

Loveland:2022:EFW

[1600] Mark Loveland, Eirik Valseth, Matt Lukac, and Clint Dawson. Extending FEniCS to work in higher dimensions using tensor product finite elements. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001922.

Wang:2022:ECS

[1601] Yihong Wang, Tinggan Yang, and Lina Chang. An edge-centered scheme for anisotropic diffusion problems with discontinuities on distorted quadrilateral meshes. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001910.

Kurniabudi:2022:IAD

[1602] Kurniabudi, Deris Stiawan, Darmawijoyo, Mohd Yazid Bin Idris, Sarjon Defit, Yaya Sudarya Triana, and Rahmat Budiarto. Improvement of attack detection performance on the Internet of Things with PSO-search and random forest. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001934.

Sitompul:2022:OAP

[1603] Yos Panagaman Sitompul, Takayuki Aoki, Seiya Watanabe, and Tomohiro Takaki. An ordered active parameter tracking method for efficient multiphase field simulations. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322001946.

Galashev:2022:NSF

[1604] Alexander Y. Galashev. Numerical simulation of functioning a silicene anode of a lithium-ion battery. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001958.

Izci:2022:MSM

[1605] Davut Izci, Serdar Ekinci, Erdal Eker, and Ayşen Demirören. Multi-strategy modified INFO algorithm: Performance analysis and application to functional electrical stimulation system. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200196X.

Kermani:2022:IGS

[1606] Ali Golzadeh Kermani, Ali Kamandi, and Ali Moeini. Integrating graph structure information and node attributes to predict protein-protein interactions. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322001971.

Nguyen-Vu:2022:EFL

[1607] Long Nguyen-Vu and Souhwan Jung. Empirical feature learning in applicationbased samples: a case study. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322001983.

Kchaou:2022:PTS

[1608] Hamdi Kchaou, Zied Kechaou, and Adel M. Alimi. A PSO task scheduling and IT2FCM fuzzy data placement strategy for scientific cloud workflows. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322001995.

Ibraheem:2022:NAS

[1609] Ghada H. Ibraheem, Mustafa Turkyilmazoglu, and M. A. Al-Jawary. Novel approximate solution for fractional differential equations by the optimal variational iteration method. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322002009.

Gambiroza:2022:DMF

[1610] Jelena Čulić Gambiroža, Toni Mastelić, Ivana Nižetić Kosović, and Mario Čagalj. Dynamic monitoring frequency for energy-efficient data collection in Internet of Things. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002010.

Yin:2022:UDM

[1611] Jia Yin, Yang hao Chan, Felipe H.

da Jornada, Diana Y. Qiu, Steven G. Louie, and Chao Yang. Using dynamic mode decomposition to predict the dynamics of a two-time non-equilibrium Green’s function. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002022.

Muller:2022:MQD

[1612] Axel Muller, Metod Saniga, Alain Giorgetti, Henri de Boutray, and Frédéric Holweck. Multi-qubit doilies: Enumeration for all ranks and classification for ranks four and five. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322002125.

Taghikhah:2022:MAA

[1613] Firouzeh Taghikhah, Alexey Voinov, Tatiana Filatova, and J. Gareth Polhill. Machine-assisted agentbased modeling: Opening the black box. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002137.

Sugiarto:2022:GCA

[1614] Hendrik Santoso Sugiarto. Graph convolution approach for labor market analysis. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002149.

Maczuga:2022:GGB

[1615] Paweł Maczuga, Albert Oliver-Serra, Anna Paszyńska, Eirik Valseth, and Maciej Paszyński. Graph-grammar based algorithm for asteroid tsunami simulations. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002150.

Wozniak:2022:SEC

[1616] Maciej Woźniak, Anna Szyszka, and Sergio Rojas. A study of efficient concurrent integration methods of B-spline basis functions in IGA-FEM. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002162.

Godzik:2022:AAH

[1617] Mateusz Godzik, Jacek Dajda, Marek Kisiel-Dorohinicki, Aleksander Byrski, Leszek Rutkowski, Patryk Orzechowski, Joost Wagenaar, and Jason H. Moore. Applying autonomous hybrid agent-based computing to difficult optimization problems. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002174.

Thies:2022:TPS

[1618] Jonas Thies, Moritz Travis Hof, Matthias Zimmermann, and Maxim Efremov. Tensor product scheme for computing bound states of the quantum mechanical three-body problem. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002186.

Toosi:2022:FAS

[1619] Ramin Toosi, Mohammadreza Sadeghi, Hossein B. Yazdi, and Mohammad Ali Akhaee. Fast and accurate spectral clustering via augmented Lagrangian. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002198.

Deng:2022:BPI

[1620] Quanling Deng, Pouria Behnoudfar, and Victor Calo. A boundarypenalized isogeometric analysis for second-order hyperbolic equations. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322002204.

Chen:2022:MCF

[1621] Yanzhe Chen and Zongxia Xie. Multichannel fusion graph neural network for multivariate time series forecasting. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002216.

Karimpour:2022:BCI

[1622] M. Karimpour and M. Rezghi. A block column iteration for nonnegative matrix factorization. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002228.

Thorimbert:2022:LMR

[1623] Y. Thorimbert, D. Lagrava, O. Malaspinas, B. Chopard, C. Coreixas, J. de Santana Neto, R. Deiterding, and J. Latt. Local mesh refinement sensor for the lattice Boltzmann method. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200223X.

Mou:2022:TLD

[1624] Hanlin Mou and Junsheng Yu. Transfer learning with DWT based clustering for blood pressure estimation of multiple patients. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002241.

Ferrada:2022:SAB

[1625] Héctor Ferrada. A sorting algorithm based on ordered block insertions. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002253.

Talpur:2022:ODN

[1626] Noureen Talpur, Said Jadid Abdulkadir, Hitham Alhussian, Mohd Hilmi Hasan, and Mohd Hafizul Afifi Abdullah. Optimizing deep neurofuzzy classifier with a novel evolutionary arithmetic optimization algorithm. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002265.

Moazeni:2022:IGT

[1627] Haniye Moazeni, Behrouz Arbab Shirani, and Seyed Reza Hejazi. An integrated game theoretic model and network data envelopment analysis to determine optimal export allocation to industrial units (a case study of the stone industry). Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002277.

Erfanian:2022:SIP

[1628] Majid Erfanian, Hamed Zeidabadi, and Omid Baghani. Solving an inverse problem for a time-fractional advection-diffusion equation with variable coefficients by rationalized Haar wavelet method. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002289.

Halder:2022:EPA

[1629] Sudip Halder, Bimal Kumar Dora, and Sunil Bhat. An Enhanced Pathfinder Algorithm based MCSA for rotor breakage detection of induction motor. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002290.

Los:2022:HLP

[1630] Marcin Łoś, Robert Schaefer, and Maciej Smołka. How to lighten parametric inverse computations if the misfit is non-convex and the forward solver needs stabilization. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002319.

Talha:2022:IPA

[1631] Adnane Talha, Anas Bouayad, and Mohammed Ouçamah Cherkaoui Malki. An improved pathfinder algorithm using opposition-based learning for tasks scheduling in cloud environment. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002320.

Sun:2022:AOM

[1632] Shuping Sun, Yaonan Tong, Biqiang Zhang, Bowen Yang, Peiguang He, Wei Song, Wenbo Yang, Yilin Wu, and Guangyu Liu. An adaptive optimization method for estimating the number of components in a Gaussian mixture model. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002332.

Tian:2022:OMM

[1633] Xiaoming Tian and Wu Jia. Optimal matching method based on rare plants in opportunistic social networks. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002344.

Abdolmaleki:2022:AAC

[1634] Nasim Abdolmaleki, Leyli Mohammad Khanli, Mahdi Hashemzadeh, and Shahin Pourbahrami. ACQC: Apollonius Circle-based Quantum Clustering. Journal of Computational Science, 64:??, October 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002368.

Anonymous:2022:EBh

[1635] Anonymous. Editorial Board. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750322002502.

Anonymous:2022:O

[1636] Anonymous. October 2022. Journal of Computational Science, 64:??, October 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Zhang:2022:MLF

[1637] Qian Zhang, Anran Huang, Lianyou Shao, Peiliang Wu, Ali Asghar Heidari, Zhennao Cai, Guoxi Liang, Huiling Chen, Fahd S. Alotaibi, Majdi Mafarja, and Jinsheng Ouyang. A machine learning framework for identifying influenza pneumonia from bacterial pneumonia for medical decision making. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002307.

Fossum:2022:VRT

[1638] Kristian Fossum, Sergey Alyaev, Jan Tveranger, and Ahmed H. Elsheikh. Verification of a real-time ensemblebased method for updating earth model based on GAN. Journal of Computational Science, 65:??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002356.

Tao:2022:LBS

[1639] Shi Tao, Liang Wang, Qing He, Jiechao Chen, and Jiahong Luo. Lattice Boltzmann simulation of complex thermal flows via a simplified immersed boundary method. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200237X.

Kumari:2022:PDG

[1640] Baby Kumari and Pankaj Singh Dholaniya. Parkinson’s disease gene prioritising using an efficient and biologically appropriate networkbased consensus strategy. Journal of Computational Science, 65: ??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322002381.

Mountris:2022:MFP

[1641] Konstantinos A. Mountris, Leiting Dong, Yue Guan, Satya N. Atluri, and Esther Pueyo. A meshless fragile points method for the solution of the monodomain model for cardiac electrophysiology simulation. Journal of Computational Science, 65:??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002393.

Zhang:2022:DBH

[1642] Zishuang Zhang, Chenxi Sun, and Zhi-Ping Liu. Discovering biomarkers of hepatocellular carcinoma from singlecell RNA sequencing data by cooperative games on gene regulatory network. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200240X.

Zhu:2022:NSD

[1643] Yingpeng Zhu, Yikun Wei, Zhengdao Wang, Rongyang Wang, Chuanyu Wu, Jianneng Chen, and Junhua Tong. Numerical simulation for deformation characteristic of tea shoot under negative pressure guidance by the immersed boundary–lattice Boltzmann method. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002411.

Tyler:2022:INM

[1644] Jack Tyler and Alexander Wittig. An improved numerical method for hyperbolic Lagrangian Coherent Structures using Differential Algebra. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002423.

Kurz:2022:DRL

[1645] Marius Kurz, Philipp Offenhäuser, Dominic Viola, Oleksandr Shcherbakov, Michael Resch, and Andrea Beck. Deep reinforcement learning for computational fluid dynamics on HPC systems. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002435.

Liu:2022:NGN

[1646] Zimian Liu, Han Qiu, Wei Guo, Junhu Zhu, and Qingxian Wang. NIE-GAT: node importance evaluation method for inter-domain routing network based on graph attention network. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002447.

Singh:2022:CFL

[1647] Priyanka Singh and Rahul Kottath. Chaos follow the leader algorithm: Application to data classification. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002459.

Cejnek:2022:PPO

[1648] Matous Cejnek and Jan Vrba. Padasip: an open-source Python toolbox for adaptive filtering. Journal of Computational Science, 65:??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002460.

Liu:2022:EDA

[1649] Xin Liu and Jason Frank. Ensemble data assimilation using optimal control in the Wasserstein metric. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032200254X.

Li:2022:IDD

[1650] Sichao Li, Jonathan Y. C. Ting, and Amanda S. Barnard. The impact of domain-driven and datadriven feature selection on the inverse design of nanoparticle catalysts. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002551.

Xie:2022:SET

[1651] Jianqiang Xie, Quanxiang Wang, and Zhiyue Zhang. Stable and efficient time second-order difference schemes for fractional Klein–Gordon–Zakharov system. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002605.

Chakrabarti:2022:SFC

[1652] Arnab Chakrabarti and Anindya S.

Chakrabarti. Sparsistent filtering of comovement networks from highdimensional data. Journal of Computational Science, 65:??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002617.

Wang:2022:EEH

[1653] Qingyu Wang, Takuji Nakashima, Chenguang Lai, Bo Hu, Xinru Du, Zhongzheng Fu, Taiga Kanehira, Yasufumi Konishi, Hiroyuki Okuizumi, and Hidemi Mutsuda. Enhanced expected hypervolume improvement criterion for parallel multi-objective optimization. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002629.

Gao:2022:PAM

[1654] Xiang Gao, Fan Zhou, Kedi Xu, Xiang Tian, and Yaowu Chen. A parallel algorithm for maximal cliques enumeration to improve hypergraph construction. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002642.

Wang:2022:TPC

[1655] Xinjie Wang, Siyuan Zhu, Yundong Guo, Peng Han, Yucheng Wang, Zhiqiang Wei, and Xiaogang Jin. TransFlowNet: a physics-constrained transformer framework for spatiotemporal super-resolution of flow simulations. Journal of Computational Science, 65:??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002654.

Anonymous:2022:EBi

[1656] Anonymous. Editorial Board. Journal of Computational Science, 65: ??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322002770.

Anonymous:2022:N

[1657] Anonymous. November 2022. Journal of Computational Science, 65:??, November 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Sgura:2023:DLB

[1658] Ivonne Sgura, Luca Mainetti, Francesco Negro, Maria Grazia Quarta, and Benedetto Bozzini. Deep-learning based parameter identification enables rationalization of battery material evolution in complex electrochemical systems. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002599.

Zhou:2023:MLP

[1659] Li Zhou, Chao Yang, Weiguo Gao, Talita Perciano, Karen M. Davies, and Nicholas K. Sauter. A machine learning pipeline for membrane segmentation of cryo-electron tomograms. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002630.

Iftikhar:2023:EGA

[1660] Babar Iftikhar, Tariq Javed, and Muhammad Arshad Siddiqu. Entropy generation analysis during MHD mixed convection flow of non-Newtonian fluid saturated inside the square cavity. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002666.

deMiras:2023:FCF

[1661] J. Ruiz de Miras, M. A. Posadas, A. J. Ibáñez-Molina, M. F. Soriano, and S. Iglesias-Parro. Fast computation of fractal dimension for 2D, 3D and 4D data. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002678.

Azad:2023:ASR

[1662] Fatemeh Azad, Maryam Zare, Mahmood Amiri, and Georgios A. Keliris. Analysis of the spike responses in the neuromorphic implementation of the two-compartmental model of hippocampal pyramidal neuron. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200268X.

Macias-Medri:2023:SMP

[1663] A. E. Macias-Medri, G. M. Viswanathan, C. E. Fiore, M. Koehler, and M. G. E. da Luz. Speedup of the Metropolis protocol via algorithmic optimization. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002691.

McLaughlin:2023:NPA

[1664] Benjamin McLaughlin and Sung Ha Kang. A new parallel adaptive clustering and its application to streaming data. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002708.

Yan:2023:CDP

[1665] Jianglong Yan, Leandro Anghinoni, Yu-Tao Zhu, Weiguang Liu, Gen Li, Qiusheng Zheng, and Liang Zhao. Characterizing data patterns with core–periphery network modeling. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200271X.

Zhang:2023:OCA

[1666] Hong Zhang and Emil M. Constantinescu. Optimal checkpointing for adjoint multistage time-stepping schemes. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002721.

Hacat:2023:ACD

[1667] Gülnur Haçat, Mine Akbas, and Aytekin Çıbık. Analysis of continuous data assimilation scheme for the Navier–Stokes equations using variational multiscale method. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322002733.

Wang:2023:NCR

[1668] Dan Wang, Feng Tian, and Daijun Wei. A new centrality ranking method for multilayer networks. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322002836.

Fahim:2023:VDB

[1669] Ahmed Fahim. A varied densitybased clustering algorithm. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750322002848.

Chatterjee:2023:DCD

[1670] Sukanya Chatterjee, Jishnu Roychowdhury, and Anilesh Dey. D-Cov19Net: a DNN based COVID-19 detection system using lung sound. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032200285X.

Sikdar:2023:NSD

[1671] Prabir Sikdar, Sunil Manohar Dash, and Kalyan Prasad Sinhamahapatra. A numerical study on the drag reduction and wake regime control of the tandem circular cylinders using splitter plates. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002861.

Lopes:2023:BAE

[1672] Tales Lopes, Victor Ströele, Regina Braga, José Maria N. David, and Michael Bauer. A broad approach to expert detection using syntactic and semantic social networks analysis in the context of Global Software Development. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002873.

Debnath:2023:MEP

[1673] Surajit Debnath, Prahlad Majumdar, Susmita Sarkar, and Uttam Ghosh. Memory effect on prey–predator dynamics: Exploring the role of fear effect, additional food and anti-predator behaviour of prey. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002885.

Daubner:2023:MFM

[1674] Simon Daubner, Martin Reder, Nishant Prajapati, Daniel Schneider, and Britta Nestler. Multiphasefield modelling of anisotropic elasticity at finite deformation in Eulerian space. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002897.

Wang:2023:NBA

[1675] Chunfeng Wang, Wenxin Song, and Peiping Shen. A new bat algorithm based on a novel topology and its convergence. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002903.

Dennunzio:2023:APS

[1676] Alberto Dennunzio, Enrico Formenti, Luciano Margara, and Sara Riva. An algorithmic pipeline for solving equations over discrete dynamical systems modelling hypothesis on real phenomena. Journal of Computational Science, 66:??, January 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002915.

Yuan:2023:DOC

[1677] Yujing Yuan and Dong Liang. Developing the optimization control of pollution constrained by the convection diffusion equations. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002927.

Anonymous:2023:EBa

[1678] Anonymous. Editorial Board. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S187775032300011X.

Anonymous:2023:Ja

[1679] Anonymous. January 2023. Journal of Computational Science, 66:??, January 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Gao:2023:MOS

[1680] Yifan Gao, Bo Yang, Shilong Wang, Guang Fu, and Peng Zhou. A multi-objective service composition method considering the interests of tri-stakeholders in cloud manufacturing based on an enhanced jellyfish search optimizer. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002939.

An:2023:FSG

[1681] Phan Thanh An, Nguyen Thi Le, Le Hong Trang, and Raymond Chi-Wing Wong. Finding shortest gentle paths on polyhedral terrains by the method of multiple shooting. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750322002940.

Spiridonov:2023:RFM

[1682] Sergey Spiridonov and Alexey A.

Shcherbakov. Reformulated Fourier Modal Method with improved near field computations. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750322002952.

Ding:2023:IRA

[1683] Jun Ding, Yinxuan Zhou, Xia Huang, Kun Song, Shiqing Lu, and Lusheng Wang. An improved RRT* algorithm for robot path planning based on path expansion heuristic sampling. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750322002964.

Lin:2023:IBH

[1684] Shiwei Lin, Ang Liu, Jianguo Wang, and Xiaoying Kong. An intelligencebased hybrid PSO-SA for mobile robot path planning in warehouse. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322002976.

Huang:2023:ECM

[1685] Dongwei Huang, Yuelin Zhao, Keqi Ye, Feng Wu, Hongwu Zhang, and Wanxie Zhong. The efficient calculation methods for stochastic nonlinear transient heat conduction problems. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322002988.

Ljubicic:2023:CPM

[1686] Karmela Ljubičić, Andro Merćep, and Zvonko Kostanjčar. Churn prediction methods based on mutual customer interdependence. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032200299X.

Stern:2023:TPR

[1687] Kevin E. Stern and Padmanabhan K. Aravind. Two-point resistances of symmetric polyhedral networks. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000017.

Guha:2023:DEO

[1688] Ritam Guha, Kushal Kanti Ghosh, Suman Kumar Bera, Ram Sarkar, and Seyedali Mirjalili. Discrete equilibrium optimizer combined with simulated annealing for feature selection. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323000029.

Varli:2023:MCE

[1689] Muhammet Varlı and Hakan Yılmaz.

Multiple classification of EEG signals and epileptic seizure diagnosis with combined deep learning. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323000030.

Peng:2023:DCD

[1690] Chao-Chung Peng and Wang Cheng-Yu. Design of constrained dynamic path planning algorithms in large-scale 3D point cloud maps for UAVs. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000042.

Sanchez-Curto:2023:IOP

[1691] J. Sánchez-Curto and P. Chamorro-Posada. The inherent overlapping in the parallel calculation of the Laplacian. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000054.

Algasov:2023:FAS

[1692] A. S. Algasov, S. A. Guda, V. I. Kolesnikov, V. V. Ilicheva, and A. V. Soldatov. Fast adaptive sampling with operation time control. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323000066.

Han:2023:HFE

[1693] Zhuang Han, Lian-Wen Sun, Xiao Yang, Xin-Tong Wu, and Yu-Bo Fan. A hierarchical 3D finite element model of osteocyte: the spectrin membrane skeleton in mechanical transmission. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000078.

Wang:2023:IIU

[1694] Chao Wang, Liangliang Ma, Lei Ma, Joel Weijia Lai, Jie Zhao, Lu Wang, and Kang Hao Cheong. Identification of influential users with cost minimization via an improved moth flame optimization. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000157.

Aksoy:2023:PMD

[1695] Necati Aksoy and Istemihan Genc. Predictive models development using gradient boosting based methods for solar power plants. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000182.

Cardoso:2023:OEK

[1696] Douglas O. Cardoso and Thalis D.

Galeno. Online evaluation of the Kolmogorov–Smirnov test on arbitrarily large samples. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000194.

Tuncel:2023:EGR

[1697] Mehmet Tunçel and Ahmet Duran. Effectiveness of grid and random approaches for a model parameter vector optimization. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000200.

Heydari:2023:NSD

[1698] M. H. Heydari, M. Razzaghi, and D. Baleanu. Numerical solution of distributed-order time fractional Klein–Gordon–Zakharov system. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750323000212.

Roozbahani:2023:CDM

[1699] Zahra Roozbahani, Jalal Rezaeenour, and Ali Katanforoush. Community detection in multi-relational directional networks. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000224.

Mycka:2023:THL

[1700] Jan Mycka, Adam Żychowski, and Jacek Mańdziuk. Toward humanlevel tonal and modal melody harmonizations. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000236.

Zheng:2023:NSI

[1701] Jiming Zheng and Jun Liu. A new scheme for identifying important nodes in complex networks based on generalized degree. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000248.

Chen:2023:LSE

[1702] Yu Chen, Ruihua Zhang, Kaijun Yang, Y. F. Yuan, and Boyuan Xu. A linearization solution for elastic-plastic torsion problems by edge-based smoothed finite element method. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032300025X.

Alsaker:2023:MRT

[1703] Melody Alsaker, Jennifer L. Mueller, and Andreas Stahel. A multithreaded real-time solution for 2D EIT reconstruction with the D-bar algorithm. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000273.

John:2023:ADA

[1704] Nicholas John and Nishant Malik. Automated discovery of analytical models for epidemic dynamics on coevolving networks. Journal of Computational Science, 67:??, March 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000285.

Anonymous:2023:EBb

[1705] Anonymous. Editorial Board. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750323000509.

Anonymous:2023:Ma

[1706] Anonymous. March 2023. Journal of Computational Science, 67:??, March 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Shubyn:2023:FLI

[1707] Bohdan Shubyn, Daniel Kostrzewa, Piotr Grzesik, Paweł Benecki, Taras Maksymyuk, Vaidy Sunderam, Jia-Hao Syu, Jerry Chun-Wei Lin, and Dariusz Mrozek. Federated learning for improved prediction of failures in Autonomous Guided Vehicles. Journal of Computational Science, 68:??, April 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323000169.

Bielak:2023:DLG

[1708] Piotr Bielak, Jakub Binkowski, Albert Sawczyn, Katsiaryna Viarenich, Daria Puchalska, and Tomasz Kajdanowicz. A deeper look at Graph Embedding RetroFitting. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032300039X.

Zychowski:2023:CPS

[1709] Adam Żychowski and Jacek Mańdziuk.

Coevolution of players strategies in security games. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000406.

Heisler:2023:MDD

[1710] Eric Heisler, Aadesh Deshmukh, Sandip Mazumder, Ponnuswamy Sadayappan, and Hari Sundar. Multidiscretization domain specific language and code generation for differential equations. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000418.

Khalili:2023:ILE

[1711] Behnam Khalili and Mahmud Ashrafizaadeh. Implementation of Lees–Edwards periodic boundary conditions for threedimensional lattice Boltzmann simulation of particle dispersions under shear flow. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032300042X.

Gao:2023:LTC

[1712] Peng Gao, Dongxing Tao, Yuan Yuan, and Shikui Dong. A low-time complexity semi-analytic Monte Carlo radiative transfer model: Application to optical characteristics of complex spatial targets. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000431.

Aseeri:2023:ERB

[1713] Ahmad O. Aseeri. Effective RNN-based forecasting methodology design for improving short-term power load forecasts: Application to large-scale powergrid time series. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000443.

Bajer:2023:PCD

[1714] Dražen Bajer. Parameter control for differential evolution by storage of successful values at an individual level. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000455.

Fujita:2023:CCC

[1715] Kohei Fujita, Takuma Yamaguchi, Yuma Kikuchi, Tsuyoshi Ichimura, Muneo Hori, and Lalith Maddegedara. Calculation of crosscorrelation function accelerated by TensorFloat-32 tensor core operations on NVIDIA’s Ampere and Hopper GPUs. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000467.

Seredynski:2023:ESO

[1716] Franciszek Seredyński, Tomasz Kulpa, and Rolf Hoffmann. Evolutionary selfoptimization of large CA-based multiagent systems. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000546.

Clempner:2023:TLA

[1717] Julio B. Clempner. A tri-level approach for computing Stackelberg Markov game equilibrium: Computational analysis. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000558.

Izgi:2023:MLD

[1718] Burhaneddin İzgi, Murat Özkaya, Nazım Kemal Üre, and Matjaž Perc. Machine learning driven extended matrix norm method for the solution of large-scale zero-sum matrix games. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000571.

Dover:2023:AAM

[1719] Kathryn Dover, Zixuan Cang, Anna Ma, Qing Nie, and Roman Vershynin. AVIDA: an alternating method for visualizing and integrating data. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000583.

Yesilkaya:2023:PCA

[1720] Bartu Yesilkaya, Ebru Sayilgan, Yilmaz Kemal Yuce, Matjaž Perc, and Yalcin Isler. Principal component analysis and manifold learning techniques for the design of brain-computer interfaces based on steady-state visually evoked potentials. Journal of Computational Science, 68:??, April 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000601.

Anonymous:2023:EBc

[1721] Anonymous. Editorial Board. Journal of Computational Science, 68:??, April 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750323000789.

Anonymous:2023:A

[1722] Anonymous. April 2023. Journal of Computational Science, 68:??, April 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

He:2023:SAF

[1723] Yi He, Judy P. Yang, and Jie Yu. Surrogate-assisted finite element model updating for detecting scour depths in a continuous bridge. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032300056X.

Wise:2023:SMV

[1724] Sarah Wise, Sveta Milusheva, Sophie Ayling, and Robert Manning Smith. Scale matters: Variations in spatial and temporal patterns of epidemic outbreaks in agent-based models. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323000595.

Salunkhe:2023:PIM

[1725] Amol Salunkhe, Dwyer Deighan, Paul E. DesJardin, and Varun Chandola. Physics informed machine learning for chemistry tabulation. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323000613.

Guo:2023:LPM

[1726] Feipeng Guo, Wei Zhou, Zifan Wang, Chunhua Ju, Shaobo Ji, and Qibei Lu. A link prediction method based on topological nearest-neighbors similarity in directed networks. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000625.

Biglarian:2023:NSE

[1727] Hassan Biglarian and Mohsen Salimi. Numerical solution of extended blackoil model incorporating capillary effects based on a high-resolution central scheme. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000637.

Phab:2023:FAC

[1728] Luca Phab, Stéphane Louise, and Renaud Sirdey. First attempts at cryptanalyzing a (toy) block cipher by means of quantum optimization approaches. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000649.

Zhu:2023:FMF

[1729] Jianqing Zhu, Juncai He, Lian Zhang, and Jinchao Xu. FV-MgNet: Fully connected V -cycle MgNet for interpretable time series forecasting. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323000650.

Phillips:2023:AIM

[1730] C. M. Phillips, E. A. B. F. Lima, C. Wu, A. M. Jarrett, Z. Zhou, N. Elshafeey, J. Ma, G. M. Rauch, and T. E. Yankeelov. Assessing the identifiability of model selection frameworks for the prediction of patient outcomes in the clinical breast cancer setting. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000662.

Georgiev:2023:CIS

[1731] Slavi G. Georgiev and Lubin G.

Vulkov. Coefficient identification in a SIS fractional-order modelling of economic losses in the propagation of COVID-19. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000674.

Xu:2023:FTT

[1732] Qin Xu, Chunmei Zhang, Huiling Chen, and Hui Yang. Finite-time topology identification of stochastic delayed coupled systems on multiweighted networks based on graphtheoretic method. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000698.

Ahmadi:2023:DHL

[1733] Bahman Ahmadi, Juan S. Giraldo, and Gerwin Hoogsteen. Dynamic Hunting Leadership optimization: Algorithm and applications. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000704.

Prottasha:2023:ILL

[1734] Nusrat Jahan Prottasha, Saydul Akbar Murad, Abu Jafar Md Muzahid, Masud Rana, Md Kowsher, Apurba Adhikary, Sujit Biswas, and Anupam Kumar Bairagi. Impact learning: a learning method from feature’s impact and competition. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000716.

Bernaschi:2023:SCN

[1735] Massimo Bernaschi, Alessandro Celestini, Marco Cianfriglia, Stefano Guarino, Giuseppe F. Italiano, Enrico Mastrostefano, and Lena Rebecca Zastrow. Seeking critical nodes in digraphs. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000728.

Paul:2023:FGA

[1736] Sri Raj Paul, Akihiro Hayashi, Kun Chen, Youssef Elmougy, and Vivek Sarkar. A Fine-grained Asynchronous Bulk Synchronous parallelism model for PGAS applications. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323000741.

Sk:2023:DEE

[1737] Nazmul Sk, Samares Pal, Prahlad Majumdar, and Bapin Mondal. Dynamics of an eco-epidemiological system: Predators get infected in two paths. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000832.

Cheng:2023:EKF

[1738] Meiling Cheng, Fangxin Fang, Ionel M.

Navon, and Christopher Pain. Ensemble Kalman filter for GAN-ConvLSTM based long lead-time forecasting. Journal of Computational Science, 69:??, May 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000844.

Bashir:2023:IOR

[1739] Rab Nawaz Bashir, Faizan Ahmed Khan, Arfat Ahmad Khan, Muhammad Tausif, Muhammad Zahid Abbas, Malik Muhammad Ali Shahid, and Nasrullah Khan. Intelligent optimization of Reference Evapotranspiration (ETo) for precision irrigation. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323000856.

Mahata:2023:IRE

[1740] Sourav Mahata and Bijoy Krishna Debnath. The impact of R&D expenditures and screening in an economic production rate (EPR) inventory model for a flawed production system with imperfect screening under an intervalvalued environment. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032300087X.

Wang:2023:CCC

[1741] Kai Wang, Xiaoping Wang, and Deli Zhang. Constructing continuous curves on point-cloud surfaces with Directed Projection Operator. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000881.

Dai:2023:IIN

[1742] Bitao Dai, Shuo Qin, Suoyi Tan, Chuchu Liu, Jianhong Mou, Hongzhong Deng, Fredrik Liljeros, and Xin Lu. Identifying influential nodes by leveraging redundant ties. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032300090X.

Anonymous:2023:EBd

[1743] Anonymous. Editorial Board. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750323001060.

Anonymous:2023:Mb

[1744] Anonymous. May 2023. Journal of Computational Science, 69:??, May 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Keivanian:2023:FAM

[1745] Farshid Keivanian, Raymond Chiong, Ali R. Kashani, and Amir H. Gandomi. A fuzzy adaptive metaheuristic algorithm for identifying sustainable, economical, and earthquakeresistant reinforced concrete cantilever retaining walls. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000388.

Fatima:2023:NGC

[1746] Ubaida Fatima, Saman Hina, and Muhammad Wasif. A novel global clustering coefficient-dependent degree centrality (GCCDC) metric for large network analysis using real-world datasets. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000686.

Kemmer:2023:CAP

[1747] Thomas Kemmer, Sebastian Hack, Bertil Schmidt, and Andreas Hildebrandt. CUDA-accelerated protein electrostatics in linear space. Journal of Computational Science, 70:??, June 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750323000820.

Muni:2023:BDM

[1748] Manoj Kumar Muni, Saroj Kumar, Chinmaya Sahu, Prasant Ranjan Dhal, Dayal R. Parhi, and Sanjay Kumar Patra. Better decision-making strategy with target seeking approach of humanoids using hybridized SOARANNfuzzy technique. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000868.

Jin:2023:GDS

[1749] Li-Li Jin and Hou-Biao Li. Greedy double subspaces coordinate descent method for solving linear least-squares problems. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000893.

Karmakar:2023:MNN

[1750] P. Karmakar and S. Das. Modeling non-Newtonian magnetized blood circulation with tri-nanoadditives in a charged artery. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000911.

Li:2023:SIA

[1751] Danyang Li and Quanling Deng. Soft IsoGeometric analysis of the bound states of a quantum three-body problem in 1D. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323000923.

Bawazeer:2023:UFS

[1752] Saleh A. Bawazeer, Saleh S. Baakeem, and A. A. Mohamad. A unified forcing scheme for the single relaxation lattice Boltzmann method. Journal of Computational Science, 70:??, June 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000947.

Cholodowicz:2023:SRN

[1753] Ewelina Cholodowicz and Przemyslaw Orlowski. Switching robust neural network control of perishable inventory with fixed shelf life products under time-varying uncertain demand. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000959.

Jaworska:2023:CAN

[1754] Irena Jaworska. Computational aspects of nonlinear and multiscale analyses by the multipoint meshless FDM. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000960.

Sahoo:2023:NES

[1755] Nirupam Sahoo and Randhir Singh. A new efficient semi-numerical method with a convergence control parameter for Lane–Emden–Fowler boundary value problem. Journal of Computational Science, 70:??, June 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001011.

Anonymous:2023:EBe

[1756] Anonymous. Editorial Board. Journal of Computational Science, 70:??, June 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S1877750323001278.

Anonymous:2023:Jb

[1757] Anonymous. June 2023. Journal of Computational Science, 70:??, June 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Bernholdt:2023:EFS

[1758] David E. Bernholdt, Mathieu Doucet, William F. Godoy, Addi Malviya-Thakur, and Gregory R. Watson. Experiential findings for sustainable software ecosystems to support experimental and observational science. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323000935.

Lenz:2023:CAR

[1759] David Lenz, Raine Yeh, Vijay Mahadevan, Iulian Grindeanu, and Tom Peterka. Customizable adaptive regularization techniques for B-spline modeling. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323000972.

Liu:2023:JWE

[1760] Hua-Yu Liu, Xiao-Wei Gao, Gui-Yong Zhang, and Kai Yang. Jacobian weighted element differential method for solid mechanics. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000984.

Campos:2023:PCE

[1761] J. O. Campos, R. M. Guedes, Y. B.

Werneck, L. P. S. Barra, R. W. dos Santos, and B. M. Rocha. Polynomial chaos expansion surrogate modeling of passive cardiac mechanics using the Holzapfel–Ogden constitutive model. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323000996.

Covaci:2023:ESS

[1762] Florina Livia Covaci. Enabling the smart supply chain ecosystems: a multi-parameter decentralized model for Supply Chain 5.0. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032300100X.

Jimenez-Ruiz:2023:UGS

[1763] Alberto Jiménez-Ruiz, Gerardo Fernández-Escribano, Miguel Cañas-Carretón, and José L. Sánchez. Using GPUs to simulate photovoltaic power plants: Special cases of performance loss. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323001023.

Easaw:2023:ECM

[1764] Nikhil Easaw, Woo Seok Lee, Prashant Singh Lohiya, Sarika Jalan, and Priodyuti Pradhan. Estimation of correlation matrices from limited time series data using machine learning. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001138.

Burkhart:2023:NRL

[1765] Michael C. Burkhart and Gabriel Ruiz. Neuroevolutionary representations for learning heterogeneous treatment effects. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032300114X.

Xue:2023:BTP

[1766] Ruixin Xue, Nafei Zhu, Jingsha He, and Lin He. Bitcoin transaction pattern recognition based on semisupervised learning. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001151.

Cafiero:2023:DDH

[1767] Florian Cafiero. Datafying diplomacy: How to enable the computational analysis and support of international negotiations. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001163.

Lytaev:2023:RND

[1768] Mikhail S. Lytaev. Reducing the numerical dispersion of the one-way Helmholtz equation via the differential evolution method. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001175.

Galam:2023:DAC

[1769] Serge Galam. The dynamics of alliances: the case of Ukraine and Russia. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001187.

Tanade:2023:EMP

[1770] Cyrus Tanade, Sarah Putney, and Amanda Randles. Establishing massively parallel models to examine the influence of cell heterogeneity on tumor growth. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001199.

Roychowdhury:2023:EMQ

[1771] Sayan Roychowdhury, Erik W. Draeger, and Amanda Randles. Establishing metrics to quantify spatial similarity in spherical and red blood cell distributions. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001205.

Helbing:2023:DDP

[1772] Dirk Helbing, Sachit Mahajan, Regula Hänggli Fricker, Andrea Musso, Carina I. Hausladen, Cesare Carissimo, Dino Carpentras, Elisabeth Stockinger, Javier Argota Sanchez-Vaquerizo, Joshua C. Yang, Mark C. Ballandies, Marcin Korecki, Rohit K. Dubey, and Evangelos Pournaras. Democracy by design: Perspectives for digitally assisted, participatory upgrades of society. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001217.

Lang:2023:NUS

[1773] Jun Lang and Chongyang Lin. Nonuniformly sampled 2D NMR spectroscopy reconstruction based on Low Rank Hankel Matrix Fast Tri-Factorization and Non-convex Factorization. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001229.

Moges:2023:SMO

[1774] Demmelash Mollalign Moges, Berhanu Guta Wordofa, and Allen Rangia Mushi. Solving multi-objective linear fractional decentralized bi-level decisionmaking problems through compensatory intuitionistic fuzzy mathematical method. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001357.

Ali:2023:MFC

[1775] N. Ali and S. Saleem. Mean flow and convective instability analysis of the BEK and related rotating boundary layer flows subject to a uniform static magnetic field. Journal of Computational Science, 71:??, July 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001369.

Anonymous:2023:EBf

[1776] Anonymous. Editorial Board. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S1877750323001485.

Anonymous:2023:Jc

[1777] Anonymous. July 2023. Journal of Computational Science, 71:??, July 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Ahmadi:2023:ITL

[1778] Zahra Ahmadi, Hoang H. Nguyen, Zijian Zhang, Dmytro Bozhkov, Daniel Kudenko, Maria Jofre, Francesco Calderoni, Noa Cohen, and Yosef Solewicz. Inductive and transductive link prediction for criminal network analysis. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001230.

Chadefaux:2023:APR

[1779] Thomas Chadefaux. An automated pattern recognition system for conflict. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001345.

Li:2023:MSF

[1780] Jinhui Li, Yanle Li, Hao Yuan, Zinan Cheng, Yong Yu, Shahid Ghafoor, and Fangyi Li. Multi-stage finite element modeling of the deformation behavior during ultrasonic-assisted incremental sheet forming. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001370.

Li:2023:DLB

[1781] Lang Li and Yu Ou. A deep learning-based side channel attack model for different block ciphers. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001382.

Morishita:2023:DAC

[1782] Y. Morishita, S. Murakami, M. Yokoyama, and G. Ueno. Data assimilation and control system for adaptive model predictive control. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001394.

Szczesna:2023:CBT

[1783] Agnieszka Szczȩsna, Dariusz Rafał Augustyn, Henryk Josiński, Katarzyna Harȩżlak, Adam Świtoński, and Paweł Kasprowski. Chaotic biomedical time signal analysis via wavelet scattering transform. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001400.

Ossandon:2023:NNC

[1784] Sebastián Ossandón and Mauricio Barrientos. Neural network control design for solid composite materials. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001412.

Cheng:2023:EDT

[1785] Dongqin Cheng. Edge-disjoint trees passing through prescribed vertices in bubble-sort star graphs. Journal of Computational Science, 72:

??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001424.

Hou:2023:EDM

[1786] Wenguo Hou, Jing Xiong, and Zeyang Xia. Efficient dynamic modeling of soft tissue deformation using a WSCintegrated order reduction method. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001436.

Erkus:2023:NCA

[1787] Ekin Can Erkuş and Vilda Purutçuoğlu. A new collective anomaly detection approach using pitch frequency and dissimilarity: Pitchy anomaly detection (PAD). Journal of Computational Science, 72: ??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001448.

Khan:2023:TRI

[1788] Sohail A. Khan, T. Hayat, and A. Alsaedi. Thermal radiation impact on chemical reactive flow of micropolar nanomaterial subject to Brownian diffusion and thermophoresis phenomenon. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001540.

Furmanczyk:2023:CFS

[1789] Konrad Furmańczyk, Kacper Paczutkowski, Marcin Dudziński, and Diana Dziewa-Dawidczyk. Classification and feature selection methods based on fitting logistic regression to PU data. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001552.

Wernli:2023:FIC

[1790] Didier Wernli. Fostering interdisciplinary collaboration in computational diplomacy: a multi-layered network approach to improve our understanding of institutional complexity and effective governance design. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001564.

Zhou:2023:HRF

[1791] Jian Zhou, Zhenyu Wang, Chuanqi Li, Wei Wei, Shiming Wang, Danial Jahed Armaghani, and Kang Peng. Hybridized random forest with population-based optimization for predicting shear properties of rock fractures. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001576.

Urbanczyk:2023:SCC

[1792] Aleksandra Urbańczyk, Piotr Kipiński, Mateusz Nabywaniec, Leszek Rutkowski, Siang Yew Chong, Xin Yao, Krzysztof Boryczko, and Aleksander Byrski. Socio-cognitive caste-based optimization. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001588.

Strobl:2023:IPS

[1793] Eric V. Strobl and Thomas A. Lasko. Identifying patient-specific root causes with the heteroscedastic noise model. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S187775032300159X.

Pacis:2023:IPM

[1794] Felix James Pacis, Adrian Ambrus, Sergey Alyaev, Rasool Khosravanian, Tron Golder Kristiansen, and Tomasz Wiktorski. Improving predictive models for rate of penetration in real drilling operations through transfer learning. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001606.

Guggilam:2023:LDA

[1795] Sreelekha Guggilam, Varun Chandola, and Abani K. Patra. Large Deviations Anomaly Detection (LAD) for collection of multivariate time series data: Applications to COVID-19 data. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001618.

Kovalchuk:2023:CP

[1796] Sergey V. Kovalchuk, Clélia de Mulatier, Derek Groen, Maciej Paszyński, Valeria V. Krzhizhanovskaya, Jack Dongarra, and Peter M. A. Sloot. The computational planet. Journal of Computational Science, 72: ??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032300162X.

Goona:2023:DSS

[1797] Nithin Kumar Goona and Saidi Reddy Parne. Distributed source scheme for Poisson equation using finite element method. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001631.

Ouertani:2023:MCV

[1798] Nasreddine Ouertani, Hajer Ben-Romdhane, Issam Nouaouri, Hamid Allaoui, and Saoussen Krichen. A multi-compartment VRP model for the health care waste transportation problem. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001643.

Gupta:2023:DEM

[1799] Shubham Gupta and Rong Su. Diversity-enhanced modified sine cosine algorithm and its application in solving engineering design problems. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001655.

Kumar:2023:QDC

[1800] Sumit Kumar, Abhinav Sharma, and Ruchir Gupta. QoS driven cost-efficient resource allocation in edge computing: a distributed game theoretic approach. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001667.

Groen:2023:FSD

[1801] Derek Groen, Diana Suleimenova, Alireza Jahani, and Yani Xue. Facilitating simulation development for global challenge response and anticipation in a timely way. Journal of Computational Science, 72: ??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323001679.

Saha:2023:ENN

[1802] Nikita Saha and Randhir Singh. An efficient new numerical algorithm for solving Emden–Fowler pantograph differential equation using Laguerre polynomials. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001680.

Tong:2023:SAA

[1803] Hua Tong, Kuanren Qian, Eni Halilaj, and Yongjie Jessica Zhang. SRLassisted AFM: Generating planar unstructured quadrilateral meshes with supervised and reinforcement learningassisted advancing front method. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001692.

Russo:2023:SFV

[1804] Luís M. S. Russo. Searching for a Feedback Vertex Set with the linkcut tree. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001709.

Antoni:2023:IMI

[1805] N. Antoni. An inverse method for the identification of parameters in beam contact problems. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001710.

Cederman:2023:CAC

[1806] Lars-Erik Cederman and Luc Girardin. Computational approaches to conflict research from modeling and data to computational diplomacy. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001722.

Zhao:2023:HRH

[1807] Xiaotong Zhao, Jingli Du, and Zhihan Wang. HCS-R-HER: Hierarchical reinforcement learning based on cross subtasks rainbow hindsight experience replay. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001734.

Wang:2023:FFA

[1808] Jian Wang, Heming Xu, Junxiang Yang, and Junseok Kim. Fractal feature analysis based on phase transitions of the Allen–Cahn and Cahn–Hilliard equations. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001746.

Akbarpour:2023:ECV

[1809] Najmeh Akbarpour, Ebrahim Akbari, and Homayun Motameni. External clustering validity index based on extended similarity measures. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032300176X.

Wang:2023:MLU

[1810] Zewen Wang, Dexiu Hu, Jie Huang, Chuang Zhao, and Zeya Zhao. A method for localization using network and analysis of data error. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001771.

Stark:2023:OSP

[1811] Justina Stark and Ivo F. Sbalzarini. An open-source pipeline for solving continuous reaction–diffusion models in image-based geometries of porous media. Journal of Computational Science, 72:??, September 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001783.

Beerman:2023:FCE

[1812] Jack T. Beerman, Gwendal G. Beaumont, and Philippe J. Giabbanelli. A framework for the comparison of errors in agent-based models using machine learning. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001795.

Wang:2023:SFF

[1813] Xiao Wang, Shufan Zou, Yi Jiang, Laiping Zhang, and Xiaogang Deng. Swin-FlowNet: Flow field oriented optimization aided by a CNN and Swin-Transformer based model. Journal of Computational Science, 72: ??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323001813.

Anonymous:2023:EBg

[1814] Anonymous. Editorial Board. Journal of Computational Science, 72: ??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750323001874.

Anonymous:2023:S

[1815] Anonymous. September 2023. Journal of Computational Science, 72:??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Anonymous:2023:PN

[1816] Anonymous. In progress (November 2023). Journal of Computational Science, 73:??, November 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Hao:2023:IWP

[1817] Yue Hao, Patricio Clark Di Leoni, Olaf Marxen, Charles Meneveau, George Em Karniadakis, and Tamer A. Zaki. Instability-wave prediction in hypersonic boundary layers with physics-informed neural operators. Journal of Computational Science, 73:??, November 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001801.

Azis:2023:NIT

[1818] Mohammad Ivan Azis, Syamsuddin Toaha, Suharman Hamzah, and Imam Solekhudin. A numerical investigation of 2D transient heat conduction problems in anisotropic FGMs with time-dependent conductivity. Journal of Computational Science, 73:??, November 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750323001825.

Prodanov:2023:CES

[1819] Emil M. Prodanov. On the cubic equation with its Siebeck–Marden– Northshield triangle and the quartic equation with its tetrahedron. Journal of Computational Science, 73:??, November 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750323001837.

Valseth:2023:SPE

[1820] Eirik Valseth, Clint Dawson, and Edward Buskey. A study of the potential effects of deepening the Corpus Christi Ship Channel on hurricane storm surge. Journal of Computational Science, 73:??, November 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001989.

Pietrzyk:2023:AUS

[1821] Kyle Pietrzyk, Giovanna Bucci, Morad Behandish, and Ilenia Battiato. Automated upscaling via symbolic computing for thermal runaway analysis in Li-ion battery modules. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323001941.

Zhu:2023:MSP

[1822] Donglin Zhu, Siwei Wang, Jiaying Shen, Changjun Zhou, Taiyong Li, and Shaoqiang Yan. A multi-strategy particle swarm algorithm with exponential noise and fitness-distance balance method for low-altitude penetration in secure space. Journal of Computational Science, 74(??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323002090.

Deng:2023:NOS

[1823] Xi Deng. A new open-source library based on novel high-resolution structure-preserving convection schemes. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323002107.

Wei:2023:DAM

[1824] Yuying Wei, Adrian Wing-Keung Law, and Chun Yang. Data assimilation with missing data in nonstationary environments for probabilistic machine learning models. Journal of Computational Science, 74(??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002119.

Ghafari:2023:ETS

[1825] R. Ghafari and N. Mansouri. An efficient task scheduling in fog computing using improved artificial hummingbird algorithm. Journal of Computational Science, 74(??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002120.

Bozorgnia:2023:GBS

[1826] Farid Bozorgnia, Morteza Fotouhi, Avetik Arakelyan, and Abderrahim Elmoataz. Graph based semi-supervised learning using spatial segregation theory. Journal of Computational Science, 74(??):??, December 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002132.

Sajadi:2023:AHM

[1827] Seyedeh Zahra Sajadi, Seyed Mojtaba Sadjadi, and Mohammad Ali Zare Chahooki. AutoMF: a hybrid matrix factorization model with deep learning to select anti-viral drugs for Covid-19. Journal of Computational Science, 74(??):??, December 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002144.

Ciuprina:2023:DFW

[1828] Gabriela Ciuprina, Daniel Ioan, and Ruth V. Sabariego. Dual full-wave formulations with Radiant Electric Circuit Element boundary conditions — application to monopole antenna modeling. Journal of Computational Science, 74(??):??, December 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002156.

Carmesin:2023:HCT

[1829] Sarah Carmesin, David Woller, David Parker, Miroslav Kulich, and Masoumeh Mansouri. The Hamiltonian Cycle and Travelling Salesperson problems with traversal-dependent edge deletion. Journal of Computational Science, 74(??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002168.

Ghosh:2023:ASO

[1830] Bappa Ghosh and Jugal Mohapatra. Analysis of a second-order numerical scheme for time-fractional partial integro-differential equations with a weakly singular kernel. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S187775032300217X.

Saxton:2023:PPE

[1831] Harry Saxton, Torsten Schenkel, Ian Halliday, and Xu Xu. Personalised parameter estimation of the cardiovascular system: Leveraging data assimilation and sensitivity analysis. Journal of Computational Science, 74(??):??, December 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002181.

Montessori:2023:TSL

[1832] Andrea Montessori, Marco Lauricella, Adriano Tiribocchi, Mihir Durve, Michele La Rocca, Giorgio Amati, Fabio Bonaccorso, and Sauro Succi. Thread-safe lattice Boltzmann for highperformance computing on GPUs. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002259.

Gdawiec:2023:OHC

[1833] Krzysztof Gdawiec, Ioannis K. Argyros, Sania Qureshi, and Amanullah Soomro. An optimal homotopy continuation method: Convergence and visual analysis. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002260.

Mao:2023:DDC

[1834] Yiqian Mao, Shan Zhong, and Hujun Yin. DRLFluent: a distributed cosimulation framework coupling deep reinforcement learning with Ansys-Fluent on high-performance computing systems. Journal of Computational Science, 74(??):??, December 2023. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323002314.

Zhou:2023:FOH

[1835] Fengying Zhou and Xiaoyong Xu. Fractional-order hybrid functions combining simulated annealing algorithm for solving fractional pantograph differential equations. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323002326.

Anonymous:2023:EB

[1836] Anonymous. Editorial Board. Journal of Computational Science, 74 (??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750323002491.

Anonymous:2023:D

[1837] Anonymous. December 2023. Journal of Computational Science, 74(??):??, December 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Szymanska:2024:MMC

[1838] Zuzanna Szymańska, Mirosław Lachowicz, Nikolaos Sfakianakis, and Mark A. J. Chaplain. Mathematical modelling of cancer invasion: Phenotypic transitioning provides insight into multifocal foci formation. Journal of Computational Science, 75(??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002351.

Trynda:2024:SCM

[1839] Jan Trynda, Maciej Woźniak, and Sergio Rojas. A study of concurrent multi-frontal solvers for modern massively parallel architectures. Journal of Computational Science, 75(??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002442.

Sinha:2024:NAS

[1840] Vikash Kumar Sinha and Prashanth Maroju. Numerical algorithm for solving real-life application problems of Lane–Emden type equation. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002454.

Baran:2024:AIP

[1841] Klaudia Baran, Mateusz Sitko, and Lukasz Madej. Analysis of the influence of a pseudo-random number generator type on the kinetics of the cellular automata recrystallization model. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002533.

Xia:2024:ICL

[1842] Huihuang Xia and Marc Kamlah. An improved Coupled Level Set and Volume of Fluid (i-CLSVoF) framework for sessile droplet evaporation. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002557.

Ibrahim:2024:MAC

[1843] Mohamed Hamza Ibrahim and Rokia Missaoui. Mining actionable concepts in concept lattice using Interestingness Propagation. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002569.

Sochala:2024:DRU

[1844] Pierre Sochala, Christophe Chiaberge, Francis Claret, and Christophe Tournassat. Dimension reduction for uncertainty propagation and global sensitivity analyses of a cesium adsorption model. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002570.

Strand:2024:MMM

[1845] Erik Strand, Filippos Tourlomousis, and Neil Gershenfeld. Mesoscale material modeling with memoryless isotropic point particles. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002582.

Toledo:2024:ECP

[1846] F. Javier Toledo, Vicente Galiano, Victoria Herranz, and Jose M. Blanes. Efficient computation of the photovoltaic single-diode model curve by means of a piecewise linear self-adaptive representation. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002594.

Garai:2024:CON

[1847] Gobinda Garai and Bankim C. Mandal. On the convergence of overlapping and non-overlapping Schwarz methods for the Cahn–Hilliard equation. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002600.

Abed-alguni:2024:IIB

[1848] Bilal H. Abed-alguni and Saqer Hamzeh AL-Jarah. IBJA: an improved binary DJaya algorithm for feature selection. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002612.

Kishi:2024:INA

[1849] Shohei Kishi, Masashi Yamakawa, Ayato Takii, Shinichi Asao, Seiichi Takeuchi, and Minsuok Kim. Impact of neck angle variation on particle and virus-laden droplet transport from lung to lung using eighth-generation airway model. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002624.

Kosztyan:2024:PIF

[1850] Zsolt T. Kosztyán and Gergely L. Novák. Project indicators and flexible project structure generators. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002636.

Rehman:2024:CAF

[1851] Ziad Ur Rehman, Salah Boulaaras, Rashid Jan, Imtiaz Ahmad, and Salma Bahramand. Computational analysis of financial system through non-integer derivative. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002648.

Saeternes:2024:APT

[1852] Erik Hide Sæternes, Andreas Thune, Alf Birger Rustad, Tor Skeie, and Xing Cai. Automated parameter tuning with accuracy control for efficient reservoir simulations. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S187775032300265X.

Dhama:2024:RSN

[1853] Soniya Dhama. A reliable scheme for nonlinear delay differential equations of pantograph-type. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002661.

Swain:2024:SAM

[1854] Sradharam Swain, Golam Mortuja Sarkar, and Bikash Sahoo. Stability analysis of MHD stagnation point flow of Casson fluid past a shrinking sheet in porous medium considering heat sink or source, thermal radiation and suction effects. Journal of Computational Science, 75(??):??, January 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002673.

Hussain:2024:EPD

[1855] Saddam Hussain and Rajesh Kumar. Elzaki projected differential transform method for multidimensional aggregation and combined aggregation-breakage equations. Journal of Computational Science, 75(??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000048.

Raul:2024:MAS

[1856] Vishal Raul and Leifur Leifsson. Multifidelity aerodynamic shape optimization for airfoil dynamic stall mitigation using manifold mapping. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000061.

Anonymous:2024:EBa

[1857] Anonymous. Editorial Board. Journal of Computational Science, 75 (??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000152.

Anonymous:2024:Ja

[1858] Anonymous. January 2024. Journal of Computational Science, 75(??):??, January 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Lin:2024:CIB

[1859] S. Lin, A. Liu, J. Wang, and X. Kong. Corrigendum to “An intelligence-based hybrid PSO-SA for mobile robot path planning in warehouse” J. Comput. Sci.

67 (2023) 101938. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323002430.

Vecchi:2024:SEI

[1860] Edoardo Vecchi, Juraj Kardoš, Malik Lechekhab, Andreas Wächter, and Olaf Schenk. Structure-exploiting interiorpoint solver for high-dimensional entropy-sparsified regression learning. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000012.

Farooq:2024:CSA

[1861] Umar Farooq, Chemseddine Maatki, Karim Kriaa, Bilel Hadrich, Muhammad Imran, Sobia Noreen, Hassan Waqas, and Ali Akgül. Characteristics of sodium alginate-based hybrid nanofluid and Darcy–Forchheimer flow induced by stretching surface with thermal radiation and Cattaneo– Christov heat flux model. Journal of Computational Science, 76 (??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000024.

Asgari:2024:DDR

[1862] Somayeh Danesh Asgari, Emran Mohammadi, Ahmad Makui, and Mostafa Jafari. Data-driven robust optimization based on position-regulated support vector clustering. Journal of Computational Science, 76 (??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000036.

Hao:2024:MIP

[1863] Xiaochen Hao, Xinqiang Wang, Jinbo Liu, Xing Wang, Zhipeng Zhang, and Yukun Ji. A multi-indicator prediction method for NOx emission concentration and ammonia escape value for cement calciner system. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032400005X.

Cengizci:2024:SFE

[1864] Süleyman Cengizci, Ömür Uğur, and Srinivasan Natesan. A stabilized finite element formulation with shock-capturing for solving advectiondominated convection–diffusion equations having time-fractional derivatives. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000073.

Arenillas:2024:DFD

[1865] Julián Arenillas, Jamie Wubben, Enrique Hernández-Orallo, and Carlos T. Calafate. D-FFP: a directional force field protocol for the efficient management of aerial conflicts between UAVs. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000085.

Skotniczny:2024:CDI

[1866] Marcin Skotniczny, Anna Paszyńska, Sergio Rojas, and Maciej Paszyński. Complexity of direct and iterative solvers on space–time formulations and time-marching schemes for hrefined grids towards singularities. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000097.

Peon-Escalante:2024:DNF

[1867] R. Peón-Escalante, K. B. Cantún-Avila, O. Carvente, A. Espinosa-Romero, and F. Peñuñuri. A dual number formulation to efficiently compute higher order directional derivatives. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000103.

Szarmach:2024:FDD

[1868] Marta Szarmach and Ireneusz Czarnowski.

A framework for damage detection in AIS data based on clustering and multi-label classification. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000115.

Castrignano:2024:MCR

[1869] Alberto Castrignanò, Roberta Bardini, Alessandro Savino, and Stefano Di Carlo. A methodology combining reinforcement learning and simulation to optimize the in silico culture of epithelial sheets. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032400019X.

Friedemann:2024:DLP

[1870] Sebastian Friedemann, Kai Keller, Yen-Sen Lu, Bruno Raffin, and Leonardo Bautista-Gomez. Dynamic load/propagate/store for data assimilation with particle filters on supercomputers. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032400022X.

Malhotra:2024:EVM

[1871] Ruchika Malhotra and Shweta Meena. Empirical validation of machine learning techniques for heterogeneous crossproject change prediction and withinproject change prediction. Journal of Computational Science, 76 (??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000231.

Dehingia:2024:DNP

[1872] Kaushik Dehingia, Salah Boulaaras, and Suman Gogoi. On the dynamics of a nutrient–plankton system with Caputo and Caputo–Fabrizio fractional operators. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000255.

Lopez-Oriona:2024:ACT

[1873] Ángel López-Oriona and José A. Vilar. Analyzing categorical time series with the R package ctsfeatures. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000267.

Amjad:2024:PIM

[1874] Maria Amjad and Mujeeb ur Rehman.

A product integration method for numerical solutions of ϕ-fractional differential equations. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000279.

Pompa:2024:PMM

[1875] Marcello Pompa, Andrea De Gaetano, Alessandro Borri, Antonella Farsetti, Simona Nanni, Alfredo Pontecorvi, and Simona Panunzi. A physiological mathematical model of the human thyroid. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000292.

Tan:2024:ECI

[1876] Rui Zhen Tan. The effects of cellular interactions on the sizes, composition, and dynamics of migrating cancer cell clusters. Journal of Computational Science, 76(??):??, March 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000309.

Lee:2024:AMS

[1877] Dongsun Lee and Chaeyoung Lee. Assessment of morphological similarities for the conservative Allen– Cahn and Cahn–Hilliard equations. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000334.

Anonymous:2024:EBb

[1878] Anonymous. Editorial Board. Journal of Computational Science, 76 (??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000413.

Anonymous:2024:M

[1879] Anonymous. March 2024. Journal of Computational Science, 76(??):??, March 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Chopard:2024:PMS

[1880] Bastien Chopard, Stephan Davishofer, Dirk Helbing, Nicolas Levrat, and Peter Sloot. Preface — from the modeling of social behavior to computational diplomacy. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750323001758.

Chen:2024:NDD

[1881] Bo Chen, Rui Tong, Xiue Gao, and Yufeng Chen. A novel dynamic decision-making method: Addressing the complexity of attribute weight and time weight. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000218.

Los:2024:SWG

[1882] Marcin Łoś, Leszek Siwik, Maciej Woźniak, Dominik Gryboś, Paweł Maczuga, Albert Oliver-Serra, Jacek Leszczyński, and Maciej Paszyński. Shock waves generators: From prevention of hail storms to reduction of the smog in urban areas — experimental verification and numerical simulations. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000310.

Wang:2024:FRM

[1883] Jing Wang, Junyang Li, Yankui Song, Yaoyao Tuo, and Chengguo Liu. FC-RRT*: a modified RRT* with rapid convergence in complex environments. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000322.

Singh:2024:RWG

[1884] Jasbir Singh, Naresh Kumar, and Ram Jiwari. A robust weak Galerkin finite element method for two parameter singularly perturbed parabolic problems on nonuniform meshes. Journal of Computational Science, 77(??):??, April 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000346.

Niemann:2024:MOP

[1885] Jan-Hendrik Niemann, Samuel Uram, Sarah Wolf, Nataša Djurdjevac Conrad, and Martin Weiser. Multilevel optimization for policy design with agent-based epidemic models. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000358.

Shahni:2024:NCA

[1886] Julee Shahni and Randhir Singh. A novel collocation approach using Chebyshev wavelets for solving fourth-order Emden–Fowler type equations. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032400036X.

Sadek:2024:GBM

[1887] Lakhlifa Sadek, Said Ounamane, Bouchra Abouzaid, and El Mostafa Sadek. The Galerkin Bell method to solve the fractional optimal control problems with inequality constraints. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000371.

Yang:2024:ACE

[1888] Junxiang Yang, Seungyoon Kang, Soobin Kwak, and Junseok Kim. The Allen–Cahn equation with a spacedependent mobility and a source term for general motion by mean curvature. Journal of Computational Science, 77(??):??, April 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000450.

Boskovic:2024:PSA

[1889] Borko Bošković, Jana Herzog, and Janez Brest. Parallel self-avoiding walks for a low-autocorrelation binary sequences problem. Journal of Computational Science, 77(??):??, April 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S187775032400053X.

Anonymous:2024:EBc

[1890] Anonymous. Editorial Board. Journal of Computational Science, 77 (??):??, April 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000668.

Anonymous:2024:A

[1891] Anonymous. April 2024. Journal of Computational Science, 77(??):??, April 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Khan:2024:NSH

[1892] Zeeshan Khan, Esraa N. Thabet, Shazia Habib, A. M. Abd-Alla, F. S. Bayones, F. M. Alharbi, and Afaf S. Alwabli. Numerical study of hydromagnetic bioconvection flow of micropolar nanofluid past an inclined stretching sheet in a porous medium with gyrotactic microorganism. Journal of Computational Science, 78(??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000498.

Shoaib:2024:MRN

[1893] Muhammad Shoaib and Tariq Javed. MHD radiative natural convective heat transfer enhancement for Casson nanofluid flow on a horizontal circular cylinder. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000504.

Torok:2024:AIG

[1894] Árpád Török and Gábor Pauer. Analyzing the impact of grid structures on traffic flow optimization in autonomous transport systems. Journal of Computational Science, 78(??):??, June 2024.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000516.

Cinaglia:2024:MBN

[1895] Pietro Cinaglia. Multilayer biological network alignment based on similarity computation via Graph Neural Networks. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000528.

Xu:2024:PCN

[1896] Liang Xu, Ziyan Liu, Yiwei Feng, and Tiegang Liu. Physics-constrained neural networks as multi-material Riemann solvers for compressible two-gas simulations. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000541.

Goswami:2024:SRM

[1897] Antriksh Goswami and Ruchir Gupta.

Stable reputation management system in P2P networks: an evolutionary perspective. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000553.

Marquardt:2024:NPD

[1898] Jan E. Marquardt, Nicolas Hafen, and Mathias J. Krause. A novel particle decomposition scheme to improve parallel performance of fully resolved particulate flow simulations. Journal of Computational Science, 78(??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000565.

Safaei:2024:AMR

[1899] F. Safaei, Z. Hashemi, and M. M. Emadi Kouchak. Analytical modeling of robustness and stochastic resilience of temporal small-world complex networks. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000577.

Korhonen:2024:GAC

[1900] Marko Korhonen, Alpo Laitinen, Gizem Ersavas Isitman, Jose L. Jimenez, and Ville Vuorinen. A GPU-accelerated computational fluid dynamics solver for assessing sheardriven indoor airflow and virus transmission by scale-resolved simulations. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000589.

Li:2024:TLE

[1901] Chuanqi Li, Jian Zhou, and Kun Du. Towards lightweight excavation: Machine learning exploration of rock size distribution prediction after tunnel blasting. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000590.

Singh:2024:PEL

[1902] Harbinder Singh, Simrandeep Singh, Jaspinder Kaur, Atipriya Sharma, Amit Gupta, and H. Singh. Performance evaluation of linear antenna array using quasi opposition modified particle swarm algorithm. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000607.

Mahadevan:2024:AMF

[1903] Vijay Mahadevan, David Lenz, Iulian Grindeanu, and Thomas Peterka. Accelerating multivariate functional approximation computation with domain decomposition techniques. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000619.

Pereira:2024:SAA

[1904] Ricardo Cardoso Pereira, Pedro Henriques Abreu, and Pedro Pereira Rodrigues. Siamese autoencoder architecture for the imputation of data missing not at random. Journal of Computational Science, 78(??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000620.

Tong:2024:HHO

[1905] Hua Tong, Eni Halilaj, and Yongjie Jessica Zhang. HybridOctree_{H}ex : Hybridoctree − basedadaptiveall − hexahedralmeshgenerationwithJacobiancontrol.JournalofComputationalScience, 78 (??) :??, June2024.CODEN????ISSN1877−

                                        7503(print), 1877−7511(electronic).URLhttpwwwsciencedirectcomsciencearticlepiiS1877750324000711.

Ke:2024:SRE

[1906] Lihua Ke, Menglong Wu, Yicheng Ye, Nanyan Hu, and Yaoyao Meng. Stability risk early warning for mine goaf: Based on D-RES and asymmetric fuzzy connection cloud model. Journal of Computational Science, 78(??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000723.

Trull:2024:PMF

[1907] Oscar Trull, J. Carlos García-Díaz, and A. Peiró-Signes. msf, a forecasting library to predict short-term electricity demand based on multiple seasonal time series. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000735.

Cunha:2024:GAC

[1908] Luís Cunha, Leandro Santiago, and Felipe Souza. Graph admissibility: Case generation and analysis by learning models. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000747.

Lei:2024:WII

[1909] Mingli Lei, Lirong Liu, and Aldo Ramirez-Arellano. Weighted information index mining of key nodes through the perspective of evidential distance. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000759.

Aref:2024:AMM

[1910] Samin Aref and Mahdi Mostajabdaveh.

Analyzing modularity maximization in approximation, heuristic, and graph neural network algorithms for community detection. Journal of Computational Science, 78(??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000760.

Elaswad:2024:TDP

[1911] Ramy Elaswad, Naser El-Sheimy, and Abdulmajeed Mohamad. Trajectory dynamics of particles in accelerated toroidal pipe: a computational study using CFD-DPM simulations. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000784.

Zhang:2024:PNM

[1912] Qingyong Zhang, Shiyang Zhou, Bingrong Xu, Zhanwei Shen, and Wanfeng Chang. PSGformer: a novel multivariate net load forecasting model for the smart grid. Journal of Computational Science, 78(??):??, June 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000814.

Anonymous:2024:EBd

[1913] Anonymous. Editorial Board. Journal of Computational Science, 78 (??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000929.

Anonymous:2024:Jb

[1914] Anonymous. June 2024. Journal of Computational Science, 78(??):??, June 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

Driscoll:2024:PSA

[1915] Simon Driscoll, Alberto Carrassi, Julien Brajard, Laurent Bertino, Marc Bocquet, and Einar Örn Ólason. Parameter sensitivity analysis of a sea ice melt pond parametrisation and its emulation using neural networks. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324000243.

Kolar-Pozun:2024:SOR

[1916] Andrej Kolar-Požun, Mitja Jančič, Miha Rot, and Gregor Kosec. Some observations regarding the RBF-FD approximation accuracy dependence on stencil size. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000772.

El-Mesady:2024:NDO

[1917] A. El-Mesady, A. A. Elsadany, A. M. S.

Mahdy, and Amr Elsonbaty. Nonlinear dynamics and optimal control strategies of a novel fractional-order lumpy skin disease model. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000796.

Saxton:2024:AIP

[1918] Harry Saxton, Xu Xu, Torsten Schenkel, and Ian Halliday. Assessing input parameter hyperspace and parameter identifiability in a cardiovascular system model via sensitivity analysis. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000802.

Vidal:2024:MAQ

[1919] Veronica Vidal, Ana Cortés, Alba Badia, and Gara Villalba. Modeling air quality at urban scale in the city of Barcelona: a matter of green resolution. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000826.

Aggarwal:2024:DCP

[1920] Manu Aggarwal and Vipul Periwal. Dory: Computation of persistence diagrams up to dimension two for Vietoris–Rips filtrations of large data sets. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000838.

Maier:2024:OES

[1921] Benjamin Maier, Dominik Göddeke, Felix Huber, Thomas Klotz, Oliver Röhrle, and Miriam Schulte. OpenDiHu: an efficient and scalable framework for biophysical simulations of the neuromuscular system. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S187775032400084X.

Nawrocki:2024:FDO

[1922] Piotr Nawrocki and Mateusz Smendowski. FinOps-driven optimization of cloud resource usage for highperformance computing using machine learning. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000851.

Milad:2024:ESD

[1923] Reham Milad, Hosam Hegazy, Jiansong Zhang, Mohamed Korashy, Said AboulHaggag, and Ahmed Ebid. Estimating the stress distribution within MERO joint using (FEM-ANN) hybrid technique. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000875.

Michopoulos:2024:BHC

[1924] J. G. Michopoulos, A. P. Iliopoulos, C. Farhat, P. Avery, G. Daeninck, J. C. Steuben, and N. A. Apetre. Bottom-up hierarchical and categorical metacomputing for automating composition and deployment of directly computable multiphysics models. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000887.

Rot:2024:SDN

[1925] Miha Rot, Mitja Jančič, and Gregor Kosec. Spatially dependent node regularity in meshless approximation of partial differential equations. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324000991.

Marques:2024:SGE

[1926] Augusto Delavald Marques, Lara Werncke Vieira, Paulo Smith Schneider, and Felipe A. C. Viana. Steam generator efficiency optimization under uncertainty through multi-fidelity modeling. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001005.

Starzec:2024:ACO

[1927] Grażyna Starzec, Mateusz Starzec, Leszek Rutkowski, Marek Kisiel-Dorohinicki, and Aleksander Byrski. Ant colony optimization using twodimensional pheromone for singleobjective transport problems. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1877750324001017.

Jastrzab:2024:RMI

[1928] Tomasz Jastrzab, Frédéric Lardeux, and Eric Monfroy. Robust models to infer flexible nondeterministic finite automata. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001029.

Zhang:2024:WVC

[1929] Xinyu Zhang, Yunjun Zhou, Wenyu Zhang, Menggang Kou, Bingyan Li, Ying Dai, and Chenglin Yang. Water vapor content prediction based on neural network model selection and optimal fusion. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001030.

Le:2024:NBF

[1930] T. T. V. Le, N. Mai-Duy, K. Le-Cao, S. Bordas, D. P. Vu, and T. Tran-Cong. A non-boundary-fitted-grid method, based on compact integrated-RBF approximations, for solving differential problems in multiply-connected domains. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001042.

Liu:2024:NHG

[1931] Yunmou Liu, Hui Du, Zhuogen Li, and Pu Chen. A new hybrid GPU-CPU sparse LDL^{T} factorization algorithm with GPU and CPU factorizing concurrently. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001054.

Akanji:2024:PPP

[1932] Lateef T. Akanji. A pressure-projection pre-conditioning multi-fractional-step method for Navier–Stokes flow in porous media. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001066.

Gryko:2024:NES

[1933] Anita Gryko and Piotr Prochor. Numerical evaluation of scaffolds as a method to restore continuity of a long bone. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324001078.

Yan:2024:EUG

[1934] Zhangmin Yan. Establishment of urban green corridor network based on neural network and landscape ecological security. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775032400108X.

Munoz:2024:CGO

[1935] Felipe Muñoz, Rafael Asenjo, Angeles Navarro, and J. Carlos Cabaleiro. CPU and GPU oriented optimizations for LiDAR data processing. Journal of Computational Science, 79(??):??, July 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750324001108.

Gao:2024:PIG

[1936] Ruisong Gao, Min Yang, and Jin Zhang. Physics-informed generatorencoder adversarial networks with latent space matching for stochastic differential equations. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S187775032400111X.

Anonymous:2024:EBe

[1937] Anonymous. Editorial Board. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1877750324001261.

Anonymous:2024:Jc

[1938] Anonymous. July 2024. Journal of Computational Science, 79(??):??, July 2024. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic).